A Complete Bibliography of Combinatorics, Probability and Computing

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

16 October 2023 Version 2.32

Title word cross-reference

(1, 2) [HJ94]. (5, 5) [CH18]. (7, 4) [Sol15]. (Δ + 1) [BBP21]. (k + 1) [oCM12]. (k − 1) [LW07]. (n, d, λ) [Ofe07]. 0 [BKHL99, Hyd23, VZ06]. 0 < q < 1 [Lar21]. 1 [BB12, BKHL99, RM92, VZ06]. 1324 [Bón14]. 2 [BLS^{+}12b, DH00, FS12, FK06, Für12, HZ17, IK09, Mát07, Nob06, Saa96, SS06b, Wu97, Yus07]. 2/3 [KLS^{+}13]. $23.00 [Sco04]. 2^{k} [NV05]. 2 × 2 [Had98, Had99]. 3 [AHL21, CP16, COMV09, EF95, FRV13, FRMZ21, GH17, GS11, HLP^{+}09, HN18, Mar13, Mat03, PRS05, RW01, RRS06, Wag06a, Wu98]. 4 [BHL^{+}15, FR01, KO04a, Mar13, Nag17, SW07a]. $49.95 [Bár04, Lea04]. 5 [BDLP20]. $65.00 [Sco04]. 7 [Bri98]. 8 [Bri98]. _{3,3} [KM12]. b [AF04]. B_{2}[g] [CV08]. B_{h} [DKL^{+}16]. β [GA19]. C_{2k} [GMT20]. C_{2k+1} [GL12b]. C_{4} [Pic11]. C_{6}^{3} [GH17]. C_{k} [EM22]. d [BCOK14a, BCOK15, BKW08, BCH22, KN03, Siv14]. Δ [FK09, KRT99]. Δ/2⌉ [MR95]. ℓ [HHWY23]. ϵ [ABFK92]. F [DP23, PR17a, Wag00]. F_{7}^{−} [Mec01]. F_{7} [Mec01]. F_{p}^{n} [Ngu11]. F_{q}[x] [Han93]. G [BFRT01, FR15, JT08b].

G(n, p) [Wol11]. G_{n,p} [CO07b]. GF (q) [Lef05]. H [AY93, AKS05, FJP13]. k [Aig05, COGL07, COW18, CF95, CDD03, DM23, DHS16, Gao14, GGH12, GJ03, HZ05, Han15, JS20, KKY06, LMMW09, LMY23, Nik01b, PV13, PS16, Ras19, Rio08, Sha11, Sub00, XY11, oCM12]. K − 4 [HLTZ17]. k = r + 1 [EGM^{+}20]. k = r + 2 [EGM^{+}20]. K^{r} [Kri94]. K^{s} [Kri94]. K_{2,s} [Dob02a]. K_{2m,2m} [DL14]. K_{4} [KM08]. K_{5} [Tho96a]. K_{N} [Fel17, CH18, Bal01b]. K_{p} [EHS^{+}94]. K_{r+1} [BCL^{+}21]. K_{r+2} [KO04b]. K_{t} [Ste22]. l [FHR09]. L^{p} [CLS13]. log n/n [Jan99]. M [Pan04, DL14, FD97, Gow96, KN14a, Sta02a]. Z [Gla16, Mun14]. Z^{2} [BCG^{+}19]. Z^{d} [ABC^{+}21, GK04]. Z_{N} [Pon13a]. Z_{p} [Pon13b]. N [BH93, GP97, Lem22, BKW08, FD97, GGX16, Hal93, KS02, KS04, Mad94, Mad95, Rei02, VS06]. n(6, 6) [BMR99]. n + D − 1 [Ham08a]. n^{−1} [VS06]. N^{2} [Heg96]. O [DKR18]. P_{7} [Mec01]. p ↑ 1 [HKNN23]. q [CST97, FHR09]. r [AH17, BCL^{+}21, BL04b, CKTV11, CFMR96, FT11, GL21a, Jin92]. r(K_{3}, G) [Bri98]. r + ε [KO04b]. s [CFMR96, FK17, GL21a]. SQS(2v) [GMT94]. t [GSS96, KK14, KM10, Tok13]. U_{2,5} [Mec01]. x + y = z^{2} [KS06]. Z^{n} [VS06]. Z_{2}^{n} [EZ12].

-Almost-Periodicity [CLS13]. -ary [FD97, Pan04]. -Assignment [LW07]. -Avoiding [Bón14]. -Bases [FK06]. -Bonds [FK09]. -Cancellative [Für12]. -Cliques [Nik01b, PV13]. -Colour [GS11]. -colourable [Sub00]. -Coloured [BLS^{+}12b, KN14a]. -Colouring [CH18, COMV09, Ras19]. -colourings [BBP21]. -Conjecture [Sol15]. -Connected [RW01, Wag06a, Wu97, Wu98, Hal93, Mad94]. -Core [Rio08]. -Cores [oCM12]. -covering [GSS96]. -critical [LMY23]. -Cross [FT11]. -Cross-Intersecting [FK17]. -CSP [SS06b]. -cube [VS06]. -Cubes [Rei02]. -Cycle [FR01, KO04a, Mar13]. -Cycles [NV05, BDLP20, DP23]. -degenerate [KN03]. -Dimensional [Siv14]. -Edge [Nag17]. -Edge-Coloured [LMMW09, Saa96]. -edge-connected [Mad95]. -Equal [Aig05]. -Extensions [HZ05]. -factorizations [HJ94]. -Factors [FS12, oCM12, GH17, HLTZ17, AY93, EM22]. -Free [AKS05, GL12b, PR17a, Pic11, BCL^{+}21, GMT20, Kri94]. -Graph [Mar13]. -Graphs [CKTV11, FRV13, GHZ19, HN18, FRMZ21]. -Hermite [GA19]. -Improper [KM10]. -in [RM92]. -increasing [GL21a]. -Independence [AF04, HHWY23, EHS^{+}94]. -Independent [BB12, FR15]. -Intersecting [Tok13, BFRT01]. -Intersection [JT08b]. -linked [KKY06, FJP13]. -Matrices [VZ06]. -nets [ABFK92]. -out [RM92]. -Partite [AH17, CKTV11, GHZ19, BCL^{+}21, Jin92]. -Partitions [XY11]. -Polymatroid [Nob06]. -Regular [Gao14, LMMW09, SW07a, CFMR96]. -SAT [COGL07, DM23, GJ03, COW18, IK09, EF95]. -Sets [Sha11, BL04b, DKL^{+}16]. -statement [Hyd23]. -Stirling [CST97]. -subdivisions [Tho96a]. -Subsystems [KRT99]. -Sums [GGX16, GGH12]. -trees [DHS16, JS20]. -tuples [GL21a]. -Uniform [BCOK14a, BCOK15, CP16, GH17, GS11, Han15, HLP^{+}09, RRS06, AHL21, PRS05, CFMR96].

-Vectors [VZ06, Wag00]. -Wide [FHR09]. -wise [Gow96]. -XOR-formulas [CDD03]. -XORSAT [PS16].

0-521-52903-4 [Sco04]. 0-521-82151-7 [Sco04].

1-cross-intersecting [FGK23]. 139.05/$159.00 [Bov04]. 14-Cycle-Free [FÖ09]. 15.99/$24.99 [Ber06]. 152pp [Ber06]. 170pp [Bár06]. 1995 [Ano99].

2-edge-coloured [BKLL22]. 2011 [SSS^{+}11]. 257pp [Big07, Ode06]. 29.00/$49.95/38 [Ode06].

3 [Lea04]. 3-540-00845-4 [Ste05b]. 3-540-66313-4 [Lea04]. 3-colouring [DEM^{+}22]. 3-Connected [FP11]. 3-Uniform [Gow06, Sch23]. 32.50 [Lea04]. 377pp [Geo08].

4 [Lea04]. 4-state [dBBR23]. 40.00/$70.00 [Ber06]. 455pp [Moh06]. 49.95 [Bár04].

50.00/$95.00 [Big07].

6-Cycles [WL01]. 61.50/$99.00/79.95 [Dud07]. 61.50/EU [McM05]. 69 [Hol04]. 69.00/$125.00/89.95 [Geo08].

73.00/$109.00 [Ste05b]. 73.00/$109.00/94.95 [Moh06].

85.55/$79.97 [McM05]. 85d [Die95].

94.95/£ [Ste05b]. 99m [Had99].

Abelian [BNMS15, GH98, GGX16, GGH12, HLW19, New23, Peb04]. Absence [Bor06]. Academic [Big07]. Achlioptas [RW16c]. Active [AHW16]. Acyclic [Bor04, Pik10, RT05]. Adaptive [CO10]. Addability [MW17]. Addable [ABMR12]. Adding [BJJ98, Ham98, KN16, Baj16]. Addition [HHP07]. Additive [Ars21, BH93, Cha07, DLM04, Wag15, RW19]. adjacency [God92]. adsorption [FP92]. Advantage [AHK10]. Adversarial [BKT13, FFV07a]. Affine [GR98, HP97, HS04, Mun14]. After [BKT13]. against [FKR^{+}03]. Albert [Mór05]. Alexander [Moh06]. Algebra [Pik01]. Algebraic [BD15, Big07, FHR09, HS98, PD17]. algebras [ESS94]. Algorithm [DV04, DFV97, DKM^{+}15, KP98, Vu18, Ann94, AST93, DFK^{+}93, FLZ19, FRS95, GMPP22, Pit93]. Algorithmic [BLGN^{+}02, FLZ17, FLZ18, Van01, Voi00]. Algorithms [CTV15, DP06, DS04, FFV07b, GG10, HS00a, LS18a, PV19, PWZ97, Sub00, Sub07, DHJM21, GJ03, NT93]. Allocation [SV03]. Allocations [DLM04, Mit99]. Almost [BKT13, CC98, CLS13, DL14, Ell11, FFO13, KM08, KMS19, MR95, Mil17, Pel94, PS92, Edw95, Gow96, Gub96, SC95]. Almost-Equidistant [KMS19]. almost-planar [Gub96]. almost-transitive [SC95]. Alon [Nao12]. Along [Can09]. Alspach [Bal01a]. Alternating [Mon14, Saa96]. alternation [Con20]. Alternative [DP15]. Amalgamated [DH94]. Amenable [BPP06]. Amidst [DJ13]. among [MS95, Mat95, Mat99]. Anagram [KLS18]. Anagram-Free [KLS18]. analogue [Law93]. analogues [CG22b]. Analyse [LCF14]. Analysing [COMV09]. Analysis [ACG^{+}04, ADH^{+}19, BD08, CP20, CTV15, DV04, DFV97, DO14, DFP92, DS04, FRS95, GLWW15, GKW14, HK14, KP11, KP19, LS18a, MM14, Ode06, PV19, PWZ97, Ars21, LW21, LS95]. Analytic [Man02b]. And/Or [CFGG04]. Andrews [Ber06]. Angel [Mát07, Bow07]. Animals [MS13, GP23]. Annealing [NS01]. Answer [HHP07]. Ante [Csi92, Sta95]. Anti [HT16, AJT03, FJZ22, JW03, JLM^{+}03, dBBR23]. anti-concentration [FJZ22]. anti-ferromagnetic [dBBR23]. anti-Ramsey [AJT03, JW03, JLM^{+}03]. Anti-Triangles [HT16]. Antiferromagnetic [GSV16, DHJM21]. any [LM23]. Apart [SS07]. Appearances [CPV01]. Application [AM10, BLGN^{+}02, FM14, MW07, Fra21, Pud03]. Applications [Alo09b, Big07, FRV13, FP10, FP14, FLZ17, FG08, KS18, Lea04, Moh06, SV03, Sie98, BDH22, Esl21, Fil20, FLZ18]. applied [GJ03]. Approach [AvLR13, COMV09, EGPN16, Hol04, Kah01, LM19, LM01, Man02b, Mar01, Alo92, DT94, ISZ22, Lyn93]. approaches [Ing93]. Approximate [BFLW05, ABIST96, FZ22, LP20]. Approximately [FK14, GJ21b, Mil23]. Approximating [Bor04, CDD03, Tim12, DFK^{+}93, Ver95]. Approximation [Bor02, CPV01, Dub08, ER99, FFV07b, Mån00, WF11, Ann94, DHJM21, GJ03, NT93, Wel94]. Approximations [ABT06, Fuc08, Sta97, Sta99b, ZD08, FH23, GSS96]. Arbitrary [Ham11, Yus04, Con20]. Arboricity [AHHT10]. Arc [BGLS04, CTY07, HMS^{+}13, SV03, Hir97b]. Arc-sine [Hir97b]. Archaeology [BCL23]. Arcs [BJGY97, BJJ98, RT05]. Area [SRT10, The04, FH23]. areas [Tho92]. Argyros [Ode06]. Arising [GLWW15, HZ14]. Arithmetic [Can09, CLS13, Ngu11, Sol06, Ver00, JLM^{+}03, RSW93]. armed [FKR^{+}03]. arrangement [KR94]. Arrangements [Bár06]. Arratia [Hol04]. Arrays [ACÖ13, KS15, GSS96]. ary [FD97, Pan04]. Ascents [BKW08]. Ashkin [Far07]. Aspects [Van01, Voi00, CK93]. Assemblies [Sta97, Sta99b]. Assignment [AS02, LW07, Par04, Wäs11]. Assignments [COW18, EES03]. Associate [FR01]. Associated [ADH^{+}19, HKL^{+}20, VW93]. association [Bie19]. Asymmetric [DMS19, HK14, MNS20]. Asymmetry [BCGR97]. Asymptotic [BP12, BCOK14a, BCOK15, BHLB^{+}13, BMW14, Cha09, Coo96, GKW14, Had98, Lyo05, MR98a, MWW02, Mon18, PR17b, Had99, Jan95, LSM19]. asymptotical [Lis96]. Asymptotically [BK16, Kuz00, Thi16, Mór92]. Asymptotics [BD15, BP01, GHO17, HM15, MPT22, Mub02, PW04, Pim11, Sta99b, Wil15, GA19]. Attachment [AMR16, BBC^{+}05, DO14, Jan19].

Author [Ano92a, Ano93a, Ano94a, Ano95a, Ano96a]. Autocorrelation [Eri97]. Autocorrelations [Mer06]. Automata [BSU15]. Automorphism [Dob06, Leh13, Sch02]. Automorphisms [Bon95, FGLP00, Sco04]. Average [Alm06, DFV97, FFV07b, GMPP22, KO04b, LV18, Mos98, Rei03, BS23, Teu03]. Average-Case [DFV97, GMPP22]. Avoidance [BS19, Fel17, Mån02a]. Avoiding [ACÖ13, AM14, BJGY97, Bón14, BM14b, EGM^{+}20, Eri97, HKL12, Jan17, MW05a, Alm93, HS95, Pin20]. Azuma [Nao12].

B [Die95, FS20]. Back [Ano92b, Ano92d, Ano92f, Ano92h, Ano93b, Ano93d, Ano93f, Ano93h, Ano94b, Ano94d, Ano94f, Ano94h, Ano95b, Ano95d, Ano95f, Ano95h, Ano96b, Ano96d, Ano96f, Ano96h, Ano09a, Ano09c, Ano09d, Ano09g, Ano09h, Ano09k, Ano09l, Ano09o, Ano10a, Ano10c, Ano10e, Ano10g, Ano10i, Ano11a, Ano11c, Ano11e, Ano11g, Ano11i, Ano11k, Ano12a, Ano12c, Ano12e, Ano12g, Ano12i, Ano13a, Ano13c, Ano13e, Ano13g, Ano13i, Ano13k, Ano14a, Ano14c, Ano14e, Ano14g, Ano14i, Ano14k, Hir97a]. backtracking [WW23]. Balanced [GHH06, Mit99, EM22, KM23]. Ball [LM19]. ballot [DR22]. Balls [Oli08]. Balls-in-Bins [Oli08]. Banach [Nao12]. Banach-Space-Valued [Nao12]. bandit [FKR^{+}03]. Bandwidth [KT13, ABE^{+}22]. Barabási [Mór05]. Barbour [Hol04]. Barely [BBW14]. barrier [BB19]. Barycentric [DM11]. Based [LM01, BCJ23]. Bases [FK06, Ars21, GJ21b, Sch94]. Basis [Sey99]. Bat [BKL^{+}06]. be [DH00, KKY06, MR95]. Beck [Buk16, Fra21]. behavior [Pit93]. behaviour [Lem22]. Beineke [Big07]. being [Mil23]. Belief [COMV09]. Below [Lo09]. Berg [Rei00]. Berge [Bah12, BBDS21, EGM^{+}20, GS11, Omi21, PS92]. Berge-cycles [Omi21]. Berge-Johnson [Bah12]. Bernoulli [BCV16, Hub16]. Berry [Fuc08]. Best [KLM02, Pau06]. Beta [DKW14]. Beta-Coalescents [DKW14]. Bethe [WF11]. Better [XY11]. Between [ESS09, KS09, PS98, RS18, SSZ15, JS21]. Betweenness [Goe10]. Bias [May05]. Biased [Bed98, GRS20]. Biasing [FG11]. Bicircular [GN06, GJ21b]. Biclique [FJKK07]. Bijections [DP19, RS18, CDTK99, JS21]. Bijective [BK11a]. Binary [AKM00, Bon00, Cal97, DG10, Fuc08, GJ99, Gra15, HK14, Hol10, KLM02, Mer06, RW01, Sta99a, Wu98, Ald92, Bez94, FGOR93, Lem94]. Binomial [BP01, GP97]. Bins [Oli08]. Bipartite [AS00, BP12, Con09, Csa07, EGM19, EKT19, FG08, GS12a, JLM23, Kah01, KT13, Moh16, NV05, Pau06, PS13, RR97, AKS03, ABD23, BDH22, Häg96, JMY22, LP23, Sli96, Tho96b, VW92]. Bipartite-ness [JLM23]. Bipartition [DD10]. biregular [BDH22, KRRS23]. Birkhäuser [Ode06]. Bisections [FHY18]. bivariate [Fil21]. BK [Més23]. BKKKL [Fri04]. BKR [KSS11]. Black [CKMP07]. blanket [And23]. Blind [DRWW10, Wäs11]. Block [RRR19, Sch94]. Blocking [GMT94, KLM06]. blocks [Neu96]. blow [EM22, Kom99]. blow-up [EM22, Kom99]. blurred [RS22]. Boards [CFKL12]. Body [KM05]. boldmath [KS06]. Bollobás [ABH22, Csa07, Häg98, KSS95, Nik09, Pin17, SSTZ21]. Boltzmann [DFLS04]. Bond [Wie03, Wie92]. Bonds [FK09]. Bonferroni [CS95, Doh02]. Bonferroni-Type [Doh02]. Book [Bár04, Bár06, Ber06, Big07, Bov04, Dud07, Geo08, Hol04, Lea04, McM05, Moh06, NRS05, Ode06, Sco04, Ste05b, CFW23]. boolean [JS21, Ahl14, AW23b, BGRS12, Eng96, ESS94, GG19, Kel11, RS18, Sav98, Spi16]. Bootstrap [BPP06, BBM09b, BBM10, CW00, MNS17, Uzz19]. born [Mon20]. Borsuk [Bár04]. both [Smi93]. Bound [BCV16, BP12, BS19, Bón14, BJ18, BJ17a, BJ17b, Fio03, Gou15, HR11, KKM18, KLS^{+}13, Mat03, MWW02, Sha11, Uzz19, Vu02, Wag06b, Wie02, Wu97, CH96, Fra19, Für96, Kur22, LSM19, Ste22, Ste23]. Boundary [Loz08]. Bounded [BPS18, BBW14, BR00, Bor06, CS10, FP08, FLS08, Geb13, Gre12, GRW03, Hli06, Lo12, McD97, Nob98, PV13, VK18, War16, BPSS21, BBP21, CS93, DEM^{+}22, Edw96, NSW23, OS20]. Bounded-Degree [CS10]. Bounded-Size [BBW14]. boundedness [DL94]. Bounding [GMP19, GLM99, KMS19]. Bounds [AP02, AS04, áBGHV18, CS98, DRWW10, DK08, Dud07, FKM01, FG08, Fuc08, GW14, HJ97, Kah97, MSS09b, Mat97, MW05b, MW07, Meh11, MP17, Mor13, Mub16, SV03, SSZ15, Sok01, Sub07, Wie03, Wit13, XY11, Alm93, Aza94, BLP^{+}21, BMM22, CF92, CH23, CS95, GSS96, HM22a, LW21, Sin92, Wie95]. Bourgain [ST08]. Boxes [DJ13, TZ21]. Branch [DKW14, Hli06, IT11]. Branch-Width [Hli06, IT11]. Branching [Rio08, Sta03]. Branko [McM05]. Breaker [BSKS11, CFKL12, DFR21]. Breaking [Fil21, Leh13]. Bridge [ABMR12, MW17, FH23]. Bridge-Addability [MW17]. Bridge-Addable [ABMR12]. Brightwell [Häg98]. broadcasting [CDP96]. Brooks [Kim95]. Brown [Lon20]. Brownian [FH23, GS93, Han93, Maa20, Oli08]. bucket [KP22]. Bunkbed [Lin11, Lin19, HKNN23]. ‘Burnside [GJ02]. Burr [HHP07].

C [McM05]. Caching [JR04]. Cambridge [Ber06, Sco04, Ano15, Ano16]. Cameron [Big07]. Can [ESS09, Mos98, MR95]. Cancellative [Für12, KS07]. Canonical [KN14a, Mon14, Sau03, Bro03, SS94]. Canonically [DFR09]. Capabilities [DZ97]. Capacity [KL16, GL21c]. capacity-preserving [GL21c]. Carathéodory [Sob18]. Card [DGHS22, Law95]. Carpets [HP97]. Cartesian [AC93, BRZK17, CT98]. Case [AS04, DFV97, FM14, FFV07b, LV18, GMPP22, VW92]. cases [EGM^{+}20]. Causal [BL12b]. Cavity [Bov04]. Cayley [BH92, Gre17, MP06a, MR06, Tim12]. Cellular [BSU15]. Centering [McD97]. Central [Bei06, Wag15, RW19]. Certain [Aus15, Aus16, BP12, Ham11]. Chain [Ald00, CLS19, EMT18, Eva04, VV97]. Chaining [Dud07]. Chains [BL10, CL05, CPV01, Cza99, Jan08, JLR15, Kah97, LW98, MR06, Mon14, NPZ17, TA97, CP20, ESS94, Ing93, Lyn93, SC95, Sin92]. Challenge [Bov04]. Channel [Gra15, KC93]. Chapters [Bár06]. characterisation [Bry92].

Characteristic [EHJ98, Sli02, Kun93]. Characterization [AS06, GK01, LM01, Gub96]. Characterizations [Fio01]. Characterizes [GdMNN11]. Charging [SS13]. chasing [Yeh20]. Chayes [BSZ22]. Check [Lef05]. cherry [AHL21]. cherry-quasirandom [AHL21]. Chip [Van01]. Chip-Firing [Van01]. Choice [Alo92, Kri00, Sob18, Vu00]. Choices [Gra98, GHSS22]. Choosability [EKT19, WL01, HW19]. Chordal [Doh02, EOZ93, GS23]. Chosen [BM14a, KS15]. Christoffel [Ben18, God92]. Chromatic [AH20, AM02, BBH^{+}16, BKT11, Bol04, Bor06, BJ18, CK07, CFRR02, DK04, DK08, FP08, FKS07, GHH06, Gre17, HJ97, Hax04, KK14, KM10, KN99, KN16, LS08, OW02, PT18, Sok01, Sok04, Tho97, Vu02, CH96, Edw95, Edw96, Hec20, Ill22, Jac93, LSM19, MW22, Sli96, Ste22]. Chvátal [KS09]. Circles [ESS09, SSZ15, SS04]. Circuit [LW02, MR11, HR95]. Circuits [AGM99, Bal01b, BF12, DR99, GJ99, Gre12, MR06, MR98a]. Circular [BGLS04]. Circumference [ML01]. Cladograms [Ald00]. Clairvoyant [Gác00]. Clarkson [Sha03]. Class [Bor04, BIJ17, DH00, JKS13, McD09, Möh07, PWZ97]. Classes [ABMR12, BPS09, BMW14, DD10, EMT18, FR01, GW18, Ham98, Loz08, Neš05, SRT10, Wil15, BCOR99, BJ94, Neš99, NSW23, Baj16, HT96]. Claw [FGJ12]. Claw-Free [FGJ12]. Clique [EOZ93, HHWY23, Mon17, HKP21, KO19, OS20, Ste23, Ver95]. Clique-factors [HHWY23, HKP21]. Cliques [DGGP14, KO04a, Lo12, Mub16, Nik01b, PRS05, PV13, Yus07, DT20, GJ21a, KS94, KS96b]. Closed [Bal03, BMW14, DD10, Ecc16, McD09, PRT16, PT18]. Cluster [BFPS11, Geo08, Bie19]. Clustered [DEM^{+}22, NSW23, HW19, KP22]. Clusters [HS16, HM94]. co [Tyo21]. co-triangle-free [Tyo21]. Coalesced [Jan08]. Coalescence [Ber08]. Coalescents [DKW14]. Cocircuit [BR00, HR95]. Cocircuits [DR99, Opo02]. Codegree [GH17, GHZ19, HLTZ17, Mar13, VK18, PSS23]. Codegrees [CP16]. Codes [AS04, CKZZ15, DZ97, Für12, MS09, DMZ23]. Coding [CH14]. Codings [HS00a]. Coefficients [BD15, OW02]. Coexistence [AMR16, BCG^{+}19]. coherent [Kun93]. Coincidences [ESS09]. Colines [Kun00]. Colleague [Fau12]. Collection [Ano99]. color [HPS93]. color-critical [HPS93]. Colour [GS11, Mat10, FLW19]. Colourability [BS09]. Colourable [EWHK98, HHK^{+}12, Ill22, Sub00]. Coloured [AP08, All08, BLS^{+}12b, CF02, DH10, DH13, DFR09, KN14a, KY12, LMMW09, ŁRS98, Tra04, ABH22, BKLL22, BR99, FLW19, Saa96]. Colourful [Sob18]. Colouring [AA07, ABD23, BNCOS10, BWY14, BEHD19, CH18, CO07a, COMV09, DP04, DD10, EWHK98, EJ14, GST14, Hax01, KP18, KK08, KK09, LV18, LMST08, NS01, Ras19, SW07a, SW07b, Sub00, Sub07, DEM^{+}22, KO19, NSW23]. Colourings [BCOE17, Far06, GKP17, HKL12, JMŠ98, KLS18, KN03, KKL12, KS18, LPRS09, Leh13, Mos98, PT09, ST12, Škr99a, Škr99b, Voi00, ABC^{+}21, BBP21, Eft20, Häg96, KM23, Ste23]. Colours [CH18, EJ14, KL17, Mos98, CLR20]. Combin [Die95, Had99].

Combinatorial [Alo99, Ano10k, Bár06, DFLS04, ESS09, EGPN16, HOO98, Hol04, Hwa98, KP19, LM19, LLMO09, Man02b, Mar01, MC04, Tyr15, Wil15, ABT99a, BBD22, DT94]. Combinatorics [Ano03, Bár04, Lea04, Erd99]. Comment [Die95]. Comments [Baj16]. Common [Dan98, GMP19, HHK^{+}12, GLLV22]. Communication [DP04, LS09a, Sha17a, Sha17b, Pud03]. Community [EMP22]. Commuting [HZ14]. comparability [KT20]. Comparable [PST01]. Comparing [Lyo17a]. comparison [KR94]. Comparisons [ACG^{+}04, FM14, KP98]. Compatible [Gác04]. Compensation [AvLR13]. Competition [BBC^{+}05]. Competition-Induced [BBC^{+}05]. Competitive [ST05]. Complement [Dob02a]. Complete [ABBL10, AP08, All08, CKTV11, CCGJ18, Dob02b, Edw05, EGM19, FHR09, FJ18, GHZ19, HJ97, HV08, Jan99, Jin92, KN14a, LTZ22, ŁRS98, MR98a, Mye02, Nik05, PS13, Prö05a, ST98, AGSW22, DH94, EKK21, LV18, Saa96]. Completing [LW07]. completion [ACM19]. Complex [FP08, Sok01, Sok04]. Complexity [Far06, GN06, GJ07, GRV15, Juk06, Kot12, LS09a, Mos02, OW02, PW16, Sav98, Sha17a, Sha17b, DT94, GMPP22, Pud03, VW92]. Component [BBF00, BCOK14b, BK06b, CF04, HN14, JS07, LPS08, MR98b, Ofe07, Rio05, Tur11]. Components [CKP18, EJ14, Hwa98, Jan03, LMST08, ŁP92, DR22, DK23, RM92]. Composite [Kut02]. Composition [HS00b]. Compositional [BG14a]. Compound [ER99, HM15, Mån00]. Compressions [Rus12]. Comput [Had99]. Computation [BFLW05]. computational [VW92]. Computer [Lea04]. Computing [BS11, GN06, Gre12, SCW96, Ano03]. Concave [AP11a, BMM22]. Concavity [Wei97]. Concentration [BMM22, Bol04, HT01, McD02, PP14, ST09, Sch04, FJZ22, Hec20]. Concerning [AC99, BD02, CCM99, FT98, Häg98, CF92]. conclusion [Mat99]. concurrent [SS94]. Condition [DH00, Sch23]. Conditional [CHJ^{+}19, MS95, Mat95, Mat99]. Conditions [DGJ08, ES05, FJP13, GHZ19, KKY06, LSS17, JLM23]. Conductance [KLM06, Mon14]. cones [DG94]. Configuration [DH06, FV17, FS05, Rio12]. Configurations [FH97]. Conflict [BNCOS10, GST14, KKL12, PT09]. Conflict-Free [BNCOS10, GST14, KKL12, PT09]. Congruence [Baj16, Ham98]. congruential [PP92]. Conjecture [BK16, BT12, CCM99, CV08, GLWW15, GS12b, Häg98, KSS11, Kaw08, KLS^{+}13, Lin11, Lin19, Mat10, PRT16, Rei00, Sol15, BDZ20, BPSS21, Fra21, FGS05, HKNN23, Hyd23, JW03, KSS95, LP20, Lon20, LP23, MNS20, Omi21, Sta95, AC99, Bal01a, Blo06, Csa07, Joh04]. Conjectures [BKN05, Pin17, Sha17a, Sha17b, SSTZ21]. Connected [BCOK14a, BCOK15, BR16, Bon00, Bri98, CY00, CRS12, CF04, FP11, FGS05, Kan19, Qin01, RW01, Wag06a, Wu97, Wu98, Hal93, Mad94, Mad95, RM92]. Connectedness [BCOR99, GRW03]. Connection [DF14, CG22a]. Connections [Nar18]. connective [Alm93]. Connectivity [ABMR12, CFR02, FV17, MW17, AO93, CF95]. Conquer [AD16]. Consecutive [GMP19]. Consistency [MPK98]. consistent [AST93]. Constant [AP02, ASS08, Cal97, CFR02, CFRR02, CS06, FS18, KO04b, Rob03, Alm93]. Constant-Degree [ASS08]. Constants [KS09, MS13]. Constellations [Cha09]. Constrained [EGPN16, LS09b, Wag06b]. constraint [COKM20]. Constraints [CDS14, Goe10, FLW19]. Constructing [Dob02a, HN20, Kut02, RSW93]. Construction [ASS08, FK08, FZ00, HZ05, KN99]. Constructions [Gre03, Kuz00]. Constructive [DN13, Mat97]. Consultant [Big07]. Contact [Can18, dFdMR94]. Contained [DF09a]. Containers [ST16]. Containing [FRR12, FS05, GW18, JLR15]. Contiguity [GJKW02, Jan95]. Contingency [Bar08, Bar10]. Continuous [Bou18, Joh22, Mar01]. Contractibility [Kri05]. Contractible [Wu03]. Contraction [KN14b]. Contrast [KS03d]. Convergence [And23, CK16, CW00, Cur12, DPPS00, DGG^{+}02, FM14, PWZ97, RW16c]. Converges [GJ02]. Convex [DT17, KM05, McM05, SS11, ABFK92, BP92, BBD22, BCH22]. Convexity [NT16]. Convolution [Hwa98]. Convolutions [Cro12]. Cooper [BCL23]. Cooper-Frieze [BCL23]. cooperation [McM05]. Coordinate [MW08, NS22]. Cop [Meh11]. Copies [Nik06, Yus17]. Cops [BBKP15, BDFM12]. Core [BCG^{+}19, DH00, GSV16, GK04, Kah01, Rio08]. Cores [FP11, oCM12]. Corners [ZK17]. Corrádi [BLS12a]. Correcting [DZ97, DMZ23]. Corrections [GG10]. Correlated [Coc08]. Correlation [AL11, áBGHV18, SW08, McD92, PP92]. Corrigendum [Ano10k, BCOK15, Had99, RW16b, Sha17a]. cospectral [HN20]. Costs [Han97, Par04]. countable [AD94]. Counter [AL11, HMS^{+}08]. Counter-Examples [HMS^{+}08]. Counter-Intuitive [AL11]. Counterexample [AC99, SSTZ21]. Counting [BR16, BFLW05, BTW12, CDM14, FK01, Far06, FK18, FK14, Gla16, GGJ^{+}21, Gow06, GL21c, Ham11, Mou00, PRS05, PS13, PST01, RS07a, SS13, Spe05, Von15, Wan10, Zah20, Ann94, BR19, Eft20, GJ21b]. counts [BCJ23, GIM21]. Coupling [Fel17]. Couplings [ABT06, BS19]. Cover [Ano92c, Ano92b, Ano92d, Ano92e, Ano92g, Ano92f, Ano92h, Ano92i, Ano93c, Ano93b, Ano93d, Ano93e, Ano93g, Ano93f, Ano93i, Ano93h, Ano94c, Ano94b, Ano94e, Ano94d, Ano94g, Ano94f, Ano94i, Ano94h, Ano95c, Ano95b, Ano95d, Ano95e, Ano95g, Ano95f, Ano95i, Ano95h, Ano96c, Ano96b, Ano96e, Ano96d, Ano96g, Ano96f, Ano96h, Ano96i, Ano09b, Ano09a, Ano09e, Ano09c, Ano09f, Ano09d, Ano09i, Ano09g, Ano09j, Ano09h, Ano09m, Ano09k, Ano09n, Ano09l, Ano09p, Ano09o, Ano10b, Ano10a, Ano10d, Ano10c, Ano10f, Ano10e, Ano10h, Ano10g, Ano10j, Ano10i, Ano11b, Ano11a, Ano11d, Ano11c, Ano11f, Ano11e, Ano11g, Ano11h, Ano11j, Ano11i, Ano11l, Ano11k, Ano12b, Ano12a, Ano12d, Ano12c, Ano12f, Ano12e]. Cover [Ano12h, Ano12g, Ano12j, Ano12i, Ano13b, Ano13a, Ano13d, Ano13c, Ano13f, Ano13e, Ano13h, Ano13g, Ano13j, Ano13i, Ano13l, Ano13k, Ano14b, Ano14a, Ano14d, Ano14c, Ano14f, Ano14e, Ano14h, Ano14g, Ano14j, Ano14i, Ano14l, Ano14k, BDNP11, GW14, Jon98, LW02, RT05, Yus12, RSS23]. Covered [SS04, MR95]. Covering [All08, CKTV11, FOT98, HLSM21, Eng96, GSS96]. Coverings [FJKK07, Kuz00, RR97, Yus97, BBDM23, FRMZ21]. Covers [Sch07]. CPC [Ano92c, Ano92b, Ano92d, Ano92e, Ano92g, Ano92f, Ano92h, Ano92i, Ano93c, Ano93b, Ano93d, Ano93e, Ano93g, Ano93f, Ano93i, Ano93h, Ano94c, Ano94b, Ano94e, Ano94d, Ano94g, Ano94f, Ano94i, Ano94h, Ano95c, Ano95b, Ano95d, Ano95e, Ano95g, Ano95f, Ano95i, Ano95h, Ano96c, Ano96b, Ano96e, Ano96d, Ano96g, Ano96f, Ano96h, Ano96i, Ano09b, Ano09a, Ano09e, Ano09c, Ano09f, Ano09d, Ano09i, Ano09g, Ano09j, Ano09h, Ano09m, Ano09k, Ano09n, Ano09l, Ano09p, Ano09o, Ano10b, Ano10a, Ano10d, Ano10c, Ano10f, Ano10e, Ano10h, Ano10g, Ano10j, Ano10i, Ano11b, Ano11a, Ano11d, Ano11c, Ano11f, Ano11e, Ano11g, Ano11h, Ano11j, Ano11i, Ano11l, Ano11k, Ano12b, Ano12a, Ano12d, Ano12c, Ano12f, Ano12e]. CPC [Ano12h, Ano12g, Ano12j, Ano12i, Ano13b, Ano13a, Ano13d, Ano13c, Ano13f, Ano13e, Ano13h, Ano13g, Ano13j, Ano13i, Ano13l, Ano13k, Ano14b, Ano14a, Ano14d, Ano14c, Ano14f, Ano14e, Ano14h, Ano14g, Ano14j, Ano14i, Ano14l, Ano14k]. Creating [BK06b]. criterion [FS20]. Critical [Ald98, BB12, FV17, FVDH20, JS07, KS08, Kun00, RRR19, VS06, Wie02, Wie03, And23, BSZ22, Cou22, DR22, HS95, HM22a, HM22b, HPS93, LMY23, Wie92, Wie95]. critically [Mad94]. Cross [FT98, FT11, FK17, FK18, SS13, Tok13, FGK23]. Cross-Graph [SS13]. Cross-Intersecting [FT98, FK18]. Crossing [BH12, KM17, MW07, Szé97, DFR21, SS94]. Cryptography [KS03d]. CSP [SS06b]. Cube [Ell11, Sam17, Gow96, VS06]. Cubes [GR98, Rei02]. Cubic [DK06, Jac07, MV06, BMS95, FRS95, KS96a, Tho92]. Cumulants [BHLB^{+}13]. CUP [Big07]. Curved [LP18]. Curves [PS98, PRT16, PD17, Zah20]. Cut [BS04, SS06b, Shi98, DG94, NT93]. Cuts [KO07]. Cutsets [FS18, Tim07]. Cuttings [Hol10]. Cycle [BLGN^{+}02, BGRS12, BJ17a, CJ96, CCGJ18, DZ97, Die05, EGM19, FR01, FÖ09, GL21b, GS11, GL12a, KP18, KO04a, LSV12, Mar13, MW13, Nik05, Ver00, ACH23, BCOR99, JW03, LL21, Saa96, BJ17b]. Cycle-Complete [CCGJ18, Nik05]. Cycles [All08, Alo09a, AFK99, BJGY97, BLS^{+}12b, BFK^{+}18, CCGJ18, CF02, Cuc07, DFRS15, FHY18, FGJ12, FLMN17, FKS10, FK05, FG15, GHZ19, GNS17, HLP^{+}09, HMS^{+}13, Jan03, JKOP19, JŁ92, KKO08, KKS15, KKS16, KO08, KLO13, KM11, ŁRS98, Mub16, NV05, SS14, Ver02, WL01, AHL21, AKPP21, BBDS21, BKLL22, BDLP20, CFM94, DT20, DP23, EGM^{+}20, GGJ^{+}21, Häg93, HLSM21, HKŁ95, Jan94, JY20, Omi21, PSS23, Sch23, Tho96b]. Cyclic [Gre03, Mon18, TA97]. Cylinder [HL07, Moh94].

D [Die95, Hol04]. D. [JKS14]. dags [BCL23]. Daisies [BLM11]. Danzer [MR19]. Darboux [Ben18, God92]. decay [SC95]. December [SSS^{+}11]. Deciding [BS09]. Decision [Hat09]. Decoding [TZ00]. Decomposable [Man02b, Von15, ABT99a]. Decompose [Yus04]. Decomposes [DL14]. Decomposing [FLW19, KKS15, KLMP19, BLP^{+}21]. Decomposition [Blo05, MR06, PW16]. Decompositions [BD02, DLM04, ES05, FS10, Mon17, VV97]. Dedicated [Bol94]. Dedication [Ano09q]. Defective [HW19]. deficiency [FHT22]. Definability [BFL^{+}07]. Defined [Blo05]. Definite [FRV13]. Degenerate [AS11a, HS16, KN03]. Degree [Alm06, ASS08, BPS18, BBC^{+}05, BPS09, BF11, BN12, Bor06, CJK97, CM97a, CFR02, CFRR02, CF04, Coo06, CS10, Day17, DHL^{+}12, DGN11, EKM13, EMT18, FGJ12, FP08, FJP13, FKS07, Geb13, GRW03, GIMM23, HMS^{+}13, JS08, KS08, KKY06, KMW12, KO04b, KLO13, LM13a, Lo12, MR08, MWW02, MR98b, Mór05, PR17a, RS00, BKLL22, BPSS21, BBP21, CS93, DEM^{+}22, Edw96, FLW19, Mad95, RW92, Sch23]. Degrees [CP16, McK11, NT18, Sal16, CDK19, FRMZ21, GIM21, Jin93, Sco92]. Delaunay [Pim11]. Delays [AA07]. Deletion [FFV07a, Gra15]. Deletions [Gra15, PV19]. Demon [Gác00]. Dense [BTW19, BK16, Bra99, EKT19, Kan19, KSS01, Kri00, KO05, McK11, Mon17, OS20, Sok04, Yus12, ANS22, Ann94, GIM21, GIMM23, Hec20]. densities [HJW22]. Density [BCGR97, BTW12, CN12, FRV13, KS03a, Kri05, Lo09, Nag17, OP18, PT18, Raz08, Sch99, BDLP20, Lam23, PSS23]. Dependence [DJR07, KCR11]. Dependency [SS06a]. Dependent [Cal97, RT05]. Depth [AGM99, CV15, Gre12]. Derivative [LM01, GA19]. Derivatives [GMP19]. Descartes [KN99]. Descent [AR01]. Describing [JS07]. Designs [FR01, Mph00]. Detachments [Bah12, Edw05]. Detect [JS02]. detection [EMP22]. Determinantal [PP14]. Determining [DP06, KS03d]. Deterministic [Bec94, DF09b, Eft20, KW10, LV13]. Deuber [HS98, Prö05a, Prö05b]. Deviation [Gra98, Kah97, Vu01, ZD08]. Deviations [AM14]. Diagnosis [PU98]. Diagonal [Wil15, CFW23]. Diagram [MP06b, Pit05]. Diagrams [Edw98, Tim12]. Diameter [BV17, CFGN15, FS10, Fra17, GSW11, RW10, ANS22, BH92]. Diameters [DKLP10, HZ14]. diamonds [AMP20, KP22]. Dice [Gri97]. Dickman [HT02]. Diestel [Woe05]. Difference [HLP08, LZ09, Mon18, Sch07]. Differences [LM19, McD97, War16]. Different [EKM13, FK01, Lyo17a]. Differential [Mit99, RW16c]. Digraph [CF04]. Digraphs [Bal03, BJJ98, FJP13, GKP17, KKS16, KO08, Alo06b, AGSW22, CFM94, Cou22, Tho95, HMS^{+}13]. Dilates [Buk08, CHS09, HR11, Pla11a, Pon13b, Sha16]. Dilemma [DGG^{+}02]. DIMANET [Ano99]. Dimension [BDG00, FJKK07, Mat03, MS13, dMR05, Lar21]. Dimensional [DF09b, LS18b, Par17, Siv14, BCH22, Lem22]. Dimensions [AS11a, APST04, BBM10, ES11, FS18, Gla16, KMSS12, Pit05, Sha11, Uzz19, HS95]. Dimer [FG08]. Dirac [CDG^{+}21, GGJ^{+}21, KKO08, KKM18, RRS06]. Dirac-Type [KKO08, RRS06]. Directed [AHHT10, Bal03, CDKM15, FKS10, Han97, Kan19, LSV12, Pin17, Tre15, AE94, CG22b, Mon20]. Dirichlet [ABT99b, ABT99a, ABT06, Ber08, Hir97a, Hir97b, Pit02].

Discrepancies [DS03]. Discrepancy [Doe00, BCPT21, Fra21]. Discrete [Alo18, Bev16, Bou18, Ell11, Gal11, GA19, NT16, ST09, Ste05b, Szé97, TZ00, BMM22, CDTK99, FSSZ23, Law93, Mar19b]. discrimination [Yeh20]. Disjoint [All08, FGJ12, FRR12, FZ00, FJ18, KM11, MR06, Ver02, Tyo21]. Disjointness [PTT21]. disk [Moh94]. Dismantling [JT08a]. Disorder [Prö05a]. Displacement [Pet13, Teu03]. Displacements [Jan08]. Disproof [Joh04, Fra21]. Dissociated [ER99]. Distance [Ald98, BV17, BEHD19, BF12, DG17, DR12, EMS07, FKR05, Fio01, KK14, KP18, Pan04, Sta97, BJ22, Big93, MR22]. Distance- [KK14]. Distance-Regular [Fio01, Big93]. Distances [APST04, ES11, EKM13, KMSS12, PD17, Rei02, SS16, EMP93, Raz20]. Distant [áBGHV18, Fio03, Jag98]. Distinct [APST04, Baj16, Dob06, DT08, ES11, Ham98, HS00b, KS02, KS04, NT18, PD17, SS16, Raz20]. Distinction [ESS09]. Distributed [DP04, GP97, Thi16]. Distribution [BCV16, BM14a, BP01, BBC^{+}05, BKW08, CV15, CDKK08, Coo06, GP97, Hir97b, ISZ22, JS08, Ngu11, Nik01a, Nik06, Pan04, RR03, ABT99b, ACH23, FGOR93, Neu96, WW23]. Distributional [FM14, TA97]. Distributions [ABL^{+}13, AS02, GK01, Gri97, Möh07, Pit02, BMM22, Jan95]. Distributive [Smy13]. Divide [AD16]. Divide-and-Conquer [AD16]. Divisible [FR01]. Do [BT11, DZ94]. Dobrushin [DGJ08]. Document [JR04]. Does [Bar10]. Dominating [CY00, HPV99, MS09, HHV21]. Domination [DW06]. Dowling [Bon95, Sch02]. Drift [LS18a, LW21]. Dual [ADH^{+}19, KSS11, Gat93, Wie92]. Dual-Pivot [ADH^{+}19]. dualities [AEM22]. Duality [BD06, BS10, FP92, Kru11, Sli02, Whi92]. Dyadic [CLS19]. Dyck [The04]. Dynamic [PWZ97]. Dynamical [DV04]. Dynamics [GLWW15, BSZ22].

easier [Ver95]. Easily [AS06, AF15]. Easy [BL12a]. Economical [Alo09a]. Eden [HM94]. Edge [AC99, All08, Alo97, AA07, AHKT20, AL06, AFST20, BKT13, CLS19, CF02, FS07, FZ00, FJ18, GW14, HKL12, KY12, KW10, LMMW09, MW17, Nag17, Nik01a, Nik06, Sta99a, WL01, Ald92, BKLL22, BDLP20, FLW19, Häg96, Mad94, Mad95, PSS23, Saa96, Smi93]. Edge-Coloured [CF02, KY12, FLW19]. edge-colourings [Häg96]. Edge-Cover [GW14]. edge-density [BDLP20]. Edge-Disjoint [FZ00, FJ18]. Edge-Expansion [Alo97, MW17]. Edge-Flip [CLS19]. Edge-Isoperimetric [AC99, FS07]. Edge-Labellings [AFST20]. Edge-statistics [AHKT20]. Edge-Weights [Sta99a, Ald92]. Edges [BD08, BJ17a, BJ17b, DL14, DKR18, FRR12, FM17, GM04, GL18, HZ17, KKS15, KKL12, KN16, KLMP19, Mos98, TW97, Wu97, BLP^{+}21, LMY23]. edit [MR22]. edited [Ste05b]. editing [LS95]. edition [McM05]. Editorial [Bol93, Bol94]. Effect [PPGSS15]. Efficient [KK09, Kub96]. efficiently [GH20b]. Eigenvalue [AS00, GLM99, Jan05, KS03e, LP18, Nik02]. Eigenvalues [CY95, CO07b, LNW21, Moh16, GA19]. Eldridge [Csa07]. Elementary [ASS08, BH93]. Elements [DR99, Wu03, Lem94]. Elimination [Alo09a]. Embeddable [CT07]. Embedding [Eva04, KT13, RRT99, Bez94, LM23, Moh94]. Embeddings [FK14, KS02, MV06]. Emission [CDS14]. empirical [WW23]. Empty [DJ13, ST98, Sku96]. Encyclopaedia [Moh06, Ste05b]. Encyclopedia [Big07]. ended [Tim23]. Endomorphisms [DŁRR98]. Ends [BS10]. Enomoto [CKZZ15]. Entropy [FG08, HN13, Kah01, Lyo10, SS97, Tao10, BMM22]. Enumerating [Bar08]. Enumeration [Cha09, Eri97, Lyo05, MR98a, MWW02, Mon18, Vat08, Coo96]. Environment [BSU15]. Environments [HS16]. Equal [Aig05, ABE^{+}21, EMP93]. Equality [TA97, Wie92]. Equations [Ano10k, LLMO09, Mit99, RW16c, BR19]. Equibipartite [BK11b]. Equidistant [KMS19]. Equidistribution [Aus15, Aus16]. Equitable [KK08, KN03, Hec20]. Equivalence [MN09, CLR20]. Erased [AP08]. Erdos [Lon20, BT94, JW03, AH17, BBM09a, BPSS21, BVH11, Bol93, Bol94, BN12, CF92, CH23, DGS15, DR22, FJW13, Fra19, GHO17, GIM12, HOO98, HK19b, HK19a, HHP07, KKM18, LP23, LPS08, Nik09, NP20b, PSST13, Sol04, ST01, Szé97, Tok13, WW23]. ergodic [ST96]. Eriksson [Ber06, CCM99]. Erratum [Aus16, BJ17b, FLZ18, Lin19]. Error [CS06, DZ97, TZ00, DMZ23]. Error-Correcting [DZ97]. Esseen [Fuc08]. Essentially [DF09a]. Estimates [Wag08, Lis96]. Estimating [HKL^{+}20]. estimation [Ing93]. Euclid [DV04]. Euclidean [GMPP22, KMS19]. Euler [Sli02]. Eulerian [BLV16, HJ94, HMS^{+}13, Lis96, MR98a]. European [Hol04]. Euros [Bár06, Dud07, Geo08, Moh06, Ode06]. Evaluating [Nob98, Nob06]. Evaluations [EMM15]. Evasion [ARS^{+}03]. Even [BJ17a, BJ17b, BNMS15, DHL^{+}12, IT11, AE94, JY20]. Even-Degree [DHL^{+}12]. Even-Order [BNMS15]. Events [CS98, DT12]. Every [DL14, MR95]. Evolution [BDNP11, JS07, Ste05a]. Evolutionary [LS18a]. Ewens [BM14a, BFJL18, GP07]. Exact [AS02, AD16, BEHD19, EES03, GG10, HLTZ17, Mub02]. Exactly [KN14a, Mil23]. examining [Kub96]. Example [AD16]. Examples [HMS^{+}08]. Exchangeable [Ber08, Ros09]. Excluded [FJW13, QSZ09]. Excluding [Die95, KS96a, TT00]. Exclusion [BLV16, GMP^{+}15]. Exist [HHK^{+}12]. Existence [PS13, PR17a, SW18]. Exit [BL10, Mar01]. expanded [PP92]. expander [CG22b]. Expanders [ASS08, JKS13, Par17, PSY13]. Expansion [Alo97, AL06, BFPS11, CS10, FVDH20, HM22b, Luk98, MW17, MS13, Ofe07, Tra04, VS06, BS92, SS94]. Expansions [Sie98, FH23]. Expectation [KK07a]. Expectations [AS02, EES03, SV03]. Expected [AGM99, CO06, Dan98, HS00b, HL07, LZ09, LM23, RRR19, SS06b, Seg00, Sie98]. Explicit [Kuz00, Sta97, Gro03]. Explosive [PSST13]. Exponent [LP00]. Exponential [Can18, HV08, Ren10, VHV06, DFK^{+}93, Fra19, GP23, GRS20]. Exponentially [LSS17]. Exponents [Kun00, JMY22, JQ23, Wie92]. Expressing [Ano10k, LLMO09]. Extended [SS03, Sha03]. Extending [ABH22]. Extension [CDKM15, HT96, Moh94, Tet94]. Extensions [HZ05, NP20b, Wal07]. Exterior [Pik01]. External [DKW14]. Extinction [Can18]. Extremal [CFS14, FG15, GT09, GW21, GR98, HPS93, KK07b, Lea04, MT19, Wol06, AC93, AO93]. Eye [KL17]. Eye-Free [KL17].

F [KM12]. face [Tho92]. Faced [DD05]. faces [LM23]. Facility [FFV07b]. Factor [áBGHV18, MA07]. Factorial [BHLB^{+}13]. Factories [Hub16]. Factorization [Han93]. factorizations [DH94, HJ94]. Factors [FS12, GH17, HLTZ17, Kri97, Lyo17b, oCM12, AY93, EM22, HKP21, HHWY23]. failures [CDP96, Smi93]. Families [AFK00, BCGR97, BJ12, BFRT01, BGHT03, Cha09, DF09a, FKM01, FR15, FHR09, FT98, FOT98, FT11, Fra17, FK17, FK18, GL09, KKK12, KS07, KW10, PT18, Rus12, RW13, Yus04, EKNS21, EL19, EFK92, Fra19, HN20]. Family [CY00, FRR12]. Fano [FS05, LPRS09]. far [Mil23]. Farr [McD92]. Fast [CFKL12, EMT18, FK08, FW18, Meh11, Mou00, ST96, FLZ19]. Faster [AAK^{+}11, LSV12, AKPP21]. Fat [HS04]. Fat-Shattering [HS04]. Fault [PU98]. Favourite [TW97]. Feasible [Alo06a]. Feedback [CTY07, HMS^{+}13, Oli08, ACH23, DGHS22]. Felsner [Bár06]. Ferrers [ABS00]. Ferromagnetic [GJ07, dBBR23]. Few [CH18, HT16, KKS15, KKL12, KM11, Nik06, PU98, BC20, Pel94, RSS23]. Fiala [Buk16]. Field [Ald98, BV17, Bov04, DG17, BSZ22, GMPP22]. Fields [GJ07, GT09, HS00a, KK01, McC03]. Final [Pic11, Mat99]. Finch [GLWW15]. Find [FM14, AST93, KP98]. Finding [AKPP21, CO06, DGGP14, DKM^{+}15, GH20b, Pau06, Pik10, Saa96, Vu18]. Finitary [HS00a]. Finite [AKM^{+}06, BL10, BCPS18, Cza99, GG10, GT09, Gre97, KK01, Leh13, MO01, MP06a, Mon14, Peb04, Smy13, Sol15, Con20, DL94, GMPP22, McC03]. Firing [Van01]. First [Alm98, AW99, AP02, BVH11, BFL^{+}07, CK16, DKLP10, Gao14, Jan05, Pim11, Ren10, Sta95, VS06, Csi92, HM22b, McM05]. First-Order [BFL^{+}07, CK16]. First-Passage [Alm98, AW99, AP02, Pim11, Ren10]. Fitness [DO14]. Five [TT00, HHV21]. Fixed [BF11, BTW12, BL12b, Eri97, Fra17, FKS07, GLS15, Gra15, PST01, AW23b, GMPP22]. FKN [Fil20]. Flaass [Joh04]. Flajolet [SSS^{+}11, Sed14]. Flat [PSY13]. Flip [CLS19]. Flow [CK07, Jac07, OW02, Sin92]. flowgraphs [Coo96]. flows [AE94, SS94]. Fluctuations [BBF00, BCJ23]. Fly [BKL^{+}06]. Fon [Joh04]. Fon-der-Flaass [Joh04]. Forbidden [BJ12, FS12, FFO13, GL09, KS03a, MP17, ABD23, BKHL99]. forces [KY20]. Forcing [RW16a, Kur22, RW16b]. Forest [FJW13, Die95]. Forests [Alo18, BK11b, HW08, Ann94, ŁP92, MR95]. Forget [LW98]. Form [HK14, MW08]. forms [Mil23]. Formula [AM10, Bev16]. Formulae [BD15, DZ94]. Formulas [COW18, Gne04, GMP^{+}15, PW17, Sav98, CDD03]. Formulation [MR11]. Forward [HS16]. Foster [Tet94]. Four [Mat10, Sha11, MS95, Mat95, Mat99]. Fourier [Ars21]. Fraction [CPT17]. Fractional [Mon17, MW22]. Fragmentation [Ber08, GP07].

Fragmentation-Coalescence [Ber08]. Frank [BT03]. Free [ABKR18, AKS05, BNCOS10, BT12, Bra99, BNMS15, CC98, CJK97, Con17, EHJ98, FGJ12, FFV07a, FÖ09, GST14, Gyo06, GL12b, HN13, Jac07, KLS18, KL17, KN99, KKL12, PT09, PR17a, Pic11, Rio05, Sch99, Ste05a, Tho97, BCL^{+}21, CE99, FGK94, GMT20, Jac93, Jin93, Kri94, MW22, Tyo21]. Freeness [BGRS12, JS02]. Freiman [GT09, Pon13a, San08]. Frequency [BL10]. Friend [Fau12]. Frieze [BCL23, DKM^{+}15]. Front [Ano92c, Ano92e, Ano92g, Ano92i, Ano93c, Ano93e, Ano93g, Ano93i, Ano94c, Ano94e, Ano94g, Ano94i, Ano95c, Ano95e, Ano95g, Ano95i, Ano96c, Ano96e, Ano96g, Ano96i, Ano09b, Ano09e, Ano09f, Ano09i, Ano09j, Ano09m, Ano09n, Ano09p, Ano10b, Ano10d, Ano10f, Ano10h, Ano10j, Ano11b, Ano11d, Ano11f, Ano11h, Ano11j, Ano11l, Ano12b, Ano12d, Ano12f, Ano12h, Ano12j, Ano13b, Ano13d, Ano13f, Ano13h, Ano13j, Ano13l, Ano14b, Ano14d, Ano14f, Ano14h, Ano14j, Ano14l]. Frozen [BBP21]. Full [GRWW21]. Function [AB00, BS11, Bev16, GSV16, Gre12, HT02, MN09, Nob06, FH23, Lem22, MR22, WF11]. Functional [HT01, LM01]. Functionals [PP14, RW19]. Functions [AR01, BD15, BGRS12, Blo05, Far07, Gra98, Hub16, HS04, IK09, Kel11, McC04, Mén18, PSY13, RS18, Sam17, Sch11, Sok01, Spi16, Wag15, Wit13, Woe05, Wol06, AY93, BMM22, God92, JS21, KR94, Lyn93, MS20, Whi92].

G [McM05]. Gallai [NP20b]. Galton [BP01, CHJ^{+}19, DGZ23, Joh22, MS20]. Galvin [Sli96]. Game [CGL16, DGG^{+}02, FKPS05, Gri97, KK09, MW11, Van01, Csi92, Pek96, Bow07, Mos02]. Games [Bec05, Bed98, BSKS11, CFKL12, FK08, Geb13, HN14, MSS09a, MSS09b, Bec94, DFR21, FKR^{+}03]. Ganter [Vu97]. Gao [Ham08a]. Gap [Wei97, BDH22, Ing93]. Gas [SS06a]. Gates [HLW19]. Gaussian [CV15, DFV97, Wol11]. GCDs [GW21]. GEM [ABT06, GK01, Hir97a, Hir97b, Pit02]. GEM/Poisson [Hir97b]. General [Fio03, GKW14, MP17, Mil17, Mon14, Vu02, ZD08, HJ94]. generalisation [ABIST96]. Generalization [AM10, GIM12, Ham08a, ST01]. Generalizations [GJ21a]. Generalized [Atk98, GJK04, GKP17, KP11, Mar13, Pan04, Rei02, CH23]. Generated [KM17, BJ22, GH20a]. Generating [BK06a, EZ12, FLS08, FFO13, Mén18, Nob06, SW99, AEM22, God92]. Generation [BFJL18, DFLS04, ŁP93]. generators [PP92]. Generic [Dud07]. Geoffrey [Geo08]. Geometric [AK14, Bár06, BDFM12, Can18, DPPS00, DT17, GRV15, Hwa98, Ing93, MS09, ST08, AEM22, EMP22, FHT23, Gat93]. Geometrical [CDTK99]. Geometries [Kun00, Sch02, Bon95]. Geometry [Bár04, Bár06, Szé97]. George [Ber06]. Getting [LSV12]. Giant [BBF00, BCOK14b, BK06b, LPS08, MR98b, Ofe07, Rio05]. Gibbs [Stu20, dBBR23]. Ginzburg [HOO98]. Girth [CT07, EWHK98, GG10, HW08, KO04b, KS18, TT00, BMS95, HRRS21, HHV21]. Given [CF04, FRR12, GL12a, HKL12, KS08, Kut02, LM13a, MWW02, MR98b, Nag17, Nik06, Pik10, BDLP20, JPP22]. Glasses [Bov04]. global [Lin93].

Good [Kuz00]. Goodness [BPS18]. Gowers [Sha17a, Sha17b]. Grabbing [MW11]. Graduate [McM05]. Graham [Sol04]. Graph [ABBL10, ABMR12, AP08, AM02, AF15, BVH11, Big07, CGL16, CO10, Die95, Die05, Dob02b, DH00, EGM19, FK08, FFV07a, FJ18, GS12a, GW18, HJ97, HV08, HT01, HL07, Jan99, JS07, Jan07, JS08, JS20, Juk06, KK07b, KKY06, KK09, KN16, LV18, LMST08, LV13, Mar13, MR08, MR98a, MN09, MW11, MR98b, MR11, Mos98, MT19, Nik01a, Nik02, Nik05, Nik06, NS01, PSST13, Pau06, PT18, Pik10, Ras19, RS00, RRT99, SS97, Sco04, SS13, SW18, Vu01, WF11, Wu97, Yus04, Zha10, AT93, Bec94, CG22a, CG22b, Eft20, FLZ19, God92, HPS93, Jan94, JŁ92, KY07, LTZ22, LM23, Lyn93, MR22, MR95, NSW23, Pek96, RW92, Saa96, SSTZ21, SS94, Teu03, WW23, Whi94]. Graph-Grabbing [MW11]. graph-theoretic [Lyn93]. Graphic [Nob06, QSZ09]. Graphical [EKM13]. Graphoids [HMS^{+}08]. graphon [BCJ23]. Graphons [HN13]. Graphs [ABJM14, ARS^{+}03, AH20, All08, ABKR18, Alo97, AKS05, ABKS09, AP11b, AFK99, AL06, AFST20, AF04, BGLS04, BLS12a, BMMM18, BTW19, Bár06, BK16, BSKS11, BLS^{+}12b, BBC^{+}05, BPS09, BDFM12, BL04a, BK06a, BFL^{+}07, BF11, BKT11, BS04, BKN05, BN12, Bor04, Bor06, BTW12, BKT13, BF01, BMW14, BT12, Bra99, Bri98, BD06, BJ17a, BJ17b, Can18, CY00, CFGN15, CT07, CJK97, CRS12, CT98, CKTV11, CM97a, CPT17, CO05, CO07a, Con17, CF02, CFR02, CFRR02, Coo06, CDG07, Csa07, CW00, CS10, CDKM15, Day17, DP06, DZ97, DR12, DH10, DH13, DPPS00, DK06, DKLP10, DD10, Dob02a, Dob06, Doh02, DGN11, DW06, Edw05, EMT06, EMS07, EWHK98, EES03, EJ14, FRV13, FHY18, FGJ12, FVDH20, FP08, FKR05, FGY00, Fio01, FH97, FP11]. Graphs [FKS10, FP10, FP14, FG08, FZ00, FKS07, FM17, FK14, GG10, GHZ19, GS12a, Geb13, GW18, GM04, GSW11, GST14, GdMNN11, Gre17, GL12b, HJ97, Häg98, HPV99, HHK^{+}12, HN18, HT16, HZ14, HN14, HW08, HKL12, IT11, Jac07, Jag98, Jan03, Jan05, JT08a, JT08b, JM13, Jon98, JMŠ98, Kah01, KK14, KLS18, KS08, KM10, Kan19, KK07b, KS03a, KL17, KKO08, KKM18, KMW12, KT13, KRS04, KSS01, KKS15, KN99, KN03, KY12, KLMP19, Kri97, Kri00, KS03e, Kru11, KO04a, KO04b, KO05, KLO13, KM11, LM13a, LW02, Leh13, Let16, LMMW09, LP18, Lo09, Lo12, LM13b, LS08, Loz08, ŁRS98, LT10, Luk98, Lyo17a, MW05a, MP06a, MSS09a, MSS09b, MR06, Mat14, McC04, McD09, MWW02, McK11, MP13, Moh06, MV06, Moh16, Mon17, Mye02, Nag17]. Graphs [NV05, Nao12, Nar18, Nik01b, Nik06, Nob98, Ofe07, PT09, PS13, PV13, PR17b, RRR19, Raz08, RSS07, Rio00, Rio05, RW10, RR97, RT05, Sal16, SS20, SS06a, Sco11, SW08, SS13, SW07a, SW07b, Shi98, SS03, Škr99b, Sta03, SW99, Ste05a, Sub00, TT00, Tho97, Tim07, Tim12, Tur11, VK18, Ver00, Vu02, Wag06a, Wan10, WL01, Woe05, Wu03, XY11, Yus97, Yus07, Yus12, Zha17, AEM22, AK94, ABE^{+}22, Alo92, AKS03, AHKT20, ANS22, ABD23, And23, Ann94, AE94, ABH22, AJT03, BH92, BS92, BCL^{+}21, BKLL22, BDZ20, BDLP20, BCJ23, Bie19, Big93, BLP^{+}21, BR99, BBP21, BPSS23, BJ94, BMS95, BDH22, BDS21, BT20, CFM92, CM93, CJ96, CY95, CLR20, CMR^{+}21, CDG^{+}21, Coo92, CF95, DT20, DR22, DP23, DL94, DH94, DEM^{+}22, DHJM21]. graphs [EOZ93, EFRS93, FRMZ21, FHT23, FHT22, FRS95, FLW19, FGK94, GRWW21, GIM21, Gro03, GMT20, GL21b, GLLV22, Gub96, Häg92, Häg93, HJ94, Häg96, Hal93, HK94, HKP21, HRRS21, HHWY23, HHV21, HN20, Hec20, HW19, HT96, HPS93, Ill22, Jac93, Jan95, JMY22, Jin92, Jin93, KO19, KSSC99, KS96a, Kim95, KRRS23, KS94, KS96b, KT20, KS03c, KC93, Kri94, Lam23, LL21, LNW21, Lin93, LMY23, Mad94, Mad95, McD93, Mon20, NS20, Neš99, OS20, PTT21, PS92, RSS23, Smi93, Ste22, Ste23, Tho92, Tho96b, Tho96a, Tim23, TZ21, Tyo21, dFdMR94]. Greedy [Ald92, FFV07b, GG10, OS14, BB19, FRS95]. Green [Bev16]. Grid [BCPS18, DF09b, Far06, CMR^{+}21]. grids [CNT23, DFR21, Nic23]. Grimmett [Geo08]. Grötzsch [Škr99a]. Grötzsch-Type [Škr99a]. Group [Eva04, FR01, GGX16, Mun14, ST05, BH92, CK93, COGHKL21, EFK92, GH20a, ŁP93, Pel94, SCW96]. Group-Divisible [FR01]. group-invariant [SCW96]. group-theoretic [CK93, EFK92]. Groups [Aus15, Aus16, BPP06, BRZK17, BNMS15, CL05, Dob06, FGLP00, GH98, GGH12, Gou15, Gow08, Gre03, Gre97, Ham11, Mon18, Peb04, Pla11a, Sch02, Sol15, ZK17, Con20, Lon20, New23]. Growing [AL09]. Growth [DH06, ES05, FS18, KP19, LPS08, MS13, DL94, Esl21, GP23]. Grünbaum [McM05]. Guaranteed [KM08]. guessing [DGHS22]. Gyárfás [CH23, Omi21].

H [Ste05b]. Hadamard [FJZ22]. Hadwiger [FGS05, Kaw08]. Hajnal [BLS12a, BKLY13, BCPT21, CDKM15, KK08, LM13b, Tre15]. Hales [Wal07]. Half [Lo09]. Hamidoune [Baj16, Pla11b]. Hamilton [AHL21, AKPP21, BFK^{+}18, CFM94, CF02, DFRS15, FLMN17, GGJ^{+}21, Häg93, Jan94, KKO08, KLO13, LSV12]. Hamiltonian [BBDS21, BJGY97, Cuc07, FK05, Häg92, Sch23, Tho96b]. Hamiltonicity [BSKS11, CFR02, Mon20]. Hamming [AK94, FVDH20, Sam17, Siv14]. Hard [BCG^{+}19, CTY07, Gác00, GSV16, GK04, JS02, Kah01, Szé97]. Hard-Core [BCG^{+}19, GSV16, GK04, Kah01]. hardback [Ber06, Sco04]. Hardness [BS09, EWHK98, Ver95]. Harmonic [Woe05]. harmonious [Edw95, Edw96]. Hashing [Jan08, Pet13, PV19, Nil94]. Hasse [MP06b]. Having [ML01, Lem94]. Heap [MPP04]. Heavily [SS04]. hedgehogs [FL20]. Height [BD08, Lew08, DGZ23]. heights [FGOR93]. Hereditary [Doe00, EMS07, KNR03, SS03]. Hermite [GA19]. Heuristic [Wit13]. Heuristics [COGL07, WW05, LW21]. Hexagonal [Wie02]. Hidden [DGGP14, Vu18]. hierarchies [DT94]. High [Alo09b, BBM10, BKT11, DGGP14, FS18, HS00a, HMS^{+}13, KN99, KLO13, LS18b, MS13, Par17, HS95, Jin93]. High-Dimensional [LS18b, Par17]. Higher [APST04, Zah20]. Highly [Kan19]. Hilbert [Ano10k, LLMO09]. Hind [CH96]. Hindman [Pud03]. Hit [NS22]. Hit-and-Run [NS22]. Hitting [ABE14, Bev16, Lew08, LW21]. Hoare [KP98]. holds [HKNN23].

Hole [CKMP07]. holes [BCH22]. Homeomorphically [CRS12]. Homogeneous [BTW12, CN06, Neš05, MV23]. Homomorphism [CN06, dMR05, Neš99]. Homomorphism-Homogeneous [CN06]. Homomorphisms [Pon13a, CG22a, FZ22]. Honour [Bol93]. Hood [PV19]. Hopcount [HV08]. Hoppe [Hir97a]. hulls [ABFK92]. Hurwitz [BP01]. Hyperbolic [MW05a]. Hypercube [BBM09b, FÖ09, Gal11, GG19, JP17, MNS17, Bez94, BS23, DK23, JPP22]. Hypercubes [BBKP15, CDP96]. hyperfinite [CG22b]. Hypergraph [BS11, Con09, CR13, HLP^{+}09, HS12, Mub02, RRT99, BB19, EGJ20]. Hypergraphs [BT11, Bah12, BBM09a, BBH^{+}16, BNCOS10, BCOK14a, BCOK14b, BCOK15, BWY14, BR16, BIJ17, CP16, CKP18, CDM14, DHL^{+}12, DN13, DFRS15, DKR18, FKP16, Für12, GW15, GH17, GHO17, Gow06, Gyo06, GL12a, HK19b, Han15, HZ17, KKL12, KKS16, KS18, LPRS09, LM16, Liu22, LM13b, PT09, PRS05, Pik99, Pik01, RRS06, RS07b, RS07a, ST12, ST16, Tal02, Vu00, AHL21, AKPP21, BBDS21, BKHL99, CFMR96, EGŁ94, GRWW21, GGJ^{+}21, GIL22, GIMM23, HK19a, HLSM21, HRRS21, JY20, McD93, New23, Sch23, Spi23, VW93]. Hypermatrices [MP17]. Hyperplanes [VZ06]. hypotheses [AST93].

I.I.D. [DZ99]. Identifying [AS04, MS09]. Identities [Ben18, BHLB^{+}13, Lyo10, CS95, God92, Kun93]. Identity [Alo09b]. II [HK19b, KS96b, LM16, Mat95, PW04]. IID [áBGHV18, Lyo17b]. III [HT96, Mat99]. Image [Ham08b, HL93]. Imperfections [BGQ97]. Imply [MP17]. Importance [BM14b]. Impossible [Prö05a]. Improper [KM10, KO19, Škr99b, Ste23]. Impropriety [Škr99a]. Improve [MW05b]. Improved [CF92, KT20, Mor13, Mub16, ST05, Sha11, SSZ15, Sin92, Ste22, Ste23, Uzz19, Wie02]. Improvement [BFPS11, Buk16, CH96, Fra19]. Improvements [BP12]. Improving [MW07]. Inapproximability [GSV16]. Incidences [ES11, PS98, SW04, SSZ15, ST08]. Inclusion [GMP^{+}15]. Inclusion-Exclusion [GMP^{+}15]. Increase [KN16, Mos98]. Increasing [DHS16, Fuc12, KS15, KW10, DZ99, Esl21, GL21a, KR94, KP22, RW19]. Increments [Nac06]. Independence [AF04, Ben18, BF01, CFRR02, Csi13, EV13, MP06b, Nik01b, PV13, Yus07, EHS^{+}94, HHWY23]. independences [MS95, Mat95, Mat99]. Independent [BB12, BMMM18, CO06, CDM14, CR13, DW06, EGŁ94, FR15, FPS05, Gal11, GG10, GLS15, GS12a, Gra98, HPV99, HHV21, HS06, JPP22, LM13a, McD02, Yus97, Zha10, Gow96, GH20b, Mór92]. Index [Ano92a, Ano93a, Ano94a, Ano95a, Ano96a, BJ18, CH14, HJ97, KK14, Nei02, Thi16, BCOR99, Sli96]. Indexed [DT12]. Indistinguishable [GG19]. Individual [Jan08]. Induced [AS06, ABKS09, BBC^{+}05, BJ12, BT12, EGM19, EKT19, HW08, LTTZ18, MP17, NT18, Rei02, SS20, SS03, DGK22, HKŁ95, How96, Sco92]. Inequalities [AW99, CT98, Doh02, FS07, FT98, HT01, LM19, Lyo10, Man10, Nik02, TZ00, CY95, Mar19b, SC95]. Inequality [Gra98, KSS11, KLM02, Nao12, Pet11, Sam17, Smy13, McD92, Més23]. Inequivalent [May05]. Inertial [The04]. Infinite [BPP06, BD06, Die05, Mén18, Tim07, DL94, Hal93, Lam23, Pin20]. infinitely [GH20a]. Influence [DJR07]. Influences [Bou18, FS07, Fri04, Hat09, Kel11]. Information [CDS14, SGO02, Sta02b, EMP22]. Inhomogeneous [Tur11]. Initial [DH06]. insertion [Aza94]. instance [Pit93]. Instances [BL12a, SS06b, GJ03]. Integer [AD16, Ber06, GLWW15, HS00b, Par04, Tok13]. Integers [DRWW10, FKR05, GR98, Heg05, LM14, Mut13, Sta02a, LPP21]. Integral [MR11]. interfaces [GP23]. Interlace [EMS07, Tra10]. Interpolation [AB00, Sal16, ABIST96]. Intersecting [AAE^{+}97, BWY14, BFRT01, BGHT03, Cza99, DF09a, FT98, FOT98, FT11, FK17, FK18, Heg05, KKK12, PRT16, Rus12, RW13, Tok13, EKNS21, EL19, EFK92, ESS94, Fra19, FGK23, Liu22]. Intersection [CS98, JS08, JT08b, JM13, Kut02, LP00, Mon14, KSSC99, Sku96, TZ21]. Interval [ABL^{+}13, BF01, Eng96, Pit02, Jac93]. Intervals [EHJ98, Jac07, Tho97]. Introduction [AB09, AB10, Bal15, BBL^{+}18, Bri09, Whi94]. Intuition [BBHKŁ16, CGL16, Bec94]. Intuitive [AL11]. Invariable [BFJL18]. Invariance [Tel02]. Invariant [BL12b, Mun14, Pit02, ABT99b, SCW96]. Inverse [Gre03, Tao10]. Inversion [Thi16]. Inversions [CHJ^{+}19]. Involutions [Bon00]. Involving [CST97, HRRS21]. Irreducible [BG14b, CRS12]. Irredundance [CM97a]. Irregular [AW23a]. Irrelevant [DH06]. ISBN [Lea04, Sco04, Ste05b]. Ising [DHJM21, GSV16, GJ07, Kot12, WF11]. Isometric [FJKK07]. Isometry [Tel02]. Isomorphic [Mec01]. Isoperimetric [AC99, BL04b, CT98, Ell11, FS07, SC95, TZ00]. Isoperimetry [Nao12]. Israel [GLWW15]. Issue [BBK^{+}03, DS04, Prö05b, Ano92c, Ano92b, Ano92d, Ano92e, Ano92g, Ano92f, Ano92h, Ano92i, Ano93c, Ano93b, Ano93d, Ano93e, Ano93g, Ano93f, Ano93i, Ano93h, Ano94c, Ano94b, Ano94e, Ano94d, Ano94g, Ano94f, Ano94i, Ano94h, Ano95c, Ano95b, Ano95d, Ano95e, Ano95g, Ano95f, Ano95i, Ano95h, Ano96c, Ano96b, Ano96e, Ano96d, Ano96g, Ano96f, Ano96h, Ano96i, Ano09b, Ano09a, Ano09e, Ano09c, Ano09f, Ano09d, Ano09i, Ano09g, Ano09j, Ano09h, Ano09m, Ano09k, Ano09n, Ano09l, Ano09p, Ano09o, Ano10b, Ano10a, Ano10d, Ano10c, Ano10f, Ano10e, Ano10h, Ano10g, Ano10j, Ano10i, Ano11b, Ano11a, Ano11d, Ano11c, Ano11f, Ano11e, Ano11g, Ano11h, Ano11j, Ano11i, Ano11l, Ano11k, Ano12b, Ano12a, Ano12d]. issue [Ano12c, Ano12f, Ano12e, Ano12h, Ano12g, Ano12j, Ano12i, Ano13b, Ano13a, Ano13d, Ano13c, Ano13f, Ano13e, Ano13h, Ano13g, Ano13j, Ano13i, Ano13l, Ano13k, Ano14b, Ano14a, Ano14d, Ano14c, Ano14f, Ano14e, Ano14h, Ano14g, Ano14j, Ano14i, Ano14l, Ano14k]. Issues [Bol94]. Iterated [Ham08b, MM14, Sch11, SC95, DGG^{+}02].

J [Bár04, Big07, Die95]. Jack [FG11]. Jackson [LP20]. Jaeger [Blo06].

Jewett [Wal07]. Jigsaw [Mar19a]. Johnson [Bah12]. Joint [HK14]. Jointly [Thi16]. Josef [Sco04]. Journals [Ano15, Ano16]. Judicious [HZ17]. Jukna [Lea04]. JumbleG [FKPS05]. Jump [BT11, BS09]. junction [Coo96]. Juntas [DF09a].

Kagomé [Wie03]. Kaibel [McM05]. Kannan [DKM^{+}15]. Karp [BF11]. Katona [CKZZ15, DGS15]. KCM [HM22a]. Kernel [MM14]. Kernels [HT01, SC95]. Kesten [Ste05b, Rei00]. Keys [FD97]. Kimmo [Ber06]. KKL [Fri04]. Klee [McM05]. Kleitman [AC99, DGS15, Den97]. knapsack [DFK^{+}93]. Kneser [AH20]. Knot [DH10]. Knowledge [DP04]. Knuth [BK11a, JKS14]. Ko [AH17, BBM09a, GHO17, HK19b, HK19a, Tok13]. Kohayakawa [MNS20, Hyd23]. Kohayakawa-Kreuter [MNS20]. Komlós [NS20]. Korado [CF92]. Kreuter [MNS20, Hyd23]. Kühn [GMT20].

L [McM05]. Labellings [AFST20]. Lacunary [Dub08]. Ladder [Ren10]. Lagrangians [BIJ17, Tal02]. Lamplighters [Woe05]. Lando [Moh06]. Laplacian [CO07b, GLM99]. Large [AM14, BMMM18, BDZ20, CT07, CJK97, CO06, EWHK98, EKK21, EJ14, Gao14, Gra98, HW08, HMS^{+}13, Kah97, KKM18, KRT99, KY12, KO04a, KO04b, KS18, Kut02, Lem22, LMST08, Lon17, Mut13, Mye02, PR17a, Sco92, Vu01, Yus12, ZD08, AHKT20, AW23b, BKLL22, DR22, HM94, HRRS21, Kri94, MW22]. Large-Deviation [ZD08]. Largest [AK14, CKP18, CF04, KS03e, Mar13, Mut13, Nik02, Tur11, ACH23]. Latin [HJ08, ACÖ13, ACM19]. Lattice [ADH^{+}19, DFV97, DG94, DPPS00, Gri97, HP97, MS13, SS06a, Sey99, Smy13, Wie02, Wie03, AK94, ABC^{+}21, AW23b, BP92, Eng96, FP92, GP23, Sch94, Wie95]. Lattices [Alm06, FK09, VV97, Bon95, CDTK99, Wie92]. Lauri [Sco04]. Law [GIM12, Hwa98, MA07, BT94, Hir97b]. Laws [Han97, SRT10, CDK19]. Layout [DPPS00]. Lazy [BBKP15]. Leader [Pin17, Woe05]. leaf [DGZ23]. leaf-height [DGZ23]. Learning [AB00, AM10, LS09a]. Least [KO04b, TT00]. Lectures [Bár04]. Left [Mar01]. Left-Continuous [Mar01]. Lehel [Omi21]. Lehmer [DV04]. Lemke [Den97]. Lemma [BFPS11, SS20, Sco11, ST01, Kom99]. Lemmas [FLZ17, FLZ18, RS07b, RS07a, FZ22]. Length [BHL^{+}15, CFI^{+}16, DKW14, Dan98, DF99, GL12a, HL07, KP18, MA07, Sch04, GL21a, Pin20]. Lengths [LZ09, Sch11, Ver00, ACH23]. Level [Cro12]. Level-Sets [Cro12]. Lévy [FH23]. liar [SW92]. Lifting [CL05]. Lifts [AL06, GJR10]. Like [Bar10, PSST13, Wan10]. Likely [LP09, Pit93]. Limit [AM14, BCOK14b, CDK19, DP19, Fuc12, Han97, MA07, SRT10, Wag15, ACH23, HKNN23, Pin20, RW19]. Limiting [WW23]. Limits [Sta97, ABT99a, Jan21, JS20, SSTZ21]. Line [BJR09, FKRD09, McC04, PWZ97, SW04, Shi98, DHJM21]. Line-of-Sight [BJR09, FKRD09]. Line-search [PWZ97]. Linear [CH14, CCGJ18, DGGP14, DHL^{+}12, DN13, Doe00, Hub16, Mar19a, NT93, Pet13, SS06b, VV97, Wit13, EFRS93, JY20, LM23, MR95, PP92]. Linear-time [NT93]. Linearity [DP15]. Lines [ESS09]. Link [VHV06, CDP96]. Linked [FJP13, KKY06]. Linkedness [KO08]. Links [Kru11]. Linus [BKS10]. Lipschitz [JS21, Mat03, PSY13, PP14, RS18]. List [HJ97, Hax01, JMŠ98, ST12, Škr99a, Škr99b, Sub07, Voi00, Vu02, Sli96, Ste22]. List-Chromatic [HJ97, Sli96]. lists [AT93]. Local [AJT03, BS92, BCOK14b, BSKS11, BFPS11, DP04, GJ07, KS03a, Lin93, LSS17, Luk98, Spi16, AO93]. Local-global [Lin93]. Locally [AFK00, CRS12, FKM01, Leh13, Vu02, Ill22, KM23]. Locating [MS09]. Locating-Dominating [MS09]. Location [FFV07b]. Lock [Bor02]. Lock-in [Bor02]. Log [Wei97, BMM22]. log-concave [BMM22]. Log-Concavity [Wei97]. Logarithmic [Hol04, Mar19b, Rob03, Sta97, Sta99b]. Logic [HLW19]. Long [AFST20, BBY08, GNS17, HMS^{+}13, DGK22, EGM^{+}20, Sta95]. Long-Range [BBY08]. Longest [ABBL10, BGLS04, BF12, KS15, Saa96, Sku96]. Look [Bar10]. Loop [AP08]. Loop-Erased [AP08]. Loose [DFRS15, FLMN17, GHZ19]. Loss [SGO02, Sta02b]. Lovász [BFPS11, CO05, FS20, Fra19]. Lowell [Big07]. Lower [AP02, Aza94, BP12, CS98, Cro12, Dud07, FG08, Gou15, HR11, Kur22, Mat03, Meh11, Wie03, Zha17, HM22a, Ste22]. lozenge [MPT22]. LRU [JR04]. Luo [NP20b]. Lyndon [MA07].

M [McM05]. Machta [BSZ22]. Major [Thi16]. Majorants [AP11a]. Majority [BBM09b, DP06, GKP17]. Makarov [Law93]. Maker [BSKS11, CFKL12, DFR21]. Maker-Breaker [BSKS11, CFKL12]. Making [BCL^{+}21]. Manipulative [BBHKŁ16]. Many [AAK^{+}11, ESS09, JQ23, JLR15, LSS17, NT18, Sha16, Tyo21, RSS23]. Map [BCGR97]. Mappings [Man02b, BDM96, BS23]. Maps [ABL^{+}13, BG14b, Cha09, GKRV18, CDK19, Lis96]. March [Pla11b, SSS^{+}11]. Markov [Ald00, BL10, CLS19, CL05, CP20, CPV01, CST97, EMT18, Eva04, FW03, HS00a, HT01, Ing93, Kah97, LW98, Lyn93, MR06, Mon14, NPZ17, SC95, Sin92, TA97]. Marriage [LP09]. Marstrand [Gla16, LM14]. Marstrand-type [Gla16]. martingale [ISZ22]. Mass [Gla16]. Match [Cur12]. Matching [AH20, HS12, Mån00, Pau06, BB19, Bry92, FRS95, Lar21, Neu96, ST96]. Matchings [AK94, BS11, FG08, GG10, GJR10, Han15, HN18, HKL12, JT08b, KY12, KKS16, PS13, Wan10, CG22a, CFMR96, EGJ20, FGS05, GRWW21, GL21c, Jan94, Més23]. Mathematical [Hol04, Moh06, Prö05a, Ste05b]. mathematician [Pla11b]. Mathematicians [Bov04]. Mathematics [Big07, Hol04, McM05]. Matousek [Bár04]. Mátraháza [Ano99]. Matrices [ABE14, Alo09b, BL10, CV10, HZ14, KK01, Lef05, Mon18, Sco11, Seg00, VZ06, Wei97, BKHL99, FLM21, FJZ22, FKS22, FSSZ23, GS93, HL93, JSS22, NP20a]. Matrix [Cha07, God92, WW23]. Matroid [BMR99, FGLP00, Mph00, Sli02, Mil94]. Matroids [Bon00, BR00, BG07, CW06, Coc08, DR99, EHJ98, EMT06, GN06, GJ99, GHH06, Hli06, ML01, May05, Mec01, Opo02, Qin01, QSZ09, RW01, SW08, Wag00, Wu98, Wu03, AO93, Bry92, GJ21b, HR95, Kor96, Kun93, Lem94]. matter [Ano92c, Ano92b, Ano92d, Ano92e, Ano92g, Ano92f, Ano92h, Ano92i, Ano93c, Ano93b, Ano93d, Ano93e, Ano93g, Ano93f, Ano93i, Ano93h, Ano94c, Ano94b, Ano94e, Ano94d, Ano94g, Ano94f, Ano94i, Ano94h, Ano95c, Ano95b, Ano95d, Ano95e, Ano95g, Ano95f, Ano95i, Ano95h, Ano96c, Ano96b, Ano96e, Ano96d, Ano96g, Ano96f, Ano96h, Ano96i, Ano09b, Ano09a, Ano09e, Ano09c, Ano09f, Ano09d, Ano09i, Ano09g, Ano09j, Ano09h, Ano09m, Ano09k, Ano09n, Ano09l, Ano09p, Ano09o, Ano10b, Ano10a, Ano10d, Ano10c, Ano10f, Ano10e, Ano10h, Ano10g, Ano10j, Ano10i, Ano11b, Ano11a, Ano11d, Ano11c, Ano11f, Ano11e, Ano11g, Ano11h, Ano11j, Ano11i, Ano11l, Ano11k, Ano12b, Ano12a, Ano12d, Ano12c, Ano12f, Ano12e]. matter [Ano12h, Ano12g, Ano12j, Ano12i, Ano13b, Ano13a, Ano13d, Ano13c, Ano13f, Ano13e, Ano13h, Ano13g, Ano13j, Ano13i, Ano13l, Ano13k, Ano14b, Ano14a, Ano14d, Ano14c, Ano14f, Ano14e, Ano14h, Ano14g, Ano14j, Ano14i, Ano14l, Ano14k]. Mauritanian [Pla11b]. Max [BS04, SS06b, NT93, Ver95]. MaxCut [AKS05]. Maxima [Spi16]. Maximal [BT12, DG17, DJ13, FH97, GS23, JLR15, Fra19]. Maximizing [GLS15, KO07]. Maximum [AE07, BNMS15, CM97a, DGN11, FRR12, GP97, GL12b, HZ05, MR08, MWW02, Mec01, Mór92, Mór05, Pet13, RRR19, RS00, Shi98, Yus17, DF96, FGK94, LPP21, LMY23]. Maximum-Size [BNMS15]. Maximum-Sized [Mec01]. McKay [LP23]. Mean [Ald98, BV17, Bov04, DG17, MW13, BSZ22, Fra21]. Mean-Field [Ald98, BV17, DG17, BSZ22]. Mean-Weight [MW13]. Means [AW99, CW00]. Measurable [DT12]. Measure [Ais13, BDG00, HT01, Mun14, Sey99, dBBR23]. Measures [Alm06, AKM^{+}06, Ber08, BL12b, DHS16, DPPS00, FG11, PP14, How96]. Median [Moh16]. Medians [Gri97]. Meeting [Ano03]. Memory [Prö05b, Sed14]. Menger [AD94]. Merge [Pit02]. Metacirculant [Dob06]. Method [AD16, BR16, DD10, ER99, FRV13, FM17, GLM99, KN14b, LCF14, MM14, Rob03, Ros09, Sal16, War16, Kub96, McD93, Wie95]. Methods [Bár04, Mon14, Ode06, Aza94]. Metrics [FW18, GRS20]. Michel [Bov04, Dud07]. mildly [DFK^{+}93]. Mills [Lem22]. Min [MW13]. Minesweeper [Mos02]. minima [Sch94]. Minimal [AKM^{+}06, CP16, HL07, IT11, Raz08, CLR20, Jin93]. Minimally [RW01, Mad95]. Minimax [MS20]. Minimization [Hal93]. Minimizing [BDLP20, GL18]. Minimum [ABBL10, Ald98, BHL^{+}15, CTY07, CP16, CFI^{+}16, Day17, DKR18, DT08, FGJ12, FM17, GH17, HLTZ17, HMS^{+}13, KLO13, LM13a, LZ09, Lo12, Mar13, MP06b, Nik01b, PR17a, Pik99, PV13, PR17b, BKLL22, DF96, Lar21]. Minimum-Weight [ABBL10, Ald98]. Minkowski [HR11]. Minor [BMW14, DD10, IT11, McD09, PT18, Tho04, KO19, RW16a, RW16b, Ste22, Ste23].

Minor-Closed [BMW14, DD10, McD09, PT18]. Minor-Minimal [IT11]. Minors [BKT11, FJW13, Mec01, Mye02, QSZ09, TT00, AGSW22, Die95, EKK21, KS96a]. minus [PSS23]. Missing [AK14, RR03, EGM^{+}20]. Mixed [BP01, LW02, RS03]. Mixing [Ald00, ABC^{+}21, BBY08, CLS19, EMT18, KLM06, Mor13, Par17, Sha17b, SW14, NS22, RSS23, Sin92, Sha17a]. Möbius [Gre12]. Model [Ald98, BV17, BCG^{+}19, DG17, DH06, DF14, EGPN16, Far07, FV17, GK04, Geo08, Gre97, Kah01, PV19, Rio12, Sok01, DF96, DHJM21, Law95, Wie95, dBBR23]. Models [BB12, Bou18, Bov04, DO14, GSV16, HM15, JS08, KP19, MC04, ZD08]. Moderately [SW18]. Modular [AP11b, Spe05, Gro03]. Modulo [FHR09, KS02, Kut02, Sta02a]. Moment [The04]. Monochromatic [All08, BKLL22, BLS^{+}12b, BHLS20, EJ14, HKL12, KS06, KM08, LL21, LPRS09, LMST08, ŁRS98]. Monographs [Hol04]. Monomer [FG08]. Monomer-Dimer [FG08]. Monotone [ABMR12, AFST20, BHL^{+}15, BSU15, LS18b, ST01, WW05, HKL^{+}20]. monotonic [GRS20]. Monotonicity [Fel17]. Most [BH01, HZ17, KKK12, KM05, Lem94, LM23]. Motion [Oli08, GS93, Han93]. Moving [SW14]. MR0723569 [Die95]. MR1680088 [Had99]. Multi [BG07, CF02, Gra98, KS15, BBD22]. Multi-Coloured [CF02]. multi-parametric [BBD22]. Multi-Path [BG07]. Multi-Row [KS15]. Multi-Way [Gra98]. Multicolour [DS03, MW18]. multicommodity [Sin92]. Multidimensional [Bei06, ST01, DFK^{+}93]. Multigraph [Jan09, Sli96]. Multipartite [LM13b, LTZ22]. Multiple [AS04, AM10, BK11b, FKP16, FM17, PW16, RSS23, CLR20, PW04]. Multiple-Orientability [FKP16]. Multiplicative [Cha07, BH93]. multiplicities [BBDM23]. multislice [Fil20]. Multivariate [BLV16, CS95, GMP19, PW04]. Multiway [DF96]. Mutual [DH06]. My [Fau12].

N [Die95]. Natural [MPK98, BB19]. Near [Han15, HKP21, Jac07, Mat14, DR22, Fra19, NP20a]. near-critical [DR22]. Near-Cubic [Jac07]. near-exponential [Fra19]. Near-Optimal [Mat14]. Near-perfect [HKP21]. near-square [NP20a]. nearest [CF95]. Nearly [EMP93, Hub16, AMP20, HJ94]. Necklaces [Peb07]. Negative [DJR07, GP97, KCR11, SW08]. Negatively [LP18]. neighborhood [AT93]. neighbour [CF95]. Neighbourhoods [FPS05]. ness [JLM23]. Nested [RW13]. nets [ABFK92]. Network [CDG07, KSV02, MW13, PPGSS15, Sie98, Tet94]. Networks [AMR16, CKMP07, FKRD09, KP19, MS09, BCL23]. nine [BMS95]. No [EJ14, GL12a, KRT99, LMST08, Mec01, Die95, Dob02a, Had99, KO19, Ste22, Ste23]. node [CDP96]. Noise [Ahl14, Bou18, HS00a]. Non [AS11a, BPP06, CPV01, CFR02, CFRR02, FK13, HHK^{+}12, KM17, Lem94, LP18, LV13, MW07, Möh07, Sey99, SW18, CP20, Esl21, Häg92, Liu22, WW23]. Non-Amenable [BPP06]. non-backtracking [WW23]. Non-binary [Lem94]. Non-Constant [CFR02, CFRR02]. Non-Crossing [KM17, MW07]. Non-Degenerate [AS11a]. Non-Deterministic [LV13]. Non-Existence [SW18]. non-Hamiltonian [Häg92]. non-increasing [Esl21]. Non-Negatively [LP18]. Non-Orthogonality [Sey99]. Non-Overlapping [CPV01]. Non-Planarity [FK13]. Non-Regenerative [Möh07]. non-reversible [CP20]. Non-Three-Colourable [HHK^{+}12]. non-trivial [Liu22]. Nonnegative [Sli02]. Nonplanar [TT00]. Nontrivial [GLM99]. Norm [Seg00]. Normal [Ais13]. Normality [Ais13, BHLB^{+}13]. Normalized [KL16]. Normally [Thi16]. norms [SCW96]. Note [Alm98, BDG00, Csi13, DFR09, FK01, Gne00, Hax01, JKS14, NV05, San08, Sol04, TA97, Tel02, Ver02, Gro03, Lon20, Raz20]. Notes [CE99]. notion [Spi23]. NP [CTY07, LV18]. NP-complete [LV18]. NP-Hard [CTY07]. Nullstellensatz [LLMO09, Alo99, Ano10k]. Number [AH20, AM02, AFK99, AF04, BHL^{+}15, BCOK14a, BCOK15, BKT11, Bol04, BMR99, Bor04, BF01, BJ17a, BJ17b, CO05, COW18, CFRR02, DKL^{+}16, DR99, DW06, DP15, DKR18, DT08, EGM19, EV13, FD97, FM14, FOT98, FRR12, FK05, FKS07, FM17, GLS15, Geb13, GMP19, GM04, GHH06, Gra15, Gre17, GJR10, GL18, GS11, GL12b, Hax04, HLP^{+}09, Hir97b, HS00b, HS12, Hwa98, IK09, JKOP19, KM10, KM12, KS06, KN99, KN16, Kri00, LP09, Let16, LS08, LPS08, MP06b, Mat97, Meh11, Mos98, Mub16, Nag17, Nik01b, PS98, PV13, PR17b, PST01, RR03, Ras19, ST98, Thi16, Vu00, Vu01, Vu02, Wol11, Wu97, Yus07, Yus17, Zha10, Ann94, BP92, BDLP20, CH96, CMR^{+}21, CNT23, DGZ23, DMZ23, DFK^{+}93, Edw95, Edw96, FJZ22, FGK94, HHWY23, HKŁ95]. number [Hec20, JW03, KY20, LPP21, LSM19, LM23, LMY23, MW22, MPT22, OS20, PP92, Ree96, SS94, Ste22, Tho96b]. Numbers [Ais13, BP12, BGH05, BIJ17, Bri98, BH12, BK11b, CCGJ18, Con09, CST97, FG15, GS12b, KMW12, Kut02, LS09b, LTTZ18, Mub02, NRS05, Nik05, Szé97, Wag06b, Alo92, AKS03, AJT03, AW23b, BT20, CH23, CFW23, EHS^{+}94, FL20, HRRS21, HR95, Jan94, KS03c, RS03]. Numerical [Gou15].

O [Baj16]. Oberwolfach [Ano03]. Observed [Ahl14]. Obstructions [Moh94]. Occupancy [CPT17, ZD08]. occurrence [Sta95]. October [Ano99]. Odd [ACÖ13, BGRS12, CC98, EGM19, FS18, FGY00, FG15, HS06, Kaw08, Peb07, BC20, KO19, Sco92, Ste23]. Odd-Cycle-Freeness [BGRS12]. Odds [Pau06]. Off [CFW23]. Off-diagonal [CFW23]. Offset [CM97b]. old [AEM22]. once [DZ94]. One [AAK^{+}11, Jan99, KP18, Lo09, Mat10, Škr99a, FKR^{+}03, PSS23, Tim23]. one-armed [FKR^{+}03]. one-ended [Tim23]. Online [BNCOS10, BFK^{+}18, MSS09a, MSS09b, RSS07, Ano15, Ano16]. Only [Opo02]. Open [HLS03]. Operations [BH01]. operators [GL21c, SCW96]. Optimal [AS11b, CKZZ15, COGHKL21, Csi92, DKM^{+}15, FZ00, Han97, Hub16, KS03d, KLO13, Mat14]. Optimality [ADH^{+}19]. Optimally [Smi93]. optimisation [BBD22]. Optimization [WW05, Wit13, Ver95]. Optimizing [JR04]. Oracles [AM10]. Orbital [CK07]. orbits [CP20]. Order [ABL^{+}13, ACÖ13, BFL^{+}07, BL12b, Bri98, BNMS15, CK16, Dob06, GIM12, Pla11a, KS03b, Zah20]. Order-Invariant [BL12b]. Ordered [DŁRR98, KO08, MPP04]. Orders [PST01, ABE^{+}21]. ordinal [DT94]. Ore [KY07]. Ore-type [KY07]. Orientability [FKP16, GW15]. Orientable [Cha09]. Orientation [Pik10, Yus17]. Orientations [AP11b, Bor04, RT05]. Oriented [GHH06, KKO08, Nar18, Bez94, Bie19, Häg93]. Orthant [HLP08]. Orthogonal [Blo05, HJ08]. Orthogonality [How96, Sey99]. Osthus [GMT20]. Other [BLM11, HJ97, JT08b, PP14, PV19, FGOR93]. Ould [Pla11b]. Outerplanar [BK06a, GdMNN11]. Overlap [CCM99, Mån02a]. Overlapping [CPV01].

P [Die95]. Packing [ABS00, Bal01b, Bal03, BKN05, BF02, BFK^{+}18, Con09, Dob02b, FLMN17, KK07b, KK09, RRT99, VK18, Yus07, Yus12, AC94, Eng96, KSS95, KY07]. Packings [DN13, KO05, KLO13, Kuz00, LM16, PW16, BDZ20]. Pair [FS12, FGK23, Sch23]. Pairings [Ham11]. Pairs [FK18, PST01, Ros09, Sta95, Wie92]. Pairwise [PRT16]. Paper [Baj16]. paperback [Ber06, Sco04]. paradox [Had99, Had98]. Paradoxical [ES05]. Parallel [DGN11, KP19]. Parameter [BCGR97, Ber08]. Parameters [MPP04, Wag15, HKL^{+}20, ISZ22]. parametric [BBD22]. Parametrized [DV04, EMT06]. Parent [AS04]. Parent-Identifying [AS04]. Parents [AS04]. Parity [Lef05]. Parity-Check [Lef05]. Parking [GP19]. Part [AE07, BG14a, HS00b, Mut13]. Partial [BE15, Cur12, FG08, PST01, Voi00, ACM19, DGHS22]. Partially [Ahl14, DŁRR98]. particular [FH23]. Partite [AH17, CKTV11, GHZ19, Mon17, Yus97, BCL^{+}21, EGŁ94, Jin92]. Partition [BS11, BHLB^{+}13, BS04, GSV16, GKW14, MR11, Sok01, WF11, HL93, HLS03, KS03b, Lem22]. Partition2 [Pit02]. Partitioning [CO10, HZ17, ŁRS98, Opo02, AC94]. Partitions [AD16, Ber06, BFP10, BKW08, DKM^{+}15, GRV15, Heg05, Jag98, KM17, MW07, Mut13, Nac06, RS07b, RS07a, Vu18, XY11, AC93, CJ96, EOZ93, GL21b, LL21, Sau03, Stu20]. Parts [áBGHV18]. Passage [Alm98, AW99, AP02, BVH11, DKLP10, Pim11, Ren10]. Passenger [Wäs11]. Passive [AHW16, JS08]. Path [ABBL10, BH01, BG07, BF12, DF99, GLM99, HV08, Let16, Rob03, Sch04, VHV06, Die95]. Path-Width [BH01, Die95]. Paths [ADH^{+}19, BGLS04, BK11b, DP15, FZ00, Jan99, Mon14, Nag17, Pin17, Ree96, The04, Wol06, BKLL22, DDF^{+}21, DGK22, Sku96]. pathwidth [NSW23]. Pattern [AM14, Jan17, Mån02a]. Pattern-Avoiding [AM14]. Patterns [ABL^{+}13, Can09, CDKK08, Jan17, LCF14, KM23, Pin20]. Paul [Bol93, Bol94]. peel [DGZ23]. Peer [CDG07]. Peer-to-Peer [CDG07]. Pegging [GSW11]. Penney [Csi92, Sta95]. Penney-ante [Sta95]. perceptron [AST93]. Percolated [Ofe07]. Percolation [Ald98, Alm98, AW99, AP02, BB12, BPP06, BBM09b, BBM10, BBY08, BVH11, BJR09, DKLP10, FVDH20, Gác04, GH20a, Lin11, Lin19, MP06a, MW05b, MW07, MNS17, MW08, Mos02, PSST13, Ren10, Siv14, Uzz19, VS06, Wie02, Wie03, Bie19, DFR21, DK23, EMP22, HS95, HM22b, Wie92, Wie95]. Perfect [AL09, BS11, BTW12, CG22a, CFMR96, GJR10, Han15, LM16, Mph00, Nil94, Pau06, PS13, HKP21, HT96, Jan94, Lar21, PS92]. Performance [FFV07b]. Period [Sch11]. Periodicity [CLS13]. Perles [McM05]. Permanents [Bar08]. Permutation [Eva04, GIM12, GJKW02, CS95, DF96, Gow96]. Permutations [ABE^{+}21, ABT06, AM14, Atk98, BM14a, BHL^{+}15, BLGN^{+}02, BDMS14, Blo05, Bón14, FW18, Gne00, Jan17, KP11, LS18b, McD02, Vat08, CK93, Fra21, Kur22, Pin20]. permuton [Maa20]. Perturbed [Alo09b, BTW19, KKS16, BPSS23, DT20]. Peter [Big07]. Petrie [AEM22]. Pevzner [Sta95]. Pfaffian [PW17]. Phase [BCG^{+}19, GK04, KS08, Rio12, COKM20, Joh22, Neu96]. Phenomena [DLM04, Lin93]. Phenomenon [Gal11]. Philippe [SSS^{+}11, Sed14]. Pivot [ADH^{+}19]. Planar [AFK99, CFGN15, EJ14, FP11, GW18, GM04, IT11, Loz08, MR08, Mén18, Škr99b, WL01, CDK19, Gub96, Lis96, Tho95, Tim23]. Planarity [FK13]. Plane [AvLR13, BFP10, Bow07, ES11, PD17, SS13, Sok04, Tho92, EMP93, Gat93, HM94, NV94, VW92, Wel94, Zah20]. Planes [LPRS09]. Plant [LCF14]. Planted [BS04]. Planting [BCOE17]. Played [CFKL12]. player [Csi92]. Playing [AHK10]. Plünnecke [Pet11]. Plus [EGM19]. Point [ABFK92, Ham08b, JS07, SW04, BCH22, HM22b]. Point-Symmetric [Ham08b]. Pointer [Yeh20]. Points [DJ13, FVDH20, PS98, PW04, SSZ15, SS16, SS04, BC20, DG94, HS95]. Poisson [AP11a, ABT99b, ABT99a, ABT06, BP01, Ber08, CPV01, ER99, Fuc08, GP07, GSS96, Hir97b, HM15, Mån00, Pit02, Sta97, Sta99b]. Poisson* [Hwa98]. Polar [Edw98]. Polarised [DM23]. Pólya [KN14b]. Polygons [DT17, SRT10, BP92]. Polyhedra [DD05]. Polyhedral [MV06]. Polymatroid [Nob06]. polymatroids [VW93, Whi92]. Polynomial [Ano10k, Ben18, BKL^{+}06, BCPS18, Bor06, CLS19, CO06, Csi13, DHJM21, EMS07, GS12a, GN06, GdMNN11, GKRV18, Hli06, LLMO09, MR11, Nob98, OP18, Tra04, Wal07, AE94, BBD22, BR99, GMPP22]. Polynomial-time [DHJM21]. Polynomials [BLV16, DH10, DH13, DK04, DK08, EHJ98, EMT06, EMM15, FP08, GMP19, Jac07, KS04, Kot12, Man10, MN09, Mph00, Nic23, OW02, Qin01, Sok01, Tho97, Tra10, Von15, Wag00, WF11, WW05, ABH22, Jac93, Kun93]. Polytopes [McM05]. Poor [Alo18]. Popular [CTV15]. Population [HM15]. Pósa [FJW13]. Posed [JKS14]. Poset [AW23b]. Position [Mil17]. Positional [Bec05, Bed98]. Positive [Bie19, DJR07, HLP08]. Possibility [DH06]. Potential [Big93]. potentials [AE94]. Potts [Sok01, dBBR23]. Potts-Model [Sok01]. Power [Mát07, GHSS22]. Powers [AM02, BN12, DDF^{+}21]. pp [Bár04, Bov04, Dud07, Hol04, Lea04, McM05, Sco04, Ste05b]. Precise [GLWW15]. Precoloring [HT96]. Predecessors [BDM96]. predicate [Coo96]. predicate-junction [Coo96]. Preface [HNZ19]. Preferential [AMR16, BBC^{+}05, DO14, Jan19]. prepared [McM05]. Prescribed [BJGY97, Day17, Sal16, CDK19, Tho92]. Preserve [BH01]. Preserving [DR12, GL21c]. Press [Ber06, Sco04]. Prime [Pla11a]. Primes [Dob06]. primitive [LPP21]. Prisoner [DGG^{+}02]. Probab [Had99]. Probabilistic [ACG^{+}04, AD16, BBHKŁ16, BR16, Hir97b, Hol04, Mat10, MC04, Alo92, Bec94, LS95]. Probabilities [BB12, BDMS14, KK01, NPZ17, PW17]. Probability [Ano03, Bor02, CS98, DGGP14, GNS17, Hat09, Jan09, Sav98, SW18, Ste05b, TZ00, Wie02, Wie03, DF96, Gat93, Nil94, Wie95]. Probably [KKK12, Rus12, RW13]. Probing [Pet13, PV19]. Problem [AC99, Aig05, Alm98, AS02, Ano99, Bah12, BGHT03, BF02, CTY07, CM97b, CN12, DGS15, DMS19, Ecc16, FRV13, FR15, FFV07b, JKS14, KM05, LW02, LP09, LV18, Mar01, MT19, MR19, RSS07, Wäs11, Zha17, ABHM21, CH23, DFK^{+}93, Für96, GMT20, KSSC99, LS95, NT93, Pit93, Pud03]. Problems [Ano10k, BL04b, BLM11, FGK23, GR98, KMSV07, LLMO09, MP17, Szé97, Tho04, CS93, COKM20, Erd99, FHT22, Gat93, GJ21a, GW21, Hal93, HLS03, HPS93, KY07, Moh94, Ver95]. Process [Can18, FM14, FFV07a, GP07, GRW03, JS07, KCR11, KSV09, Pic11, Sta99b, ABT99b, Esl21, JŁ92, GJ02]. Processes [AP11a, áBGHV18, Ber08, BLV16, CST97, Dud07, Oli08, PSST13, Rio08, RW16c, RW00, Sta97, FW03, RW92]. Product [Dob06, Fri04, Häg98, Hat09, Kel11, AC93, Mar19b]. Products [BDMS14, CT98, DH13, HQ05, HT01, Sta02a, Wil15, AC94]. Profile [DG10, BJ22, Ill22]. Profiles [Fuc12]. Programming [CH14]. Progressions [CLS13, Ngu11, Sol06, Ver00, JLM^{+}03, RSW93]. Projecting [HLP08]. Projective [Kun00]. Proof [Alo09b, BK11a, DP15, HS98, KK08, KSS95, NS16, Rei00, Die95, Omi21, Sli96]. Proofs [BD02, CFS14, Pin17, DGK22, FKS22]. Propagation [COMV09]. Properties [AF15, BMW14, BJ94, Cro12, FHR09, HM94, HN13, Heg05, JM13, KNR03, KN99, SS03, WL01, ABC^{+}21, AE94, DT20, FSSZ23, HKL^{+}20, MNS20]. Property [FJW13, FW18, LV13, NT16, TA97, DKR18, FS20]. Proportion [Had98, Had99]. proposal [Pit93]. Propp [HS00a]. Pseudo [SS04, Lar21]. Pseudo-Circles [SS04]. pseudo-dimension [Lar21]. Pseudographs [GJKW02]. Pseudorandom [Con17, EGJ20, HKP21]. Pseudorandomness [AKM^{+}06]. Pursuit [ARS^{+}03]. Pursuit-Evasion [ARS^{+}03]. Puzzles [Mar19a].

Quadratic [Spi16]. Quadruples [Aus15, Aus16, FK06]. Quadtrees [Cur12]. Quantitative [Aus15, Aus16, Bou18, FS20]. Quantum [BFLW05, Jan07]. Quarter [AvLR13]. Quartets [AKM00]. Quasi [AP11b, Aus15, Aus16, Mye02, NRS05, SS03, Sta03]. Quasi-Random [AP11b, Aus15, Aus16, Mye02, SS03]. Quasi-Randomness [NRS05]. Quasi-Transitive [Sta03]. Quasirandom [BRZK17, Gow08, LM16, ZK17, AHL21, Kur22]. Quasirandomness [Gow06]. Queries [Cur12]. Query [GRV15, ABHM21]. Question [GLWW15, HHP07, Sol04, ST08]. questions [AKS03]. Quickly [SW99]. QuickSelect [FM14, HT02]. Quicksort [ACG^{+}04, ADH^{+}19].

Radius [Gou15]. Rado [AH17, BBM09a, GHO17, HK19b, HK19a, Tok13]. Raffaele [Sco04]. Railways [GJK04]. Rainbow [AH17, JLM^{+}03, KMSV07, KY12, PS13, GRWW21, GJ21a]. Ramanujan [BL04a]. Ramsey [NS16, Ode06, AKS03, ABKS09, AJT03, AW23b, BPS18, BH93, BGH05, BT03, BBK^{+}03, BMR99, Bri98, Bro03, CH23, CS93, CP16, CLR20, CMR^{+}21, CCGJ18, Con09, CFW23, CNT23, DT20, DR12, DP15, EFRS93, EHS^{+}94, FL20, FKR^{+}03, FW03, Geb13, Gro03, GS11, GS12b, HRRS21, HKŁ95, HLP^{+}09, HR95, JKOP19, JW03, JM13, JLM^{+}03, KMW12, KN14a, KT20, KS03c, Lam23, Let16, LS09b, MSS09a, MSS09b, MNS20, Mub16, NV94, Neš05, NRS05, Nik05, Pek96, RS03, SS14, TW98, Wag06b]. Ramsey-Minimal [CP16]. Ramsey-type [NV94, AKS03, HRRS21, KT20]. Ramsey-Unsaturated [SS14]. Random [AP11a, ABE14, ABJM14, AR01, AP08, ABKR18, AS02, AL11, AP11b, AAK^{+}11, AL06, AFST20, AK14, AS11b, AM10, ABT06, ACH23, AGM99, AF04, Aus15, Aus16, BBM09a, BLS12a, BGQ97, Bar08, Bar10, BCOK14b, BSKS11, BPS09, BVH11, Blo05, BFP10, BK06a, BFL^{+}07, BF11, BS04, BSU15, BKT13, BF01, BD08, BF12, CC98, CK93, Can18, CL05, CFGN15, CDKK08, CFKL12, CGL16, CO05, COGL07, COW18, CDS14, CKP18, CF02, CFR02, CFRR02, CF04, CS06, CFI^{+}16, CV10, CW00, Cur12, DHS16, DF14, DPPS00, DKLP10, DF99, DF09b, DLM04, DFLS04, DW06, DJ13, ER99, EES03, EGPN16, Eva04, FFV07a, FP11, FKP16, FZ00, FKS07, FK13, FK14, Gal11, GW15, GJK04, GHO17, GW14, Geo08, GM04, GIM12, Goe10, GS93, GP19, Gou15]. Random [GP97, Gre17, Gre97, GJR10, Häg98, HS00a, HK19b, Han97, HN13, HZ14, HS00b, HJ14, HS16, Hol10, Jan95, Jan99, Jan03, Jan05, JS07, Jan07, JT08a, Jan09, Jan17, Jan19, JS08, Jon98, KK01, KS08, KM10, KLM06, KRS04, KKS15, Kri97, Kri00, KS03e, KSV09, KM11, LP00, Leh13, Let16, Lew08, LMMW09, LS08, MO01, Mar01, MSS09a, MSS09b, Mar19a, MW13, McC04, MR08, McD09, McK11, Mer06, MP13, MR98b, Mór05, MS09, Mut13, Mye02, Nac06, NT16, Nei02, OS14, PSST13, Par04, Pit05, PV19, Pon13a, RR03, RRR19, Ras19, Rei02, RS00, Rio00, Rio05, RW10, RW00, RW92, Sal16, SV03, SS06b, Seg00, SW07a, SW07b, Sie98, SS03, Sob18, Sta99a, Sta03, Sta02a, Sta97, Sta99b, Sta02b, SW18, SW99]. Random [Sub00, TW97, TW98, Tur11, VZ06, Vu00, Vu01, Wag08, Woe05, Yus04, Zha17, Ald92, AKPP21, ABE^{+}22, And23, ABT99a, BH92, BBDS21, BDM96, BB19, BDZ20, BCJ23, BJ94, BCL23, BDH22, BHLS20, CFM92, CM93, CDP96, COKM20, CDG^{+}21, Coo92, CFM94, CF95, CFMR96, Cou22, CDD03, DM23, DFP92, DP23, DF96, EF95, EKK21, FP92, FLM21, FKS22, FSSZ23, FRS95, GHSS22, GJ03, Gow96, GIM21, GIL22, GIMM23, HK19a, Hec20, How96, JSS22, Jan94, Jan21, JS20, KSSC99, KRRS23, Kor96, LL21, LM23, ŁP92, ŁP93, MR22, MS95, Mat95, Mat99, McD93, Més23, Mon20, MV23, NS20, New23, NP20a, PP92, Pin20, Pit93, RSS23, Teu03, Tim23, WW23, Whi94, NS16]. Random-Cluster [Geo08]. Random-Free [HN13]. Random-Walks [Gre97]. Randomised [Wel94, Ann94]. Randomized [ARS^{+}03, DP06, GG10, PPGSS15, WW05, Wit13, DPR21, LW21]. Randomly [BTW19, FJ18, KS15, KKS16, Nar18, RSS07, Bie19, BPSS23, DT20]. Randomness [NRS05]. Range [BBY08]. Rank [Alo09b, CV10, Nob06, ABH22, CG22a, FLM21]. Rapid [KM05]. rate [BT94]. rates [GP23, Sin92]. Ratio [AM14, KLM02]. Ratios [LP18]. Rayleigh [CW06, PP14]. read [DZ94]. read-once [DZ94]. Real [DK04, DK08, OP18]. Realistic [CTV15]. Realizations [EKM13]. Reals [BDG00, DRWW10]. Reciprocity [AE94]. recognition [SWŻ96]. Recognizing [VW93]. recolouring [McD93]. Reconstructibility [Peb04]. Reconstructing [AT93, Peb07]. Reconstruction [Sco04, CM93, Mil94]. Records [Hol10, Fil21]. Recovering [SS97]. recovery [FSSZ23]. Rectangles [HJ08, SS04, AC94]. Recurrence [JKS14, Law95]. Recursive [DF99, Fuc08, Jan19, Lew08, BCL23, Esl21]. reduced [Sch94]. Reduction [DFV97]. Refined [HM22a, Sta99b]. Reformulation [Mat10]. Refutation [COGL07]. Regenerative [Möh07]. Regime [BBW14]. Regions [FP08]. registers [ACH23, SWŻ96]. Regular [AP11b, BSKS11, BL04a, CPT17, COW18, CFR02, CFRR02, CDG07, Cuc07, DKM^{+}15, DW06, DFRS15, Fio01, FG08, FZ00, GG10, Gao14, HN14, HW08, KO05, LMMW09, Lo09, QSZ09, RS07b, RS07a, ST12, SW07a, SW07b, SW99, Wan10, Zha10, Big93, CDG^{+}21, CFM94, CFMR96, GIL22, GL21b, HJ94, HN20, Jan95, dBBR23]. Regularity [FLZ17, FLZ18, Gow06, RS07b, Sco11, FLZ19, HL93, HLS03]. Reimer [Smy13]. Related [AP11a, BP01, Cha09, HHP07, Jan07, Sta97, AKS03, Bon95, CE99]. Relation [KS09]. Relational [CN06]. Relations [Ham08b, MPK98]. Relative [DH10, DH13]. Relaxed [BS09]. Reliability [KSV02, OW02, Sie98, Wag00]. Reliable [CDP96, PU98, Smi93]. Remark [Ale97, ST98]. Removable [GJ99]. Removal [BKT13, FZ22, SS20]. Removing [Mos98]. Rényi [BVH11, DR22, DMS19, LPS08, PSST13, WW23]. Rényi-Like [PSST13]. replica [COKM20]. Representation [GP07, PWZ97]. Representations [May05]. Residues [Sta02a]. Resilience [BSKS11, FLM21]. resilient [Mon20]. Resistance [GLM99, Tel02]. Resolution [GLWW15]. Restricted [AW99, ACM19, BLGN^{+}02, Häg96, HLS00, HHP07, Nik01b, Vat08]. Restrictions [FOT98, RW92]. Result [Den97, Ecc16, Fra17, Had98, KL17, KKO08, Vu01, Had99, Mil94]. Results [BCGR97, CFS14, DK04, GG10, HHP07, LM01, Mub02, ST05, Sob18, AO93, BH93, Erd99, FGK23, GJ03, JLM^{+}03, KT20, TZ21]. Retain [AHK10]. Reversal [LW98]. Reversible [Ber08, CP20, Ing93]. Reversing [BJJ98]. Review [Bár04, Bár06, Ber06, Big07, Bov04, Dud07, Geo08, Hol04, Lea04, McM05, Moh06, Ode06, Sco04, Ste05b]. Revisited [CFGG04, Fri04, LS18a, RS03, Sha03]. Revolutionaries [MP13]. ribbon [AEM22]. Richard [Hol04, Fau12]. Richardson [DH06]. Richter [PRT16]. Riemann [FH23]. Riffle [SGO02]. Right [MA07]. Rings [DP04]. Riordan [ABH22, SSTZ21]. Robber [Meh11]. Robbers [BBKP15, BDFM12]. Robertson [Die95]. Robin [Big07, PV19]. Robust [DO14, Sob18]. Robustness [DPR21]. Rödl [Kuz00]. Roichman [Nao12]. roommates [Pit93]. Root [Csi13, Fra21]. root-mean-squared [Fra21]. rooted [BJ22]. Roots [CK07, CK16, KS04, PT18, Sok04]. Rotating [Alo06a]. Rotational [Edw98]. Rough [Tel02]. Row [KS15]. Rowlinson [CPT17]. RP [Saa96]. Rudich [KSS11]. Rules [BBW14]. Rumour [PPGSS15, AMP20, DPR21]. Run [NS22]. Russo [AM10]. Ruzsa [GJ21a].

safe [Pel94]. Same [FFO13]. Sample [AK14]. Sampled [Wag08]. Samplers [DFLS04, BBD22]. samples [DZ99]. Sampling [BFP10, BM14b, BFJL18, CDG07, EMT18, Gne04, GRV15, GRS20, KCR11, MR06, Möh07]. Sandwiching [KRRS23]. Sankoff [KS09]. Sárközy [McC03, Omi21]. sat [EF95, COGL07, COW18, DM23, GJ03, IK09]. satisfaction [COKM20]. Satisfiability [JKS14, PS16, CDD03]. Satisfiable [KSV09]. Satisfying [COW18]. Saturated [Day17, JP17, Pik99, Pik01]. Saturation [MNS17]. Sauer [KK07b]. Scale [FFV07a, Rio05, ABT99b]. Scale-Free [FFV07a, Rio05]. scale-invariant [ABT99b]. Scan [DGJ08]. Scapellato [Sco04]. scattered [KS03b]. Sceneries [GG19]. scenery [How96]. Schedules [Alo06a]. Schelp [Fau12, GS12b, Omi21]. Scheme [GKW14]. Schemes [KS03d, SS13, Vat08]. Schinzel [CV08]. Schmidt [CV08]. Schur [BGH05]. Schütte [BGHT03]. Science [Lea04]. Sciences [Moh06, Ste05b]. Scott [BT12]. Search [DG10, DRWW10, FD97, Fuc08, Hol10, Pan04, WW05, Wit13, Ald92, LW21, PWZ97]. Searches [Sta99a]. Searching [CKMP07]. secants [BC20]. second [McM05]. Secret [KS03d]. Secretary [KLM02]. Segment [McC04]. segments [PTT21]. Selecting [SS04]. Selection [AS11b, Gne00, Erd99]. selections [ABFK92]. Self [BM14b, FKM01, HP97, MW05a, Alm93, HS95]. Self-Affine [HP97]. Self-Avoiding [BM14b, MW05a, Alm93]. self-avoiding-walk [HS95]. Self-Similarity [FKM01]. Semi [BWY14, FRV13, HMS^{+}08]. Semi-Definite [FRV13]. Semi-Graphoids [HMS^{+}08]. Semi-Strong [BWY14]. Semicomplete [BJJ98]. Semidefinite [CH14]. Semirandom [CO07a]. Sensitivity [Ahl14]. separable [Maa20]. Separation [BDMS14, EKT19]. Separator [FP10, FP14]. Separators [Mat14]. Sequence [AS11b, BKS10, BPS09, BF11, Con17, CF04, Dub08, EKM13, FKS07, GLWW15, KS08, Mån00, MR98b, GL21a, Ham95]. Sequences [Ahl14, AKM^{+}06, Can09, DFR09, EMT18, FK01, Gác04, KS09, McD97, Mer06, PW04, Tok13, TW98, And23, CG22b, Eft20, GIMM23]. Sequential [AS11b, Gne00, FP92]. Ser [Die95]. Sergei [Moh06]. Series [DGN11, KP19]. Series-Parallel [DGN11, KP19]. Set [AFK00, BJ12, BE15, CTY07, Edw98, FKM01, GT09, HZ05, JLR15, KS07, KMS19, Kut02, Nao12, Rei03, SS97, Sub07, AD94, AC93, FGK23, Kur22, PP92, Whi92]. Sets [BMMM18, Bei06, BDG00, BGHT03, BTW12, BL12b, Cal97, CC98, CV08, CO06, CDM14, Cro12, CR13, DKL^{+}16, DŁRR98, EZ12, FT11, FLS08, FFO13, Gal11, GG10, GLS15, GS12a, HLS99, HPV99, HLP08, HMS^{+}13, KRT99, LM13a, MPK98, Mil17, Sch99, SS11, Sha11, Sol06, Wol06, Zha10, BC20, BL04b, BCH22, CE99, Con20, GMT94, HHV21, JPP22, LPP21, RSW93, Sku96, Tyr15]. setting [EMP22, EFK92]. Seven [Edw98]. Seven-Set [Edw98]. Several [CS98, CPV01, KS09, KO07]. Seymour [Die95]. Shadows [BE15]. Shannon [KL16, Tao10]. Shape [DHS16, MPT22]. Shapes [ABS00, DP19]. Sharing [KS03d]. Sharp [BCV16, BLP^{+}21, Bol04, BNMS15, Hec20, KKM18, KSV02]. Shattering [HS04]. Shephard [McM05]. shift [ACH23, SWŻ96]. Shor [Sha03]. Short [AE07, BD02, CFS14, DGK22, FHY18, FKS10, KK08, NS16, Sli96, Die95]. Shortcutting [Tho95]. Shortening [Wol06]. Shortest [HV08, VHV06]. shrink [DZ94]. Shuffle [Mor13]. shuffles [DFP92]. Shuffling [SGO02, Sta02b, Law95]. Sided [Mar01]. Sierpiński [Teu03]. sieves [Gro03]. Sight [BJR09, FKRD09]. Sign [Kun93]. Sign-coherent [Kun93]. Signed [QSZ09]. Signed-Graphic [QSZ09]. Signing [BL04a]. Signings [GHH06]. Silently [BCOE17]. Similarity [FKM01, CP20]. Simon [Hol04]. Simonovits [JW03]. Simple [AFK99, BR00, DD05, GdMNN11, HJ97, Jan09, LP00, ST16, TW97, Vu18, WW05, FKS22, FGOR93, FRS95, Teu03]. Simplex [MR09]. simplified [Fra21]. Simplifying [GMP^{+}15]. Simply [KM17, BJ22]. Simpson [Had99, Had98]. Simulated [NS01]. Simulating [CS06]. Simulation [AD16]. Simultaneously [KO07]. sine [Hir97b]. Singular [PW04, VZ06, JSS22]. Singularity [FKS22, KK01, MM14]. Sipser [BF11]. Site [Siv14, Wie02, DK23, HM22b, Wie95]. Size [AE07, BS19, BBW14, BR00, BKW08, BNMS15, CF04, DG10, DP15, Ecc16, FRR12, Fuc12, GLS15, Geb13, Ham08b, HR11, HKL12, HZ17, HS12, JKOP19, KMS19, Mat03, MR98b, Mut13, Pic11, Pik99, RRR19, Rei03, Ros09, Sch04, VHV06, CMR^{+}21, CNT23, EFRS93, HKŁ95, JPP22, Kur22]. Size-Ramsey [DP15, JKOP19, CMR^{+}21, CNT23, HKŁ95]. Sized [Mec01]. Sizes [ABKS09, BG14a, Fuc08, HS00b, JS07, JR04, Kut02, Lyo17a, CFM92]. Skeletons [Mén18]. Skew [HZ14, MPT22]. Skew-Symmetric [HZ14]. Slices [BG14b]. Slicing [KM05]. Slow [Gác04]. Slowly [GJ02]. Small [AvLR13, CY00, DR99, FS10, GSW11, GRV15, HLS99, KT13, Kor96, KRT99, LM19, ML01, Nao12, Opo02, PRS05, Rio05, Sol06, Vu01, Wag15, BS23, Con20, DK23, HR95, RSW93]. Small-Set [Nao12]. Smallest [AS00, Csi13, GLM99, Sku96, BMS95, JSS22]. Smooth [BG14a, JLM23]. smoother [Spi23]. Sobolev [CY95, Mar19b, Rob03]. Society [Hol04]. sofic [Tim23]. Solid [Pit05]. Solution [MT19]. Solutions [KS06, LP09, Mar19a, Ras19, RW16c, BR19, DFK^{+}93]. Solving [SS06b]. Some [AO93, Bár06, BGH05, BMW14, CTV15, CFS14, Cro12, DPPS00, Fio01, FFV07b, FT98, Gat93, GJ03, Gre03, HR95, Lis96, Man10, MPP04, Mub02, Nik02, Sav98, BH93, CK93, EFK92, Ver95]. Sorting [CTV15]. Sós [BPSS21, JW03, Lon20]. source [AD94]. Sources [CV15, ST96]. Space [BH12, CKZZ15, Die05, DT08, KMS19, Nao12, SW04, PTT21]. Spaces [Cha07, Cza99, Fri04, Hat09, Kel11, San08, AC93, Mar19b, MV23]. Spacing [FH97]. Spanned [VZ06]. Spanning [BKT13, CRS12, CKTV11, CFI^{+}16, DP23, FJ18, GIL22, HL07, JKS13, Kan19, KT13, KSS01, LZ09, Lyo05, LPS08, PW17, PR17a, RSS07, Rio00, Yus17, ABE^{+}22, ANS22, Jan94]. Sparse [AF04, BLS12a, BKHL99, BFL^{+}07, Bor04, Con09, CV10, FK08, FSSZ23, HK14, JT08a, Kan19, KS03e, Lef05, Ngu11, RW10, SS20, SS06b, Sco11, SW18, Vu02, Yus97, ACM19, BR19, BDZ20, Coo92, EGŁ94, FKS22, HKP21, HW19, Kim95, KS03c, RW16a, RW16b, SSTZ21]. Special [BBK^{+}03, DS04, Prö05b]. Specified [McK11, GIM21]. Spectral [BDH22, Chu05, COMV09, CO10, Fio01, Fio03, Gou15, Gre03, Nik09, Wei97, Ing93, WW23]. Speculated [KS09]. Speed [AL09]. Spencer [KK07b]. Sperner [AE07, DGS15, EFK92, FR15, FJKK07]. Sphere [CT07]. Spheres [AS11a, SS04]. Spies [MP13]. Spin [Bov04]. Spiros [Ode06]. Split [CHJ^{+}19, Jan19, Pit02]. Split-and-Merge [Pit02]. Splitter [DK06]. Splitting [Alo06b]. Spread [ABJM14, CS93, Spi23]. Spreading [PPGSS15, AMP20, DPR21, DL94]. Springer [Bár04, Bov04, Dud07, Geo08, Lea04, McM05, Moh06, Ste05b]. Springer-Verlag [Dud07, Moh06]. Square [BF02, HP97, NP20a, Wie95]. squared [Fra21]. Squares [ACÖ13, BRZK17, ACM19]. Srinivasan [KCR11]. Stability [Bou18, BIJ17, Ecc16, Fra17, GHO17, HN18, MR09]. Stable [BL12a, HZ05, LP09, Pit93]. Stack [Atk98]. Staircase [HJ14, SRT10]. Standard [MA07, MPT22]. Star [AHHT10, GRW03, GS12b, RW00]. stars [Ree96]. Stasys [Lea04]. state [dBBR23]. statement [Hyd23]. stationary [ST96]. Statistics [BM14a, GP97, GKW14, KP11, AHKT20]. Stefan [Bár06]. Stein [ER99, Ros09]. Steiner [DG17, EV13, KM05, Pan04]. Step [Ros09]. Steps [AvLR13]. Stevo [Ode06]. Stirling [CST97, KP11]. Stochastic [BV17, Bor02, DG17, Dud07, PWZ97, Wei97, KR94]. Stone [Nik09, BN12]. Straight [ESS09]. stranded [ABH22]. Strategies [CFKL12]. strategy [Csi92, Pek96]. stretch [BS23]. strict [Bry92]. String [AKM00, Eri97, FP10, FP14, Mat14, AMP20, LS95, ST96]. Strings [CCM99, Eri97, Mån02a]. Stripes [GS12b]. Strong [AGSW22, BWY14, BJ18, COGL07, CW00, Cur12, FJKK07, Hax04, LS08, PP14, LSM19]. Stronger [BJ18]. Strongly [CV15, CF04, Fio01, SS14, RM92]. Structural [WL01]. Structure [BLGN^{+}02, Bra99, Can09, Cha07, KL17, LCF14, PR17b, Häg92, JPP22, Neš99, OS20]. structured [Nic23]. Structures [BG14a, CN06, DD05, DFLS04, Hol04, Hwa98, Man02b, Neš05, Spe05, Ste05b, Tyr15, ABT99a, GJ03]. Studying [Mit99]. Sub [Wol11]. Sub-Gaussian [Wol11]. Subcritical [BPS09, BBW14, GW18, Tur11]. Subcube [JM13]. Subcubic [Moh16]. Subdivision [BT12, DM11]. Subdivisions [Had98, KO04b, LSS17, Had99, JQ23, Tho96a]. Subgraph [FK13, Gao14, GIM21, HJW22, SW18, ABHM21, BCJ23, CFM92, KSSC99]. Subgraphs [ABKR18, AS00, AS06, ABKS09, BKT13, CY00, DHL^{+}12, EKT19, Far06, FS10, FS12, FFO13, FÖ09, HMS^{+}13, JP17, Kan19, KS03a, KN14a, LMMW09, McK11, NT18, PR17a, Rei02, Rio00, ST98, Sie98, SS03, Vu01, AW23a, ABD23, Eft20, EKK21, EM22, GS23, GMT20, Jin92, Kri94, Liu22, LTZ22, MW22, Sco92]. sublinear [HHWY23]. Submap [BCGR97]. submatroids [Kor96]. Subposet [BJ12, GL09, MP17]. Subquadratic [PW16]. Subsequence [AS11b, Ham03, HQ05]. Subsequences [BHL^{+}15, Dan98, KS15, LS18b, ST01, DZ99]. Subset [HLS99, Heg96, Tra04]. Subsets [BNMS15, Ell11, Fio03, GL09, KKK12, LM14, Mat03, Ngu11, Pon13a]. Subspace [BBDM23]. Substitution [Spe05, Wie95]. Subsystems [KRT99]. Subtournaments [Lon17]. Subtree [DG10, Fuc08, Fuc12]. successive [Sch94]. Sufficient [DH00, KKY06]. sufficiently [GIMM23]. Sum [BNMS15, CC98, HLS99, HR11, Sch99, Sha16, CE99]. Sum-Free [BNMS15, CC98, Sch99, CE99]. Sums [BCV16, Buk08, CHS09, EZ12, GH98, GGX16, GGH12, HLS00, Ham03, KS02, LM19, Pla11a, Pon13b, Wag08, Ham95, Heg96]. Sumset [Tao10]. Sumsets [CLS13, SS11, Sol06]. Sunflowers [MW18]. Super [Can18]. Super-Exponential [Can18]. Supercritical [BG14a, DKLP10, DK23]. Supersaturation [JY20]. Supersequences [Dan98]. supposedly [Ver95]. Surface [Gou15, HJW22]. Surfaces [Cha09, Kru11, Moh06, Sli02, GRS20]. Surjectivity [NP20a]. Surprise [NPZ17]. SVD [Vu18]. Swap [EKM13, EMT18]. Swap-Distances [EKM13]. switched [MV23]. Symbol [FM14]. Symbolic [LCF14]. symbols [KY20]. Symmetric [Blo05, Ham08b, HZ14, HM15, MN09, Qin01, VV97, BH92, COKM20, EKNS21, JSS22, ŁP93, Mil23]. symmetrical [BS92]. Symmetrization [FM17, KM05]. Symmetry [Edw98, FGLP00, MW05b, SRT10]. Synchronous [CKMP07]. System [FRR12]. Systematic [DGJ08]. Systems [AAE^{+}97, Ano10k, AE07, BE15, EV13, FS05, FPS05, JLR15, KNR03, KRT99, Kut02, LLMO09, SS97, BR19, FGK23]. Szarek [Ale97]. Szekeres [ST01]. Szemerédi [KK08, BKLY13, BCPT21, CDKM15, GJ21a, LM13b, Sco11, Tre15].

T. [GLWW15]. Table [Bar10, Had98, Had99]. Tableaux [AR01, HJ14, MPT22]. Tables [Bar08]. Tail [LW21, SV03, Wag08, Zha17]. Tails [Wol11]. Talagrand [Ale97, Bov04, Dud07]. Tale [ABT06]. Tame [CV15]. tangencies [Zah20]. Targets [SW14]. Task [Gác00]. Tavaré [Hol04]. Teasing [SS07]. technique [Sha03]. Techniques [CO10]. Teller [Far07]. Tensor [DH13]. tensors [CG22a]. Term [Ngu11]. Terms [VS06, HM22b]. Testable [AS06, AF15]. Testing [AHW16, BGRS12, CS10, FW18, LV13, ST05, COGHKL21, Pel94]. Tests [PU98, Pel94, PP92]. Tetrahedra [DT08]. tetrahedron [FRMZ21]. Texts [McM05, RR03]. th [KS04, CF95]. Their [GMP19, Moh06, Dan98, FLZ17, FLZ18, MO01, OW02]. Theorem [AH17, BLS12a, BKLY13, Bár04, Bei06, BK11a, BN12, CDKM15, DGS15, FP10, FP14, GT09, HOO98, Ham08a, HN18, KK07b, KN14a, KRS04, LM14, LM13b, Nik09, Rei03, RRS06, Škr99a, Tre15, Vu97, Wal07, AD94, ABE^{+}22, BCPT21, Bro03, CF92, CM93, CDG^{+}21, Die95, Fil20, Kim95, KS03b, Law93, McC03, NS20, NV94, NP20b, Pud03, RW19, Sli96, Tet94, Ale97, Buk16, Chu05, FJKK07, HS98, KK08, NS16, San08]. Theorems [ABE14, AM14, BCOK14b, DPPS00, DK06, Fuc12, Gla16, JT08b, Wag15, BH93, DR22, EHS^{+}94, RS22]. theoretic [CK93, EFK92, Lyn93]. Theory [AW99, Big07, BBK^{+}03, DH10, Die95, Gre03, GT09, Hir97b, Jan07, Tao10, TW98, Big93, Bro03, DT94, FW03, Whi94]. theta [BT20]. thickness [Coo92]. Thin [AFK00, FKM01]. Thomassen [PRT16]. Thorp [Mor13]. Three [AS11a, APST04, ABT06, DT08, ES11, EJ14, Gne04, HHK^{+}12, HMS^{+}08, Jan99, KMSS12, Ngu11, SS16, SW92, VS06, Fra21, HM22b, Lem94, Pin20, Ree96]. Three-Space [DT08]. Three-Term [Ngu11]. Threshold [AY93, BNMS15, DLM04, Gal11, GH17, GNS17, HLTZ17, KC93, KSV02, Lyn93, Mar19a, MW05b, MW07, McC04, PS16, CDD03, HK94]. Thresholds [BBH^{+}16, FKP16, GW15, KK07a, MO01, SW92]. Tied [TW97]. Tight [AHL21, BK16, DRWW10, HLP^{+}09, Mub16, Wit13, AKPP21, AMP20, HLSM21, PSS23]. Tiling [GHZ19, HLSM21, NS20]. Tilings [BMMM18, BTW19, CLS19, MPT22]. Time [ABE14, Ald00, AP02, BDNP11, BKL^{+}06, BBY08, Bev16, Can18, CO06, DGGP14, Jon98, Lew08, LW98, Mor13, SS06b, Wit13, DFK^{+}93, DHJM21, NS22, NT93]. Times [AW99, Jan99, LSV12, Pim11, Ren10, SW14, And23, LW21, Mór92]. Todorcevic [Ode06]. Toll [Wag15]. Top [Sta02b, DFP92]. Topics [Big07, Sco04]. Topological [Bár04, EMM15, KS94, KS96b, KO04a, Sli02, Whi94]. Topologies [MO01]. Topology [PPGSS15]. Torus [Alo09a, DF14, ST09, Siv14]. Total [DKW14, DF99, JMŠ98, Sta97, Sta99b, CH96]. Tour [TA97]. Tournament [AL11, BGHT03, CGL16, FK05, Yus17]. Tournaments [BJGY97, BGQ97, CTY07, Cuc07, BHLS20, DDF^{+}21]. Trails [AFST20, Bal03]. Transcendental [MT19]. Transformations [Pit02]. Transforms [MC04]. transience [Law95]. Transition [GK04, Rio12, Neu96, SCW96]. transitions [Joh22]. Transitive [Luk98, Sta03, SC95]. Transmission [CDS14]. Transmissions [Alo06a]. Transportation [ST09]. Transversal [EM22, KY20]. Transversals [HS06, Yus97, EGŁ94, GH20b]. Tree [BD02, CFI^{+}16, CKMP07, DL14, FM14, GP19, Han97, HL07, Jan21, KS02, KW10, Lew08, Lyo10, Mór05, Nob98, PW17, RSS07, Sta99a, Wag15, Wan10, Ald92, ANS22, Bez94, Esl21, ISZ22, dBBR23]. Tree-Decompositions [BD02]. Tree-Like [Wan10]. Tree-Width [Nob98]. treedepth [NSW23]. Trees [Ald98, BPS18, BPP06, BP01, BFL^{+}07, BEHD19, CHJ^{+}19, CFGG04, CRS12, CDKK08, DG17, DF99, Dob02a, Dob02b, DT12, FD97, FJ18, Fuc08, Fuc12, Hat09, Hol10, Jan19, JKS13, KSS01, KLM02, LZ09, Lyo05, LPS08, Lyo17b, MS13, MPP04, Nei02, Pan04, SS07, VHV06, Yus04, BJ22, BPSS21, BHLS20, DHS16, DG10, DGZ23, DF96, Edw95, Edw96, Esl21, FGOR93, FW03, GIL22, Jan94, Jan21, JS20, Joh22, KP22, Kub96, MS20, RW19].

Triads [Wu98]. trialities [AEM22]. Trials [BCV16]. Triangle [ABKR18, BMMM18, BT12, Bra99, CJK97, Con17, FRMZ21, Gyo06, Jin93, JS02, KN99, Kri05, Kri97, MW22, Ste05a, Yus12, BDZ20, Tyo21, dFdMR94]. Triangle-degrees [FRMZ21]. Triangle-Free [ABKR18, BT12, Bra99, CJK97, Con17, Gyo06, KN99, Ste05a, Jin93, MW22]. Triangle-Freeness [JS02]. Triangle-Tilings [BMMM18]. Triangles [BRZK17, BPSS23, GL12b, HT16, KLMP19, Lo09, PR17b, Raz08, Wol11, BLP^{+}21, FGK94, LNW21, Tyo21]. Triangular [FM17, GL18, Yeh20]. Triangulation [Mén18]. Triangulations [DT17, Pim11]. Trie [CV15]. Tries [BD08, Sch04]. Tripartite [HN18, GLLV22]. Triple [FRR12, FS05, FPS05, KNR03]. tripling [Con20]. trivial [Liu22]. TSP [Aza94]. tuner [BBD22]. Tuning [BBD22]. tuples [GL21a]. Turán [EHS^{+}94, AKS03, BP12, BT94, BLM11, BIJ17, BT20, BK11b, Chu05, CCGJ18, CN12, EGM19, FRV13, GIM12, GJ21a, JMY22, JQ23, JM13, KMSV07, KM12, KL17, KRS04, LTTZ18, Mat97, Mub02, PSS23, TZ21]. Turán-Type [KL17, TZ21]. Tutte [ABH22, BR99, DH10, DH13, EMT06, EMM15, Far07, GN06, GdMNN11, GKRV18, Hli06, Man10, Mph00, Nob98, OP18, Qin01, Tra04, VW92, Wel94]. Tuza [BK16, BDZ20]. Tverberg [Sob18]. Twisted [FGY00]. Two [All08, AE07, BH01, BD02, BCGR97, BS09, Ber08, BKN05, DH06, DD05, Dob06, DF09b, DK04, FR01, Gri97, GS12b, Jan99, KL17, LSV12, ŁRS98, Mar01, MN09, Pin17, SS07, Tho04, GHSS22, Lem22, Wie92]. Two-Colourability [BS09]. Two-Coloured [ŁRS98]. Two-Dimensional [DF09b, Lem22]. Two-Edge-Coloured [All08]. Two-Faced [DD05]. Two-Parameter [Ber08]. Two-Part [AE07]. Two-Sided [Mar01]. Two-Type [DH06]. Type [Ben18, DH06, Doh02, KL17, KKO08, KKM18, RRS06, Škr99a, AKS03, Gla16, HRRS21, KT20, KY07, RS22, TZ21, NV94]. types [KS03b]. Typical [KL17, War16]. Typically [PSY13].

Ulam [Bár04]. ULD [FK09]. ULD-Lattices [FK09]. Unavoidable [KM23, Lon17, BDS21]. Unbiased [CLS19]. Unbounded [DH06]. Uncapacitated [FFV07b]. Unicyclic [Jan03]. Uniform [BCV16, BCOK14a, BCOK15, CP16, DKR18, FOT98, FK17, GH17, Gow06, GS11, Han15, HLP^{+}09, KKL12, KS18, Mén18, RRS06, AHL21, ANS22, CFMR96, GIL22, GIMM23, JŁ92, PRS05, RSW93, Sch23]. Uniformly [Alo18, BK06a]. Unimodal [AS11b]. Unimodular [Mou00, Tim23]. Union [Ecc16, KL16, PW16, EL19]. Union-Closed [Ecc16]. Unions [CS98, FLS08]. Unique [RW16c]. Uniquely [EWHK98]. Uniqueness [Mar19a, dBBR23]. Unit [DT08, ESS09, KMSS12]. Unity [KS04]. Univariate [Von15]. Universal [BDS21, DFR09, FHT23, JKS13, Sch99, HK94]. Universality [HT16, HM22a]. University [Ber06, Sco04]. Unknown [Pik10]. Unlabelled [Hwa98, Stu20, JS20]. Unreliable [ACG^{+}04]. unrooted [BJ22]. unsatisfiable [GJ03]. Unsaturated [SS14]. Unusually [DR22]. Upper [Alm93, AP02, BS19, Bón14, Dud07, MSS09b, Mat97, MWW02, Uzz19, Vu02, Wag06b, Wie02, Wie03, Wu97, CF92, CH23, CS95, CH96, Für96, GSS96, Lam23]. Urn [MC04, Hir97a]. Urns [KN14b]. Used [FM14]. Using [AST93, Bár04, Gre12, MW05b, ST09, Eft20, GRS20, LW21].

V [McM05]. Valid [ABIST96]. Value [Ald98, AK14, Sie98, JSS22]. Valued [Nao12]. Values [AKM^{+}06, VS06, FH23]. Vanish [KS04]. Variable [JR04, LW21]. Variables [AK14, ER99, GP97, Hat09, Kel11, TW98, Wag08, MS95, Mat95, Mat99]. Variation [Sta97, Sta99b]. Variational [Zha17]. variations [LS95]. Variety [PW04]. VC [BDG00]. VC-Dimension [BDG00]. Vector [Cza99, San08]. Vectors [Cal97, Can09, Dub08, VZ06, EKNS21, Wag00]. Venn [Edw98]. Verlag [Bár06, Dud07, Moh06]. Version [AH17, BKLY13, LM13b, BCPT21, Bro03, LP20, LP23]. Versions [Tre15]. Versus [GS12b, Mub16, Tim12, AW23b]. Vertex [Coo06, DD10, Fio03, Hax01, Jag98, JS08, Luk98, Ver02, Wu98, CDK19, HPS93, Mad94, Sau03, Smi93]. Vertex-Bipartition [DD10]. Vertex-Disjoint [Ver02]. vertex-edge-critically [Mad94]. Vertex-Transitive [Luk98]. Vertex-Triads [Wu98]. vertices [Mad95]. Very [FK01]. Via [DKLP10, KN14b, MM14, Ars21, Bar08, BFPS11, BR16, BIJ17, CH14, CP20, CO10, DR22, DP19, DT94, Doh02, ER99, FJZ22, GT09, GL21c, Hir97a, HT01, JQ23, KK09, PWZ97, RW16c, Sob18, Yeh20]. Vicarious [LT10]. view [RS22]. Vieweg [Bár06]. Viola’ [Sha17a, Sha17b]. Virtual [DH10]. Visibility [Alo18]. Visual [KS03d]. Vol [McM05, Ste05b]. Volume [Ano92a, Ano93a, Ano94a, Ano95a, Ano96a, Ano92c, Ano92b, Ano92d, Ano92e, Ano92g, Ano92f, Ano92h, Ano92i, Ano93c, Ano93b, Ano93d, Ano93e, Ano93g, Ano93f, Ano93i, Ano93h, Ano94c, Ano94b, Ano94e, Ano94d, Ano94g, Ano94f, Ano94i, Ano94h, Ano95c, Ano95b, Ano95d, Ano95e, Ano95g, Ano95f, Ano95i, Ano95h, Ano96c, Ano96b, Ano96e, Ano96d, Ano96g, Ano96f, Ano96h, Ano96i, Ano09b, Ano09a, Ano09e, Ano09c, Ano09f, Ano09d, Ano09i, Ano09g, Ano09j, Ano09h, Ano09m, Ano09k, Ano09n, Ano09l, Ano09p, Ano09o, Ano10b, Ano10a, Ano10d, Ano10c, Ano10f, Ano10e, Ano10h, Ano10g, Ano10j, Ano10i, Ano11b, Ano11a, Ano11d, Ano11c, Ano11f, Ano11e, Ano11g, Ano11h, Ano11j, Ano11i, Ano11l, Ano11k, Ano12b]. volume [Ano12a, Ano12d, Ano12c, Ano12f, Ano12e, Ano12h, Ano12g, Ano12j, Ano12i, Ano13b, Ano13a, Ano13d, Ano13c, Ano13f, Ano13e, Ano13h, Ano13g, Ano13j, Ano13i, Ano13l, Ano13k, Ano14b, Ano14a, Ano14d, Ano14c, Ano14f, Ano14e, Ano14h, Ano14g, Ano14j, Ano14i, Ano14l, Ano14k]. Volumes [DT08, Mén18]. vs [LS09a].

W [Big07]. Waerden [RS22, Tyr15]. Waiters [BBHKŁ16]. Waiting [BKL^{+}06, Mór92]. Walk [AP08, CS06, Eva04, GW14, God92, LP00, OS14, TW97, Wag08, HS95, Teu03]. Walkers [EGPN16]. Walks [AP11a, AvLR13, AAK^{+}11, BM14b, CL05, DF09b, Gou15, Gre97, Häg98, Jon98, KLM06, MW05a, Mar01, NT16, Sta03, Wan10, Woe05, Alm93, And23, GHSS22, How96, MV23, RSS23]. Walter [Prö05a, Prö05b]. Ward [GLWW15]. Watson [BP01, CHJ^{+}19, DGZ23, Joh22, MS20]. Way [Gra98, Yus04]. WDM [AHHT10]. weak [ABFK92, FLZ19]. Weakening [Kaw08]. Weakly [Pik01, Wag06a, ABH22]. Web [Coo06]. Web-Graphs [Coo06]. Weight [ABBL10, Ald98, Cal97, HV08, MW13, VHV06]. Weighted [BKLY13, BR19, FJ18, HQ05, Ham08a, HJ14, KL17, Mut13, RSS07, Tok13, Tra10, Ham95, MPT22]. Weights [BD08, HV08, Jan99, KW10, Sta99a, VHV06, Ald92]. Welcome [Ano15, Ano16]. Welsh [Alm98]. wheel [FGK94]. wheel-free [FGK94]. Which [BH01]. Whitney [Far07, Man10]. Whole [Sok04]. whose [Dob02a]. Wide [FHR09]. Widom [CPT17]. Width [BH01, Hli06, IT11, Nob98, Die95]. Wiener [Nei02]. Wilson [Big07, AEM22, HS00a]. Window [FV17, JS07, Cou22]. Winkler [Gác04]. Winning [FK08, Pek96]. Wins [Mát07]. wise [Gow96]. Without [BMMM18, BJ17a, BJ17b, FHY18, FP08, FKS10, KP18, KO04a, LPRS09, Mar13, NV05, PV19, Kri94, Liu22, Mye02, WL01]. Word [MA07]. Words [CPV01, RR03, Thi16, Sta95]. Work [Prö05a]. Workshop [Ano99]. Worst [FM14]. Worst-Case [FM14]. written [McM05].

XOR [CDD03]. XORSAT [PS16]. xvii [Lea04].

Y. [Baj16]. Yahya [Pla11b]. Yang [Lem22]. Yields [GSW11]. Young [AR01].

Zarankiewicz [Für96]. Zero [EHJ98, FG11, GH98, Jac07, Tho97, Jac93]. Zero-Free [EHJ98, Jac07, Tho97, Jac93]. Zeros [ADH^{+}19, BCPS18, Bor06, DK04, DK08, FP08, GMP19, OP18, Sok01, Wag00]. zeta [FH23]. Ziegler [McM05]. Zigzag [DD05]. Ziv [HOO98]. Zvonkin [Moh06].

References

Alon:2007:ECD

[AA07] Noga Alon and Vera Asodi. Edge colouring with delays. Combinatorics, Probability and Computing, 16(2):173–191, March 2007. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

            (electronic). URL http://journals.cambridge.org/action/
            displayIssue?jid=CPC&volumeId=16&issueId=02.

Ahlswede:1997:IS

[AAE^{+}97] R. Ahlswede, N. Alon, P. L. Erdős, M. Ruszinkó, and L. A. Székely. Intersecting systems. Combinatorics, Probability and Computing, 6(2):127–137, June 1997. CODEN CPCOFG.

ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=6&issueId=02.

Alon:2011:MRW

[AAK^{+}11] Noga Alon, Chen Avin, Michal Koucký, Gady Kozma, Zvi Lotker, and Mark R. Tuttle. Many random walks are faster than one. Combinatorics, Probability and Computing, 20(4):481–502, July 2011. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anthony:2000:FLI

[AB00] Martin Anthony and Peter L. Bartlett. Function learning from interpolation. Combinatorics, Probability and Computing, 9(3):213–225, May 2000. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http://journals.

cambridge.org/action/displayIssue?jid=CPC&volumeId=9&
issueId=03.

Alon:2009:I

[AB09] Noga Alon and Bela Bollobas. Introduction. Combinatorics, Probability and Computing, 18(1–2):1, March 2009. CODEN CP-COFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Alon:2010:I

[AB10] Noga Alon and Béla Bollobás. Introduction. Combinatorics, Probability and Computing, 19(5–6):641, November 2010. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Addario-Berry:2010:LMW

[ABBL10] Louigi Addario-Berry, Nicolas Broutin, and Gábor Lugosi. The longest minimum-weight path in a complete graph. Combinatorics, Probability and Computing, 19(1):1–19, January 2010. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Alon:2021:MPC

[ABC^{+}21] Noga Alon, Raimundo Briceño, Nishant Chandgotia, Alexander Magazinov, and Yinon Spinka. Mixing properties of colourings of the Z^{d} lattice. Combinatorics, Probability and Computing, 30(3): 360–373, May 2021. CODEN CPCOFG. ISSN 0963-5483 (print),

1469-2163 (electronic). URL https://www.cambridge.org/
core/journals/combinatorics-probability-and-computing/

article/mixing-properties-of-colourings-of-the-d-lattice/
70C1934D455244CE425346DA38793052.

Anderson:2023:CGF

[ABD23] James Anderson, Anton Bernshteyn, and Abhishek Dhawan. Colouring graphs with forbidden bipartite subgraphs. Combinatorics, Probability and Computing, 32(1):45–67, January 2023. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/colouring-
graphs-with-forbidden-bipartite-subgraphs/CEA9F0C38B954C6343406893669AA47C.

Addario-Berry:2014:HTT

[ABE14] Louigi Addario-Berry and Laura Eslava. Hitting time theorems for random matrices. Combinatorics, Probability and Computing, 23(5):635–669, September 2014. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). Special issue on Honouring the Memory of Philippe Flajolet - Part 1.

Acan:2021:PEO

[ABE^{+}21] Huseyin Acan, Charles Burnette, Sean Eberhard, Eric Schmutz, and James Thomas. Permutations with equal orders. Combinatorics, Probability and Computing, 30(5):800–810, September 2021. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL https://www.cambridge.org/core/
journals/combinatorics-probability-and-computing/article/
permutations-with-equal-orders/E73A1AEDF5DD5CDC16EA2A5427C87A0E.

Allen:2022:SBT

[ABE^{+}22] Peter Allen, Julia Böttcher, Julia Ehrenmüller, Jakob Schnitzer, and Anusch Taraz. A spanning bandwidth theorem in random graphs. Combinatorics, Probability and Computing, 31(4):598– 628, July 2022. CODEN CPCOFG. ISSN 0963-5483 (print),

1469-2163 (electronic). URL https://www.cambridge.org/
core/journals/combinatorics-probability-and-computing/
article/spanning-bandwidth-theorem-in-random-graphs/
CFC205325B6FB60E6A420F82F6BF8A06.

Alon:1992:PSW

[ABFK92] Noga Alon, Imre Bárány, Zoltán Füredi, and Daniel J. Kleitman. Point selections and weak ϵ-nets for convex hulls. Combinatorics, Probability and Computing, 1(3):189–200, September 1992. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Backhausz:2018:CBD

[áBGHV18] ágnes Backhausz, Balázs Gerencsér, Viktor Harangi, and Máté Vizer. Correlation bounds for distant parts of factor of IID processes. Combinatorics, Probability and Computing, 27(1):1–20, January 2018. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL https://www.cambridge.org/core/
product/6BB584C6997C00A89484AC2CE76EFC32.

Avohou:2022:ETB

[ABH22] Remi C. Avohou, Joseph Ben Geloun, and Mahouton N. Hounkonnou. Extending the Tutte and Bollobás–Riordan polynomials to rank 3 weakly coloured stranded graphs. Combinatorics, Probability and Computing, 31(3):507–549, May 2022. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/extending-
the-tutte-and-bollobasriordan-polynomials-to-rank-3-
weakly-coloured-stranded-graphs/8F31EB8200D21D8C997D1D35407A94C0.

Alweiss:2021:SQP

[ABHM21] Ryan Alweiss, Chady Ben Hamida, Xiaoyu He, and Alexander Moreira. On the subgraph query problem. Combinatorics, Probability and Computing, 30(1):1–16, January 2021. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/on-the-
subgraph-query-problem/9CF0AC38D8234446F39D8245A4A037B7.

Anthony:1996:VGA

[ABIST96] Martin Anthony, Peter Bartlett, Yuval Ishai, and John Shawe-Taylor. Valid generalisation from approximate interpolation. Combinatorics, Probability and Computing, 5(3):191–214, September 1996. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Addario-Berry:2014:SRG

[ABJM14] Louigi Addario-Berry, Svante Janson, and Colin McDiarmid. On the spread of random graphs. Combinatorics, Probability and Computing, 23(4):477–504, July 2014. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Allen:2018:TFS

[ABKR18] Peter Allen, Julia Böttcher, Yoshiharu Kohayakawa, and Barnaby Roberts. Triangle-free subgraphs of random graphs. Combinatorics, Probability and Computing, 27(2):141–161, March 2018. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/trianglefree-
subgraphs-of-random-graphs/E12FFAC201496968403DC5A98BBF04F5.

Alon:2009:SIS

[ABKS09] Noga Alon, József Balogh, Alexandr Kostochka, and Wojciech Samotij. Sizes of induced subgraphs of Ramsey graphs. Combinatorics, Probability and Computing, 18(4):459–476, July 2009. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Abrams:2013:DOP

[ABL^{+}13] Aaron Abrams, Eric Babson, Henry Landau, Zeph Landau, and James Pommersheim. Distributions of order patterns of interval maps. Combinatorics, Probability and Computing, 22(3):319– 341, ???? 2013. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Addario-Berry:2012:CBA

[ABMR12] L. Addario-Berry, C. Mcdiarmid, and B. Reed. Connectivity for bridge-addable monotone graph classes. Combinatorics, Probability and Computing, 21(6):803–815, November 2012. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Alon:2000:PFS

[ABS00] Noga Alon, Miklós Bóna, and Joel Spencer. Packing Ferrers shapes. Combinatorics, Probability and Computing, 9 (3):205–211, May 2000. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http://journals.

cambridge.org/action/displayIssue?jid=CPC&volumeId=9&
issueId=03.

Arratia:1999:PDL

[ABT99a] Richard Arratia, A. D. Barbour, and Simon Tavaré. On Poisson– Dirichlet limits for random decomposable combinatorial structures. Combinatorics, Probability and Computing, 8(3):193–208, May 1999. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL http://journals.cambridge.org/
action/displayIssue?jid=CPC&volumeId=8&issueId=03.

Arratia:1999:PDD

[ABT99b] Richard Arratia, A. D. Barbour, and Simon Tavaré. The Poisson–Dirichlet distribution and the scale-invariant Poisson process. Combinatorics, Probability and Computing, 8(5): 407–416, September 1999. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http://journals.

cambridge.org/action/displayIssue?jid=CPC&volumeId=8&
issueId=05.

Arratia:2006:TTC

[ABT06] Richard Arratia, A. D. Barbour, and Simon Tavaré. A tale of three couplings: Poisson–Dirichlet and GEM approximations for random permutations. Combinatorics, Probability and Computing, 15(1–2):31–62, January 2006. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=15&issueId=01.

Ahlswede:1993:ESP

[AC93] Rudolf Ahlswede and Ning Cai. On extremal set partitions in Cartesian product spaces. Combinatorics, Probability and Computing, 2(3):211–220, September 1993. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Ahlswede:1994:PPP

[AC94] Rudolf Ahlswede and Ning Cai. On partitioning and packing products with rectangles. Combinatorics, Probability and Computing, 3(4):429–434, December 1994. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Ahlswede:1999:CKC

[AC99] Rudolf Ahlswede and Ning Cai. A counterexample to Kleitman’s Conjecture concerning an edge-isoperimetric problem. Combinatorics, Probability and Computing, 8(4):301–305, July 1999. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=8&issueId=04. Random

graphs and combinatorial structures (Oberwolfach, 1997).

Alonso:2004:QUC

[ACG^{+}04] Laurent Alonso, Philippe Chassaing, Florent Gillet, Svante Janson, Edward M. Reingold, and René Schott. Quicksort with unreliable comparisons: a probabilistic analysis. Combinatorics, Probability and Computing, 13(4–5):419–449, July 2004. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=13&issueId=04.

Arratia:2023:RFS

[ACH23] Richard A. Arratia, E. Rodney Canfield, and Alfred W. Hales. Random feedback shift registers and the limit distribution for largest cycle lengths. Combinatorics, Probability and Computing, 32(4):559–593, July 2023. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/random-
feedback-shift-registers-and-the-limit-distribution-
for-largest-cycle-lengths/71EF5787EDBBD2A83DF6331DC98620FC.

Andren:2019:RCS

[ACM19] Lina J. Andrén, Carl Johan Casselgren, and Klas Markström. Restricted completion of sparse partial Latin squares. Combinatorics, Probability and Computing, 28(5):675–695, September 2019. CODEN CPCOFG. ISSN 0963-5483 (print),

1469-2163 (electronic). URL /core/journals/combinatorics-
probability-and-computing/article/restricted-completion-
of-sparse-partial-latin-squares/E865D1C0D0A778062F66F730558246C1.

Andren:2013:AAO

[ACÖ13] Lina J. Andrén, Carl Johan Casselgren, and Lars-Daniel Öhman. Avoiding arrays of odd order by Latin squares. Combinatorics, Probability and Computing, 22(2):184–212, ???? 2013. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Aharoni:1994:MTC

[AD94] Ron Aharoni and Reinhard Diestel. Menger’s theorem for a countable source set. Combinatorics, Probability and Computing, 3(2):145–156, June 1994. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Arratia:2016:PDC

[AD16] Richard Arratia and Stephen DeSalvo. Probabilistic divide-and-conquer: A new exact simulation method, with integer partitions as an example. Combinatorics, Probability and Computing, 25(3): 324–351, May 2016. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Aumuller:2019:DPQ

[ADH^{+}19] Martin Aumüller, Martin Dietzfelbinger, Clemens Heuberger, Daniel Krenn, and Helmut Prodinger. Dual-pivot Quicksort: Optimality, analysis and zeros of associated lattice paths. Combinatorics, Probability and Computing, 28(4):485–518, July 2019. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/dualpivot-
quicksort-optimality-analysis-and-zeros-of-associated-
lattice-paths/180377D10918C761E5FA26BE1ACE58BC.

Arrowsmith:1994:RPP

[AE94] D. K. Arrowsmith and J. W. Essam. Reciprocity and polynomial properties for even flows and potentials on directed graphs. Combinatorics, Probability and Computing, 3(1):1–11, March 1994. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Aydinian:2007:AMS

[AE07] Harout Aydinian and Péter L. Erdős. All maximum size twopart Sperner systems: In short. Combinatorics, Probability and Computing, 16(4):553–555, July 2007. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=16&issueId=04.

Abrams:2022:NDO

[AEM22] Lowell Abrams and Joanna A. Ellis-Monaghan. New dualities from old: generating geometric, Petrie, and Wilson dualities and trialities of ribbon graphs. Combinatorics, Probability and Computing, 31(4):574–597, July 2022. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/new-
dualities-from-old-generating-geometric-petrie-and-wilson-
dualities-and-trialities-of-ribbon-graphs/CC19789CBD9372C87A970D388751B23B.

Atkinson:2004:INS

[AF04] Geoffrey Atkinson and Alan Frieze. On the b-independence number of sparse random graphs. Combinatorics, Probability and Computing, 13(3):295–309, May 2004. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=13&issueId=03.

Alon:2015:ETG

[AF15] Noga Alon and Jacob Fox. Easily testable graph properties. Combinatorics, Probability and Computing, 24(4):646–657, July 2015. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). Special issue on Oberwolfach Special Issue Part 1.

Alt:1999:NSC

[AFK99] Helmut Alt, Ulrich Fuchs, and Klaus Kriegel. On the number of simple cycles in planar graphs. Combinatorics, Probability and Computing, 8(5):397–405, September 1999. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=8&issueId=05.

Alon:2000:LTS

[AFK00] Noga Alon, Emanuela Fachini, and János Körner. Locally thin set families. Combinatorics, Probability and Computing, 9(6): 481–488, November 2000. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http://journals.

cambridge.org/action/displayIssue?jid=CPC&volumeId=9&
issueId=06.

Angel:2020:LMT

[AFST20] Omer Angel, Asaf Ferber, Benny Sudakov, and Vincent Tassion. Long monotone trails in random edge-labellings of random graphs. Combinatorics, Probability and Computing, 29(1):22–30, January 2020. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL https://www.cambridge.org/core/
journals/combinatorics-probability-and-computing/article/
long-monotone-trails-in-random-edgelabellings-of-random-
graphs/7DA12B9041407DFC8C86D01F21EAE25F.

Arya:1999:EDR

[AGM99] Sunil Arya, Mordecai J. Golin, and Kurt Mehlhorn. On the expected depth of random circuits. Combinatorics, Probability and Computing, 8(3):209–228, May 1999. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=8&issueId=03.

Axenovich:2022:SCM

[AGSW22] Maria Axenovich, António Girão, Richard Snyder, and Lea Weber. Strong complete minors in digraphs. Combinatorics, Probability and Computing, 31(3):489–506, May 2022. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/strong-
complete-minors-in-digraphs/8700C8DFB647BB036B70141BD5491C3E.

Aharoni:2017:RPV

[AH17] Ron Aharoni and David Howard. A rainbow r-partite version of the Erdős–Ko–Rado theorem. Combinatorics, Probability and Computing, 26(3):321–337, May 2017. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/product/
645E07E1715A0CA2E98E64D039E1451C.

Alishahi:2020:CNM

[AH20] Meysam Alishahi and Hajiabolhassan Hossein. On the chromatic number of matching Kneser graphs. Combinatorics, Probability and Computing, 29(1):1–21, January 2020. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/on-the-
chromatic-number-of-matching-kneser-graphs/520DDDC21F77F9DA847AAED4752C7837.

Amini:2010:WDS

[AHHT10] Omid Amini, Frédéric Havet, Florian Huc, and Stéphan Thomassé. WDM and directed star arboricity. Combinatorics, Probability and Computing, 19(2):161–182, March 2010. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Alon:2010:PRA

[AHK10] Noga Alon, Dan Hefetz, and Michael Krivelevich. Playing to retain the advantage. Combinatorics, Probability and Computing, 19(4):481–491, July 2010. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Alon:2020:ESL

[AHKT20] Noga Alon, Dan Hefetz, Michael Krivelevich, and Mykhaylo Tyomkyn. Edge-statistics on large graphs. Combinatorics, Probability and Computing, 29(2):163–189, March 2020. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/edgestatistics-
on-large-graphs/440578B759F245F3F417172975C3F511.

Ahlberg:2014:POB

[Ahl14] Daniel Ahlberg. Partially observed Boolean sequences and noise sensitivity. Combinatorics, Probability and Computing, 23(3): 317–330, May 2014. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Aigner-Horev:2021:THC

[AHL21] Elad Aigner-Horev and Gil Levy. Tight Hamilton cycles in cherry-quasirandom 3-uniform hypergraphs. Combinatorics, Probability and Computing, 30(3):412–443, May 2021. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/tight-
hamilton-cycles-in-cherryquasirandom-3uniform-hypergraphs/
DB476B5F977F5C6EC9F947498CD2D1FB.

Alon:2016:APT

[AHW16] Noga Alon, Rani Hod, and Amit Weinstein. On active and passive testing. Combinatorics, Probability and Computing, 25(1): 1–20, January 2016. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). Special issue on Oberwolfach Special Issue Part 2.

Aigner:2005:EP

[Aig05] Martin Aigner. The k-equal problem. Combinatorics, Probability and Computing, 14(1–2):17–24, January 2005. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic).  URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=14&issueId=01.

Aistleitner:2013:NNN

[Ais13] Christoph Aistleitner. Normal numbers and the normality measure. Combinatorics, Probability and Computing, 22(3):342–345, ???? 2013. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). Disproof by counterexample of conjecture in [AKM^{+}06].

Axenovich:2003:LAR

[AJT03] Maria Axenovich, Tao Jiang, and Zsolt Tuza. Local anti-Ramsey numbers of graphs. Combinatorics, Probability and Computing, 12(5–6):495–511, November 2003. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). Special issue on Ramsey theory.

Aigner:1994:MLG

[AK94] Martin Aigner and Regina Klimmek. Matchings in lattice graphs and Hamming graphs. Combinatorics, Probability and Computing, 3(2):157–166, June 1994. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Archibald:2014:LMV

[AK14] Margaret Archibald and Arnold Knopfmacher. The largest missing value in a sample of geometric random variables. Combinatorics, Probability and Computing, 23(5):670–685, September 2014. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). Special issue on Honouring the Memory of Philippe Flajolet - Part 1.

Alon:2000:SQB

[AKM00] Noga Alon, János Körner, and Angelo Monti. String quartets in binary. Combinatorics, Probability and Computing, 9 (5):381–390, September 2000. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http://journals.

cambridge.org/action/displayIssue?jid=CPC&volumeId=9&
issueId=05.

Alon:2006:MPF

[AKM^{+}06] N. Alon, Y. Kohayakawa, C. Mauduit, C. G. Moreira, and V. Rödl. Measures of pseudorandomness for finite sequences: Minimal values. Combinatorics, Probability and Computing, 15(1–2):1–29, January 2006. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&

volumeId=15&issueId=01. See [Ais13] for a disproof of a conjecture of this paper.

Allen:2021:FTH

[AKPP21] Peter Allen, Christoph Koch, Olaf Parczyk, and Yury Person. Finding tight Hamilton cycles in random hypergraphs faster. Combinatorics, Probability and Computing, 30(2):239–257, March 2021. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL https://www.cambridge.org/core/
journals/combinatorics-probability-and-computing/article/
finding-tight-hamilton-cycles-in-random-hypergraphs-
faster/614DD0BE59B687A582E97C88311EF1FC.

Alon:2003:TNB

[AKS03] Noga Alon, Michael Krivelevich, and Benny Sudakov. Turán numbers of bipartite graphs and related Ramsey-type questions. Combinatorics, Probability and Computing, 12(5–6):477– 494, November 2003. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). Special issue on Ramsey theory.

Alon:2005:MBF

[AKS05] Noga Alon, Michael Krivelevich, and Benny Sudakov. Max-Cut in h-free graphs. Combinatorics, Probability and Computing, 14(5–6):629–647, November 2005. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=14&issueId=05.

Amit:2006:RLG

[AL06] Alon Amit and Nathan Linial. Random lifts of graphs: Edge expansion. Combinatorics, Probability and Computing, 15(3):317– 332, May 2006. CODEN CPCOFG. ISSN 0963-5483 (print),

1469-2163 (electronic). URL http://journals.cambridge.
org/action/displayIssue?jid=CPC&volumeId=15&issueId=
03.

Albert:2009:GPS

[AL09] M. H. Albert and S. A. Linton. Growing at a perfect speed. Combinatorics, Probability and Computing, 18(3):301–308, May 2009. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Alm:2011:CIC

[AL11] Sven Erick Alm and Svante Linusson. A counter-intuitive correlation in a random tournament. Combinatorics, Probability and Computing, 20(1):1–9, January 2011. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Aldous:1992:GSB

[Ald92] David Aldous. Greedy search on the binary tree with random edge-weights. Combinatorics, Probability and Computing, 1(4): 281–293, December 1992. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Aldous:1998:CVP

[Ald98] David Aldous. On the critical value for ‘percolation’ of minimumweight trees in the mean-field distance model. Combinatorics, Probability and Computing, 7(1):1–10, March 1998. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic).  URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=7&issueId=01.

Aldous:2000:MTM

[Ald00] David J. Aldous. Mixing time for a Markov chain on cladograms. Combinatorics, Probability and Computing, 9(3):191–204, May 2000. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=9&issueId=03.

Alesker:1997:RST

[Ale97] S. Alesker. A remark on the Szarek–Talagrand Theorem. Combinatorics, Probability and Computing, 6(2):139–144, June 1997. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=6&issueId=02.

Allen:2008:CTE

[All08] Peter Allen. Covering two-edge-coloured complete graphs with two disjoint monochromatic cycles. Combinatorics, Probability and Computing, 17(4):471–486, July 2008. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Alm:1993:UBC

[Alm93] Sven Erick Alm. Upper bounds for the connective constant of self-avoiding walks. Combinatorics, Probability and Computing, 2(2):115–136, June 1993. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Alm:1998:NPW

[Alm98] Sven Erick Alm. A note on a problem by Welsh in firstpassage percolation. Combinatorics, Probability and Computing, 7(1):11–15, March 1998. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http://journals.

cambridge.org/action/displayIssue?jid=CPC&volumeId=7&
issueId=01.

Alm:2006:MAD

[Alm06] Sven Erick Alm. On measures of average degree for lattices. Combinatorics, Probability and Computing, 15(4):477–488, July 2006. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=15&issueId=04.

Alon:1992:CNG

[Alo92] Noga Alon. Choice numbers of graphs: a probabilistic approach. Combinatorics, Probability and Computing, 1(2):107–114, June 1992. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Alon:1997:EEG

[Alo97] Noga Alon. On the edge-expansion of graphs. Combinatorics, Probability and Computing, 6(2):145–152, June 1997. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=6&issueId=02.

Alon:1999:CN

[Alo99] Noga Alon. Combinatorial Nullstellensatz. Combinatorics, Probability and Computing, 8(1–2):7–29, January 1999. CODEN CP-COFG. ISSN 0963-5483 (print), 1469-2163 (electronic). Recent trends in combinatorics (Mátraháza, 1995).

Alon:2006:FSR

[Alo06a] Noga Alon. Feasible schedules for rotating transmissions. Combinatorics, Probability and Computing, 15(5):783–787, September 2006. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=15&issueId=05.

Alon:2006:SD

[Alo06b] Noga Alon. Splitting digraphs. Combinatorics, Probability and Computing, 15(6):933–937, November 2006. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=15&issueId=06.

Alon:2009:EEC

[Alo09a] Noga Alon. Economical elimination of cycles in the torus. Combinatorics, Probability and Computing, 18(5):619–627, September 2009. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Alon:2009:PIM

[Alo09b] Noga Alon. Perturbed identity matrices have high rank: Proof and applications. Combinatorics, Probability and Computing, 18 (1–2):3–15, March 2009. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Alon:2018:UDF

[Alo18] Noga Alon. Uniformly discrete forests with poor visibility. Combinatorics, Probability and Computing, 27(4):442–448, July 2018. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/uniformly-
discrete-forests-with-poor-visibility/603918F93D8502F042E3E1F9F5EAC11F.

Alon:2002:CNG

[AM02] Noga Alon and Bojan Mohar. The chromatic number of graph powers. Combinatorics, Probability and Computing, 11(1):1–10, January 2002. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL http://journals.cambridge.org/
action/displayIssue?jid=CPC&volumeId=11&issueId=01.

Arpe:2010:AGR

[AM10] Jan Arpe and Elchanan Mossel. Application of a generalization of Russo’s formula to learning from multiple random oracles. Combinatorics, Probability and Computing, 19(2):183–199, March 2010. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Atapour:2014:LDR

[AM14] Mahshid Atapour and Neal Madras. Large deviations and ratio limit theorems for pattern-avoiding permutations. Combinatorics, Probability and Computing, 23(2):161–200, March 2014. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Angel:2020:SDN

[AMP20] Omer Angel, Abbas Mehrabian, and Yuval Peres. The string of diamonds is nearly tight for rumour spreading. Combinatorics, Probability and Computing, 29(2):190–199, March 2020. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/string-
of-diamonds-is-nearly-tight-for-rumour-spreading/B7CDFEC04F314FF83405691BD73BDBAC.

Antunovic:2016:CPA

[AMR16] Tonći Antunović, Elchanan Mossel, and Miklós Z. Rácz. Coexistence in preferential attachment networks. Combinatorics, Probability and Computing, 25(6):797–822, November 2016. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/product/
070E46AE919E2E7C67BD9554F20411C9.

Andriopoulos:2023:CBT

[And23] George Andriopoulos. Convergence of blanket times for sequences of random walks on critical random graphs. Combinatorics, Probability and Computing, 32(3):478–515, May 2023. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/convergence-
of-blanket-times-for-sequences-of-random-walks-on-critical-
random-graphs/ECF4CCE3B6D32192281982F1A6E3CA30.

Annan:1994:RAA

[Ann94] J. D. Annan. A randomised approximation algorithm for counting the number of forests in dense graphs. Combinatorics, Probability and Computing, 3(3):273–283, September 1994. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:1992:AIV

[Ano92a] Anonymous. Author index to volume 1. Combinatorics, Probability and Computing, 1(4):383–384, December 1992. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:1992:CVIb

[Ano92b] Anonymous. CPC volume 1 issue 1 cover and back matter. Combinatorics, Probability and Computing, 1(1):b1–b3, March 1992. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:1992:CVIa

[Ano92c] Anonymous. CPC volume 1 issue 1 cover and front matter. Combinatorics, Probability and Computing, 1(1):f1–f2, March 1992. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:1992:CVIc

[Ano92d] Anonymous. CPC volume 1 issue 2 cover and back matter. Combinatorics, Probability and Computing, 1(2):b1–b6, June 1992. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:1992:CVId

[Ano92e] Anonymous. CPC volume 1 issue 2 cover and front matter. Combinatorics, Probability and Computing, 1(2):f1–f2, June 1992. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:1992:CVIf

[Ano92f] Anonymous. CPC volume 1 issue 3 cover and back matter. Combinatorics, Probability and Computing, 1(3):b1–b6, September 1992. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:1992:CVIe

[Ano92g] Anonymous. CPC volume 1 issue 3 cover and front matter. Combinatorics, Probability and Computing, 1(3):f1–f2, September 1992. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:1992:CVIg

[Ano92h] Anonymous. CPC volume 1 issue 4 cover and back matter. Combinatorics, Probability and Computing, 1(4):b1–b2, December 1992. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:1992:CVIh

[Ano92i] Anonymous. CPC volume 1 issue 4 cover and front matter. Combinatorics, Probability and Computing, 1(4):f1–f2, December 1992. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:1993:AIV

[Ano93a] Anonymous. Author index to volume 2. Combinatorics, Probability and Computing, 2(4):531–532, December 1993. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:1993:CVIb

[Ano93b] Anonymous. CPC volume 2 issue 1 cover and back matter. Combinatorics, Probability and Computing, 2(1):b1, March 1993. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:1993:CVIa

[Ano93c] Anonymous. CPC volume 2 issue 1 cover and front matter. Combinatorics, Probability and Computing, 2(1):f1–f2, March 1993. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:1993:CVIc

[Ano93d] Anonymous. CPC volume 2 issue 2 cover and back matter. Combinatorics, Probability and Computing, 2(2):b1–b2, June 1993. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:1993:CVId

[Ano93e] Anonymous. CPC volume 2 issue 2 cover and front matter. Combinatorics, Probability and Computing, 2(2):f1–f2, June 1993. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:1993:CVIf

[Ano93f] Anonymous. CPC volume 2 issue 3 cover and back matter. Combinatorics, Probability and Computing, 2(3):b1–b2, September 1993. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:1993:CVIe

[Ano93g] Anonymous. CPC volume 2 issue 3 cover and front matter. Combinatorics, Probability and Computing, 2(3):f1–f2, September 1993. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:1993:CVIh

[Ano93h] Anonymous. CPC volume 2 issue 4 cover and back matter. Combinatorics, Probability and Computing, 2(4):b1–b2, December 1993. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:1993:CVIg

[Ano93i] Anonymous. CPC volume 2 issue 4 cover and front matter. Combinatorics, Probability and Computing, 2(4):f1–f2, December 1993. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:1994:AIV

[Ano94a] Anonymous. Author index to volume 3. Combinatorics, Probability and Computing, 3(4):557–558, December 1994. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:1994:CVIb

[Ano94b] Anonymous. CPC volume 3 issue 1 cover and back matter. Combinatorics, Probability and Computing, 3(1):b1, March 1994. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:1994:CVIa

[Ano94c] Anonymous. CPC volume 3 issue 1 cover and front matter. Combinatorics, Probability and Computing, 3(1):f1–f2, March 1994. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:1994:CVId

[Ano94d] Anonymous. CPC volume 3 issue 2 cover and back matter. Combinatorics, Probability and Computing, 3(2):b1–b2, June 1994. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:1994:CVIc

[Ano94e] Anonymous. CPC volume 3 issue 2 cover and front matter. Combinatorics, Probability and Computing, 3(2):f1–f3, June 1994. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:1994:CVIf

[Ano94f] Anonymous. CPC volume 3 issue 3 cover and back matter. Combinatorics, Probability and Computing, 3(3):b1, September 1994. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:1994:CVIe

[Ano94g] Anonymous. CPC volume 3 issue 3 cover and front matter. Combinatorics, Probability and Computing, 3(3):f1–f2, September 1994. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:1994:CVIh

[Ano94h] Anonymous. CPC volume 3 issue 4 cover and back matter. Combinatorics, Probability and Computing, 3(4):b1, December 1994. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:1994:CVIg

[Ano94i] Anonymous. CPC volume 3 issue 4 cover and front matter. Combinatorics, Probability and Computing, 3(4):f1–f2, December 1994. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:1995:AIV

[Ano95a] Anonymous. Author index to volume 4. Combinatorics, Probability and Computing, 4(4):443–444, December 1995. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:1995:CVIb

[Ano95b] Anonymous. CPC volume 4 issue 1 cover and back matter. Combinatorics, Probability and Computing, 4(1):b1–b2, March 1995. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:1995:CVIa

[Ano95c] Anonymous. CPC volume 4 issue 1 cover and front matter. Combinatorics, Probability and Computing, 4(1):f1–f2, March 1995. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:1995:CVIc

[Ano95d] Anonymous. CPC volume 4 issue 2 cover and back matter. Combinatorics, Probability and Computing, 4(2):b1–b4, June 1995. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:1995:CVId

[Ano95e] Anonymous. CPC volume 4 issue 2 cover and front matter. Combinatorics, Probability and Computing, 4(2):f1–f2, June 1995. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:1995:CVIf

[Ano95f] Anonymous. CPC volume 4 issue 3 cover and back matter. Combinatorics, Probability and Computing, 4(3):b1–b3, September 1995. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:1995:CVIe

[Ano95g] Anonymous. CPC volume 4 issue 3 cover and front matter. Combinatorics, Probability and Computing, 4(3):f1–f2, September 1995. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:1995:CVIh

[Ano95h] Anonymous. CPC volume 4 issue 4 cover and back matter. Combinatorics, Probability and Computing, 4(4):b1–b2, December 1995. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:1995:CVIg

[Ano95i] Anonymous. CPC volume 4 issue 4 cover and front matter. Combinatorics, Probability and Computing, 4(4):f1–f2, December 1995. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:1996:AIV

[Ano96a] Anonymous. Author index to volume 5. Combinatorics, Probability and Computing, 5(4):443–444, December 1996. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:1996:CVIb

[Ano96b] Anonymous. CPC volume 5 issue 1 cover and back matter. Combinatorics, Probability and Computing, 5(1):b1–b4, March 1996. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:1996:CVIa

[Ano96c] Anonymous. CPC volume 5 issue 1 cover and front matter. Combinatorics, Probability and Computing, 5(1):f1–f2, March 1996. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:1996:CVId

[Ano96d] Anonymous. CPC volume 5 issue 2 cover and back matter. Combinatorics, Probability and Computing, 5(2):b1–b3, June 1996. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:1996:CVIc

[Ano96e] Anonymous. CPC volume 5 issue 2 cover and front matter. Combinatorics, Probability and Computing, 5(2):f1–f2, June 1996. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:1996:CVIf

[Ano96f] Anonymous. CPC volume 5 issue 3 cover and back matter. Combinatorics, Probability and Computing, 5(3):b1–b5, September 1996. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:1996:CVIe

[Ano96g] Anonymous. CPC volume 5 issue 3 cover and front matter. Combinatorics, Probability and Computing, 5(3):f1–f2, September 1996. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:1996:CVIg

[Ano96h] Anonymous. CPC volume 5 issue 4 cover and back matter. Combinatorics, Probability and Computing, 5(4):b1–b2, December 1996. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:1996:CVIh

[Ano96i] Anonymous. CPC volume 5 issue 4 cover and front matter. Combinatorics, Probability and Computing, 5(4):f1–f2, December 1996. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:1999:PCD

[Ano99] Anonymous. Problem collection of the DIMANET Mátraháza Workshop, 22–28 October 1995. Combinatorics, Probability and Computing, 8(1–2):185–192, January 1999. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). Recent trends in combinatorics (Mátraháza, 1995).

Anonymous:2003:OMC

[Ano03] Anonymous. The Oberwolfach Meeting on Combinatorics, Probability and Computing. Combinatorics, Probability and Computing, 12(3):223, May 2003. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:/

/journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=12&issueId=03.

Anonymous:2009:CVIb

[Ano09a] Anonymous. CPC volume 18 issue 1 cover and back matter. Combinatorics, Probability and Computing, 18(1–2):b1–b6, March 2009. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:2009:CVIa

[Ano09b] Anonymous. CPC volume 18 issue 1 cover and front matter. Combinatorics, Probability and Computing, 18(1–2):f1–f2, March 2009. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:2009:CVId

[Ano09c] Anonymous. CPC volume 18 issue 3 cover and back matter. Combinatorics, Probability and Computing, 18(3):b1–b4, May 2009. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:2009:CVIf

[Ano09d] Anonymous. CPC volume 18 issue 3 cover and back matter. Combinatorics, Probability and Computing, 18(3):b1–b4, May 2009. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:2009:CVIc

[Ano09e] Anonymous. CPC volume 18 issue 3 cover and front matter. Combinatorics, Probability and Computing, 18(3):f1–f2, May 2009. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:2009:CVIe

[Ano09f] Anonymous. CPC volume 18 issue 3 cover and front matter. Combinatorics, Probability and Computing, 18(3):f1–f2, May 2009. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:2009:CVIh

[Ano09g] Anonymous. CPC volume 18 issue 4 cover and back matter. Combinatorics, Probability and Computing, 18(4):b1–b5, July 2009. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:2009:CVIj

[Ano09h] Anonymous. CPC volume 18 issue 4 cover and back matter. Combinatorics, Probability and Computing, 18(4):b1–b5, July 2009. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:2009:CVIg

[Ano09i] Anonymous. CPC volume 18 issue 4 cover and front matter. Combinatorics, Probability and Computing, 18(4):f1–f2, July 2009. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:2009:CVIi

[Ano09j] Anonymous. CPC volume 18 issue 4 cover and front matter. Combinatorics, Probability and Computing, 18(4):f1–f2, July 2009. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:2009:CVIl

[Ano09k] Anonymous. CPC volume 18 issue 5 cover and back matter. Combinatorics, Probability and Computing, 18(5):b1–b4, September 2009. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:2009:CVIn

[Ano09l] Anonymous. CPC volume 18 issue 5 cover and back matter. Combinatorics, Probability and Computing, 18(5):b1–b4, September 2009. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:2009:CVIk

[Ano09m] Anonymous. CPC volume 18 issue 5 cover and front matter. Combinatorics, Probability and Computing, 18(5):f1–f3, September 2009. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:2009:CVIm

[Ano09n] Anonymous. CPC volume 18 issue 5 cover and front matter. Combinatorics, Probability and Computing, 18(5):f1–f3, September 2009. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:2009:CVIp

[Ano09o] Anonymous. CPC volume 18 issue 6 cover and back matter. Combinatorics, Probability and Computing, 18(6):b1–b5, November 2009. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:2009:CVIo

[Ano09p] Anonymous. CPC volume 18 issue 6 cover and front matter. Combinatorics, Probability and Computing, 18(6):f1–f2, November 2009. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:2009:D

[Ano09q] Anonymous. Dedication. Combinatorics, Probability and Computing, 18(1–2):2, March 2009. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:2010:CVIb

[Ano10a] Anonymous. CPC volume 19 issue 1 cover and back matter. Combinatorics, Probability and Computing, 19(1):b1–b3, January 2010. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:2010:CVIa

[Ano10b] Anonymous. CPC volume 19 issue 1 cover and front matter. Combinatorics, Probability and Computing, 19(1):f1–f2, January 2010. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:2010:CVId

[Ano10c] Anonymous. CPC volume 19 issue 2 cover and back matter. Combinatorics, Probability and Computing, 19(2):b1–b2, March 2010. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:2010:CVIc

[Ano10d] Anonymous. CPC volume 19 issue 2 cover and front matter. Combinatorics, Probability and Computing, 19(2):f1–f2, March 2010. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:2010:CVIf

[Ano10e] Anonymous. CPC volume 19 issue 3 cover and back matter. Combinatorics, Probability and Computing, 19(3):b1–b2, May 2010. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:2010:CVIe

[Ano10f] Anonymous. CPC volume 19 issue 3 cover and front matter. Combinatorics, Probability and Computing, 19(3):f1–f2, May 2010. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:2010:CVIh

[Ano10g] Anonymous. CPC volume 19 issue 4 cover and back matter. Combinatorics, Probability and Computing, 19(4):b1–b3, July 2010. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:2010:CVIg

[Ano10h] Anonymous. CPC volume 19 issue 4 cover and front matter. Combinatorics, Probability and Computing, 19(4):f1–f2, July 2010. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:2010:CVIj

[Ano10i] Anonymous. CPC volume 19 issue 5–6 cover and back matter. Combinatorics, Probability and Computing, 19(5–6):b1–b4, November 2010. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:2010:CVIi

[Ano10j] Anonymous. CPC volume 19 issue 5–6 cover and front matter. Combinatorics, Probability and Computing, 19(5–6):f1–f2, November 2010. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:2010:ECP

[Ano10k] Anonymous. Expressing combinatorial problems by systems of polynomial equations and Hilbert’s Nullstellensatz — corrigendum. Combinatorics, Probability and Computing, 19(1):159, January 2010. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:2011:CVIb

[Ano11a] Anonymous. CPC volume 20 issue 1 cover and back matter. Combinatorics, Probability and Computing, 20(1):b1–b2, January 2011. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:2011:CVIa

[Ano11b] Anonymous. CPC volume 20 issue 1 cover and front matter. Combinatorics, Probability and Computing, 20(1):f1–f2, January 2011. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:2011:CVId

[Ano11c] Anonymous. CPC volume 20 issue 2 cover and back matter. Combinatorics, Probability and Computing, 20(2):b1–b2, March 2011. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:2011:CVIc

[Ano11d] Anonymous. CPC volume 20 issue 2 cover and front matter. Combinatorics, Probability and Computing, 20(2):f1–f2, March 2011. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:2011:CVIf

[Ano11e] Anonymous. CPC volume 20 issue 3 cover and back matter. Combinatorics, Probability and Computing, 20(3):b1–b2, May 2011. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:2011:CVIe

[Ano11f] Anonymous. CPC volume 20 issue 3 cover and front matter. Combinatorics, Probability and Computing, 20(3):f1–f2, May 2011. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:2011:CVIg

[Ano11g] Anonymous. CPC volume 20 issue 4 cover and back matter. Combinatorics, Probability and Computing, 20(4):b1–b2, July 2011. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:2011:CVIh

[Ano11h] Anonymous. CPC volume 20 issue 4 cover and front matter. Combinatorics, Probability and Computing, 20(4):f1–f2, July 2011. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:2011:CVIj

[Ano11i] Anonymous. CPC volume 20 issue 5 cover and back matter. Combinatorics, Probability and Computing, 20(5):b1–b4, September 2011. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:2011:CVIi

[Ano11j] Anonymous. CPC volume 20 issue 5 cover and front matter. Combinatorics, Probability and Computing, 20(5):f1–f2, September 2011. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:2011:CVIl

[Ano11k] Anonymous. CPC volume 20 issue 6 cover and back matter. Combinatorics, Probability and Computing, 20(6):b1–b4, November 2011. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:2011:CVIk

[Ano11l] Anonymous. CPC volume 20 issue 6 cover and front matter. Combinatorics, Probability and Computing, 20(6):f1–f2, November 2011. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:2012:CVIb

[Ano12a] Anonymous. CPC volume 21 issue 1–2 cover and back matter. Combinatorics, Probability and Computing, 21(1–2):b1–b8, ???? 2012. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:2012:CVIa

[Ano12b] Anonymous. CPC volume 21 issue 1–2 cover and front matter. Combinatorics, Probability and Computing, 21(1–2):f1–f2, ???? 2012. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:2012:CVId

[Ano12c] Anonymous. CPC volume 21 issue 3 cover and back matter. Combinatorics, Probability and Computing, 21(3):b1–b2, May 2012. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:2012:CVIc

[Ano12d] Anonymous. CPC volume 21 issue 3 cover and front matter. Combinatorics, Probability and Computing, 21(3):f1–f2, May 2012. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:2012:CVIf

[Ano12e] Anonymous. CPC volume 21 issue 4 cover and back matter. Combinatorics, Probability and Computing, 21(4):b1–b3, July 2012. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:2012:CVIe

[Ano12f] Anonymous. CPC volume 21 issue 4 cover and front matter. Combinatorics, Probability and Computing, 21(4):f1–f2, July 2012. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:2012:CVIh

[Ano12g] Anonymous. CPC volume 21 issue 5 cover and back matter. Combinatorics, Probability and Computing, 21(5):b1–b3, September 2012. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:2012:CVIg

[Ano12h] Anonymous. CPC volume 21 issue 5 cover and front matter. Combinatorics, Probability and Computing, 21(5):f1–f2, September 2012. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:2012:CVIj

[Ano12i] Anonymous. CPC volume 21 issue 6 cover and back matter. Combinatorics, Probability and Computing, 21(6):b1–b2, November 2012. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:2012:CVIi

[Ano12j] Anonymous. CPC volume 21 issue 6 cover and front matter. Combinatorics, Probability and Computing, 21(6):f1–f2, November 2012. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:2013:CVIb

[Ano13a] Anonymous. CPC volume 22 issue 1 cover and back matter. Combinatorics, Probability and Computing, 22(1):b1–b4, ???? 2013. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:2013:CVIa

[Ano13b] Anonymous. CPC volume 22 issue 1 cover and front matter. Combinatorics, Probability and Computing, 22(1):f1–f2, ???? 2013. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:2013:CVId

[Ano13c] Anonymous. CPC volume 22 issue 2 cover and back matter. Combinatorics, Probability and Computing, 22(2):b1–b6, ???? 2013. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:2013:CVIc

[Ano13d] Anonymous. CPC volume 22 issue 2 cover and front matter. Combinatorics, Probability and Computing, 22(2):f1–f2, ???? 2013. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:2013:CVIf

[Ano13e] Anonymous. CPC volume 22 issue 3 cover and back matter. Combinatorics, Probability and Computing, 22(3):b1–b6, ???? 2013. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:2013:CVIe

[Ano13f] Anonymous. CPC volume 22 issue 3 cover and front matter. Combinatorics, Probability and Computing, 22(3):f1–f2, ???? 2013. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:2013:CVIh

[Ano13g] Anonymous. CPC volume 22 issue 4 cover and back matter. Combinatorics, Probability and Computing, 22(4):b1–b3, July 2013. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:2013:CVIg

[Ano13h] Anonymous. CPC volume 22 issue 4 cover and front matter. Combinatorics, Probability and Computing, 22(4):f1–f2, July 2013. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:2013:CVIj

[Ano13i] Anonymous. CPC volume 22 issue 5 cover and back matter. Combinatorics, Probability and Computing, 22(5):b1–b9, September 2013. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:2013:CVIi

[Ano13j] Anonymous. CPC volume 22 issue 5 cover and front matter. Combinatorics, Probability and Computing, 22(5):f1–f2, September 2013. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:2013:CVIl

[Ano13k] Anonymous. CPC volume 22 issue 6 cover and back matter. Combinatorics, Probability and Computing, 22(6):b1–b8, November 2013. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:2013:CVIk

[Ano13l] Anonymous. CPC volume 22 issue 6 cover and front matter. Combinatorics, Probability and Computing, 22(6):f1–f2, November 2013. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:2014:CVIb

[Ano14a] Anonymous. CPC volume 23 issue 1 cover and back matter. Combinatorics, Probability and Computing, 23(1):b1–b2, January 2014. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:2014:CVIa

[Ano14b] Anonymous. CPC volume 23 issue 1 cover and front matter. Combinatorics, Probability and Computing, 23(1):f1–f2, January 2014. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:2014:CVId

[Ano14c] Anonymous. CPC volume 23 issue 2 cover and back matter. Combinatorics, Probability and Computing, 23(2):b1–b7, March 2014. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:2014:CVIc

[Ano14d] Anonymous. CPC volume 23 issue 2 cover and front matter. Combinatorics, Probability and Computing, 23(2):f1–f2, March 2014. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:2014:CVIf

[Ano14e] Anonymous. CPC volume 23 issue 3 cover and back matter. Combinatorics, Probability and Computing, 23(3):b1–b2, May 2014. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:2014:CVIe

[Ano14f] Anonymous. CPC volume 23 issue 3 cover and front matter. Combinatorics, Probability and Computing, 23(3):f1–f2, May 2014. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:2014:CVIh

[Ano14g] Anonymous. CPC volume 23 issue 4 cover and back matter. Combinatorics, Probability and Computing, 23(4):b1–b8, July 2014. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:2014:CVIg

[Ano14h] Anonymous. CPC volume 23 issue 4 cover and front matter. Combinatorics, Probability and Computing, 23(4):f1–f2, July 2014. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:2014:CVIj

[Ano14i] Anonymous. CPC volume 23 issue 5 cover and back matter. Combinatorics, Probability and Computing, 23(5):b1–b8, September 2014. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). Special issue on Honouring the Memory of Philippe Flajolet - Part 1.

Anonymous:2014:CVIi

[Ano14j] Anonymous. CPC volume 23 issue 5 cover and front matter. Combinatorics, Probability and Computing, 23(5):f1–f2, September 2014. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). Special issue on Honouring the Memory of Philippe Flajolet - Part 1.

Anonymous:2014:CVIl

[Ano14k] Anonymous. CPC volume 23 issue 6 cover and back matter. Combinatorics, Probability and Computing, 23(6):b1–b8, November 2014. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). Special issue on Honouring the Memory of Philippe Flajolet - Part 2.

Anonymous:2014:CVIk

[Ano14l] Anonymous. CPC volume 23 issue 6 cover and front matter. Combinatorics, Probability and Computing, 23(6):f1–f2, November 2014. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). Special issue on Honouring the Memory of Philippe Flajolet - Part 2.

Anonymous:2015:WCJ

[Ano15] Anonymous. Welcome to Cambridge Journals Online. Combinatorics, Probability and Computing, 24(6), ???? 2015. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Anonymous:2016:WCJa

[Ano16] Anonymous. Welcome to Cambridge Journals Online. Combinatorics, Probability and Computing, 25(1), ???? 2016. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). Special issue on Oberwolfach Special Issue Part 2.

Alon:2022:DUS

[ANS22] Noga Alon, Asaf Nachmias, and Matan Shalev. The diameter of the uniform spanning tree of dense graphs. Combinatorics, Probability and Computing, 31(6):1010–1030, November 2022. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/diameter-
of-the-uniform-spanning-tree-of-dense-graphs/FF6B7FCAE2C832F2FCDC6953DE1DE93A.

Akkari:1993:SLE

[AO93] Safwan Akkari and James Oxley. Some local extremal connectivity results for matroids. Combinatorics, Probability and Computing, 2(4):367–384, December 1993. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Alm:2002:LUB

[AP02] Sven Erick Alm and Robert Parviainen. Lower and upper bounds for the time constant of first-passage percolation. Combinatorics, Probability and Computing, 11(5):433–445, September 2002. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=11&issueId=05.

Alappattu:2008:CLE

[AP08] Jomy Alappattu and Jim Pitman. Coloured loop-erased random walk on the complete graph. Combinatorics, Probability and Computing, 17(6):727–740, November 2008. CODEN CP-COFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Abramson:2011:CMR

[AP11a] Josh Abramson and Jim Pitman. Concave majorants of random walks and related Poisson processes. Combinatorics, Probability and Computing, 20(5):651–682, September 2011. CODEN CP-COFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Alon:2011:MOR

[AP11b] Noga Alon and Pawel Pralat. Modular orientations of random and quasi-random regular graphs. Combinatorics, Probability and Computing, 20(3):321–329, May 2011. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Aronov:2004:DDT

[APST04] Boris Aronov, János Pach, Micha Sharir, and Gábor Tardos. Distinct distances in three and higher dimensions. Combinatorics, Probability and Computing, 13(3):283–293, May 2004. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=13&issueId=03.

Adin:2001:DFR

[AR01] Ron M. Adin and Yuval Roichman. Descent functions and random Young tableaux. Combinatorics, Probability and Computing, 10(3):187–201, May 2001. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=10&issueId=03.

Adler:2003:RPE

[ARS^{+}03] Micah Adler, Harald Räcke, Naveen Sivadasan, Christian Sohler, and Berthold Vöcking. Randomized pursuit-evasion in graphs. Combinatorics, Probability and Computing, 12(3):225–244, May 2003. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=12&issueId=03. Combina-

torics, probability and computing (Oberwolfach, 2001).

Arsovski:2021:ABF

[Ars21] Bodan Arsovski. Additive bases via Fourier analysis. Combinatorics, Probability and Computing, 30(6):930–941, November 2021. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL https://www.cambridge.org/core/
journals/combinatorics-probability-and-computing/article/
additive-bases-via-fourier-analysis/DBD0AB8EE3250C164D4739F1FB19C84E.

Alon:2000:BSS

[AS00] Noga Alon and Benny Sudakov. Bipartite subgraphs and the smallest eigenvalue. Combinatorics, Probability and Computing, 9(1):1–12, January 2000. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http://journals.

cambridge.org/action/displayIssue?jid=CPC&volumeId=9&
issueId=01.

Alm:2002:EED

[AS02] Sven Erick Alm and Gregory B. Sorkin. Exact expectations and distributions for the random assignment problem. Combinatorics, Probability and Computing, 11(3):217–248, May 2002. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=11&issueId=03.

Alon:2004:NBP

[AS04] Noga Alon and Uri Stav. New bounds on parent-identifying codes: The case of multiple parents. Combinatorics, Probability and Computing, 13(6):795–807, November 2004. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic).  URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=13&issueId=06.

Alon:2006:CET

[AS06] Noga Alon and Asaf Shapira. A characterization of easily testable induced subgraphs. Combinatorics, Probability and Computing, 15(6):791–805, November 2006. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=15&issueId=06.

Apfelbaum:2011:NDS

[AS11a] Roel Apfelbaum and Micha Sharir. Non-degenerate spheres in three dimensions. Combinatorics, Probability and Computing, 20(4):503–512, July 2011. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Arlotto:2011:OSS

[AS11b] Alessandro Arlotto and J. Michael Steele. Optimal sequential selection of a unimodal subsequence of a random sequence. Combinatorics, Probability and Computing, 20(6):799–814, November 2011. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Alon:2008:ECC

[ASS08] Noga Alon, Oded Schwartz, and Asaf Shapira. An elementary construction of constant-degree expanders. Combinatorics, Probability and Computing, 17(3):319–327, May 2008. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=17&issueId=03.

Anthony:1993:UPA

[AST93] Martin Anthony and John Shawe-Taylor. Using the perceptron algorithm to find consistent hypotheses. Combinatorics, Probability and Computing, 2(4):385–387, December 1993. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Aigner:1993:RGN

[AT93] Martin Aigner and Eberhard Triesch. Reconstructing a graph from its neighborhood lists. Combinatorics, Probability and Computing, 2(2):103–113, June 1993. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Atkinson:1998:GSP

[Atk98] M. D. Atkinson. Generalized stack permutations. Combinatorics, Probability and Computing, 7(3):239–246, September 1998. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=7&issueId=03.

Austin:2015:QEC

[Aus15] Tim Austin. Quantitative equidistribution for certain quadruples in quasi-random groups. Combinatorics, Probability and Computing, 24(2):376–381, March 2015. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). See erratum [Aus16].

Austin:2016:QEC

[Aus16] Tim Austin. Quantitative equidistribution for certain quadruples in quasi-random groups: Erratum. Combinatorics, Probability and Computing, 25(3):484–485, May 2016. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). See [Aus15].

Adan:2013:CAW

[AvLR13] Ivo J. B. F. Adan, Johan S. H. van Leeuwaarden, and Kilian Raschel. The compensation approach for walks with small steps in the quarter plane. Combinatorics, Probability and Computing, 22(2):161–183, ???? 2013. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Alm:1999:IMR

[AW99] Sven Erick Alm and John C. Wierman. Inequalities for means of restricted first-passage times in percolation theory. Combinatorics, Probability and Computing, 8(4):307–315, July 1999. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=8&issueId=04. Random

graphs and combinatorial structures (Oberwolfach, 1997).

Alon:2023:IS

[AW23a] Noga Alon and Fan Wei. Irregular subgraphs. Combinatorics, Probability and Computing, 32(2):269–283, March 2023. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/irregular-
subgraphs/9133DEC8E9706E1E20485881B4D3A643.

Axenovich:2023:PRN

[AW23b] Maria Axenovich and Christian Winter. Poset Ramsey numbers: large Boolean lattice versus a fixed poset. Combinatorics, Probability and Computing, 32(4):638–653, July 2023. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/poset-
ramsey-numbers-large-boolean-lattice-versus-a-fixed-
poset/747A9639BC766BCC4DEAE190E288E185.

Alon:1993:TFF

[AY93] Noga Alon and Raphael Yuster. Threshold functions for Hfactors. Combinatorics, Probability and Computing, 2(2):137– 144, June 1993. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Azar:1994:LBI

[Aza94] Yossi Azar. Lower bounds for insertion methods for TSP. Combinatorics, Probability and Computing, 3(3):285–292, September 1994. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Bahmanian:2012:DHB

[Bah12] M. A. Bahmanian. Detachments of hypergraphs I: The bergejohnson problem. Combinatorics, Probability and Computing, 21 (4):483–495, July 2012. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Bajnok:2016:CHP

[Baj16] Béla Bajnok. Comments on Y. O. Hamidoune’s paper ‘Adding Distinct Congruence Classes’. Combinatorics, Probability and Computing, 25(5):641–644, September 2016. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/product/
77C4ABDB26D16EE905651217C0569C95. See [Ham98].

Balister:2001:AC

[Bal01a] P. N. Balister. On the Alspach Conjecture. Combinatorics, Probability and Computing, 10(2):95–125, March 2001. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic).  URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=10&issueId=02.

Balister:2001:PC

[Bal01b] Paul Balister. Packing circuits into K_{N}. Combinatorics, Probability and Computing, 10(6):463–499, November 2001. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic).  URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=10&issueId=06.

Balister:2003:PDD

[Bal03] Paul Balister. Packing digraphs with directed closed trails. Combinatorics, Probability and Computing, 12(1):1–15, January 2003. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=12&issueId=01.

Balister:2015:I

[Bal15] Paul Balister. Introduction. Combinatorics, Probability and Computing, 24(4):584, July 2015. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). Special issue on Oberwolfach Special Issue Part 1.

Barany:2004:BRU

[Bár04] Imre Bárány. Book review: “Using the Borsuk–Ulam Theorem: Lectures on Topological Methods in Combinatorics and Geometry,” by J Matoušek, Springer, 2003, 196 pp., £30.50, $49.95, EU 49.95. Combinatorics, Probability and Computing, 13(2):281–282, March 2004. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=13&issueId=02.

Barany:2006:BRG

[Bár06] Imre Bárány. Book review: “Geometric Graphs and Arrangements: Some Chapters from Combinatorial Geometry”, by Stefan Felsner, Vieweg Verlag, 2004, 170pp., 29.90 Euros. Combinatorics, Probability and Computing, 15(6):941–942, November 2006. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=15&issueId=06.

Barvinok:2008:ECT

[Bar08] Alexander Barvinok. Enumerating contingency tables via random permanents. Combinatorics, Probability and Computing, 17(1):1–19, January 2008. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:/

/journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=17&issueId=01.

Barvinok:2010:WDR

[Bar10] Alexander Barvinok. What does a random contingency table look like? Combinatorics, Probability and Computing, 19(4):517–539, July 2010. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Balister:2012:CPI

[BB12] Paul Balister and Béla Bollobás. Critical probabilities of 1independent percolation models. Combinatorics, Probability and Computing, 21(1–2):11–22, ???? 2012. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Bennett:2019:NBR

[BB19] Patrick Bennett and Tom Bohman. A natural barrier in random greedy hypergraph matching. Combinatorics, Probability and Computing, 28(6):816–825, November 2019. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL /core/journals/combinatorics-probability-
and-computing/article/natural-barrier-in-random-greedy-
hypergraph-matching/9ECB54E77217E02E4D5A1D87603F33A0.

Berger:2005:DDC

[BBC^{+}05] N. Berger, C. Borgs, J. T. Chayes, R. M. D’Souza, and R. D. Kleinberg. Degree distribution of competition-induced preferential attachment graphs. Combinatorics, Probability and Computing, 14(5–6):697–721, November 2005. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=14&issueId=05.

Bendkowski:2022:TCO

[BBD22] Maciej Bendkowski, Olivier Bodini, and Sergey Dovgal. Tuning as convex optimisation: a polynomial tuner for multiparametric combinatorial samplers. Combinatorics, Probability and Computing, 31(5):765–811, September 2022. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/tuning-
as-convex-optimisation-a-polynomial-tuner-for-multiparametric-
combinatorial-samplers/2A79BF69E8CC80011E6AADF81BFC0291.

Bishnoi:2023:SCM

[BBDM23] Anurag Bishnoi, Simona Boyadzhiyska, Shagnik Das, and Tamás Mészáros. Subspace coverings with multiplicities. Combinatorics, Probability and Computing, 32(5):782–795, September 2023. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/subspace-
coverings-with-multiplicities/35FD9151D023FE344F47C9CBFB908410.

Bal:2021:HBC

[BBDS21] Deepak Bal, Ross Berkowitz, Pat Devlin, and Mathias Schacht. Hamiltonian Berge cycles in random hypergraphs. Combinatorics, Probability and Computing, 30(2):228–238, March 2021. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/hamiltonian-
berge-cycles-in-random-hypergraphs/23FFCED26E2B4B74BBF1F3BAB41F94F3.

Barraez:2000:FGC

[BBF00] D. Barraez, S. Boucheron, and W. Fernandez de la Vega. On the fluctuations of the giant component. Combinatorics, Probability and Computing, 9(4):287–304, July 2000. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=9&issueId=04.

Balogh:2016:CTH

[BBH^{+}16] József Balogh, Jane Butterfield, Ping Hu, John Lenz, and Dhruv Mubayi. On the chromatic thresholds of hypergraphs. Combinatorics, Probability and Computing, 25(2):172–212, March 2016. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Bednarska-Bzdega:2016:MWP

[BBHKŁ16] Małgorzata Bednarska-Bzdega, Dan Hefetz, Michael Krivelevich, and Tomasz Łuczak. Manipulative waiters with probabilistic intuition. Combinatorics, Probability and Computing, 25(6): 823–849, November 2016. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL https://www.cambridge.

org/core/product/CB15B3CDBC0C064EA8C701BE78F1A98A.

Bollobas:2003:SIR

[BBK^{+}03] Béla Bollobás, Graham Brightwell, Yoshiharu Kohayakawa, Imre Leader, and Alex Scott. Special issue on Ramsey theory. Combinatorics, Probability and Computing, 12(5–6):467, November 2003. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Bal:2015:LCR

[BBKP15] Deepak Bal, Anthony Bonato, William B. Kinnersley, and Pawel Pralat. Lazy cops and robbers on hypercubes. Combinatorics, Probability and Computing, 24(6):829–837, November 2015. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Balister:2018:I

[BBL^{+}18] Paul Balister, Béla Bollobás, Imre Leader, Rob Morris, and Oliver Riordan. Introduction. Combinatorics, Probability and Computing, 27(4):441, July 2018. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL https://www.

cambridge.org/core/journals/combinatorics-probability-
and-computing/article/introduction/0B2BB359AEE24F7A6B7C8A6C0C34AB10.

Balogh:2009:EKR

[BBM09a] József Balogh, Tom Bohman, and Dhruv Mubayi. Erdős–Ko– Rado in random hypergraphs. Combinatorics, Probability and Computing, 18(5):629–646, September 2009. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Balogh:2009:MBP

[BBM09b] József Balogh, Béla Bollobás, and Robert Morris. Majority bootstrap percolation on the hypercube. Combinatorics, Probability and Computing, 18(1–2):17–51, March 2009. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Balogh:2010:BPH

[BBM10] József Balogh, Béla Bollobás, and Robert Morris. Bootstrap percolation in high dimensions. Combinatorics, Probability and Computing, 19(5–6):643–692, November 2010. CODEN CP-COFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Bonamy:2021:FCB

[BBP21] Marthe Bonamy, Nicolas Bousquet, and Guillem Perarnau. Frozen (Δ + 1)-colourings of bounded degree graphs. Combinatorics, Probability and Computing, 30(3):330–343, May 2021. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/frozen-
1colourings-of-bounded-degree-graphs/6B29091FC990E0F109BCCCE22D16C841.

Bhamidi:2014:BSR

[BBW14] Shankar Bhamidi, Amarjit Budhiraja, and Xuan Wang. Bounded-size rules: The barely subcritical regime. Combinatorics, Probability and Computing, 23(4):505–538, July 2014. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Benjamini:2008:LRP

[BBY08] Itai Benjamini, Noam Berger, and Ariel Yadin. Long-range percolation mixing time. Combinatorics, Probability and Computing, 17(4):487–494, July 2008. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Ball:2020:SPF

[BC20] Simeon Ball and Bence Csajbók. On sets of points with few odd secants. Combinatorics, Probability and Computing, 29(1):31–43, January 2020. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL https://www.cambridge.org/core/
journals/combinatorics-probability-and-computing/article/
on-sets-of-points-with-few-odd-secants/8833E150CC39E02DF802D2F5D2EE837C.

Blanca:2019:PCH

[BCG^{+}19] Antonio Blanca, Yuxuan Chen, David Galvin, Dana Randall, and Prasad Tetali. Phase coexistence for the hard-core model on Z^{2}. Combinatorics, Probability and Computing, 28(1):1–22, January 2019. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL https://www.cambridge.org/core/
journals/combinatorics-probability-and-computing/article/
phase-coexistence-for-the-hardcore-model-on-2/9B652165B36865C568285FD7A37D8B59.

Bender:1997:SDA

[BCGR97] Edward A. Bender, E. Rodney Canfield, Zhicheng Gao, and L. Bruce Ricmond. Submap density and asymmetry results for two parameter map families. Combinatorics, Probability and Computing, 6(1):17–25, March 1997. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=6&issueId=01.

Bukh:2022:CHD

[BCH22] Boris Bukh, Ting-Wei Chao, and Ron Holzman. On convex holes in d dimensional point sets. Combinatorics, Probability and Computing, 31(1):101–108, January 2022. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/on-convex-
holes-in-ddimensional-point-sets/4FB2995DE76EE1D45F468AD952E8E952.

Bhattacharya:2023:FSC

[BCJ23] Bhaswar B. Bhattacharya, Anirban Chatterjee, and Svante Janson. Fluctuations of subgraph counts in graphon based random graphs. Combinatorics, Probability and Computing, 32(3):428– 464, May 2023. CODEN CPCOFG. ISSN 0963-5483 (print),

1469-2163 (electronic). URL https://www.cambridge.org/
core/journals/combinatorics-probability-and-computing/
article/fluctuations-of-subgraph-counts-in-graphon-based-
random-graphs/0A7EAECDD0474BE1C44C67B874ACA62D.

Balogh:2021:MFG

[BCL^{+}21] József Balogh, Felix Christian Clemen, Mikhail Lavrov, Bernard Lidický, and Florian Pfender. Making K_{r+1}-free graphs r-partite. Combinatorics, Probability and Computing, 30(4):609–618, July 2021. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL https://www.cambridge.org/core/
journals/combinatorics-probability-and-computing/article/
making-kr1free-graphs-rpartite/4B3AD0723DC446B87610E28F70A376D8.

Briend:2023:ARR

[BCL23] Simon Briend, Francisco Calvillo, and Gábor Lugosi. Archaeology of random recursive dags and Cooper-Frieze random networks. Combinatorics, Probability and Computing, 32(6):859– 873, November 2023. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL https://www.cambridge.

org/core/journals/combinatorics-probability-and-computing/
article/archaeology-of-random-recursive-dags-and-cooperfrieze-
random-networks/1426DB76F9F93BEC18D3CD25565184E8.

Bapst:2017:PCS

[BCOE17] Victor Bapst, Amin Coja-Oghlan, and Charilaos Efthymiou. Planting colourings silently. Combinatorics, Probability and Computing, 26(3):338–366, May 2017. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL https://www.

cambridge.org/core/product/2BB714622DD89AC2FDB8D3479F52BC14.

Behrisch:2014:ANC

[BCOK14a] Michael Behrisch, Amin Coja-Oghlan, and Mihyun Kang. The asymptotic number of connected d-uniform hypergraphs. Combinatorics, Probability and Computing, 23(3):367–385, May 2014. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). See corrigendum [BCOK15].

Behrisch:2014:LLT

[BCOK14b] Michael Behrisch, Amin Coja-Oghlan, and Mihyun Kang. Local limit theorems for the giant component of random hypergraphs. Combinatorics, Probability and Computing, 23(3):331–366, May 2014. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Behrisch:2015:CAN

[BCOK15] Michael Behrisch, Amin Coja-Oghlan, and Mihyun Kang. Corrigendum: “The Asymptotic Number of Connected d-Uniform Hypergraphs”. Combinatorics, Probability and Computing, 24 (1):373–375, January 2015. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). See [BCOK14a].

Bender:1999:CCC

[BCOR99] E. A. Bender, P. J. Cameron, A. M. Odlyzko, and L. B. Richmond. Connectedness, classes and cycle index. Combinatorics, Probability and Computing, 8(1–2):31–43, January 1999. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). Recent trends in combinatorics (Mátraháza, 1995).

Bishnoi:2018:ZPF

[BCPS18] Anurag Bishnoi, Pete L. Clark, Aditya Potukuchi, and John R. Schmitt. On zeros of a polynomial in a finite grid. Combinatorics, Probability and Computing, 27(3):310–333, May 2018. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/on-zeros-
of-a-polynomial-in-a-finite-grid/D9CB84B64CFBAAC44F3BB2A1265B5A2D.

Balogh:2021:DVH

[BCPT21] József Balogh, Béla Csaba, András Pluhár, and Andrew Treglown. A discrepancy version of the Hajnal–Szemerédi theorem. Combinatorics, Probability and Computing, 30(3):444–459, May 2021. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL https://www.cambridge.org/core/
journals/combinatorics-probability-and-computing/article/
discrepancy-version-of-the-hajnalszemeredi-theorem/1793166873B96794297CBEAC0C754A60.

Baillon:2016:SUB

[BCV16] Jean-Bernard Baillon, Roberto Cominetti, and José Vaisman. A sharp uniform bound for the distribution of sums of Bernoulli trials. Combinatorics, Probability and Computing, 25(3):352–361, May 2016. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Bellenbaum:2002:TSP

[BD02] Patrick Bellenbaum and Reinhard Diestel. Two short proofs concerning tree-decompositions. Combinatorics, Probability and Computing, 11(6):541–547, November 2002. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=11&issueId=06.

Bruhn:2006:DIG

[BD06] Henning Bruhn and Reinhard Diestel. Duality in infinite graphs. Combinatorics, Probability and Computing, 15(1–2):75–90, January 2006. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL http://journals.cambridge.org/
action/displayIssue?jid=CPC&volumeId=15&issueId=01.

Broutin:2008:AHT

[BD08] N. Broutin and L. Devroye. An analysis of the height of tries with random weights on the edges. Combinatorics, Probability and Computing, 17(2):161–202, March 2008. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=17&issueId=02.

Banderier:2015:FAC

[BD15] Cyril Banderier and Michael Drmota. Formulae and asymptotics for coefficients of algebraic functions. Combinatorics, Probability and Computing, 24(1):1–53, January 2015. CODEN CPCOFG.

ISSN 0963-5483 (print), 1469-2163 (electronic). Special issue on Honouring the Memory of Philippe Flajolet - Part 3.

Beveridge:2012:CRG

[BDFM12] Andrew Beveridge, Andrzej Dudek, Alan Frieze, and Tobias Müller. Cops and robbers on geometric graphs. Combinatorics, Probability and Computing, 21(6):816–834, November 2012. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Ben-David:2000:NVD

[BDG00] Shai Ben-David and Leonid Gurvits. A note on VC-dimension and measure of sets of reals. Combinatorics, Probability and Computing, 9(5):391–405, September 2000. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=9&issueId=05.

Brito:2022:SGR

[BDH22] Gerandy Brito, Ioana Dumitriu, and Kameron Decker Harris. Spectral gap in random bipartite biregular graphs and applications. Combinatorics, Probability and Computing, 31(2):229–267, March 2022. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL https://www.cambridge.org/core/
journals/combinatorics-probability-and-computing/article/
spectral-gap-in-random-bipartite-biregular-graphs-and-
applications/FF0544B9780460143733B7A5317D6CC1.

Bennett:2020:MNC

[BDLP20] Patrick Bennett, Andrzej Dudek, Bernard Lidický, and Oleg Pikhurko. Minimizing the number of 5-cycles in graphs with given edge-density. Combinatorics, Probability and Computing, 29(1):44–67, January 2020. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/minimizing-
the-number-of-5cycles-in-graphs-with-given-edgedensity/
3740B1846281CAA0AA7FF6E396EBA949.

Baron:1996:PRM

[BDM96] Gerd Baron, Michael Drmota, and Ljuben Mutafchiev. Predecessors in random mappings. Combinatorics, Probability and Computing, 5(4):317–335, December 1996. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Bernardi:2014:SPP

[BDMS14] Olivier Bernardi, Rosena R. X. Du, Alejandro H. Morales, and Richard P. Stanley. Separation probabilities for products of permutations. Combinatorics, Probability and Computing, 23(2): 201–222, March 2014. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Barlow:2011:ECT

[BDNP11] Martin T. Barlow, Jian Ding, Asaf Nachmias, and Yuval Peres. The evolution of the cover time. Combinatorics, Probability and Computing, 20(3):331–345, May 2011. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Bucic:2021:UUG

[BDS21] Matija Bucić, Nemanja Draganić, and Benny Sudakov. Universal and unavoidable graphs. Combinatorics, Probability and Computing, 30(6):942–955, November 2021. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/universal-
and-unavoidable-graphs/AF73AF5A5324830C149338577258AFFC.

Bennett:2020:LTP

[BDZ20] Patrick Bennett, Andrzej Dudek, and Shira Zerbib. Large triangle packings and Tuza’s conjecture in sparse random graphs. Combinatorics, Probability and Computing, 29(5):757–779, September 2020. CODEN CPCOFG. ISSN 0963-5483 (print),

1469-2163 (electronic). URL https://www.cambridge.org/
core/journals/combinatorics-probability-and-computing/
article/large-triangle-packings-and-tuzas-conjecture-
in-sparse-random-graphs/BDB72B1BCA1D362D08E87EE25D1BCA15.

Bollobas:2015:PSS

[BE15] Béla Bollobás and Tom Eccles. Partial shadows of set systems. Combinatorics, Probability and Computing, 24(5):825–828, September 2015. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Beck:1994:DGG

[Bec94] József Beck. Deterministic graph games and a probabilistic intuition. Combinatorics, Probability and Computing, 3(1):13–26, March 1994. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Beck:2005:PG

[Bec05] József Beck. Positional games. Combinatorics, Probability and Computing, 14(5–6):649–696, November 2005. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic).  URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=14&issueId=05.

Bednarska:1998:BPG

[Bed98] Małgorzata Bednarska. On biased positional games. Combinatorics, Probability and Computing, 7(4):339–351, December 1998. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=7&issueId=04.

Bousquet:2019:EDC

[BEHD19] Nicolas Bousquet, Louis Esperet, Ararat Harutyunyan, and Rémi De Joannis De Verclos. Exact distance colouring in trees. Combinatorics, Probability and Computing, 28(2):177–186, March 2019. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/exact-
distance-colouring-in-trees/3B14559A9DA7A5ACD438C555186CD136.

Beiglbock:2006:MCS

[Bei06] M. Beiglböck. A multidimensional central sets theorem. Combinatorics, Probability and Computing, 15(6):807–814, November 2006. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=15&issueId=06.

Bencs:2018:CDT

[Ben18] Ferenc Bencs. Christoffel–Darboux type identities for the independence polynomial. Combinatorics, Probability and Computing, 27(5):716–724, September 2018. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL https://www.

cambridge.org/core/journals/combinatorics-probability-
and-computing/article/christoffeldarboux-type-identities-
for-the-independence-polynomial/0E6F178300E28D1259B6CC5DFE6CE642.

Berndt:2006:BRI

[Ber06] Bruce C. Berndt. Book review: “Integer Partitions”, by George E. Andrews and Kimmo Eriksson, Cambridge University Press 2004, 152pp., £15.99/$24.99 (paperback); £40.00/$70.00 (hardback). Combinatorics, Probability and Computing, 15 (5):789–790, September 2006. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:/

/journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=15&issueId=05.

Bertoin:2008:TPP

[Ber08] Jean Bertoin. Two-parameter Poisson–Dirichlet measures and reversible exchangeable fragmentation-coalescence processes. Combinatorics, Probability and Computing, 17(3):329–337, May 2008. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=17&issueId=03.

Beveridge:2016:HTF

[Bev16] Andrew Beveridge. A hitting time formula for the discrete Green’s function. Combinatorics, Probability and Computing, 25(3):362–379, May 2016. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Bezrukov:1994:OEB

[Bez94] Sergej L. Bezrukov. On oriented embedding of the binary tree into the hypercube. Combinatorics, Probability and Computing, 3(1):27–38, March 1994. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Boucheron:2001:INR

[BF01] S. Boucheron and W. Fernandez de la Vega. On the independence number of random interval graphs. Combinatorics, Probability and Computing, 10(5):385–396, September 2001. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=10&issueId=05.

Boucheron:2002:SPP

[BF02] S. Boucheron and W. Fernandez de la Vega. On a square packing problem. Combinatorics, Probability and Computing, 11(2):113– 127, March 2002. CODEN CPCOFG. ISSN 0963-5483 (print),

1469-2163 (electronic). URL http://journals.cambridge.
org/action/displayIssue?jid=CPC&volumeId=11&issueId=
02.

Bohman:2011:KSR

[BF11] Tom Bohman and Alan Frieze. Karp–Sipser on random graphs with a fixed degree sequence. Combinatorics, Probability and Computing, 20(5):721–741, September 2011. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Broutin:2012:LPD

[BF12] Nicolas Broutin and Omar Fawzi. Longest path distance in random circuits. Combinatorics, Probability and Computing, 21(6): 856–881, November 2012. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Brito:2018:ESI

[BFJL18] Gerandy Brito, Christopher Fowler, Matthew Junge, and Avi Levy. Ewens sampling and invariable generation. Combinatorics, Probability and Computing, 27(6):853–891, November 2018. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/ewens-
sampling-and-invariable-generation/149BD838A147A4E585BC3E0FDFE3BE9F.

Briggs:2018:PHC

[BFK^{+}18] Joseph Briggs, Alan Frieze, Michael Krivelevich, Po-Shen Loh, and Benny Sudakov. Packing Hamilton cycles online. Combinatorics, Probability and Computing, 27(4):475–495, July 2018. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/packing-
hamilton-cycles-online/94DCA7AE4B328047C54E77FB39A0B632.

Bohman:2007:FOD

[BFL^{+}07] Tom Bohman, Alan Frieze, Tomasz Luczak, Oleg Pikhurko, Clifford Smyth, Joel Spencer, and Oleg Verbitsky. Firstorder definability of trees and sparse random graphs. Combinatorics, Probability and Computing, 16(3):375–400, May 2007. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=16&issueId=03.

Bordewich:2005:ACQ

[BFLW05] M. Bordewich, M. Freedman, L. Lovász, and D. Welsh. Approximate counting and quantum computation. Combinatorics, Probability and Computing, 14(5–6):737–754, November 2005. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=14&issueId=05.

Bodini:2010:RSP

[BFP10] Olivier Bodini, Éric Fusy, and Carine Pivoteau. Random sampling of plane partitions. Combinatorics, Probability and Computing, 19(2):201–226, March 2010. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Bissacot:2011:ILL

[BFPS11] Rodrigo Bissacot, Roberto Fernández, Aldo Procacci, and Benedetto Scoppola. An improvement of the Lovász local lemma via cluster expansion. Combinatorics, Probability and Computing, 20(5):709–719, September 2011. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Bohman:2001:IF

[BFRT01] Tom Bohman, Alan Frieze, Miklós Ruszinkó, and Luboš Thoma. G-intersecting families. Combinatorics, Probability and Computing, 10(5):367–384, September 2001. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=10&issueId=05.

Bonin:2007:MPM

[BG07] Joseph E. Bonin and Omer Giménez. Multi-path matroids. Combinatorics, Probability and Computing, 16(2):193–217, March 2007. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=16&issueId=02.

Bender:2014:PSS

[BG14a] Edward A. Bender and Zhicheng Gao. Part sizes of smooth supercritical compositional structures. Combinatorics, Probability and Computing, 23(5):686–716, September 2014. CODEN CP-COFG. ISSN 0963-5483 (print), 1469-2163 (electronic). Special issue on Honouring the Memory of Philippe Flajolet - Part 1.

Bouttier:2014:IMS

[BG14b] J. Bouttier and E. Guitter. On irreducible maps and slices. Combinatorics, Probability and Computing, 23(6):914–972, November 2014. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). Special issue on Honouring the Memory of Philippe Flajolet - Part 2.

Bode:2005:SRS

[BGH05] Jens-P. Bode, Hans-Dietrich O. F. Gronau, and Heiko Harborth. Some Ramsey Schur numbers. Combinatorics, Probability and Computing, 14(1–2):25–30, January 2005. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=14&issueId=01.

Borowiecki:2003:STP

[BGHT03] Mieczysław Borowiecki, Jarosław Grytczuk, Mariusz Hałuszczak, and Zsolt Tuza. Schütte’s tournament problem and intersecting families of sets. Combinatorics, Probability and Computing, 12(4):359–364, July 2003. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=12&issueId=04.

Balister:2004:LPC

[BGLS04] Paul N. Balister, Ervin Györi, Jenö Lehel, and Richard H. Schelp. Longest paths in circular arc graphs. Combinatorics, Probability and Computing, 13(3):311–317, May 2004. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=13&issueId=03.

Barbour:1997:IRT

[BGQ97] A. D. Barbour, Anant P. Godbole, and Jinghua Qian. Imperfections in random tournaments. Combinatorics, Probability and Computing, 6(1):1–15, March 1997. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=6&issueId=01.

Bhattacharyya:2012:TOC

[BGRS12] Arnab Bhattacharyya, Elena Grigorescu, Prasad Raghavendra, and Asaf Shapira. Testing odd-cycle-freeness in Boolean functions. Combinatorics, Probability and Computing, 21(6):835–855, November 2012. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Babai:1992:DRC

[BH92] L. Babai and G. L. Hetyei. On the diameter of random Cayley graphs of the symmetric group. Combinatorics, Probability and Computing, 1(3):201–208, September 1992. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Bergelson:1993:AMR

[BH93] Vitaly Bergelson and Neil Hindman. Additive and multiplicative Ramsey theorems in N—some elementary results. Combinatorics, Probability and Computing, 2(3):221–241, September 1993. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Barat:2001:OWP

[BH01] János Barát and Péter Hajnal. Operations which preserve path-width at most two. Combinatorics, Probability and Computing, 10(4):277–291, July 2001. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=10&issueId=04.

Bukh:2012:SCN

[BH12] Boris Bukh and Alfredo Hubard. Space crossing numbers. Combinatorics, Probability and Computing, 21(3):358–373, May 2012. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Balogh:2015:MNM

[BHL^{+}15] József Balogh, Ping Hu, Bernard Lidický, Oleg Pikhurko, Balázs Udvari, and Jan Volec. Minimum number of monotone subsequences of length 4 in permutations. Combinatorics, Probability and Computing, 24(4):658–679, July 2015. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). Special issue on Oberwolfach Special Issue Part 1.

Bobecka:2013:ANT

[BHLB^{+}13] Konstancja Bobecka, Pawel Hitczenko, Fernando López-Blázquez, Grzegorz Rempala, and Jacek Wesolowski. Asymptotic normality through factorial cumulants and partition identities. Combinatorics, Probability and Computing, 22(2):213–240, ???? 2013.

CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Bucic:2020:MTR

[BHLS20] Matija Bucić, Sven Heberle, Shoham Letzter, and Benny Sudakov. Monochromatic trees in random tournaments. Combinatorics, Probability and Computing, 29(3):318–345, May 2020. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/monochromatic-
trees-in-random-tournaments/10C4DDD1B10B931397BB8D02177CAD00.

Bienvenu:2019:PAO

[Bie19] François Bienvenu. Positive association of the oriented percolation cluster in randomly oriented graphs. Combinatorics, Probability and Computing, 28(6):811–815, November 2019. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL /core/journals/combinatorics-probability-
and-computing/article/positive-association-of-the-oriented-
percolation-cluster-in-randomly-oriented-graphs/920D8AD2010B132FA32A25B1CC49A4E1.

Biggs:1993:PTD

[Big93] Norman L. Biggs. Potential theory on distance-regular graphs. Combinatorics, Probability and Computing, 2(3):243– 255, September 1993. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Biggs:2007:BRT

[Big07] Norman Biggs. Book review: “Topics in Algebraic Graph Theory, by Lowell W. Beineke and Robin J. Wilson (Academic Consultant: Peter J. Cameron), Encyclopedia of Mathematics and its Applications 102, CUP 2005, 257pp., £50.00/$95.00”. Combinatorics, Probability and Computing, 16(1):171–172, January 2007. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=16&issueId=01.

Brandt:2017:STN

[BIJ17] Axel Brandt, David Irwin, and Tao Jiang. Stability and Turán numbers of a class of hypergraphs via Lagrangians. Combinatorics, Probability and Computing, 26(3):367–405, May 2017. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/product/
D297DB8319BED747B9F41BCA1B0EA911.

Brand:1994:PCR

[BJ94] Neal Brand and Steve Jackson. Properties of classes of random graphs. Combinatorics, Probability and Computing, 3(4): 435–454, December 1994. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Boehnlein:2012:SFF

[BJ12] Edward Boehnlein and Tao Jiang. Set families with a forbidden induced subposet. Combinatorics, Probability and Computing, 21(4):496–511, July 2012. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Bukh:2017:BNE

[BJ17a] Boris Bukh and Zilin Jiang. A bound on the number of edges in graphs without an even cycle. Combinatorics, Probability and Computing, 26(1):1–15, January 2017. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/product/
02680F879444253FC3A96D5ABA2ACE7F. See erratum [BJ17b].

Bukh:2017:EBN

[BJ17b] Boris Bukh and Zilin Jiang. Erratum for ‘A Bound on the Number of Edges in Graphs Without an Even Cycle’. Combinatorics, Probability and Computing, 26(6):952–953, November 2017. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL https://www.cambridge.org/core/
product/F39FB0CF429D08CEDEE075C48B7A7865. See [BJ17a].

Bruhn:2018:SBS

[BJ18] Henning Bruhn and Felix Joos. A stronger bound for the strong chromatic index. Combinatorics, Probability and Computing, 27 (1):21–43, January 2018. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL https://www.cambridge.

org/core/product/33093E14462B979409A32806F17CDE78.

BerzunzaOjeda:2022:DPR

[BJ22] Gabriel Berzunza Ojeda and Svante Janson. The distance profile of rooted and unrooted simply generated trees. Combinatorics, Probability and Computing, 31(3):368–410, May 2022.

CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/distance-
profile-of-rooted-and-unrooted-simply-generated-trees/
B32482FCA00F9FCF73A32330A3396DF3.

Bang-Jensen:1997:HCA

[BJGY97] Jørgen Bang-Jensen, Gregory Gutin, and Anders Yeo. Hamiltonian cycles avoiding prescribed arcs in tournaments. Combinatorics, Probability and Computing, 6(3):255–261, September 1997. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=6&issueId=03.

Bang-Jensen:1998:ARA

[BJJ98] Jørgen Bang-Jensen and Tibor Jordán. Adding and reversing arcs in semicomplete digraphs. Combinatorics, Probability and Computing, 7(1):17–25, March 1998. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=7&issueId=01.

Bollobas:2009:LSP

[BJR09] Béla Bollobás, Svante Janson, and Oliver Riordan. Line-of-sight percolation. Combinatorics, Probability and Computing, 18(1– 2):83–106, March 2009. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Bodirsky:2006:GOG

[BK06a] Manuel Bodirsky and Mihyun Kang. Generating outerplanar graphs uniformly at random. Combinatorics, Probability and Computing, 15(3):333–343, May 2006. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=15&issueId=03.

Bohman:2006:CGC

[BK06b] Tom Bohman and David Kravitz. Creating a giant component. Combinatorics, Probability and Computing, 15(4):489–511, July 2006. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=15&issueId=04.

Bidkhori:2011:BPT

[BK11a] Hoda Bidkhori and Shaunak Kishore. A bijective proof of a theorem of Knuth. Combinatorics, Probability and Computing, 20(1):11–25, January 2011. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Bushaw:2011:TNM

[BK11b] Neal Bushaw and Nathan Kettle. Turán numbers of multiple paths and equibipartite forests. Combinatorics, Probability and Computing, 20(6):837–853, November 2011. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Baron:2016:TCA

[BK16] Jacob D. Baron and Jeff Kahn. Tuza’s conjecture is asymptotically tight for dense graphs. Combinatorics, Probability and Computing, 25(5):645–667, September 2016. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/product/
14C03274ADF24F51444023CD4B3892F9.

Bertram-Kretzberg:1999:SMF

[BKHL99] Claudia Bertram-Kretzberg, Thomas Hofmeister, and Hanno Lefmann. Sparse 0–1 matrices and forbidden hypergraphs. Combinatorics, Probability and Computing, 8(5):417–427, September 1999. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=8&issueId=05.

Benjamini:2006:WBF

[BKL^{+}06] Itai Benjamini, Gady Kozma, László Lovász, Dan Romik, and Gábor Tardos. Waiting for a bat to fly by (in polynomial time). Combinatorics, Probability and Computing, 15(5):673–683, September 2006. CODEN CPCOFG. ISSN 0963-5483 (print),

1469-2163 (electronic). URL http://journals.cambridge.
org/action/displayIssue?jid=CPC&volumeId=15&issueId=
05.

Balogh:2022:MPC

[BKLL22] József Balogh, Alexandr Kostochka, Mikhail Lavrov, and Xujun Liu. Monochromatic paths and cycles in 2-edge-coloured graphs with large minimum degree. Combinatorics, Probability and Computing, 31(1):109–122, January 2022. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/monochromatic-
paths-and-cycles-in-2edgecoloured-graphs-with-large-
minimum-degree/46D67972642591FEAE29E9A51B2E9149.

Balogh:2013:TWV

[BKLY13] Jozsef Balogh, Graeme Kemkes, Choongbum Lee, and Stephen J. Young. Towards a weighted version of the Hajnal–Szemerédi theorem. Combinatorics, Probability and Computing, 22(3):346– 350, ???? 2013. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Bollobas:2005:TCP

[BKN05] Béla Bollobás, Alexandr Kostochka, and Kittikorn Nakprasit. On two conjectures on packing of graphs. Combinatorics, Probability and Computing, 14(5–6):723–736, November 2005. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=14&issueId=05.

Balister:2010:LS

[BKS10] Paul Balister, Steve Kalikow, and Amites Sarkar. The Linus sequence. Combinatorics, Probability and Computing, 19(1):21– 46, January 2010. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Bohme:2011:MGH

[BKT11] Thomas Böhme, Alexandr Kostochka, and Andrew Thomason. Minors in graphs with high chromatic number. Combinatorics, Probability and Computing, 20(4):513–518, July 2011. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Bottcher:2013:ASS

[BKT13] Julia Böttcher, Yoshiharu Kohayakawa, and Anusch Taraz. Almost spanning subgraphs of random graphs after adversarial edge removal. Combinatorics, Probability and Computing, 22(5): 639–683, September 2013. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Brennan:2008:DAS

[BKW08] Charlotte Brennan, Arnold Knopfmacher, and Stephan Wagner. The distribution of ascents of size d or more in partitions of n.

Combinatorics, Probability and Computing, 17(4):495–509, July 2008. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Bilu:2004:RSR

[BL04a] Yonatan Bilu and Nathan Linial. Ramanujan signing of regular graphs. Combinatorics, Probability and Computing, 13(6):911–912, November 2004. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=13&issueId=06.

Bollobas:2004:IPS

[BL04b] Béla Bollobás and Imre Leader. Isoperimetric problems for rsets. Combinatorics, Probability and Computing, 13(2):277–279, March 2004. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL http://journals.cambridge.org/
action/displayIssue?jid=CPC&volumeId=13&issueId=02.

Beveridge:2010:EFM

[BL10] Andrew Beveridge and László Lovász. Exit frequency matrices for finite Markov chains. Combinatorics, Probability and Computing, 19(4):541–560, July 2010. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Bilu:2012:SIE

[BL12a] Yonatan Bilu and Nathan Linial. Are stable instances easy? Combinatorics, Probability and Computing, 21(5):643–660, September 2012. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Brightwell:2012:OIM

[BL12b] Graham Brightwell and Malwina Luczak. Order-invariant measures on fixed causal sets. Combinatorics, Probability and Computing, 21(3):330–357, May 2012. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Beals:2002:PRC

[BLGN^{+}02] Robert Beals, Charles R. Leedham-Green, Alice C. Niemeyer, Cheryl E. Praeger, and Ákos Seress. Permutations with restricted cycle structure and an algorithmic application. Combinatorics, Probability and Computing, 11(5):447–464, September 2002. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=11&issueId=05.

Bollobas:2011:DOT

[BLM11] Béla Bollobás, Imre Leader, and Claudia Malvenuto. Daisies and other Turán problems. Combinatorics, Probability and Computing, 20(5):743–747, September 2011. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Bloznelis:2005:ODS

[Blo05] M. Bloznelis. Orthogonal decomposition of symmetric functions defined on random permutations. Combinatorics, Probability and Computing, 14(3):249–268, May 2005. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=14&issueId=03.

Blokhuis:2006:JC

[Blo06] Aart Blokhuis. Jaeger’s Conjecture. Combinatorics, Probability and Computing, 15(4):627–629, July 2006. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=15&issueId=04.

Blumenthal:2021:SBD

[BLP^{+}21] Adam Blumenthal, Bernard Lidický, Yanitsa Pehova, Florian Pfender, Oleg Pikhurko, and Jan Volec. Sharp bounds for decomposing graphs into edges and triangles. Combinatorics, Probability and Computing, 30(2):271–287, March 2021. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/sharp-
bounds-for-decomposing-graphs-into-edges-and-triangles/
09035E84588903B1D116E515DF036FBB.

Balogh:2012:CHT

[BLS12a] József Balogh, Choongbum Lee, and Wojciech Samotij. Corrádi and Hajnal’s theorem for sparse random graphs. Combinatorics, Probability and Computing, 21(1–2):23–55, ???? 2012. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Benevides:2012:MCC

[BLS^{+}12b] F. S. Benevides, T. Luczak, A. Scott, J. Skokan, and M. White. Monochromatic cycles in 2-coloured graphs. Combinatorics, Probability and Computing, 21(1–2):57–87, ???? 2012. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Branden:2016:MEP

[BLV16] P. Brändén, M. Leander, and M. Visontai. Multivariate Eulerian polynomials and exclusion processes. Combinatorics, Probability and Computing, 25(4):486–499, July 2016. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Bakshajeva:2014:SPC

[BM14a] Tatjana Bakshajeva and Eugenijus Manstavicius. On statistics of permutations chosen from the Ewens distribution. Combinatorics, Probability and Computing, 23(6):889–913, November 2014. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). Special issue on Honouring the Memory of Philippe Flajolet - Part 2.

Bousquet-Melou:2014:ISS

[BM14b] Mireille Bousquet-Mélou. On the importance sampling of selfavoiding walks. Combinatorics, Probability and Computing, 23 (5):725–748, September 2014. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). Special issue on Honouring the Memory of Philippe Flajolet - Part 1.

Bobkov:2022:CFE

[BMM22] Sergey G. Bobkov, Arnaud Marsiglietti, and James Melbourne. Concentration functions and entropy bounds for discrete log-concave distributions. Combinatorics, Probability and Computing, 31(1):54–72, January 2022. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/concentration-
functions-and-entropy-bounds-for-discrete-logconcave-
distributions/6FE5494747FA2B309E956C93537527EE.

Balogh:2018:TTG

[BMMM18] József Balogh, Andrew McDowell, Theodore Molla, and Richard Mycroft. Triangle-tilings in graphs without large independent sets. Combinatorics, Probability and Computing, 27(4):449–474, July 2018. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL https://www.cambridge.org/core/
journals/combinatorics-probability-and-computing/article/

triangletilings-in-graphs-without-large-independent-
sets/A9BAC608FDBAE53B950C797CC7530F39.

Bonin:1999:MRN

[BMR99] Joseph E. Bonin, Jennifer McNulty, and Talmage James Reid. The matroid Ramsey number n(6, 6). Combinatorics, Probability and Computing, 8(3):229–235, May 1999. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=8&issueId=03.

Brinkmann:1995:SCG

[BMS95] Gunnar Brinkmann, Brendan D. McKay, and Carsten Saager. The smallest cubic graphs of girth nine. Combinatorics, Probability and Computing, 4(4):317–329, December 1995. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Bousquet-Melou:2014:APS

[BMW14] Mireille Bousquet-Mélou and Kerstin Weller. Asymptotic properties of some minor-closed classes of graphs. Combinatorics, Probability and Computing, 23(5):749–795, September 2014. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). Special issue on Honouring the Memory of Philippe Flajolet - Part 1.

Bollobas:2012:DPG

[BN12] Béla Bollobás and Vladimir Nikiforov. Degree powers in graphs: The Erdős–Stone theorem. Combinatorics, Probability and Computing, 21(1–2):89–105, ???? 2012. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Bar-Noy:2010:OCF

[BNCOS10] A. Bar-Noy, P. Cheilaris, S. Olonetsky, and S. Smorodinsky. Online conflict-free colouring for hypergraphs. Combinatorics, Probability and Computing, 19(4):493–516, July 2010. CODEN CP-COFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Bushaw:2015:STM

[BNMS15] Neal Bushaw, Maurício Collares Neto, Robert Morris, and Paul Smith. The sharp threshold for maximum-size sum-free subsets in even-order Abelian groups. Combinatorics, Probability and Computing, 24(4):609–640, July 2015. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). Special issue on Oberwolfach Special Issue Part 1.

Bollobas:1993:EHP

[Bol93] Béla Bollobás. Editorial: In honour of Paul Erdős. Combinatorics, Probability and Computing, 2(2):101, June 1993. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Bollobas:1994:EID

[Bol94] Béla Bollobás. Editorial: Issues dedicated to Paul Erdős. Combinatorics, Probability and Computing, 3(2):i, June 1994. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Bollobas:2004:HSC

[Bol04] Béla Bollobás. How sharp is the concentration of the chromatic number? Combinatorics, Probability and Computing, 13(1):115– 117, January 2004. CODEN CPCOFG. ISSN 0963-5483 (print),

1469-2163 (electronic). URL http://journals.cambridge.
org/action/displayIssue?jid=CPC&volumeId=13&issueId=
01.

Bonin:1995:ADL

[Bon95] Joseph E. Bonin. Automorphisms of Dowling lattices and related geometries. Combinatorics, Probability and Computing, 4(1):1– 9, March 1995. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Bonin:2000:ICB

[Bon00] Joseph E. Bonin. Involutions of connected binary matroids. Combinatorics, Probability and Computing, 9(4):305–308, July 2000. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=9&issueId=04.

Bona:2014:NUB

[Bón14] Miklós Bóna. A new upper bound for 1324-avoiding permutations. Combinatorics, Probability and Computing, 23(5):717–724, September 2014. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). Special issue on Honouring the Memory of Philippe Flajolet - Part 1.

Borkar:2002:LPS

[Bor02] V. S. Borkar. On the lock-in probability of stochastic approximation. Combinatorics, Probability and Computing, 11(1):11–20, January 2002. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL http://journals.cambridge.org/
action/displayIssue?jid=CPC&volumeId=11&issueId=01.

Bordewich:2004:ANA

[Bor04] Magnus Bordewich. Approximating the number of acyclic orientations for a class of sparse graphs. Combinatorics, Probability and Computing, 13(1):1–16, January 2004. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=13&issueId=01.

Borgs:2006:AZC

[Bor06] Christian Borgs. Absence of zeros for the chromatic polynomial on bounded degree graphs. Combinatorics, Probability and Computing, 15(1–2):63–74, January 2006. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=15&issueId=01.

Bouyrie:2018:QNS

[Bou18] Raphaël Bouyrie. On quantitative noise stability and influences for discrete and continuous models. Combinatorics, Probability and Computing, 27(3):334–357, May 2018. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/on-quantitative-
noise-stability-and-influences-for-discrete-and-continuous-
models/3F36799B3E41F18CBCE5ACACB18BF2D6.

Bovier:2004:BRS

[Bov04] Anton Bovier. Book review: “Spin Glasses: A Challenge for Mathematicians – Cavity and Mean Field Models”, by Michel Talagrand, Springer, 2003, 586 pp., EU 139.05/$159.00. Combinatorics, Probability and Computing, 13(6):913–915, November 2004. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=13&issueId=06.

Bowditch:2007:AGP

[Bow07] Brian H. Bowditch. The Angel Game in the plane. Combinatorics, Probability and Computing, 16(3):345–362, May 2007. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=16&issueId=03.

Barany:1992:NCL

[BP92] Imre Bárány and János Pach. On the number of convex lattice polygons. Combinatorics, Probability and Computing, 1(4): 295–302, December 1992. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Bennies:2001:AHB

[BP01] Jürgen Bennies and Jim Pitman. Asymptotics of the Hurwitz binomial distribution related to mixed Poisson Galton–Watson trees. Combinatorics, Probability and Computing, 10(3):203–211, May 2001. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL http://journals.cambridge.org/
action/displayIssue?jid=CPC&volumeId=10&issueId=03.

Ball:2012:AIL

[BP12] Simeon Ball and Valentina Pepe. Asymptotic improvements to the lower bound of certain bipartite Turán numbers. Combinatorics, Probability and Computing, 21(3):323–329, May 2012. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Balogh:2006:BPI

[BPP06] József Balogh, Yuval Peres, and Gábor Pete. Bootstrap percolation on infinite trees and non-amenable groups. Combinatorics, Probability and Computing, 15(5):715–730, September 2006. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=15&issueId=05.

Bernasconi:2009:DSR

[BPS09] Nicla Bernasconi, Konstantinos Panagiotou, and Angelika Steger. The degree sequence of random graphs from subcritical classes. Combinatorics, Probability and Computing, 18(5):647– 681, September 2009. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Balla:2018:RGB

[BPS18] Igor Balla, Alexey Pokrovskiy, and Benny Sudakov. Ramsey goodness of bounded degree trees. Combinatorics, Probability and Computing, 27(3):289–309, May 2018. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/ramsey-
goodness-of-bounded-degree-trees/D4361A3055FE5FBE06BC3357DF28E4A2.

Besomi:2021:ESC

[BPSS21] Guido Besomi, Matías Pavez-Signé, and Maya Stein. On the Erdős–Sós conjecture for trees with bounded degree. Combinatorics, Probability and Computing, 30(5):741–761, September 2021. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL https://www.cambridge.org/core/
journals/combinatorics-probability-and-computing/article/
on-the-erdossos-conjecture-for-trees-with-bounded-degree/
22B7ECB15D4B56C66BCA9565C4828AE9.

Bottcher:2023:TRP

[BPSS23] Julia Böttcher, Olaf Parczyk, Amedeo Sgueglia, and Jozef Skokan. Triangles in randomly perturbed graphs. Combinatorics, Probability and Computing, 32(1):91–121, January 2023. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/triangles-
in-randomly-perturbed-graphs/4B5F54F3B87F50F254C862604C26E7C6.

Bollobas:1999:TPC

[BR99] Béla Bollobás and Oliver Riordan. A Tutte polynomial for coloured graphs. Combinatorics, Probability and Computing, 8 (1–2):45–93, January 1999. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). Recent trends in combinatorics (Mátraháza, 1995).

Bonin:2000:SMB

[BR00] Joseph E. Bonin and Talmage James Reid. Simple matroids with bounded cocircuit size. Combinatorics, Probability and Computing, 9(5):407–419, September 2000. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=9&issueId=05.

Bollobas:2016:CCH

[BR16] Béla Bollobás and Oliver Riordan. Counting connected hypergraphs via the probabilistic method. Combinatorics, Probability and Computing, 25(1):21–75, January 2016. CODEN CPCOFG.

ISSN 0963-5483 (print), 1469-2163 (electronic). Special issue on Oberwolfach Special Issue Part 2.

Barvinok:2019:WCS

[BR19] Alexander Barvinok and Guus Regts. Weighted counting of solutions to sparse systems of equations. Combinatorics, Probability and Computing, 28(5):696–719, September 2019. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL /core/journals/combinatorics-probability-
and-computing/article/weighted-counting-of-solutions-
to-sparse-systems-of-equations/DED5ECF9DA97244AB61FBAAA53154310.

Brandt:1999:SDT

[Bra99] Stephan Brandt. On the structure of dense triangle-free graphs. Combinatorics, Probability and Computing, 8(3):237–245, May 1999. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=8&issueId=03.

Brinkmann:1998:ARN

[Bri98] Gunnar Brinkmann. All Ramsey numbers r(K_{3}, G) for connected graphs of order 7 and 8. Combinatorics, Probability and Computing, 7(2):129–140, June 1998. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http://journals.

cambridge.org/action/displayIssue?jid=CPC&volumeId=7&
issueId=02.

Brightwell:2009:I

[Bri09] Graham Brightwell. Introduction. Combinatorics, Probability and Computing, 18(5):617–618, September 2009. CODEN CP-COFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Brown:2003:CVT

[Bro03] Tom C. Brown. On the canonical version of a theorem in Ramsey theory. Combinatorics, Probability and Computing, 12(5–6): 513–514, November 2003. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). Special issue on Ramsey theory.

Bryant:1992:CSM

[Bry92] Victor Bryant. A characterisation of strict matching matroids. Combinatorics, Probability and Computing, 1(3):209–217, September 1992. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Bergelson:2017:TCS

[BRZK17] V. Bergelson, D. Robertson, and P. Zorin-Kranich. Triangles in Cartesian squares of quasirandom groups. Combinatorics, Probability and Computing, 26(2):161–182, March 2017. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/product/
3D14BB454708AE0CDACDCDF6B809A419.

Babai:1992:LES

[BS92] László Babai and Márió Szegedy. Local expansion of symmetrical graphs. Combinatorics, Probability and Computing, 1(1):1–11, March 1992. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Bollobas:2004:MCR

[BS04] B. Bollobás and A. D. Scott. Max cut for random graphs with a planted partition. Combinatorics, Probability and Computing, 13(4–5):451–474, July 2004. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=13&issueId=04.

Berke:2009:DRT

[BS09] R. Berke and T. Szabó. Deciding relaxed two-colourability: a hardness jump. Combinatorics, Probability and Computing, 18 (1–2):53–81, March 2009. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Bruhn:2010:DE

[BS10] Henning Bruhn and Maya Stein. Duality of ends. Combinatorics, Probability and Computing, 19(1):47–60, January 2010. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Barvinok:2011:CPF

[BS11] Alexander Barvinok and Alex Samorodnitsky. Computing the partition function for perfect matchings in a hypergraph. Combinatorics, Probability and Computing, 20(6):815–835, November 2011. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Bates:2019:UBS

[BS19] Erik Bates and Lisa Sauermann. An upper bound on the size of avoidance couplings. Combinatorics, Probability and Computing, 28(3):325–334, May 2019. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL https://www.cambridge.

org/core/journals/combinatorics-probability-and-computing/
article/an-upper-bound-on-the-size-of-avoidance-couplings/
49169F4D710A2657F4FA70E3E345AE03.

Boczkowski:2023:MHS

[BS23] Lucas Boczkowski and Igor Shinkar. On mappings on the hypercube with small average stretch. Combinatorics, Probability and Computing, 32(2):334–348, March 2023. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/on-mappings-
on-the-hypercube-with-small-average-stretch/6849E6C388735541481A91AEC1F775B9.

Ben-Shimon:2011:LRH

[BSKS11] Sonny Ben-Shimon, Michael Krivelevich, and Benny Sudakov. Local resilience and hamiltonicity maker-breaker games in random regular graphs. Combinatorics, Probability and Computing, 20(2):173–211, March 2011. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Bollobas:2015:MCA

[BSU15] Béla Bollobás, Paul Smith, and Andrew Uzzell. Monotone cellular automata in a random environment. Combinatorics, Probability and Computing, 24(4):687–722, July 2015. CODEN CP-COFG. ISSN 0963-5483 (print), 1469-2163 (electronic). Special issue on Oberwolfach Special Issue Part 1.

Blanca:2022:CMF

[BSZ22] Antonio Blanca, Alistair Sinclair, and Xusheng Zhang. The critical mean-field Chayes–Machta dynamics. Combinatorics, Probability and Computing, 31(6):924–975, November 2022. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/critical-
meanfield-chayesmachta-dynamics/DE97A2ADF20A29F4A30B7C400DD79455.

Barbour:1994:RET

[BT94] A. D. Barbour and Simon Tavaré. A rate for the Erdős–Turán law. Combinatorics, Probability and Computing, 3(2):167–176, June 1994. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Bollobas:2003:FR

[BT03] Béla Bollobás and Andrew Thomason. Frank Ramsey. Combinatorics, Probability and Computing, 12(5–6):469–475, November 2003. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). Special issue on Ramsey theory.

Baber:2011:HDJ

[BT11] Rahil Baber and John Talbot. Hypergraphs do jump. Combinatorics, Probability and Computing, 20(2):161–171, March 2011. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Bousquet:2012:SIS

[BT12] Nicolas Bousquet and Stéphan Thomassé. Scott’s induced subdivision conjecture for maximal triangle-free graphs. Combinatorics, Probability and Computing, 21(4):512–514, July 2012. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Bukh:2020:TNT

[BT20] Boris Bukh and Michael Tait. Turán numbers of theta graphs. Combinatorics, Probability and Computing, 29(4):495–507, July 2020. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL https://www.cambridge.org/core/
journals/combinatorics-probability-and-computing/article/
turan-numbers-of-theta-graphs/FE576F751EEBC3B87AB516A99B2AE0B3.

Bottcher:2012:PGF

[BTW12] Julia Böttcher, Anusch Taraz, and Andreas Würfl. Perfect graphs of fixed density: Counting and homogeneous sets. Combinatorics, Probability and Computing, 21(5):661–682, September 2012. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Balogh:2019:TRP

[BTW19] József Balogh, Andrew Treglown, and Adam Zsolt Wagner. Tilings in randomly perturbed dense graphs. Combinatorics, Probability and Computing, 28(2):159–176, March 2019. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/tilings-
in-randomly-perturbed-dense-graphs/EA3A3FF4E53B807CF4E9584DA1A3B7CE.

Bukh:2008:SD

[Buk08] Boris Bukh. Sums of dilates. Combinatorics, Probability and Computing, 17(5):627–639, September 2008. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Bukh:2016:IBF

[Buk16] Boris Bukh. An improvement of the Beck–Fiala Theorem. Combinatorics, Probability and Computing, 25(3):380–398, May 2016. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Bhamidi:2017:DSM

[BV17] Shankar Bhamidi and Remco Van Der Hofstad. Diameter of the stochastic mean-field model of distance. Combinatorics, Probability and Computing, 26(6):797–825, November 2017. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/product/
0B669E2EB71EBCCC9DE9539A1EAFD0B5.

Bhamidi:2011:FPP

[BVH11] Shankar Bhamidi, Remco Van Der Hofstad, and Gerard Hooghiemstra. First passage percolation on the Erdős–Rényi random graph. Combinatorics, Probability and Computing, 20 (5):683–707, September 2011. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Blais:2014:SSC

[BWY14] Eric Blais, Amit Weinstein, and Yuichi Yoshida. Semi-strong colouring of intersecting hypergraphs. Combinatorics, Probability and Computing, 23(1):1–7, January 2014. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Calkin:1997:DSC

[Cal97] Neil J. Calkin. Dependent sets of constant weight binary vectors. Combinatorics, Probability and Computing, 6(3): 263–271, September 1997. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http://journals.

cambridge.org/action/displayIssue?jid=CPC&volumeId=6&
issueId=03.

Candela:2009:SAA

[Can09] P. Candela. Structure along arithmetic patterns in sequences of vectors. Combinatorics, Probability and Computing, 18(6):

861–870, November 2009. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Can:2018:SEE

[Can18] Van Hao Can. Super-exponential extinction time of the contact process on random geometric graphs. Combinatorics, Probability and Computing, 27(2):162–185, March 2018. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/superexponential-
extinction-time-of-the-contact-process-on-random-geometric-
graphs/C21FF83C60F8BAA5CD0DD834E4076169.

Calkin:1998:AOR

[CC98] Neil J. Calkin and P. J. Cameron. Almost odd random sumfree sets. Combinatorics, Probability and Computing, 7(1):27–32, March 1998. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL http://journals.cambridge.org/
action/displayIssue?jid=CPC&volumeId=7&issueId=01.

Collier-Cartaino:2018:LTN

[CCGJ18] Clayton Collier-Cartaino, Nathan Graber, and Tao Jiang. Linear Turán numbers of linear cycles and cycle-complete Ramsey numbers. Combinatorics, Probability and Computing, 27(3):358–386, May 2018. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL https://www.cambridge.org/core/
journals/combinatorics-probability-and-computing/article/
linear-turan-numbers-of-linear-cycles-and-cyclecomplete-
ramsey-numbers/2BF318AC92FF263554B771C594938B36.

Cakir:1999:CEC

[CCM99] Isa Cakir, Ourania Chryssaphinou, and Marianne Månsson. On a conjecture by Eriksson concerning overlap in strings. Combinatorics, Probability and Computing, 8(5):429–440, September 1999. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=8&issueId=05.

Creignou:2003:AST

[CDD03] Nadia Creignou, Hervé Daudé, and Olivier Dubois. Approximating the satisfiability threshold for random k-XOR-formulas. Combinatorics, Probability and Computing, 12(2):113–126, March 2003. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL http://journals.cambridge.org/
action/displayIssue?jid=CPC&volumeId=12&issueId=02.

Cooper:2007:SRG

[CDG07] Colin Cooper, Martin Dyer, and Catherine Greenhill. Sampling regular graphs and a peer-to-peer network. Combinatorics, Probability and Computing, 16(4):557–593, July 2007. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=16&issueId=04.

Condon:2021:DTR

[CDG^{+}21] Padraig Condon, Alberto Espuny Díaz, António Girão, Daniela Kühn, and Deryk Osthus. Dirac’s theorem for random regular graphs. Combinatorics, Probability and Computing, 30(1):17–36, January 2021. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL https://www.cambridge.org/core/
journals/combinatorics-probability-and-computing/article/
diracs-theorem-for-random-regular-graphs/FB9A338AD26B9F0BF44F524649956F1E.

Collet:2019:LLP

[CDK19] G. Collet, M. Drmota, and L. D. Klausner. Limit laws of planar maps with prescribed vertex degrees. Combinatorics, Probability and Computing, 28(4):519–541, July 2019. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/limit-
laws-of-planar-maps-with-prescribed-vertex-degrees/64A212D62A428BC44DCCD864C85AA42F.

Chyzak:2008:DPR

[CDKK08] Frédéric Chyzak, Michael Drmota, Thomas Klausner, and Gerard Kok. The distribution of patterns in random trees. Combinatorics, Probability and Computing, 17(1):21–59, January 2008. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=17&issueId=01.

Czygrinow:2015:EHS

[CDKM15] Andrzej Czygrinow, Louis DeBiasio, H. A. Kierstead, and Theodore Molla. An extension of the Hajnal–Szemerédi theorem to directed graphs. Combinatorics, Probability and Computing, 24(5):754–773, September 2015. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Cooper:2014:CIS

[CDM14] Jeff Cooper, Kunal Dutta, and Dhruv Mubayi. Counting independent sets in hypergraphs. Combinatorics, Probability and Computing, 23(4):539–550, July 2014. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Chlebus:1996:RBH

[CDP96] Bogdan S. Chlebus, Krzysztof Diks, and Andrzej Pelc. Reliable broadcasting in hypercubes with random link and node failures. Combinatorics, Probability and Computing, 5(4):337–350, December 1996. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Comets:2014:ITU

[CDS14] Francis Comets, François Delarue, and René Schott. Information transmission under random emission constraints. Combinatorics, Probability and Computing, 23(6):973–1009, November 2014. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). Special issue on Honouring the Memory of Philippe Flajolet - Part 2.

Carstens:1999:GBD

[CDTK99] Hans-Georg Carstens, Walter A. Deuber, Wolfgang Thumser, and Elke Koppenrade. Geometrical bijections in discrete lattices. Combinatorics, Probability and Computing, 8(1–2):109–129, January 1999. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). Recent trends in combinatorics (Mátraháza, 1995).

Cameron:1999:NSF

[CE99] Peter J. Cameron and Paul Erdős. Notes on sum-free and related sets. Combinatorics, Probability and Computing, 8(1–2):95–107, January 1999. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). Recent trends in combinatorics (Mátraháza, 1995).

Calderbank:1992:IUB

[CF92] A. R. Calderbank and P. Frankl. Improved upper bounds concerning the Erdős–Korado theorem. Combinatorics, Probability and Computing, 1(2):115–122, June 1992. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Cooper:1995:CRT

[CF95] Colin Cooper and Alan Frieze. On the connectivity of random kth nearest neighbour graphs. Combinatorics, Probability and Computing, 4(4):343–362, December 1995. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Cooper:2002:MCH

[CF02] Colin Cooper and Alan Frieze. Multi-coloured Hamilton cycles in random edge-coloured graphs. Combinatorics, Probability and Computing, 11(2):129–133, March 2002. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=11&issueId=02.

Cooper:2004:SLS

[CF04] Colin Cooper and Alan Frieze. The size of the largest strongly connected component of a random digraph with a given degree sequence. Combinatorics, Probability and Computing, 13(3):319– 337, May 2004. CODEN CPCOFG. ISSN 0963-5483 (print),

1469-2163 (electronic). URL http://journals.cambridge.
org/action/displayIssue?jid=CPC&volumeId=13&issueId=
03.

Chauvin:2004:ATR

[CFGG04] B. Chauvin, P. Flajolet, D. Gardy, and B. Gittenberger. And/or trees revisited. Combinatorics, Probability and Computing, 13(4–5):475–497, July 2004. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=13&issueId=04.

Chapuy:2015:DRP

[CFGN15] Guillaume Chapuy, Éric Fusy, Omer Gim

nez, and Marc Noy. On the diameter of random planar graphs. Combinatorics, Probability and Computing, 24(1):145–178, January 2015. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). Special issue on Honouring the Memory of Philippe Flajolet - Part 3.

Cooper:2016:LRM

[CFI^{+}16] Colin Cooper, Alan Frieze, Nate Ince, Svante Janson, and Joel Spencer. On the length of a random minimum spanning tree. Combinatorics, Probability and Computing, 25(1):89–107, January 2016. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). Special issue on Oberwolfach Special Issue Part 2.

Clemens:2012:FSM

[CFKL12] Dennis Clemens, Asaf Ferber, Michael Krivelevich, and Anita Liebenau. Fast strategies in maker-breaker games played on random boards. Combinatorics, Probability and Computing, 21(6): 897–915, November 2012. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Calkin:1992:SSR

[CFM92] Neil Calkin, Alan Frieze, and Brendan D. McKay. On subgraph sizes in random graphs. Combinatorics, Probability and Computing, 1(2):123–134, June 1992. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Cooper:1994:HCR

[CFM94] Colin Cooper, Alan Frieze, and Michael Molloy. Hamilton cycles in random regular digraphs. Combinatorics, Probability and Computing, 3(1):39–49, March 1994. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Cooper:1996:PMR

[CFMR96] Colin Cooper, Alan Frieze, Michael Molloy, and Bruce Reed. Perfect matchings in random r-regular, s-uniform hypergraphs. Combinatorics, Probability and Computing, 5(1):1–14, March 1996. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Cooper:2002:RRGa

[CFR02] Colin Cooper, Alan Frieze, and Bruce Reed. Random regular graphs of non-constant degree: Connectivity and Hamiltonicity. Combinatorics, Probability and Computing, 11(3):249–261, May 2002. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=11&issueId=03.

Cooper:2002:RRGb

[CFRR02] Colin Cooper, Alan Frieze, Bruce Reed, and Oliver Riordan. Random regular graphs of non-constant degree: Independence and chromatic number. Combinatorics, Probability and Computing, 11(4):323–341, July 2002. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=11&issueId=04.

Conlon:2014:SPS

[CFS14] David Conlon, Jacob Fox, and Benny Sudakov. Short proofs of some extremal results. Combinatorics, Probability and Computing, 23(1):8–28, January 2014. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Conlon:2023:DBR

[CFW23] David Conlon, Jacob Fox, and Yuval Wigderson. Offdiagonal book Ramsey numbers. Combinatorics, Probability and Computing, 32(3):516–545, May 2023. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/offdiagonal-
book-ramsey-numbers/56B267C2FD057C2E646E9AEC5026D5FE.

Cai:2022:PMR

[CG22a] Jin-Yi Cai and Artem Govorov. Perfect matchings, rank of connection tensors and graph homomorphisms. Combinatorics, Probability and Computing, 31(2):268–303, March 2022. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/perfect-
matchings-rank-of-connection-tensors-and-graph-homomorphisms/
46898FD6A40484251A432661320116F4.

Csoka:2022:DAE

[CG22b] Endre Csóka and Lukasz Grabowski. On directed analogues of expander and hyperfinite graph sequences. Combinatorics, Probability and Computing, 31(2):184–197, March 2022. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/on-directed-
analogues-of-expander-and-hyperfinite-graph-sequences/
44BE06135C05015A20E2E9BF825CE954.

Clemens:2016:RGI

[CGL16] Dennis Clemens, Heidi Gebauer, and Anita Liebenau. The random graph intuition for the tournament game. Combinatorics, Probability and Computing, 25(1):76–88, January 2016. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). Special issue on Oberwolfach Special Issue Part 2.

Chetwynd:1996:IHU

[CH96] Amanda Chetwynd and Roland Häggkvist. An improvement of Hind’s upper bound on the total chromatic number. Combinatorics, Probability and Computing, 5(2):99–104, June 1996. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Chlamtac:2014:LIC

[CH14] Eden Chlamtác and Ishay Haviv. Linear index coding via semidefinite programming. Combinatorics, Probability and Computing, 23(2):223–247, March 2014. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Cameron:2018:CFC

[CH18] Alex Cameron and Emily Heath. A (5, 5)-colouring of K_{n} with few colours. Combinatorics, Probability and Computing, 27(6): 892–912, November 2018. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL https://www.cambridge.

org/core/journals/combinatorics-probability-and-computing/
article/55colouring-of-kn-with-few-colours/C3871E492827BDEE82E4F94C11F2B10D.

Cameron:2023:NUB

[CH23] Alex Cameron and Emily Heath. New upper bounds for the Erdős–Gyárfás problem on generalized Ramsey numbers. Combinatorics, Probability and Computing, 32(2):349–362, March 2023. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/new-
upper-bounds-for-the-erdosgyarfas-problem-on-generalized-
ramsey-numbers/1CCC980AD15C6AA943B7171023B9CB05.

Chang:2007:AMS

[Cha07] Mei-Chu Chang. Additive and multiplicative structure in matrix spaces. Combinatorics, Probability and Computing, 16(2):219– 238, March 2007. CODEN CPCOFG. ISSN 0963-5483 (print),

1469-2163 (electronic). URL http://journals.cambridge.
org/action/displayIssue?jid=CPC&volumeId=16&issueId=
02.

Chapuy:2009:AEC

[Cha09] Guillaume Chapuy. Asymptotic enumeration of constellations and related families of maps on orientable surfaces. Combinatorics, Probability and Computing, 18(4):477–516, July 2009. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Cai:2019:IST

[CHJ^{+}19] Xing Shi Cai, Cecilia Holmgren, Svante Janson, Tony Johansson, and Fiona Skerman. Inversions in split trees and conditional Galton–Watson trees. Combinatorics, Probability and Computing, 28(3):335–364, May 2019. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/inversions-
in-split-trees-and-conditional-galtonwatson-trees/49BA76E23B14724A60D6D5904D910BCA.

Cilleruelo:2009:SD

[CHS09] Javier Cilleruelo, Yahya O. Hamidoune, and Oriol Serra. On sums of dilates. Combinatorics, Probability and Computing, 18 (6):871–880, November 2009. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Chung:2005:STT

[Chu05] Fan Chung. A spectral Turán Theorem. Combinatorics, Probability and Computing, 14(5–6):755–767, November 2005. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic).  URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=14&issueId=05.

Chen:1996:CPG

[CJ96] C. C. Chen and G. P. Jin. Cycle partitions in graphs. Combinatorics, Probability and Computing, 5(2):95–97, June 1996. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Chen:1997:TFG

[CJK97] C. C. Chen, G. P. Jin, and K. M. Koh. Triangle-free graphs with large degree. Combinatorics, Probability and Computing, 6 (4):381–396, December 1997. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http://journals.

cambridge.org/action/displayIssue?jid=CPC&volumeId=6&
issueId=04.

Cameron:1993:RPS

[CK93] Peter J. Cameron and William M. Kantor. Random permutations: some group-theoretic aspects. Combinatorics, Probability and Computing, 2(3):257–262, September 1993. CODEN CP-COFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Cameron:2007:OCF

[CK07] Peter J. Cameron and K. K. Kayibi. Orbital chromatic and flow roots. Combinatorics, Probability and Computing, 16(3):401–407, May 2007. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL http://journals.cambridge.org/
action/displayIssue?jid=CPC&volumeId=16&issueId=03.

Christofides:2016:FOC

[CK16] Demetres Christofides and Daniel Král’. First-order convergence and roots. Combinatorics, Probability and Computing, 25(2): 213–221, March 2016. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Czyzowicz:2007:SBH

[CKMP07] Jurek Czyzowicz, Dariusz Kowalski, Euripides Markou, and Andrzej Pelc. Searching for a black hole in synchronous tree networks. Combinatorics, Probability and Computing, 16(4):595– 619, July 2007. CODEN CPCOFG. ISSN 0963-5483 (print),

1469-2163 (electronic). URL http://journals.cambridge.
org/action/displayIssue?jid=CPC&volumeId=16&issueId=
04.

Cooley:2018:LCR

[CKP18] Oliver Cooley, Mihyun Kang, and Yury Person. Largest components in random hypergraphs. Combinatorics, Probability and Computing, 27(5):741–762, September 2018. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/largest-
components-in-random-hypergraphs/7DD8B2411550207E3D92F8745D5AE47D.

Cioaba:2011:CCG

[CKTV11] Sebastian M. Cioaba, André Kündgen, Craig M. Timmons, and Vladislav V. Vysotsky. Covering complete r-graphs with spanning complete r-partite r-graphs. Combinatorics, Probability and Computing, 20(4):519–527, July 2011. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Chee:2015:OCE

[CKZZ15] Yeow Meng Chee, Han Mao Kiah, Hui Zhang, and Xiande Zhang. Optimal codes in the Enomoto–Katona space. Combinatorics, Probability and Computing, 24(2):382–406, March 2015. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Chan:2005:LMC

[CL05] Onn Chan and T. K. Lam. Lifting Markov chains to random walks on groups. Combinatorics, Probability and Computing, 14(3):269–273, May 2005. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=14&issueId=03.

Clemens:2020:MRG

[CLR20] Dennis Clemens, Anita Liebenau, and Damian Reding. On minimal Ramsey graphs and Ramsey equivalence in multiple colours. Combinatorics, Probability and Computing, 29(4):537–554, July 2020. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL https://www.cambridge.org/core/
journals/combinatorics-probability-and-computing/article/
on-minimal-ramsey-graphs-and-ramsey-equivalence-in-multiple-
colours/2F34FFC7F40E288D808EF2C4BA59D79F.

Croot:2013:APS

[CLS13] Ernie Croot, Izabella Laba, and Olof Sisask. Arithmetic progressions in sumsets and L^{p}-almost-periodicity. Combinatorics, Probability and Computing, 22(3):351–365, ???? 2013. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Cannon:2019:PME

[CLS19] S. Cannon, D. A. Levin, and A. Stauffer. Polynomial mixing of the edge-flip Markov chain for unbiased dyadic tilings. Combinatorics, Probability and Computing, 28(3):365–387, May 2019.

CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/polynomial-
mixing-of-the-edgeflip-markov-chain-for-unbiased-dyadic-
tilings/8972300E3CCFAFFDC28D15CE208357D8.

Cameron:1993:TRR

[CM93] Peter J. Cameron and Cleide Martins. A theorem on reconstruction of random graphs. Combinatorics, Probability and Computing, 2(1):1–9, March 1993. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Cockayne:1997:IMD

[CM97a] E. J. Cockayne and C. M. Mynhardt. Irredundance and maximum degree in graphs. Combinatorics, Probability and Computing, 6(2):153–157, June 1997. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http://journals.

cambridge.org/action/displayIssue?jid=CPC&volumeId=6&
issueId=02.

Cowen:1997:OP

[CM97b] Lenore Cowen and Rudolph Mathar. The offset problem. Combinatorics, Probability and Computing, 6(2):159–164, June 1997. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=6&issueId=02.

Clemens:2021:SRN

[CMR^{+}21] Dennis Clemens, Meysam Miralaei, Damian Reding, Mathias Schacht, and Anusch Taraz. On the size-Ramsey number of grid graphs. Combinatorics, Probability and Computing, 30(5): 670–685, September 2021. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL https://www.cambridge.

org/core/journals/combinatorics-probability-and-computing/
article/on-the-sizeramsey-number-of-grid-graphs/BA31C4B7C5407A492438C6DED05F381A.

Cameron:2006:HHR

[CN06] Peter J. Cameron and Jaroslav Nešetril. Homomorphismhomogeneous relational structures. Combinatorics, Probability and Computing, 15(1–2):91–103, January 2006. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic).  URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=15&issueId=01.

Csikvari:2012:DTP

[CN12] Péter Csikvári and Zoltán Lóránt Nagy. The density Turán problem. Combinatorics, Probability and Computing, 21(4):531–553, July 2012. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Conlon:2023:SRN

[CNT23] David Conlon, Rajko Nenadov, and Milos Trujić. On the size-Ramsey number of grids. Combinatorics, Probability and Computing, 32(6):874–880, November 2023. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/on-the-
sizeramsey-number-of-grids/2C4A8CAD6D4A1343AC542D8795875984.

Coja-Oghlan:2005:LNR

[CO05] Amin Coja-Oghlan. The Lovász number of random graphs. Combinatorics, Probability and Computing, 14(4):439–465, July 2005. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=14&issueId=04.

Coja-Oghlan:2006:FLI

[CO06] Amin Coja-Oghlan. Finding large independent sets in polynomial expected time. Combinatorics, Probability and Computing, 15(5):731–751, September 2006. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=15&issueId=05.

Coja-Oghlan:2007:CSG

[CO07a] Amin Coja-Oghlan. Colouring semirandom graphs. Combinatorics, Probability and Computing, 16(4):515–552, July 2007. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=16&issueId=04.

Coja-Oghlan:2007:LEG

[CO07b] Amin Coja-Oghlan. On the Laplacian eigenvalues of G_{n,p}. Combinatorics, Probability and Computing, 16(6):923–946, November 2007. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=16&issueId=06.

Coja-Oghlan:2010:GPA

[CO10] Amin Coja-Oghlan. Graph partitioning via adaptive spectral techniques. Combinatorics, Probability and Computing, 19(2): 227–284, March 2010. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Cocks:2008:CM

[Coc08] Clifford C. Cocks. Correlated matroids. Combinatorics, Probability and Computing, 17(4):511–518, July 2008. CODEN CP-COFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Coja-Oghlan:2021:OGT

[COGHKL21] Amin Coja-Oghlan, Oliver Gebhard, Max Hahn-Klimroth, and Philipp Loick. Optimal group testing. Combinatorics, Probability and Computing, 30(6):811–848, November 2021. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/optimal-
group-testing/6AACA31D8DB0A45F594471E0B731C7DB.

Coja-Oghlan:2007:SRH

[COGL07] Amin Coja-Oghlan, Andreas Goerdt, and André Lanka. Strong refutation heuristics for random k-SAT. Combinatorics, Probability and Computing, 16(1):5–28, January 2007. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic).  URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=16&issueId=01.

Coja-Oghlan:2020:RSP

[COKM20] Amin Coja-Oghlan, Tobias Kapetanopoulos, and Noela Müller. The replica symmetric phase of random constraint satisfaction problems. Combinatorics, Probability and Computing, 29(3):346– 422, May 2020. CODEN CPCOFG. ISSN 0963-5483 (print),

1469-2163 (electronic). URL https://www.cambridge.org/
core/journals/combinatorics-probability-and-computing/
article/replica-symmetric-phase-of-random-constraint-
satisfaction-problems/5815BB497BE83C3370F796CF8CD9C9FB.

Coja-Oghlan:2009:SAA

[COMV09] Amin Coja-Oghlan, Elchanan Mossel, and Dan Vilenchik. A spectral approach to analysing belief propagation for 3colouring. Combinatorics, Probability and Computing, 18(6):

881–912, November 2009. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Conlon:2009:HPS

[Con09] David Conlon. Hypergraph packing and sparse bipartite Ramsey numbers. Combinatorics, Probability and Computing, 18(6): 913–923, November 2009. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Conlon:2017:STF

[Con17] David Conlon. A sequence of triangle-free pseudorandom graphs. Combinatorics, Probability and Computing, 26(2):195– 200, March 2017. CODEN CPCOFG. ISSN 0963-5483 (print),

1469-2163 (electronic). URL https://www.cambridge.org/
core/product/78853F75AE8CD1BDF8B6B86A2E091A3B.

Conant:2020:FSS

[Con20] Gabriel Conant. On finite sets of small tripling or small alternation in arbitrary groups. Combinatorics, Probability and Computing, 29(6):807–829, November 2020. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/on-finite-
sets-of-small-tripling-or-small-alternation-in-arbitrary-
groups/BC08B81A497706752074E5EE497AF233.

Cooper:1992:TSR

[Coo92] Colin Cooper. On the thickness of sparse random graphs. Combinatorics, Probability and Computing, 1(4):303–309, December 1992. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Cooper:1996:AEP

[Coo96] C. Cooper. Asymptotic enumeration of predicate-junction flowgraphs. Combinatorics, Probability and Computing, 5(3):215– 226, September 1996. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Cooper:2006:DVD

[Coo06] Colin Cooper. Distribution of vertex degree in Web-graphs. Combinatorics, Probability and Computing, 15(5):637–661, September 2006. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=15&issueId=05.

Coulson:2022:CWR

[Cou22] Matthew Coulson. The critical window in random digraphs. Combinatorics, Probability and Computing, 31(3):411–429, May 2022. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL https://www.cambridge.org/core/
journals/combinatorics-probability-and-computing/article/
critical-window-in-random-digraphs/E112B68627A8DDDC1D1ED28FE3D044E8.

Coja-Oghlan:2018:NSA

[COW18] Amin Coja-Oghlan and Nick Wormald. The number of satisfying assignments of random regular k-SAT formulas. Combinatorics, Probability and Computing, 27(4):496–530, July 2018. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/number-
of-satisfying-assignments-of-random-regular-ksat-formulas/
4495BF86B69F1C02F7F9A435B9DFE3CE.

Clemens:2016:MDC

[CP16] Dennis Clemens and Yury Person. Minimum degrees and codegrees of Ramsey-minimal 3-uniform hypergraphs. Combinatorics, Probability and Computing, 25(6):850–869, November 2016. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL https://www.cambridge.org/core/
product/83583E57BB628995F6C29CB359FDE3BC.

Choi:2020:ANR

[CP20] Michael C. H. Choi and Pierre Patie. Analysis of nonreversible Markov chains via similarity orbits. Combinatorics, Probability and Computing, 29(4):508–536, July 2020. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/analysis-
of-nonreversible-markov-chains-via-similarity-orbits/
EC0DCE3956D8A46E61FDBFB5A17DDDC2.

Cohen:2017:WRO

[CPT17] Emma Cohen, Will Perkins, and Prasad Tetali. On the Widom– Rowlinson occupancy fraction in regular graphs. Combinatorics, Probability and Computing, 26(2):183–194, March 2017.

CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/product/
3C2AF2F1637B9337F54A2DEA32D3EFEA.

Chryssaphinou:2001:PAN

[CPV01] Ourania Chryssaphinou, Stavros Papastavridis, and Eutichia Vaggelatou. Poisson approximation for the non-overlapping appearances of several words in Markov chains. Combinatorics, Probability and Computing, 10(4):293–308, July 2001. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=10&issueId=04.

Cutler:2013:HIS

[CR13] Jonathan Cutler and A. J. Radcliffe. Hypergraph independent sets. Combinatorics, Probability and Computing, 22(1):9–20, ???? 2013. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Croot:2012:SPL

[Cro12] Ernie Croot. Some properties of lower level-sets of convolutions. Combinatorics, Probability and Computing, 21(4):515–530, July 2012. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Chen:2012:HIS

[CRS12] Guantao Chen, Han Ren, and Songling Shan. Homeomorphically irreducible spanning trees in locally connected graphs. Combinatorics, Probability and Computing, 21(1–2):107–111, ???? 2012. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Chen:1993:RPB

[CS93] G. Chen and R. H. Schelp. Ramsey problems with bounded degree spread. Combinatorics, Probability and Computing, 2(3): 263–269, September 1993. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Chen:1995:MIP

[CS95] Tuhao Chen and E. Seneta. Multivariate identities, permutation and Bonferroni upper bounds. Combinatorics, Probability and Computing, 4(4):331–342, December 1995. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Chen:1998:LBP

[CS98] Tuhao Chen and E. Seneta. Lower bounds for the probability of intersection of several unions of events. Combinatorics, Probability and Computing, 7(4):353–364, December 1998. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=7&issueId=04.

Cooper:2006:SRW

[CS06] Joshua N. Cooper and Joel Spencer. Simulating a random walk with constant error. Combinatorics, Probability and Computing, 15(6):815–822, November 2006. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=15&issueId=06.

Czumaj:2010:TEB

[CS10] Artur Czumaj and Christian Sohler. Testing expansion in bounded-degree graphs. Combinatorics, Probability and Computing, 19(5–6):693–709, November 2010. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Csaba:2007:BEC

[Csa07] B. Csaba. On the Bollobás–Eldridge Conjecture for bipartite graphs. Combinatorics, Probability and Computing, 16(5):661–691, September 2007. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=16&issueId=05.

Csirik:1992:OSF

[Csi92] János A. Csirik. Optimal strategy for the first player in the Penney Ante game. Combinatorics, Probability and Computing, 1(4):311–321, December 1992. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Csikvari:2013:NSR

[Csi13] Péter Csikvári. Note on the smallest root of the independence polynomial. Combinatorics, Probability and Computing, 22(1): 1–8, ???? 2013. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Crippa:1997:MPI

[CST97] D. Crippa, K. Simon, and P. Trunz. Markov processes involving q-Stirling numbers. Combinatorics, Probability and Computing, 6(2):165–178, June 1997. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http://journals.

cambridge.org/action/displayIssue?jid=CPC&volumeId=6&
issueId=02.

Chung:1998:IIC

[CT98] F. R. K. Chung and Prasad Tetali. Isoperimetric inequalities for Cartesian products of graphs. Combinatorics, Probability and Computing, 7(2):141–148, June 1998. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=7&issueId=02.

Charbit:2007:GLG

[CT07] Pierre Charbit and Stéphan Thomassé. Graphs with large girth not embeddable in the sphere. Combinatorics, Probability and Computing, 16(6):829–832, November 2007. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=16&issueId=06.

Clement:2015:TRA

[CTV15] J. Clément, T. H. Nguyen Thi, and B. Vallée. Towards a realistic analysis of some popular sorting algorithms. Combinatorics, Probability and Computing, 24(1):104–144, January 2015. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). Special issue on Honouring the Memory of Philippe Flajolet - Part 3.

Charbit:2007:MFA

[CTY07] Pierre Charbit, Stéphan Thomassé, and Anders Yeo. The minimum feedback arc set problem is NP-hard for tournaments. Combinatorics, Probability and Computing, 16(1):1–4, January 2007. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=16&issueId=01.

Cuckler:2007:HCR

[Cuc07] Bill Cuckler. Hamiltonian cycles in regular tournaments. Combinatorics, Probability and Computing, 16(2):239–249, March 2007. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=16&issueId=02.

Curien:2012:SCP

[Cur12] Nicolas Curien. Strong convergence of partial match queries in random quadtrees. Combinatorics, Probability and Computing, 21(5):683–694, September 2012. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Cilleruelo:2008:SCS

[CV08] Javier Cilleruelo and Carlos Vinuesa. B_{2}[g] sets and a conjecture of Schinzel and Schmidt. Combinatorics, Probability and Computing, 17(6):741–747, November 2008. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Costello:2010:RRS

[CV10] Kevin P. Costello and Van Vu. On the rank of random sparse matrices. Combinatorics, Probability and Computing, 19(3):321– 342, May 2010. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Cesaratto:2015:GDT

[CV15] Eda Cesaratto and Brigitte Vallée. Gaussian distribution of trie depth for strongly tame sources. Combinatorics, Probability and Computing, 24(1):54–103, January 2015. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). Special issue on Honouring the Memory of Philippe Flajolet - Part 3.

Csorgo:2000:RGS

[CW00] Sándor Csörgo and Wei Biao Wu. Random graphs and the strong convergence of bootstrap means. Combinatorics, Probability and Computing, 9(4):315–347, July 2000. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=9&issueId=04.

Choe:2006:RM

[CW06] Youngbin Choe and David G. Wagner. Rayleigh matroids. Combinatorics, Probability and Computing, 15(5):765–781, September 2006. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=15&issueId=05.

Chung:1995:EGS

[CY95] F. R. K. Chung and S.-T. Yau. Eigenvalues of graphs and Sobolev inequalities. Combinatorics, Probability and Computing, 4(1): 11–25, March 1995. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Caro:2000:DFG

[CY00] Yair Caro and Raphael Yuster. Dominating a family of graphs with small connected subgraphs. Combinatorics, Probability and Computing, 9(4):309–313, July 2000. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=9&issueId=04.

Czabarka:1999:ICF

[Cza99] Éva Czabarka. Intersecting chains in finite vector spaces. Combinatorics, Probability and Computing, 8(6):509–528, November 1999. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=8&issueId=06.

Dancik:1998:CSS

[Dan98] Vlado Dancík. Common subsequences and supersequences and their expected length. Combinatorics, Probability and Computing, 7(4):365–373, December 1998. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=7&issueId=04.

Day:2017:SGP

[Day17] A. Nicholas Day. Saturated graphs of prescribed minimum degree. Combinatorics, Probability and Computing, 26(2):201–207, March 2017. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL https://www.cambridge.org/core/
product/A1FC8D7941DF0B67A3BA4DF9F90DE6B4.

deBoer:2023:UGM

[dBBR23] David de Boer, Pjotr Buys, and Guus Regts. Uniqueness of the Gibbs measure for the 4-state anti-ferromagnetic Potts model on the regular tree. Combinatorics, Probability and Computing, 32(1):158–182, January 2023. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/uniqueness-
of-the-gibbs-measure-for-the-4state-antiferromagnetic-
potts-model-on-the-regular-tree/0EE86171BCA21ABE510C8CE98C733C41.

Deza:2005:ZSS

[DD05] Michel Deza and Mathieu Dutour. Zigzag structures of simple two-faced polyhedra. Combinatorics, Probability and Computing, 14(1–2):31–57, January 2005. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=14&issueId=01.

Ding:2010:VBM

[DD10] Guoli Ding and Stan Dziobiak. Vertex-bipartition method for colouring minor-closed classes of graphs. Combinatorics, Probability and Computing, 19(4):579–591, July 2010. CODEN CP-COFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Draganic:2021:PPT

[DDF^{+}21] Nemanja Draganić, François Dross, Jacob Fox, António Girão, Frédéric Havet, Dániel Korándi, William Lochet, David Munhá Correia, Alex Scott, and Benny Sudakov. Powers of paths in tournaments. Combinatorics, Probability and Computing, 30(6):894–898, November 2021. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/powers-
of-paths-in-tournaments/3CA48CC1DB5AA4C7A2BB4F9F056A6933.

Dujmovic:2022:CCG

[DEM^{+}22] Vida Dujmović, Louis Esperet, Pat Morin, Bartosz Walczak, and David R. Wood. Clustered 3-colouring graphs of bounded degree. Combinatorics, Probability and Computing, 31(1):123–135, January 2022. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL https://www.cambridge.org/core/
journals/combinatorics-probability-and-computing/article/
clustered-3colouring-graphs-of-bounded-degree/378EA59B7067BA10EBA1D2EEF5BE835E.

Denley:1997:RLK

[Den97] Tristan Denley. On a result of Lemke and Kleitman. Combinatorics, Probability and Computing, 6(1):39–43, March 1997. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=6&issueId=01.

Dobrow:1996:MTM

[DF96] Robert P. Dobrow and James Allen Fill. Multiway trees of maximum and minimum probability under the random permutation model. Combinatorics, Probability and Computing, 5(4):351–371, December 1996. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Dobrow:1999:TPL

[DF99] Robert P. Dobrow and James Allen Fill. Total path length for random recursive trees. Combinatorics, Probability and Computing, 8(4):317–333, July 1999. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http://journals.

cambridge.org/action/displayIssue?jid=CPC&volumeId=8&

issueId=04. Random graphs and combinatorial structures (Oberwolfach, 1997).

Dinur:2009:IFE

[DF09a] Irit Dinur and Ehud Friedgut. Intersecting families are essentially contained in juntas. Combinatorics, Probability and Computing, 18(1–2):107–122, March 2009. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Doerr:2009:DRW

[DF09b] Benjamin Doerr and Tobias Friedrich. Deterministic random walks on the two-dimensional grid. Combinatorics, Probability and Computing, 18(1–2):123–144, March 2009. CODEN CP-COFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Devroye:2014:RCM

[DF14] Luc Devroye and Nicolas Fraiman. The random connection model on the torus. Combinatorics, Probability and Computing, 23(5):796–804, September 2014. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). Special issue on Honouring the Memory of Philippe Flajolet - Part 1.

deFraysseix:1994:TCG

[dFdMR94] Hubert de Fraysseix, Patrice Ossona de Mendez, and Pierre Rosenstiehl. On triangle contact graphs. Combinatorics, Probability and Computing, 3(2):233–246, June 1994. CODEN CP-COFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Dyer:1993:MET

[DFK^{+}93] Martin Dyer, Alan Frieze, Ravi Kannan, Ajai Kapoor, Ljubomir Perkovic, and Umesh Vazirani. A mildly exponential time algorithm for approximating the number of solutions to a multidimensional knapsack problem. Combinatorics, Probability and Computing, 2(3):271–284, September 1993. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Duchon:2004:BSR

[DFLS04] Philippe Duchon, Philippe Flajolet, Guy Louchard, and Gilles Schaeffer. Boltzmann samplers for the random generation of combinatorial structures. Combinatorics, Probability and Computing, 13(4–5):577–625, July 2004. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=13&issueId=04.

Diaconis:1992:ATR

[DFP92] Persi Diaconis, James Allen Fill, and Jim Pitman. Analysis of top to random shuffles. Combinatorics, Probability and Computing, 1(2):135–155, June 1992. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Dudek:2009:NUC

[DFR09] Andrzej Dudek, Peter Frankl, and Vojtech Rödl. A note on universal and canonically coloured sequences. Combinatorics, Probability and Computing, 18(5):683–689, September 2009. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Day:2021:MBP

[DFR21] A. Nicholas Day and Victor Falgas-Ravry. Maker–Breaker percolation games I: crossing grids. Combinatorics, Probability and Computing, 30(2):200–227, March 2021. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/makerbreaker-
percolation-games-i-crossing-grids/291ED756915B0FF38068DFB3C22A87EF.

Dudek:2015:LHC

[DFRS15] Andrzej Dudek, Alan Frieze, Andrzej Ruciński, and Matas Sileikis. Loose Hamilton cycles in regular hypergraphs. Combinatorics, Probability and Computing, 24(1):179–194, January 2015. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). Special issue on Honouring the Memory of Philippe Flajolet - Part 3.

Daude:1997:ACA

[DFV97] Hervé Daudé, Philippe Flajolet, and Brigitte Vallée. An averagecase analysis of the Gaussian algorithm for lattice reduction. Combinatorics, Probability and Computing, 6(4):397–433, December 1997. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL http://journals.cambridge.org/
action/displayIssue?jid=CPC&volumeId=6&issueId=04.

Deza:1994:LPC

[DG94] Michel Deza and Viatcheslav Grishukhin. Lattice points of cut cones. Combinatorics, Probability and Computing, 3(2):191–214, June 1994. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Dennert:2010:SSP

[DG10] Florian Dennert and Rudolf Grübel. On the subtree size profile of binary search trees. Combinatorics, Probability and Computing, 19(4):561–578, July 2010. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Davidson:2017:MST

[DG17] A. Davidson and A. Ganesh. Maximal Steiner trees in the stochastic mean-field model of distance. Combinatorics, Probability and Computing, 26(6):826–838, November 2017. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/product/
CB14B2899251F310E8BF1EF558CA7D6F.

Dyer:2002:CIP

[DGG^{+}02] Martin Dyer, Leslie Ann Goldberg, Catherine Greenhill, Gabriel Istrate, and Mark Jerrum. Convergence of the Iterated Prisoner’s Dilemma game. Combinatorics, Probability and Computing, 11(2):135–147, March 2002. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=11&issueId=02.

Dekel:2014:FHC

[DGGP14] Yael Dekel, Ori Gurel-Gurevich, and Yuval Peres. Finding hidden cliques in linear time with high probability. Combinatorics, Probability and Computing, 23(1):29–49, January 2014. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Diaconis:2022:CGP

[DGHS22] Persi Diaconis, Ron Graham, Xiaoyu He, and Sam Spiro. Card guessing with partial feedback. Combinatorics, Probability and Computing, 31(1):1–20, January 2022. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/card-
guessing-with-partial-feedback/E89E90B78CFCE8E223E721297EA78FF6.

Dyer:2008:DCS

[DGJ08] Martin Dyer, Leslie Ann Goldberg, and Mark Jerrum. Dobrushin conditions and systematic scan. Combinatorics, Probability and Computing, 17(6):761–779, November 2008. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Draganic:2022:SPL

[DGK22] Nemanja Draganić, Stefan Glock, and Michael Krivelevich. Short proofs for long induced paths. Combinatorics, Probability and Computing, 31(5):870–878, September 2022. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/short-
proofs-for-long-induced-paths/2067DBBDA97D8270DAD793853A2B955C.

Drmota:2011:MDS

[DGN11] Michael Drmota, Omer Giménez, and Marc Noy. The maximum degree of series-parallel graphs. Combinatorics, Probability and Computing, 20(4):529–570, July 2011. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Das:2015:STP

[DGS15] Shagnik Das, Wenying Gan, and Benny Sudakov. Sperner’s theorem and a problem of Erdős, Katona and Kleitman. Combinatorics, Probability and Computing, 24(4):585–608, July 2015. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). Special issue on Oberwolfach Special Issue Part 1.

Devroye:2023:PNL

[DGZ23] Luc Devroye, Marcel K. Goh, and Rosie Y. Zhao. On the peel number and the leaf-height of Galton–Watson trees. Combinatorics, Probability and Computing, 32(1):68–90, January 2023. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/on-the-
peel-number-and-the-leafheight-of-galtonwatson-trees/
631190C0A22C46DD848D98D26A5AC5CD.

Dugdale:1994:AFC

[DH94] J. K. Dugdale and A. J. W. Hilton. Amalgamated factorizations of complete graphs. Combinatorics, Probability and Computing, 3(2):215–232, June 1994. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Dugdale:2000:SCG

[DH00] J. K. Dugdale and A. J. W. Hilton. A sufficient condition for a graph to be the core of a class 2 graph. Combinatorics, Probability and Computing, 9(2):97–104, March 2000. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=9&issueId=02.

Deijfen:2006:ICI

[DH06] Maria Deijfen and Olle Häggström. The initial configuration is irrelevant for the possibility of mutual unbounded growth in the two-type Richardson model. Combinatorics, Probability and Computing, 15(3):345–353, May 2006. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=15&issueId=03.

Diao:2010:RTP

[DH10] Y. Diao and G. Hetyei. Relative Tutte polynomials for coloured graphs and virtual knot theory. Combinatorics, Probability and Computing, 19(3):343–369, May 2010. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Diao:2013:RTP

[DH13] Y. Diao and G. Hetyei. Relative Tutte polynomials of tensor products of coloured graphs. Combinatorics, Probability and Computing, 22(6):801–828, November 2013. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Dyer:2021:PTA

[DHJM21] Martin Dyer, Marc Heinrich, Mark Jerrum, and Haiko Müller. Polynomial-time approximation algorithms for the antiferromagnetic Ising model on line graphs. Combinatorics, Probability and Computing, 30(6):905–921, November 2021. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/polynomialtime-
approximation-algorithms-for-the-antiferromagnetic-ising-
model-on-line-graphs/0245A08F316EEE9C02F5588657215437.

Dellamonica:2012:EDS

[DHL^{+}12] D. Dellamonica, P. Haxell, T. Luczak, D. Mubayi, B. Nagle, Y. Person, V. Rödl, M. Schacht, and J. Verstraëte. On evendegree subgraphs of linear hypergraphs. Combinatorics, Probability and Computing, 21(1–2):113–127, ???? 2012. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Darrasse:2016:SMR

[DHS16] Alexis Darrasse, Hsien-Kuei Hwang, and Michèle Soria. Shape measures of random increasing k-trees. Combinatorics, Probability and Computing, 25(5):668–699, September 2016. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/product/
30BBE834DA01F94EB58ADF53F5F22367.

Diestel:1995:GMS

[Die95] Reinhard Diestel. Graph minors. I. A short proof of the pathwidth theorem. comment on: “Graph minors. I. Excluding a forest” [J. Combin. Theory Ser. B 35 (1983), no. 1, 39–61; MR0723569 (85d:05148)] by N. Robertson and P. D. Seymour. Combinatorics, Probability and Computing, 4(1):27– 30, March 1995. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Diestel:2005:CSI

[Die05] Reinhard Diestel. The cycle space of an infinite graph. Combinatorics, Probability and Computing, 14(1–2):59–79, January 2005. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=14&issueId=01.

Dumitrescu:2013:MEB

[DJ13] Adrian Dumitrescu and Minghui Jiang. Maximal empty boxes amidst random points. Combinatorics, Probability and Computing, 22(4):477–498, July 2013. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Dubhashi:2007:PIN

[DJR07] Devdatt Dubhashi, Johan Jonasson, and Desh Ranjan. Positive influence and negative dependence. Combinatorics, Probability and Computing, 16(1):29–41, January 2007. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=16&issueId=01.

Dong:2004:TRR

[DK04] F. M. Dong and K. M. Koh. Two results on real zeros of chromatic polynomials. Combinatorics, Probability and Computing, 13(6):809–813, November 2004. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=13&issueId=06.

Ding:2006:STC

[DK06] Guoli Ding and Jinko Kanno. Splitter theorems for cubic graphs. Combinatorics, Probability and Computing, 15(3):355–375, May 2006. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=15&issueId=03.

Dong:2008:BRZ

[DK08] F. M. Dong and K. M. Koh. Bounds for the real zeros of chromatic polynomials. Combinatorics, Probability and Computing, 17(6):749–759, November 2008. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Diskin:2023:SSP

[DK23] Sahar Diskin and Michael Krivelevich. Supercritical site percolation on the hypercube: small components are small. Combinatorics, Probability and Computing, 32(3):422–427, May 2023. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/supercritical-

site-percolation-on-the-hypercube-small-components-are-
small/28269664D9C55BFE2DED35BA90E471DD.

Dellamonica:2016:NS

[DKL^{+}16] Domingos Dellamonica, Jr., Yoshiharu Kohayakawa, Sang June Lee, Vojtech Rödl, and Wojciech Samotij. On the number of B_{h}- sets. Combinatorics, Probability and Computing, 25(1):108–129, January 2016. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). Special issue on Oberwolfach Special Issue Part 2.

Ding:2010:DSR

[DKLP10] Jian Ding, Jeong Han Kim, Eyal Lubetzky, and Yuval Peres. Diameters in supercritical random graphs via first passage percolation. Combinatorics, Probability and Computing, 19(5–6): 729–751, November 2010. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Dellamonica:2015:OAF

[DKM^{+}15] Domingos Dellamonica, Jr., Subrahmanyam Kalyanasundaram, Daniel M. Martin, Vojtech Rödl, and Asaf Shapira. An optimal algorithm for finding Frieze–Kannan regular partitions. Combinatorics, Probability and Computing, 24(2):407–437, March 2015. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Duffus:2018:MNE

[DKR18] Dwight Duffus, Bill Kay, and Vojtech Rödl. The minimum number of edges in uniform hypergraphs with Property O. Combinatorics, Probability and Computing, 27(4):531–538, July 2018. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/minimum-
number-of-edges-in-uniform-hypergraphs-with-property-
o/87F2ABCAD0DA50CEB5092D62D2856B59.

Dahmer:2014:TEB

[DKW14] Iulia Dahmer, Götz Kersting, and Anton Wakolbinger. The total external branch length of beta-coalescents. Combinatorics, Probability and Computing, 23(6):1010–1027, November 2014. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). Special issue on Honouring the Memory of Philippe Flajolet - Part 2.

Diestel:1994:GIG

[DL94] Reinhard Diestel and Imre Leader. The growth of infinite graphs: boundedness and finite spreading. Combinatorics, Probability and Computing, 3(1):51–55, March 1994. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Drmota:2014:AET

[DL14] M. Drmota and A. Lladó. Almost every tree with m edges decomposes K_{2m,2m}. Combinatorics, Probability and Computing, 23(1):50–65, January 2014. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Dubois:2004:ADR

[DLM04] Olivier Dubois, Guy Louchard, and Jacques Mandler. Additive decompositions, random allocations, and threshold phenomena. Combinatorics, Probability and Computing, 13(4–5):537–575, July 2004. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL http://journals.cambridge.org/
action/displayIssue?jid=CPC&volumeId=13&issueId=04.

Duffus:1998:EPO

[DŁRR98] Dwight Duffus, Tomasz Łuczak, Vojtěch Rödl, and Andrzej Ruciński. Endomorphisms of partially ordered sets. Combinatorics, Probability and Computing, 7(1):33–46, March 1998. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=7&issueId=01.

Diaconis:2011:BS

[DM11] Persi Diaconis and Laurent Miclo. On barycentric subdivision. Combinatorics, Probability and Computing, 20(2):213–237, March 2011. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Danielsson:2023:PRS

[DM23] Joel Larsson Danielsson and Klas Markström. Polarised random k-SAT. Combinatorics, Probability and Computing, 32(6): 885–899, November 2023. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL https://www.cambridge.

org/core/journals/combinatorics-probability-and-computing/
article/polarised-random-ksat/0A68B966E7BD0401414FD390E2A59B80.

deMendez:2005:HD

[dMR05] Patrice Ossona de Mendez and Pierre Rosenstiehl. Homomorphism and dimension. Combinatorics, Probability and Computing, 14(5–6):861–872, November 2005. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=14&issueId=05.

Drmota:2019:ARP

[DMS19] M. Drmota, A. Magner, and W. Szpankowski. Asymmetric Rényi problem. Combinatorics, Probability and Computing, 28(4):542– 573, July 2019. CODEN CPCOFG. ISSN 0963-5483 (print),

1469-2163 (electronic). URL https://www.cambridge.org/
core/journals/combinatorics-probability-and-computing/
article/asymmetric-renyi-problem/E62EA1F7567705E822A4E226AA1557DC.

Dong:2023:NEC

[DMZ23] Dingding Dong, Nitya Mani, and Yufei Zhao. On the number of error correcting codes. Combinatorics, Probability and Computing, 32(5):819–832, September 2023. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/on-the-
number-of-error-correcting-codes/7BDE7E79815F0145DC142CC108756FBF.

Dizona:2013:CPL

[DN13] Jill Dizona and Brendan Nagle. Constructive packings by linear hypergraphs. Combinatorics, Probability and Computing, 22(6): 829–858, November 2013. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Dereich:2014:RAP

[DO14] Steffen Dereich and Marcel Ortgiese. Robust analysis of preferential attachment models with fitness. Combinatorics, Probability and Computing, 23(3):386–411, May 2014. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Dobson:2002:CTG

[Dob02a] Edward Dobson. Constructing trees in graphs whose complement has no K_{2,s}. Combinatorics, Probability and Computing, 11(4): 343–347, July 2002. CODEN CPCOFG. ISSN 0963-5483 (print),

1469-2163 (electronic). URL http://journals.cambridge.
org/action/displayIssue?jid=CPC&volumeId=11&issueId=
04.

Dobson:2002:PTC

[Dob02b] Edward Dobson. Packing trees into the complete graph. Combinatorics, Probability and Computing, 11(3):263–272, May 2002. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=11&issueId=03.

Dobson:2006:AGM

[Dob06] Edward Dobson. Automorphism groups of metacirculant graphs of order a product of two distinct primes. Combinatorics, Probability and Computing, 15(1–2):105–130, January 2006. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=15&issueId=01.

Doerr:2000:LHD

[Doe00] Benjamin Doerr. Linear and hereditary discrepancy. Combinatorics, Probability and Computing, 9(4):349–354, July 2000. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=9&issueId=04.

Dohmen:2002:BTI

[Doh02] Klaus Dohmen. Bonferroni-type inequalities via chordal graphs. Combinatorics, Probability and Computing, 11(4):349–351, July 2002. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=11&issueId=04.

Dessmark:2004:DCC

[DP04] Anders Dessmark and Andrzej Pelc. Distributed colouring and communication in rings with local knowledge. Combinatorics, Probability and Computing, 13(2):123–136, March 2004. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=13&issueId=02.

DeMarco:2006:RAD

[DP06] Gianluca De Marco and Andrzej Pelc. Randomized algorithms for determining the majority on graphs. Combinatorics, Probability and Computing, 15(6):823–834, November 2006. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=15&issueId=06.

Dudek:2015:APL

[DP15] Andrzej Dudek and Pawel Pralat. An alternative proof of the linearity of the size-Ramsey number of paths. Combinatorics, Probability and Computing, 24(3):551–555, May 2015. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

DeSalvo:2019:LSB

[DP19] Stephen DeSalvo and Igor Pak. Limit shapes via bijections. Combinatorics, Probability and Computing, 28(2):187–240, March 2019. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL https://www.cambridge.org/core/
journals/combinatorics-probability-and-computing/article/
limit-shapes-via-bijections/AE20CFC9B4E4ED0E2832A29BE9F29CFA.

Diaz:2023:SCR

[DP23] Alberto Espuny Díaz and Yury Person. Spanning F -cycles in random graphs. Combinatorics, Probability and Computing, 32(5): 833–850, September 2023. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL https://www.cambridge.

org/core/journals/combinatorics-probability-and-computing/
article/spanning-fcycles-in-random-graphs/19EAA37AA3CE031CC1A927724A1C7931.

Diaz:2000:CTS

[DPPS00] Josep Díaz, Mathew D. Penrose, Jordi Petit, and María Serna. Convergence theorems for some layout measures on random lattice and random geometric graphs. Combinatorics, Probability and Computing, 9(6):489–511, November 2000. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=9&issueId=06.

Daknama:2021:RRR

[DPR21] Rami Daknama, Konstantinos Panagiotou, and Simon Reisser. Robustness of randomized rumour spreading. Combinatorics, Probability and Computing, 30(1):37–78, January 2021. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/robustness-
of-randomized-rumour-spreading/385AB29A7B1A2E44CC16EDE50718844F.

Denley:1999:NEM

[DR99] Tristan Denley and Talmage James Reid. On the number of elements in matroids with small circuits or cocircuits. Combinatorics, Probability and Computing, 8(6):529–537, November 1999. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=8&issueId=06.

Dellamonica:2012:DPR

[DR12] Domingos Dellamonica and Vojtech Rödl. Distance preserving Ramsey graphs. Combinatorics, Probability and Computing, 21 (4):554–581, July 2012. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

DeAmbroggio:2022:ULC

[DR22] Umberto De Ambroggio and Matthew I. Roberts. Unusually large components in near-critical Erdős–Rényi graphs via ballot theorems. Combinatorics, Probability and Computing, 31(5): 840–869, September 2022. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL https://www.cambridge.

org/core/journals/combinatorics-probability-and-computing/
article/unusually-large-components-in-nearcritical-erdosrenyi-
graphs-via-ballot-theorems/CC5D53F6C6C281DDB013B9C03EFBF88F.

Dietzfelbinger:2010:TBB

[DRWW10] Martin Dietzfelbinger, Jonathan E. Rowe, Ingo Wegener, and Philipp Woelfel. Tight bounds for blind search on the integers and the reals. Combinatorics, Probability and Computing, 19(5– 6):711–728, November 2010. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Doerr:2003:MD

[DS03] Benjamin Doerr and Anand Srivastav. Multicolour discrepancies. Combinatorics, Probability and Computing, 12(4):365–399, July 2003. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=12&issueId=04.

Drmota:2004:SIA

[DS04] Michael Drmota and Wojciech Szpankowski. Special issue on analysis of algorithms. Combinatorics, Probability and Computing, 13(4–5):415–417, July 2004. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=13&issueId=04.

Deuber:1994:CAC

[DT94] Walter A. Deuber and Wolfgang Thumser. A combinatorial approach to complexity theory via ordinal hierarchies. Combinatorics, Probability and Computing, 3(2):177–190, June 1994. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Dumitrescu:2008:NTM

[DT08] Adrian Dumitrescu and Csaba D. Tóth. On the number of tetrahedra with minimum, unit, and distinct volumes in threespace. Combinatorics, Probability and Computing, 17(2):203– 224, March 2008. CODEN CPCOFG. ISSN 0963-5483 (print),

1469-2163 (electronic). URL http://journals.cambridge.
org/action/displayIssue?jid=CPC&volumeId=17&issueId=
02.

Dodos:2012:MEI

[DT12] Vassilis Kanellopoulos Pandelis Dodos and Konstantinos Tyros. Measurable events indexed by trees. Combinatorics, Probability and Computing, 21(3):374–411, May 2012. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Dumitrescu:2017:CPG

[DT17] Adrian Dumitrescu and Csaba D. Tóth. Convex polygons in geometric triangulations. Combinatorics, Probability and Computing, 26(5):641–659, September 2017. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/product/
8615247CB0E1D7DDFAF61649F342DE71.

Das:2020:RPR

[DT20] Shagnik Das and Andrew Treglown. Ramsey properties of randomly perturbed graphs: cliques and cycles. Combinatorics, Probability and Computing, 29(6):830–867, November 2020. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/

combinatorics-probability-and-computing/article/ramsey-
properties-of-randomly-perturbed-graphs-cliques-and-
cycles/C368E54DCC99B5C0651FE8403BB05FEF.

Dubickas:2008:ALS

[Dub08] Arturas Dubickas. An approximation by lacunary sequence of vectors. Combinatorics, Probability and Computing, 17(3):339– 345, May 2008. CODEN CPCOFG. ISSN 0963-5483 (print),

1469-2163 (electronic). URL http://journals.cambridge.
org/action/displayIssue?jid=CPC&volumeId=17&issueId=
03.

Dudley:2007:BRG

[Dud07] Richard Dudley. Book review: “The Generic Chaining: Upper and Lower Bounds of Stochastic Processes”, by Michel Talagrand, Springer-Verlag, 2005, 222 pp., £61.50/$99.00/79.95 Euros. Combinatorics, Probability and Computing, 16(4):659–660, July 2007. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL http://journals.cambridge.org/
action/displayIssue?jid=CPC&volumeId=16&issueId=04.

Daireaux:2004:DAP

[DV04] Benoît Daireaux and Brigitte Vallée. Dynamical analysis of the parametrized Lehmer–Euclid algorithm. Combinatorics, Probability and Computing, 13(4–5):499–536, July 2004. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=13&issueId=04.

Duckworth:2006:IDN

[DW06] W. Duckworth and N. C. Wormald. On the independent domination number of random regular graphs. Combinatorics, Probability and Computing, 15(4):513–522, July 2006. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=15&issueId=04.

Dubiner:1994:HDR

[DZ94] Moshe Dubiner and Uri Zwick. How do read-once formulae shrink? Combinatorics, Probability and Computing, 3(4):455– 469, December 1994. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Decreusefond:1997:ECC

[DZ97] Laurent Decreusefond and Gilles Zémor. On the error-correcting capabilities of cycle codes of graphs. Combinatorics, Probability and Computing, 6(1):27–38, March 1997. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=6&issueId=01.

Deuschel:1999:ISD

[DZ99] Jean-Dominique Deuschel and Ofer Zeitouni. On increasing subsequences of I.I.D. samples. Combinatorics, Probability and Computing, 8(3):247–263, May 1999. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http://journals.

cambridge.org/action/displayIssue?jid=CPC&volumeId=8&
issueId=03.

Eccles:2016:SRU

[Ecc16] Tom Eccles. A stability result for the union-closed size problem. Combinatorics, Probability and Computing, 25(3):399–418, May 2016. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Edwards:1995:HCN

[Edw95] Keith Edwards. The harmonious chromatic number of almost all trees. Combinatorics, Probability and Computing, 4(1):31– 46, March 1995. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Edwards:1996:HCN

[Edw96] Keith Edwards. The harmonious chromatic number of bounded degree trees. Combinatorics, Probability and Computing, 5(1): 15–28, March 1996. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Edwards:1998:SSV

[Edw98] A. W. F. Edwards. Seven-set Venn diagrams with rotational and polar symmetry. Combinatorics, Probability and Computing, 7(2):149–152, June 1998. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http://journals.

cambridge.org/action/displayIssue?jid=CPC&volumeId=7&
issueId=02.

Edwards:2005:DCG

[Edw05] Keith Edwards. Detachments of complete graphs. Combinatorics, Probability and Computing, 14(3):275–310, May 2005. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=14&issueId=03.

Eriksson:2003:EER

[EES03] Henrik Eriksson, Kimmo Eriksson, and Jonas Sjöstrand. Exact expectations for random graphs and assignments. Combinatorics, Probability and Computing, 12(4):401–412, July 2003. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=12&issueId=04.

ElMaftouhi:1995:RS

[EF95] A. El Maftouhi and W. Fernandez de la Vega. On random 3sat. Combinatorics, Probability and Computing, 4(3):189–195, September 1995. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Erdos:1992:GTS

[EFK92] Péter L. Erdős, Ulrich Faigle, and Walter Kern. A grouptheoretic setting for some intersecting Sperner families. Combinatorics, Probability and Computing, 1(4):323–334, December 1992. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Erdos:1993:RSL

[EFRS93] Paul Erdős, R. J. Faudree, C. C. Rousseau, and R. H. Schelp. Ramsey size linear graphs. Combinatorics, Probability and Computing, 2(4):389–399, December 1993. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Efthymiou:2020:DCG

[Eft20] Charilaos Efthymiou. Deterministic counting of graph colourings using sequences of subgraphs. Combinatorics, Probability and Computing, 29(4):555–586, July 2020. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/deterministic-
counting-of-graph-colourings-using-sequences-of-subgraphs/
F60A60F621EDA6286FEB0B640EFBDB97.

Ehard:2020:PHM

[EGJ20] Stefan Ehard, Stefan Glock, and Felix Joos. Pseudorandom hypergraph matchings. Combinatorics, Probability and Computing, 29(6):868–885, November 2020. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/pseudorandom-
hypergraph-matchings/AE0DE17C95900E7DD3F221E2C17FE203.

Erdos:1994:ITS

[EGŁ94] Paul Erdős, András Gyárfás, and Tomasz Łuczak. Independent transversals in sparse partite hypergraphs. Combinatorics, Probability and Computing, 3(3):293–296, September 1994. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Ergemlidze:2019:TNI

[EGM19] Beka Ergemlidze, Ervin Győri, and Abhishek Methuku. Turán number of an induced complete bipartite graph plus an odd cycle. Combinatorics, Probability and Computing, 28(2):241–252, March 2019. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL https://www.cambridge.org/core/
journals/combinatorics-probability-and-computing/article/
turan-number-of-an-induced-complete-bipartite-graph-
plus-an-odd-cycle/83B2DBA6BD60C6778D3B01BB44E42E95.

Ergemlidze:2020:ALB

[EGM^{+}20] Beka Ergemlidze, Ervin Győri, Abhishek Methuku, Nika Salia, Casey Tompkins, and Oscar Zamora. Avoiding long Berge cycles: the missing cases k = r + 1 and k = r + 2. Combinatorics, Probability and Computing, 29(3):423–435, May 2020. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/avoiding-
long-berge-cycles-the-missing-cases-k-r-1-and-k-r-2/
2746031EF10465255DDC3A1E36CCFD76.

Espinasse:2016:CAM

[EGPN16] T. Espinasse, N. Guillotin-Plantard, and P. Nadeau. A combinatorial approach to a model of constrained random walkers. Combinatorics, Probability and Computing, 25(2):222–235, March 2016. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Edwards:1998:ZFI

[EHJ98] Hugh Edwards, Robert Hierons, and Bill Jackson. The zero-free intervals for characteristic polynomials of matroids. Combinatorics, Probability and Computing, 7(2):153–165, June 1998. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=7&issueId=02.

Erdos:1994:TRT

[EHS^{+}94] P. Erdős, A. Hajnal, M. Simonovits, V. T. Sós, and E. Szemerédi. Turán–Ramsey theorems and K_{p}-independence numbers. Combinatorics, Probability and Computing, 3(3):297–325, September 1994. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Esperet:2014:CPG

[EJ14] Louis Esperet and Gwenaël Joret. Colouring planar graphs with three colours and no large monochromatic components. Combinatorics, Probability and Computing, 23(4):551–570, July 2014. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Erde:2021:LCM

[EKK21] Joshua Erde, Mihyun Kang, and Michael Krivelevich. Large complete minors in random subgraphs. Combinatorics, Probability and Computing, 30(4):619–630, July 2021. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/large-
complete-minors-in-random-subgraphs/20B3183EE801111DB0723C0DC024E10E.

Erdos:2013:SDD

[EKM13] Péter L. Erdős, Zoltán Király, and István Miklós. On the swapdistances of different realizations of a graphical degree sequence. Combinatorics, Probability and Computing, 22(3):366–383, ???? 2013. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Eberhard:2021:SIF

[EKNS21] Sean Eberhard, Jeff Kahn, Bhargav Narayanan, and Sophie Spirkl. On symmetric intersecting families of vectors. Combinatorics, Probability and Computing, 30(6):899–904, November 2021. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL https://www.cambridge.org/core/
journals/combinatorics-probability-and-computing/article/
on-symmetric-intersecting-families-of-vectors/BD56D62DBC2186D38F97F72A3C6D26D7.

Esperet:2019:SCD

[EKT19] Louis Esperet, Ross J. Kang, and Stéphan Thomassé. Separation choosability and dense bipartite induced subgraphs. Combinatorics, Probability and Computing, 28(5):720–732, September 2019. CODEN CPCOFG. ISSN 0963-5483 (print),

1469-2163 (electronic). URL /core/journals/combinatorics-
probability-and-computing/article/separation-choosability-
and-dense-bipartite-induced-subgraphs/265127482211C42CA56294C8468D0FEF.

Ellis:2019:UIF

[EL19] David Ellis and Noam Lifshitz. On the union of intersecting families. Combinatorics, Probability and Computing, 28(6):826–839, November 2019. CODEN CPCOFG. ISSN 0963-5483 (print),

1469-2163 (electronic). URL /core/journals/combinatorics-
probability-and-computing/article/on-the-union-of-intersecting-
families/1F6C8FDA3DF75E447412E89D4687B3A2.

Ellis:2011:AIS

[Ell11] David Ellis. Almost isoperimetric subsets of the discrete cube. Combinatorics, Probability and Computing, 20(3):363–380, May 2011. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Ergemlidze:2022:TFS

[EM22] Beka Ergemlidze and Theodore Molla. Transversal C_{k}-factors in subgraphs of the balanced blow-up of C_{k}. Combinatorics, Probability and Computing, 31(6):1031–1047, November 2022. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/transversal-
ckfactors-in-subgraphs-of-the-balanced-blowup-of-ck/
543CA9541355EFFDFE1F11822FA78268.

Ellis-Monaghan:2015:ETT

[EMM15] J. Ellis-Monaghan and I. Moffatt. Evaluations of topological Tutte polynomials. Combinatorics, Probability and Computing, 24(3):556–583, May 2015. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Erdos:1993:NED

[EMP93] Paul Erdős, Endre Makai, and János Pach. Nearly equal distances in the plane. Combinatorics, Probability and Computing, 2(4):401–408, December 1993. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Eldan:2022:CDP

[EMP22] Ronen Eldan, Dan Mikulincer, and Hester Pieters. Community detection and percolation of information in a geometric setting. Combinatorics, Probability and Computing, 31(6):1048–1069, November 2022. CODEN CPCOFG. ISSN 0963-5483 (print),

1469-2163 (electronic). URL https://www.cambridge.org/
core/journals/combinatorics-probability-and-computing/
article/community-detection-and-percolation-of-information-
in-a-geometric-setting/CF3636A72A9B62A40AFDA0CA1A1A18B6.

Ellis-Monaghan:2007:DHG

[EMS07] Joanna A. Ellis-Monaghan and Irasema Sarmiento. Distance hereditary graphs and the interlace polynomial. Combinatorics, Probability and Computing, 16(6):947–973, November 2007. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=16&issueId=06.

Ellis-Monaghan:2006:PTP

[EMT06] Joanna A. Ellis-Monaghan and Lorenzo Traldi. Parametrized Tutte polynomials of graphs and matroids. Combinatorics, Probability and Computing, 15(6):835–854, November 2006. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic).  URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=15&issueId=06.

Erdos:2018:NCD

[EMT18] Péter L. Erdős, István Miklós, and Zoltán Toroczkai. New classes of degree sequences with fast mixing swap Markov chain sampling. Combinatorics, Probability and Computing, 27(2):186–207, March 2018. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL https://www.cambridge.org/core/
journals/combinatorics-probability-and-computing/article/
new-classes-of-degree-sequences-with-fast-mixing-swap-
markov-chain-sampling/AD6E839043739A421C482D486853D2DA.

Engel:1996:IPC

[Eng96] Konrad Engel. Interval packing and covering in the Boolean lattice. Combinatorics, Probability and Computing, 5(4):373–384, December 1996. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Erdos:1993:CPC

[EOZ93] Paul Erdős, Edward T. Ordman, and Yechezkel Zalcstein. Clique partitions of chordal graphs. Combinatorics, Probability and Computing, 2(4):409–415, December 1993. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Eichelsbacher:1999:CPA

[ER99] Peter Eichelsbacher and Małgorzata Roos. Compound Poisson approximation for dissociated random variables via Stein’s method. Combinatorics, Probability and Computing, 8(4): 335–346, July 1999. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http://journals.

cambridge.org/action/displayIssue?jid=CPC&volumeId=8&

issueId=04. Random graphs and combinatorial structures (Oberwolfach, 1997).

Erdos:1999:SPR

[Erd99] Paul Erdős. A selection of problems and results in combinatorics. Combinatorics, Probability and Computing, 8(1–2):1–6, January 1999. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). Recent trends in combinatorics (Mátraháza, 1995).

Eriksson:1997:AES

[Eri97] Kimmo Eriksson. Autocorrelation and the enumeration of strings avoiding a fixed string. Combinatorics, Probability and Computing, 6(1):45–48, March 1997. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http://journals.

cambridge.org/action/displayIssue?jid=CPC&volumeId=6&
issueId=01.

Elek:2005:PDG

[ES05] Gábor Elek and Vera T. Sós. Paradoxical decompositions and growth conditions. Combinatorics, Probability and Computing, 14(1–2):81–105, January 2005. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=14&issueId=01.

Elekes:2011:ITD

[ES11] György Elekes and Micha Sharir. Incidences in three dimensions and distinct distances in the plane. Combinatorics, Probability and Computing, 20(4):571–608, July 2011. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Eslava:2021:NIT

[Esl21] Laura Eslava. A non-increasing tree growth process for recursive trees and applications. Combinatorics, Probability and Computing, 30(1):79–104, January 2021. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/nonincreasing-
tree-growth-process-for-recursive-trees-and-applications/
041AA200905EF06843C3CBE02161A3E6.

Erdos:1994:ICB

[ESS94] Péter L. Erdős, Ákos Seress, and László A. Székely. On intersecting chains in Boolean algebras. Combinatorics, Probability and Computing, 3(1):57–62, March 1994. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Elekes:2009:CDB

[ESS09] György Elekes, Miklós Simonovits, and Endre Szabó. A combinatorial distinction between unit circles and straight lines: How many coincidences can they have? Combinatorics, Probability and Computing, 18(5):691–705, September 2009. CODEN CP-COFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Eustis:2013:INS

[EV13] Alex Eustis and Jacques Verstraëte. On the independence number of Steiner systems. Combinatorics, Probability and Computing, 22(2):241–252, ???? 2013. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Evans:2004:EMC

[Eva04] Steven N. Evans. Embedding a Markov chain into a random walk on a permutation group. Combinatorics, Probability and Computing, 13(3):339–351, May 2004. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=13&issueId=03.

Emden-Weinert:1998:UCG

[EWHK98] Thomas Emden-Weinert, Stefan Hougardy, and Bernd Kreuter. Uniquely colourable graphs and the hardness of colouring graphs of large girth. Combinatorics, Probability and Computing, 7 (4):375–386, December 1998. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http://journals.

cambridge.org/action/displayIssue?jid=CPC&volumeId=7&
issueId=04.

Even-Zohar:2012:SGS

[EZ12] Chaim Even-Zohar. On sums of generating sets in Z_{2}^{n}. Combinatorics, Probability and Computing, 21(6):916–941, November 2012. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Farr:2006:CCC

[Far06] G. E. Farr. The complexity of counting colourings of subgraphs of the grid. Combinatorics, Probability and Computing, 15(3):377– 383, May 2006. CODEN CPCOFG. ISSN 0963-5483 (print),

1469-2163 (electronic). URL http://journals.cambridge.
org/action/displayIssue?jid=CPC&volumeId=15&issueId=
03.

Farr:2007:ATM

[Far07] G. E. Farr. On the Ashkin–Teller model and Tutte–Whitney functions. Combinatorics, Probability and Computing, 16(2):251– 260, March 2007. CODEN CPCOFG. ISSN 0963-5483 (print),

1469-2163 (electronic). URL http://journals.cambridge.
org/action/displayIssue?jid=CPC&volumeId=16&issueId=
02.

Faudree:2012:MFC

[Fau12] Ralph J. Faudree. My friend and colleague, Richard Schelp. Combinatorics, Probability and Computing, 21(1–2):1–9, ???? 2012. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Fill:1997:NAS

[FD97] James Allen Fill and Robert P. Dobrow. The number of mary search trees on n keys. Combinatorics, Probability and Computing, 6(4):435–453, December 1997. CODEN CPCOFG.

ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=6&issueId=04.

Feldheim:2017:MAC

[Fel17] Ohad N. Feldheim. Monotonicity of avoidance coupling on K_{N}. Combinatorics, Probability and Computing, 26(1):16–23, January 2017. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL https://www.cambridge.org/core/
product/91774A32FA58E8C44D7298BF47496307.

Fujita:2013:FSG

[FFO13] Shinya Fujita, Michitaka Furuya, and Kenta Ozeki. Forbidden subgraphs generating almost the same sets. Combinatorics, Probability and Computing, 22(5):733–748, September 2013. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Flaxman:2007:ADS

[FFV07a] Abraham D. Flaxman, Alan M. Frieze, and Juan Vera. Adversarial deletion in a scale-free random graph process. Combinatorics, Probability and Computing, 16(2):261–270, March 2007. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=16&issueId=02.

Flaxman:2007:ACP

[FFV07b] Abraham D. Flaxman, Alan M. Frieze, and Juan Carlos Vera. On the average case performance of some greedy approximation algorithms for the uncapacitated facility location problem. Combinatorics, Probability and Computing, 16(5):713–732, September 2007. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=16&issueId=05.

Friedland:2008:LBP

[FG08] Shmuel Friedland and Leonid Gurvits. Lower bounds for partial matchings in regular bipartite graphs and applications to the monomer-dimer entropy. Combinatorics, Probability and Computing, 17(3):347–361, May 2008. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=17&issueId=03.

Fulman:2011:ZBJ

[FG11] Jason Fulman and Larry Goldstein. Zero biasing and jack measures. Combinatorics, Probability and Computing, 20(5):753–762, September 2011. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Furedi:2015:ENO

[FG15] Zoltan Füredi and David S. Gunderson. Extremal numbers for odd cycles. Combinatorics, Probability and Computing, 24(4): 641–645, July 2015. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). Special issue on Oberwolfach Special Issue Part 1.

Faudree:2012:MDD

[FGJ12] Ralph J. Faudree, Ronald J. Gould, and Michael S. Jacobson. Minimum degree and disjoint cycles in claw-free graphs. Combinatorics, Probability and Computing, 21(1–2):129–139, ???? 2012. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Furedi:1994:MNT

[FGK94] Zoltán Füredi, Michel X. Goemans, and Daniel J. Kleitman. On the maximum number of triangles in wheel-free graphs. Combinatorics, Probability and Computing, 3(1):63–75, March 1994. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Furedi:2023:PRC

[FGK23] Zoltán Füredi, András Gyárfás, and Zoltán Király. Problems and results on 1-cross-intersecting set pair systems. Combinatorics, Probability and Computing, 32(4):691–702, July 2023. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/problems-
and-results-on-1crossintersecting-set-pair-systems/F76EC4E32A48E5434F04F482B2BD3A46.

Fern:2000:MAS

[FGLP00] Lori Fern, Gary Gordon, Jason Leasure, and Sharon Pronchik. Matroid automorphisms and symmetry groups. Combinatorics, Probability and Computing, 9(2):105–123, March 2000. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic).  URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=9&issueId=02.

Flajolet:1993:DHB

[FGOR93] Philippe Flajolet, Zhicheng Gao, Andrew Odlyzko, and Bruce Richmond. The distribution of heights of binary trees and other simple trees. Combinatorics, Probability and Computing, 2(2): 145–156, June 1993. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Furedi:2005:CMH

[FGS05] Zoltán Füredi, András Gyárfás, and Gábor Simonyi. Connected matchings and Hadwiger’s conjecture. Combinatorics, Probability and Computing, 14(3):435–438, May 2005. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic).  URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=14&issueId=03.

Fiol:2000:TOG

[FGY00] M. A. Fiol, E. Garriga, and J. L. A. Yebra. On twisted odd graphs. Combinatorics, Probability and Computing, 9 (3):227–240, May 2000. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http://journals.

cambridge.org/action/displayIssue?jid=CPC&volumeId=9&
issueId=03.

Firby:1997:MSC

[FH97] Peter Firby and Julie Haviland. Maximal spacing configurations in graphs. Combinatorics, Probability and Computing, 6 (4):455–463, December 1997. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http://journals.

cambridge.org/action/displayIssue?jid=CPC&volumeId=6&
issueId=04.

Foster:2023:BBE

[FH23] James Foster and Karen Habermann. Brownian bridge expansions for Lévy area approximations and particular values of the Riemann zeta function. Combinatorics, Probability and Computing, 32(3):370–397, May 2023. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/brownian-
bridge-expansions-for-levy-area-approximations-and-particular-
values-of-the-riemann-zeta-function/6C1D65580D064B415E14AF19DAE55517.

Felszeghy:2009:APM

[FHR09] Bálint Felszeghy, Gábor Heged’us, and Lajos Rónyai. Algebraic properties of modulo q complete l-wide families. Combinatorics, Probability and Computing, 18(3):309–333, May 2009. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Freschi:2022:DPG

[FHT22] Andrea Freschi, Joseph Hyde, and Andrew Treglown. On deficiency problems for graphs. Combinatorics, Probability and Computing, 31(3):478–488, May 2022. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL https://www.

cambridge.org/core/journals/combinatorics-probability-
and-computing/article/on-deficiency-problems-for-graphs/
749905010B8382DBC4E8C420A1A713FF.

Frati:2023:UGG

[FHT23] Fabrizio Frati, Michael Hoffmann, and Csaba D. Tóth. Universal geometric graphs. Combinatorics, Probability and Computing, 32(5):742–761, September 2023. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL https://www.

cambridge.org/core/journals/combinatorics-probability-
and-computing/article/universal-geometric-graphs/CC9F25476CCC26533B14FE910E707D8F.

Fan:2018:BGS

[FHY18] Genghua Fan, Jianfeng Hou, and Xingxing Yu. Bisections of graphs without short cycles. Combinatorics, Probability and Computing, 27(1):44–59, January 2018. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/product/
DCE64149EA6A004047A95F4550968FFE.

Filmus:2020:FTM

[Fil20] Yuval Filmus. FKN theorem for the multislice, with applications. Combinatorics, Probability and Computing, 29(2):200–212, March 2020. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL https://www.cambridge.org/core/
journals/combinatorics-probability-and-computing/article/
fkn-theorem-for-the-multislice-with-applications/FB5F9A27D51E18D7804F4B371257118F.

Fill:2021:BBR

[Fil21] James Allen Fill. Breaking bivariate records. Combinatorics, Probability and Computing, 30(1):105–123, January 2021. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/breaking-
bivariate-records/02A030814A8C56F1C334EF9CBA868261.

Fiol:2001:SSC

[Fio01] M. A. Fiol. Some spectral characterizations of strongly distance-regular graphs. Combinatorics, Probability and Computing, 10(2):127–135, March 2001. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=10&issueId=02.

Fiol:2003:GSB

[Fio03] M. A. Fiol. A general spectral bound for distant vertex subsets. Combinatorics, Probability and Computing, 12(1):17–26, January 2003. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=12&issueId=01.

Frieze:2018:EDS

[FJ18] Alan Frieze and Tony Johansson. On edge-disjoint spanning trees in a randomly weighted complete graph. Combinatorics, Probability and Computing, 27(2):228–244, March 2018. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/on-edgedisjoint-
spanning-trees-in-a-randomly-weighted-complete-graph/
C2FA711CCC2218FFA03C064CBE932F37.

Froncek:2007:SID

[FJKK07] Dalibor Froncek, Janja Jerebic, Sandi Klavzar, and Petr Kovár. Strong isometric dimension, biclique coverings, and Sperner’s Theorem. Combinatorics, Probability and Computing, 16(2):271– 275, March 2007. CODEN CPCOFG. ISSN 0963-5483 (print),

1469-2163 (electronic). URL http://journals.cambridge.
org/action/displayIssue?jid=CPC&volumeId=16&issueId=
02.

Ferrara:2013:DCL

[FJP13] Michael Ferrara, Michael Jacobson, and Florian Pfender. Degree conditions for H-linked digraphs. Combinatorics, Probability and Computing, 22(5):684–699, September 2013. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Fiorini:2013:EFM

[FJW13] Samuel Fiorini, Gwenaël Joret, and David R. Wood. Excluded forest minors and the Erdős–Pósa property. Combinatorics, Probability and Computing, 22(5):700–721, September 2013. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Ferber:2022:NHM

[FJZ22] Asaf Ferber, Vishesh Jain, and Yufei Zhao. On the number of Hadamard matrices via anti-concentration. Combinatorics, Probability and Computing, 31(3):455–477, May 2022. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/on-the-
number-of-hadamard-matrices-via-anticoncentration/887EFBBF79B804BCDD942029283D4CD7.

Fachini:2001:NCV

[FK01] Emanuela Fachini and János Körner. A note on counting very different sequences. Combinatorics, Probability and Computing, 10(6):501–504, November 2001. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=10&issueId=06.

Friedgut:2005:NHC

[FK05] Ehud Friedgut and Jeff Kahn. On the number of Hamiltonian cycles in a tournament. Combinatorics, Probability and Computing, 14(5–6):769–781, November 2005. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=14&issueId=05.

Furedi:2006:BQ

[FK06] Zoltán Füredi and Gyula O. H. Katona. 2-bases of quadruples. Combinatorics, Probability and Computing, 15(1–2):131–141, January 2006. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL http://journals.cambridge.org/
action/displayIssue?jid=CPC&volumeId=15&issueId=01.

Feldheim:2008:WFS

[FK08] Ohad N. Feldheim and Michael Krivelevich. Winning fast in sparse graph construction games. Combinatorics, Probability and Computing, 17(6):781–791, November 2008. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Felsner:2009:ULB

[FK09] Stefan Felsner and Kolja B. Knauer. ULD-lattices and Δbonds. Combinatorics, Probability and Computing, 18(5):707– 724, September 2009. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Frieze:2013:NPR

[FK13] Alan Frieze and Michael Krivelevich. On the non-planarity of a random subgraph. Combinatorics, Probability and Computing, 22(5):722–732, September 2013. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Furer:2014:ACE

[FK14] Martin Fürer and Shiva Prasad Kasiviswanathan. Approximately counting embeddings into random graphs. Combinatorics, Probability and Computing, 23(6):1028–1056, November 2014. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). Special issue on Honouring the Memory of Philippe Flajolet - Part 2.

Frankl:2017:UCI

[FK17] Peter Frankl and Andrey Kupavskii. Uniform s-cross-intersecting families. Combinatorics, Probability and Computing, 26(4):517– 524, July 2017. CODEN CPCOFG. ISSN 0963-5483 (print),

1469-2163 (electronic). URL https://www.cambridge.org/
core/product/35ABB02D3123D9B84A84FF5159A437EA.

Frankl:2018:CIP

[FK18] Peter Frankl and Andrey Kupavskii. Counting intersecting and pairs of cross-intersecting families. Combinatorics, Probability and Computing, 27(1):60–68, January 2018. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/product/
CADB0D089786311C5B15CA4122531ADE.

Fachini:2001:SSB

[FKM01] Emanuela Fachini, János Körner, and Angelo Monti. Selfsimilarity bounds for locally thin set families. Combinatorics, Probability and Computing, 10(4):309–315, July 2001. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic).  URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=10&issueId=04.

Fountoulakis:2016:MOT

[FKP16] Nikolaos Fountoulakis, Megha Khosla, and Konstantinos Panagiotou. The multiple-orientability thresholds for random hypergraphs. Combinatorics, Probability and Computing, 25(6): 870–908, November 2016. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL https://www.cambridge.

org/core/product/A6ED335A0C90E5DBEC6D1DFBCF0D674B.

Frieze:2005:GJ

[FKPS05] Alan Frieze, Michael Krivelevich, Oleg Pikhurko, and Tibor Szabó. The game of JumbleG. Combinatorics, Probability and Computing, 14(5–6):783–793, November 2005. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic).  URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=14&issueId=05.

Friedgut:2003:RGA

[FKR^{+}03] Ehud Friedgut, Yoshiharu Kohayakawa, Vojtěch Rödl, Andrzej Ruciński, and Prasad Tetali. Ramsey games against a onearmed bandit. Combinatorics, Probability and Computing, 12 (5–6):515–545, November 2003. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). Special issue on Ramsey theory.

Ferrara:2005:DGI

[FKR05] M. Ferrara, Y. Kohayakawa, and V. Rödl. Distance graphs on the integers. Combinatorics, Probability and Computing, 14(1–2):107–131, January 2005. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=14&issueId=01.

Frieze:2009:LSN

[FKRD09] Alan Frieze, Jon Kleinberg, R. Ravi, and Warren Debany. Lineof-sight networks. Combinatorics, Probability and Computing, 18 (1–2):145–163, March 2009. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Frieze:2007:CNR

[FKS07] Alan Frieze, Michael Krivelevich, and Cliff Smyth. On the chromatic number of random graphs with a fixed degree sequence. Combinatorics, Probability and Computing, 16(5):733–746, September 2007. CODEN CPCOFG. ISSN 0963-5483 (print),

1469-2163 (electronic). URL http://journals.cambridge.
org/action/displayIssue?jid=CPC&volumeId=16&issueId=
05.

Fox:2010:DGS

[FKS10] Jacob Fox, Peter Keevash, and Benny Sudakov. Directed graphs without short cycles. Combinatorics, Probability and Computing, 19(2):285–301, March 2010. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Ferber:2022:SSR

[FKS22] Asaf Ferber, Matthew Kwan, and Lisa Sauermann. Singularity of sparse random matrices: simple proofs. Combinatorics, Probability and Computing, 31(1):21–28, January 2022. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/singularity-
of-sparse-random-matrices-simple-proofs/C6661A163329C87519145BE42BAA8E52.

Fox:2020:RNH

[FL20] Jacob Fox and Ray Li. On Ramsey numbers of hedgehogs. Combinatorics, Probability and Computing, 29(1):101–112, January 2020. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL https://www.cambridge.org/core/
journals/combinatorics-probability-and-computing/article/
on-ramsey-numbers-of-hedgehogs/B0DC14A250A162A462CA9CC65560F84F.

Ferber:2021:RRR

[FLM21] Asaf Ferber, Kyle Luh, and Gweneth McKinley. Resilience of the rank of random matrices. Combinatorics, Probability and Computing, 30(2):163–174, March 2021. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/resilience-
of-the-rank-of-random-matrices/7FA16339531066EDE2BCABC6B46E58FB.

Ferber:2017:PLH

[FLMN17] Asaf Ferber, Kyle Luh, Daniel Montealegre, and Oanh Nguyen. Packing loose Hamilton cycles. Combinatorics, Probability and Computing, 26(6):839–849, November 2017. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/product/
C3605ADC21B383BA7ADF896CFFB1BC02.

Frein:2008:GAS

[FLS08] Yannick Frein, Benjamin Lévêque, and András Seb’o. Generating all sets with bounded unions. Combinatorics, Probability and Computing, 17(5):641–660, September 2008. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Fujita:2019:DEC

[FLW19] Shinya Fujita, Ruonan Li, and Guanghui Wang. Decomposing edge-coloured graphs under colour degree constraints. Combinatorics, Probability and Computing, 28(5):755–767, September 2019. CODEN CPCOFG. ISSN 0963-5483 (print),

1469-2163 (electronic). URL /core/journals/combinatorics-
probability-and-computing/article/decomposing-edgecoloured-
graphs-under-colour-degree-constraints/9FCF9A373C0FFABB4B1978CA63D3C903.

Fox:2017:RLT

[FLZ17] Jacob Fox, László Miklós Lovász, and Yufei Zhao. On regularity lemmas and their algorithmic applications. Combinatorics, Probability and Computing, 26(4):481–505, July 2017. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/product/
73EB6DAA7B1CFA2518C13B451F46EF1C. See erratum [FLZ18].

Fox:2018:ERL

[FLZ18] Jacob Fox, László Miklós Lovász, and Yufei Zhao. Erratum: On Regularity Lemmas and their Algorithmic Applications. Combinatorics, Probability and Computing, 27(5):851–852, September 2018. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL https://www.cambridge.org/core/
journals/combinatorics-probability-and-computing/article/
erratum-on-regularity-lemmas-and-their-algorithmic-applications/
2DFE5362C10D46B374D639F1BBD3B26C. See [FLZ17].

Fox:2019:FNA

[FLZ19] Jacob Fox, László Miklós Lovász, and Yufei Zhao. A fast new algorithm for weak graph regularity. Combinatorics, Probability and Computing, 28(5):777–790, September 2019. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL /core/journals/combinatorics-
probability-and-computing/article/fast-new-algorithm-
for-weak-graph-regularity/E17D06308910088718F987EB1C8A6560.

Fill:2014:QTP

[FM14] James Allen Fill and Jason Matterer. QuickSelect tree process convergence, with an application to distributional convergence for the number of symbol comparisons used by worst-case find. Combinatorics, Probability and Computing, 23(5):805–828, September 2014. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). Special issue on Honouring the Memory of Philippe Flajolet - Part 1.

Furedi:2017:MNT

[FM17] Zoltán Füredi and Zeinab Maleki. The minimum number of triangular edges and a symmetrization method for multiple graphs. Combinatorics, Probability and Computing, 26(4):525–535, July 2017. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL https://www.cambridge.org/core/
product/609A04D622AA730BCECA86D0017837B0.

Furedi:2009:CFS

[FÖ09] Zoltán Füredi and Lale Özkahya. On 14-cycle-free subgraphs of the hypercube. Combinatorics, Probability and Computing, 18 (5):725–729, September 2009. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Frankl:1998:UIF

[FOT98] P. Frankl, K. Ota, and N. Tokushige. Uniform intersecting families with covering number restrictions. Combinatorics, Probability and Computing, 7(1):47–56, March 1998. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=7&issueId=01.

Fan:1992:DRS

[FP92] Y. Fan and J. K. Percus. Duality for random sequential adsorption on a lattice. Combinatorics, Probability and Computing, 1 (3):219–222, September 1992. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Fernandez:2008:RCZ

[FP08] Roberto Fernández and Aldo Procacci. Regions without complex zeros for chromatic polynomials on graphs with bounded degree. Combinatorics, Probability and Computing, 17(2):225–238, March 2008. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL http://journals.cambridge.org/
action/displayIssue?jid=CPC&volumeId=17&issueId=02.

Fox:2010:STS

[FP10] Jacob Fox and János Pach. A separator theorem for string graphs and its applications. Combinatorics, Probability and Computing, 19(3):371–390, May 2010. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Fountoulakis:2011:CCR

[FP11] Nikolaos Fountoulakis and Konstantinos Panagiotou. 3connected cores in random planar graphs. Combinatorics, Probability and Computing, 20(3):381–412, May 2011. CODEN CP-COFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Fox:2014:ANS

[FP14] Jacob Fox and János Pach. Applications of a new separator theorem for string graphs. Combinatorics, Probability and Computing, 23(1):66–74, January 2014. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Furedi:2005:TSI

[FPS05] Zoltán Füredi, Oleg Pikhurko, and Miklós Simonovits. On triple systems with independent neighbourhoods. Combinatorics, Probability and Computing, 14(5–6):795–813, November 2005. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=14&issueId=05.

Fu:2001:CGD

[FR01] H. L. Fu and C. A. Rodger. 4-cycle group-divisible designs with two associate classes. Combinatorics, Probability and Computing, 10(4):317–343, July 2001. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=10&issueId=04.

Falgas-Ravry:2015:SPI

[FR15] Victor Falgas-Ravry. Sperner’s problem for G-independent families. Combinatorics, Probability and Computing, 24(3):528–550, May 2015. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Frankl:2017:SRF

[Fra17] Peter Frankl. A stability result for families with fixed diameter. Combinatorics, Probability and Computing, 26(4):506–516, July 2017. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL https://www.cambridge.org/core/
product/D090AB65516766E67F1488A47945006D.

Frankl:2019:NEI

[Fra19] Peter Frankl. A near-exponential improvement of a bound of Erdős and Lovász on maximal intersecting families. Combinatorics, Probability and Computing, 28(5):733–739, September 2019. CODEN CPCOFG. ISSN 0963-5483 (print),

1469-2163 (electronic). URL /core/journals/combinatorics-
probability-and-computing/article/nearexponential-improvement-
of-a-bound-of-erdos-and-lovasz-on-maximal-intersecting-
families/FFD1BAC368E23B845516A49BC1FC7441.

Franks:2021:SDB

[Fra21] Cole Franks. A simplified disproof of Beck’s three permutations conjecture and an application to root-mean-squared discrepancy. Combinatorics, Probability and Computing, 30(3):398–411, May 2021. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL https://www.cambridge.org/core/
journals/combinatorics-probability-and-computing/article/
simplified-disproof-of-becks-three-permutations-conjecture-
and-an-application-to-rootmeansquared-discrepancy/4025C9CD6D1BB769B47625F4D647B731.

Friedgut:2004:IPS

[Fri04] Ehud Friedgut. Influences in product spaces: KKL and BKKKL revisited. Combinatorics, Probability and Computing, 13(1):17– 29, January 2004. CODEN CPCOFG. ISSN 0963-5483 (print),

1469-2163 (electronic). URL http://journals.cambridge.
org/action/displayIssue?jid=CPC&volumeId=13&issueId=
01.

Falgas-Ravry:2021:TDG

[FRMZ21] Victor Falgas-Ravry, Klas Markström, and Yi Zhao. Triangledegrees in graphs and tetrahedron coverings in 3-graphs. Combinatorics, Probability and Computing, 30(2):175–199, March 2021. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/triangledegrees-
in-graphs-and-tetrahedron-coverings-in-3graphs/B671A519BCD972559E885DB6E4B2B50C.

Frankl:2012:MNE

[FRR12] Peter Frankl, Vojtech Rödl, and Andrzej Ruciński. On the maximum number of edges in a triple system not containing a disjoint family of a given size. Combinatorics, Probability and Computing, 21(1–2):141–148, ???? 2012. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Frieze:1995:ASG

[FRS95] Alan Frieze, A. J. Radcliffe, and Stephen Suen. Analysis of a simple greedy matching algorithm on random cubic graphs. Combinatorics, Probability and Computing, 4(1):47–66, March 1995. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Falgas-Ravry:2013:ASD

[FRV13] Victor Falgas-Ravry and Emil R. Vaughan. Applications of the semi-definite method to the Turán density problem for 3-graphs. Combinatorics, Probability and Computing, 22(1):21–54, ???? 2013. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Furedi:2005:TSC

[FS05] Zoltán Füredi and Miklós Simonovits. Triple systems not containing a Fano configuration. Combinatorics, Probability and Computing, 14(4):467–484, July 2005. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=14&issueId=04.

Falik:2007:EII

[FS07] Dvir Falik and Alex Samorodnitsky. Edge-isoperimetric inequalities and influences. Combinatorics, Probability and Computing, 16(5):693–712, September 2007. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=16&issueId=05.

Fox:2010:DSS

[FS10] Jacob Fox and Benny Sudakov. Decompositions into subgraphs of small diameter. Combinatorics, Probability and Computing, 19 (5–6):753–774, November 2010. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Fujisawa:2012:PFS

[FS12] Jun Fujisawa and Akira Saito. A pair of forbidden subgraphs and 2-factors. Combinatorics, Probability and Computing, 21(1– 2):149–158, ???? 2012. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Feldheim:2018:GCO

[FS18] Ohad Noy Feldheim and Yinon Spinka. The growth constant of odd cutsets in high dimensions. Combinatorics, Probability and Computing, 27(2):208–227, March 2018. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/growth-
constant-of-odd-cutsets-in-high-dimensions/F27A3FE99E1C9EBC27346DF9ABE2C5E6.

Ferber:2020:QLC

[FS20] Asaf Ferber and Asaf Shapira. A quantitative Lovász criterion for Property B. Combinatorics, Probability and Computing, 29(6): 956–960, November 2020. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL https://www.cambridge.

org/core/journals/combinatorics-probability-and-computing/
article/quantitative-lovasz-criterion-for-property-b/
70A24EAB2FF2D5928265C7B54C87B31D.

Ferber:2023:SRP

[FSSZ23] Asaf Ferber, Ashwin Sah, Mehtaab Sawhney, and Yizhe Zhu. Sparse recovery properties of discrete random matrices. Combinatorics, Probability and Computing, 32(2):316–325, March 2023. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/sparse-
recovery-properties-of-discrete-random-matrices/654DC865362DE107A143C1C98E4F3732.

Frankl:1998:SIC

[FT98] P. Frankl and N. Tokushige. Some inequalities concerning crossintersecting families. Combinatorics, Probability and Computing, 7(3):247–260, September 1998. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http://journals.

cambridge.org/action/displayIssue?jid=CPC&volumeId=7&
issueId=03.

Frankl:2011:CIF

[FT11] Peter Frankl and Norihide Tokushige. On r-cross intersecting families of sets. Combinatorics, Probability and Computing, 20 (5):749–752, September 2011. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Fuchs:2008:SSR

[Fuc08] Michael Fuchs. Subtree sizes in recursive trees and binary search trees: Berry–Esseen bounds and Poisson approximations. Combinatorics, Probability and Computing, 17(5):661–680, September 2008. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Fuchs:2012:LTS

[Fuc12] Michael Fuchs. Limit theorems for subtree size profiles of increasing trees. Combinatorics, Probability and Computing, 21 (3):412–441, May 2012. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Furedi:1996:UBZ

[Für96] Zoltán Füredi. An upper bound on Zarankiewicz’ problem. Combinatorics, Probability and Computing, 5(1):29–33, March 1996. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Furedi:2012:CHC

[Für12] Zoltán Füredi. 2-cancellative hypergraphs and codes. Combinatorics, Probability and Computing, 21(1–2):159–177, ???? 2012. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Federico:2017:CWC

[FV17] Lorenzo Federico and Remco Van Der Hofstad. Critical window for connectivity in the configuration model. Combinatorics, Probability and Computing, 26(5):660–680, September 2017. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL https://www.cambridge.org/core/
product/F06C1CA390A309CF8975C7419D364533.

Federico:2020:EPC

[FVDH20] Lorenzo Federico, Remco Van Der Hofstad, Frank Den Hollander, and Tim Hulshof. Expansion of percolation critical points for Hamming graphs. Combinatorics, Probability and Computing, 29(1):68–100, January 2020. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/expansion-
of-percolation-critical-points-for-hamming-graphs/09C142CBCB3ED00646355E97200FD3A8.

Furstenberg:2003:MPR

[FW03] Hillel Furstenberg and Benjamin Weiss. Markov processes and Ramsey theory for trees. Combinatorics, Probability and Computing, 12(5–6):547–563, November 2003. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). Special issue on Ramsey theory.

Fox:2018:FPT

[FW18] Jacob Fox and Fan Wei. Fast property testing and metrics for permutations. Combinatorics, Probability and Computing, 27(4): 539–579, July 2018. CODEN CPCOFG. ISSN 0963-5483 (print),

1469-2163 (electronic). URL https://www.cambridge.org/
core/journals/combinatorics-probability-and-computing/
article/fast-property-testing-and-metrics-for-permutations/
A1847C895C28605E0BFB1015C43D1DDB.

Frieze:2000:OCE

[FZ00] Alan M. Frieze and Lei Zhao. Optimal construction of edgedisjoint paths in random regular graphs. Combinatorics, Probability and Computing, 9(3):241–263, May 2000. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic).  URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=9&issueId=03.

Fox:2022:RLA

[FZ22] Jacob Fox and Yufei Zhao. Removal lemmas and approximate homomorphisms. Combinatorics, Probability and Computing, 31 (4):721–736, July 2022. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL https://www.cambridge.

org/core/journals/combinatorics-probability-and-computing/
article/removal-lemmas-and-approximate-homomorphisms/
A0618B680623DD624744FA294C9F2504.

Goel:2019:DDA

[GA19] Gopal Goel and Andrew Ahn. Discrete derivative asymptotics of the β-Hermite eigenvalues. Combinatorics, Probability and Computing, 28(5):657–674, September 2019. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL /core/journals/combinatorics-probability-
and-computing/article/discrete-derivative-asymptotics-
of-the-hermite-eigenvalues/D33D217D66D21E8B064414A70C7AA620.

Gacs:2000:CDH

[Gác00] Peter Gács. The clairvoyant demon has a hard task. Combinatorics, Probability and Computing, 9(5):421–424, September 2000. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=9&issueId=05.

Gacs:2004:CSS

[Gác04] Peter Gács. Compatible sequences and a slow Winkler percolation. Combinatorics, Probability and Computing, 13(6):815–856, November 2004. CODEN CPCOFG. ISSN 0963-5483 (print),

1469-2163 (electronic). URL http://journals.cambridge.
org/action/displayIssue?jid=CPC&volumeId=13&issueId=
06.

Galvin:2011:TPR

[Gal11] David Galvin. A threshold phenomenon for random independent sets in the discrete hypercube. Combinatorics, Probability and Computing, 20(1):27–51, January 2011. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Gao:2014:FRS

[Gao14] Pu Gao. The first k-regular subgraph is large. Combinatorics, Probability and Computing, 23(3):412–433, May 2014. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Gates:1993:SDP

[Gat93] John Gates. Some dual problems of geometric probability in the plane. Combinatorics, Probability and Computing, 2(1):11–23, March 1993. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Goodall:2011:TPC

[GdMNN11] A. J. Goodall, A. de Mier, S. D. Noble, and M. Noy. The Tutte polynomial characterizes simple outerplanar graphs. Combinatorics, Probability and Computing, 20(4):609–616, July 2011.

CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Gebauer:2013:SRN

[Geb13] Heidi Gebauer. Size Ramsey number of bounded degree graphs for games. Combinatorics, Probability and Computing, 22(4): 499–516, July 2013. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Georgii:2008:BRR

[Geo08] Hans-Otto Georgii. Book review: “The Random-Cluster Model”, by Geoffrey Grimmett, Springer, 2006, 377pp., £69.00/$125.00/89.95 Euros. Combinatorics, Probability and Computing, 17(2):317–318, March 2008. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=17&issueId=02.

Gamarnik:2010:RGA

[GG10] David Gamarnik and David A. Goldberg. Randomized greedy algorithms for independent sets and matchings in regular graphs: Exact results and finite girth corrections. Combinatorics, Probability and Computing, 19(1):61–85, January 2010. CODEN CP-COFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Gross:2019:ISB

[GG19] Renan Gross and Uri Grupel. Indistinguishable sceneries on the Boolean hypercube. Combinatorics, Probability and Computing, 28(1):46–60, January 2019. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL https://www.cambridge.

org/core/journals/combinatorics-probability-and-computing/
article/indistinguishable-sceneries-on-the-boolean-hypercube/
8C4E5A41F6492C2F6D80BEC49F79FD0F.

Girard:2012:SAG

[GGH12] Benjamin Girard, Simon Griffiths, and Yahya Ould Hamidoune. k-sums in Abelian groups. Combinatorics, Probability and Computing, 21(4):582–596, July 2012. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Glock:2021:CHC

[GGJ^{+}21] Stefan Glock, Stephen Gould, Felix Joos, Daniela Kühn, and Deryk Osthus. Counting Hamilton cycles in Dirac hypergraphs.

Combinatorics, Probability and Computing, 30(4):631–653, July 2021. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL https://www.cambridge.org/core/
journals/combinatorics-probability-and-computing/article/
counting-hamilton-cycles-in-dirac-hypergraphs/DED5BD050BAD9AC439BAFD59CC81C80A.

Gao:2016:SAG

[GGX16] Weidong Gao, David J. Grynkiewicz, and Xingwu Xia. On nsums in an Abelian group. Combinatorics, Probability and Computing, 25(3):419–435, May 2016. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Gao:1998:ZSA

[GH98] W. D. Gao and Y. O. Hamidoune. Zero sums in Abelian groups. Combinatorics, Probability and Computing, 7(3): 261–263, September 1998. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http://journals.

cambridge.org/action/displayIssue?jid=CPC&volumeId=7&
issueId=03.

Gao:2017:MCT

[GH17] Wei Gao and Jie Han. Minimum codegree threshold for C_{6}^{3}-factors in 3-uniform hypergraphs. Combinatorics, Probability and Computing, 26(4):536–559, July 2017. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/product/
811610BD9AEEA2BCFA699D97D9D89CA8.

Georgakopoulos:2020:PIG

[GH20a] Agelos Georgakopoulos and John Haslegrave. Percolation on an infinitely generated group. Combinatorics, Probability and Computing, 29(4):587–615, July 2020. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/percolation-
on-an-infinitely-generated-group/7AEFCF8A59FADB48F27A3E959CFC2E80.

Graf:2020:FIT

[GH20b] Alessandra Graf and Penny Haxell. Finding independent transversals efficiently. Combinatorics, Probability and Computing, 29(5):780–806, September 2020. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL https://www.

cambridge.org/core/journals/combinatorics-probability-

and-computing/article/finding-independent-transversals-
efficiently/F2503D2F735E5ADC2D4640938183CE64.

Goddyn:2006:BSC

[GHH06] Luis Goddyn, Petr Hlinený, and Winfried Hochstättler. Balanced signings and the chromatic number of oriented matroids. Combinatorics, Probability and Computing, 15(4):523–539, July 2006. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=15&issueId=04.

Gauy:2017:EKR

[GHO17] Marcelo M. Gauy, Hiêp Hàn, and Igor C. Oliveira. Erdős– Ko–Rado for random hypergraphs: Asymptotics and stability. Combinatorics, Probability and Computing, 26(3):406–422, May 2017. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL https://www.cambridge.org/core/
product/0BF9A0CD5AA0B89DDE292D291FC35A59.

Georgakopoulos:2022:PTC

[GHSS22] Agelos Georgakopoulos, John Haslegrave, Thomas Sauerwald, and John Sylvester. The power of two choices for random walks. Combinatorics, Probability and Computing, 31(1):73–100, January 2022. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL https://www.cambridge.org/core/
journals/combinatorics-probability-and-computing/article/
power-of-two-choices-for-random-walks/6792931BB35F640A3DF2844D33BC8AF2.

Gao:2019:CCT

[GHZ19] Wei Gao, Jie Han, and Yi Zhao. Codegree conditions for tiling complete k -partite k -graphs and loose cycles. Combinatorics, Probability and Computing, 28(6):840–870, November 2019. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL /core/journals/combinatorics-probability-
and-computing/article/codegree-conditions-for-tiling-
complete-kpartite-kgraphs-and-loose-cycles/40A58319656D7A7EC15932E3DEA82E1B.

Greenhill:2022:STR

[GIL22] Catherine Greenhill, Mikhail Isaev, and Gary Liang. Spanning trees in random regular uniform hypergraphs. Combinatorics, Probability and Computing, 31(1):29–53, January 2022. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/

combinatorics-probability-and-computing/article/spanning-
trees-in-random-regular-uniform-hypergraphs/78D9710FC7CC7605CE9FFCB842964CBD.

Gnedin:2012:GET

[GIM12] Alexander Gnedin, Alexander Iksanov, and Alexander Marynych. A generalization of the Erdős–Turán law for the order of random permutation. Combinatorics, Probability and Computing, 21(5): 715–733, September 2012. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Greenhill:2021:SCD

[GIM21] Catherine Greenhill, Mikhail Isaev, and Brendan D. McKay. Subgraph counts for dense random graphs with specified degrees. Combinatorics, Probability and Computing, 30(3):460–497, May 2021. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL https://www.cambridge.org/core/
journals/combinatorics-probability-and-computing/article/
subgraph-counts-for-dense-random-graphs-with-specified-
degrees/9D52A45D91060244A52FEE668F5C6E93.

Greenhill:2023:DSS

[GIMM23] Catherine Greenhill, Mikhail Isaev, Tamás Makai, and Brendan D. McKay. Degree sequences of sufficiently dense random uniform hypergraphs. Combinatorics, Probability and Computing, 32(2):183–224, March 2023. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL https://www.

cambridge.org/core/journals/combinatorics-probability-
and-computing/article/degree-sequences-of-sufficiently-
dense-random-uniform-hypergraphs/67BBA22E7C0A384E92CD13443AD2CBAE.

Goddyn:1999:RCB

[GJ99] Luis A. Goddyn and Bill Jackson. Removable circuits in binary matroids. Combinatorics, Probability and Computing, 8 (6):539–545, November 1999. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http://journals.

cambridge.org/action/displayIssue?jid=CPC&volumeId=8&
issueId=06.

Goldberg:2002:BPC

[GJ02] Leslie Ann Goldberg and Mark Jerrum. The ‘Burnside Process’ converges slowly. Combinatorics, Probability and Computing, 11(1):21–34, January 2002. CODEN CPCOFG.

ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=11&issueId=01.

Goerdt:2003:SRR

[GJ03] Andreas Goerdt and Tomasz Jurdziński. Some results on random unsatisfiable k-SAT instances and approximation algorithms applied to random structures. Combinatorics, Probability and Computing, 12(3):245–267, May 2003. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&

volumeId=12&issueId=03. Combinatorics, probability and computing (Oberwolfach, 2001).

Goldberg:2007:CFI

[GJ07] Eslie Ann Goldberg and Mark Jerrum. The complexity of ferromagnetic Ising with local fields. Combinatorics, Probability and Computing, 16(1):43–61, January 2007. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=16&issueId=01.

Gowers:2021:GRS

[GJ21a] W. T. Gowers and Barnabás Janzer. Generalizations of the Ruzsa–Szemerédi and rainbow Turán problems for cliques. Combinatorics, Probability and Computing, 30(4):591–608, July 2021. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/generalizations-
of-the-ruzsaszemeredi-and-rainbow-turan-problems-for-
cliques/3ACA5B930E52081F0717810A454B4DB6.

Guo:2021:ACB

[GJ21b] Heng Guo and Mark Jerrum. Approximately counting bases of bicircular matroids. Combinatorics, Probability and Computing, 30 (1):124–135, January 2021. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL https://www.cambridge.

org/core/journals/combinatorics-probability-and-computing/
article/approximately-counting-bases-of-bicircular-matroids/
1503038473E8C262382C9D231718F2DF.

Garmo:2004:GRR

[GJK04] Hans Garmo, Svante Janson, and Michal Karonski. On generalized random railways. Combinatorics, Probability and Computing, 13(1):31–35, January 2004. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=13&issueId=01.

Greenhill:2002:PPC

[GJKW02] Catherine Greenhill, Svante Janson, Jeong Han Kim, and Nicholas C. Wormald. Permutation pseudographs and contiguity. Combinatorics, Probability and Computing, 11(3):273–298, May 2002. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=11&issueId=03.

Greenhill:2010:NPM

[GJR10] Catherine Greenhill, Svante Janson, and Andrzej Ruciński. On the number of perfect matchings in random lifts. Combinatorics, Probability and Computing, 19(5–6):791–817, November 2010. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Gnedin:2001:CGD

[GK01] Alexander Gnedin and Sergei Kerov. A characterization of GEM distributions. Combinatorics, Probability and Computing, 10(3): 213–217, May 2001. CODEN CPCOFG. ISSN 0963-5483 (print),

1469-2163 (electronic). URL http://journals.cambridge.
org/action/displayIssue?jid=CPC&volumeId=10&issueId=
03.

Galvin:2004:PTH

[GK04] David Galvin and Jeff Kahn. On phase transition in the hard-core model on Z^{d}. Combinatorics, Probability and Computing, 13(2):137–164, March 2004. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=13&issueId=02.

Girao:2017:GMC

[GKP17] António Girão, Teeradej Kittipassorn, and Kamil Popielarz. Generalized majority colourings of digraphs. Combinatorics, Probability and Computing, 26(6):850–855, November 2017. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/product/
8560D452B38B085A45C17E6662563527.

Goodall:2018:TPM

[GKRV18] Andrew Goodall, Thomas Krajewski, Guus Regts, and Lluís Vena. A Tutte polynomial for maps. Combinatorics, Probability and Computing, 27(6):913–945, November 2018. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/tutte-
polynomial-for-maps/9C9A2E83973A585E7F163723D846DE28.

Grabner:2014:GAS

[GKW14] Peter J. Grabner, Arnold Knopfmacher, and Stephan Wagner. A general asymptotic scheme for the analysis of partition statistics. Combinatorics, Probability and Computing, 23(6):1057– 1086, November 2014. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). Special issue on Honouring the Memory of Philippe Flajolet - Part 2.

Griggs:2009:FSF

[GL09] Jerrold R. Griggs and Linyuan Lu. On families of subsets with a forbidden subposet. Combinatorics, Probability and Computing, 18(5):731–748, September 2009. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Gyori:2012:HNC

[GL12a] Ervin Győri and Nathan Lemons. Hypergraphs with no cycle of a given length. Combinatorics, Probability and Computing, 21 (1–2):193–201, ???? 2012. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Gyori:2012:MNT

[GL12b] Ervin Győri and Hao Li. The maximum number of triangles in C_{2k+1}-free graphs. Combinatorics, Probability and Computing, 21(1–2):187–191, ???? 2012. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Gruslys:2018:MNT

[GL18] Vytautas Gruslys and Shoham Letzter. Minimizing the number of triangular edges. Combinatorics, Probability and Computing, 27(4):580–622, July 2018. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL https://www.cambridge.

org/core/journals/combinatorics-probability-and-computing/
article/minimizing-the-number-of-triangular-edges/5B9368B4340E69D25AE0EC5661DDA156.

Gowers:2021:LIS

[GL21a] W. T. Gowers and J. Long. The length of an s-increasing sequence of r-tuples. Combinatorics, Probability and Computing, 30(5):686–721, September 2021. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL https://www.

cambridge.org/core/journals/combinatorics-probability-
and-computing/article/length-of-an-sincreasing-sequence-
of-rtuples/7301418D47DB1ECD6BE71C20E8A98D0A.

Gruslys:2021:CPR

[GL21b] Vytautas Gruslys and Shoham Letzter. Cycle partitions of regular graphs. Combinatorics, Probability and Computing, 30(4): 526–549, July 2021. CODEN CPCOFG. ISSN 0963-5483 (print),

1469-2163 (electronic). URL https://www.cambridge.org/
core/journals/combinatorics-probability-and-computing/
article/cycle-partitions-of-regular-graphs/38EA443A8ED4AE63A812431A410DFF2F.

Gurvits:2021:CMC

[GL21c] Leonid Gurvits and Jonathan Leake. Counting matchings via capacity-preserving operators. Combinatorics, Probability and Computing, 30(6):956–981, November 2021. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/counting-
matchings-via-capacitypreserving-operators/9CF63D280E8003E749CEE81A9C2859C7.

Glasscock:2016:MTT

[Gla16] Daniel Glasscock. Marstrand-type theorems for the counting and mass dimensions in Z. Combinatorics, Probability and Computing, 25(5):700–743, September 2016. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/product/
B04A5D22570ED534A737BF11BB90E191.

Grzesik:2022:TCG

[GLLV22] Andrzej Grzesik, Joonkyung Lee, Bernard Lidický, and Jan Volec. On tripartite common graphs. Combinatorics, Probability and Computing, 31(5):907–923, September 2022. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/on-tripartite-
common-graphs/1B3EF2A2954FEAF3EED8F9209D5A6D86.

Guattery:1999:PRM

[GLM99] S. Guattery, T. Leighton, and G. L. Miller. The path resistance method for bounding the smallest nontrivial eigenvalue of a Laplacian. Combinatorics, Probability and Computing, 8 (5):441–460, September 1999. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http://journals.

cambridge.org/action/displayIssue?jid=CPC&volumeId=8&
issueId=05.

Gan:2015:MNI

[GLS15] Wenying Gan, Po-Shen Loh, and Benny Sudakov. Maximizing the number of independent sets of a fixed size. Combinatorics, Probability and Computing, 24(3):521–527, May 2015. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Gaither:2015:RWQ

[GLWW15] Jeffrey Gaither, Guy Louchard, Stephan Wagner, and Mark Daniel Ward. Resolution of T. Ward’s question and the Israel–Finch conjecture: Precise analysis of an integer sequence arising in dynamics. Combinatorics, Probability and Computing, 24(1):195– 215, January 2015. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). Special issue on Honouring the Memory of Philippe Flajolet - Part 3.

Gerke:2004:NER

[GM04] Stefanie Gerke and Colin McDiarmid. On the number of edges in random planar graphs. Combinatorics, Probability and Computing, 13(2):165–183, March 2004. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=13&issueId=02.

Goaoc:2015:SIE

[GMP^{+}15] Xavier Goaoc, Jiří Matousek, Pavel Paták, Zuzana Safernová, and Martin Tancer. Simplifying inclusion-exclusion formulas. Combinatorics, Probability and Computing, 24(2):438–456, March 2015. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Geil:2019:BNC

[GMP19] O. Geil and U. Martínez-Peñas. Bounding the number of common zeros of multivariate polynomials and their consecutive derivatives. Combinatorics, Probability and Computing, 28(2):253–279, March 2019. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL https://www.cambridge.org/core/
journals/combinatorics-probability-and-computing/article/
bounding-the-number-of-common-zeros-of-multivariate-
polynomials-and-their-consecutive-derivatives/678AC5C9D8B85601FDD844BA63126937.

Gimenez:2022:ACC

[GMPP22] Nardo Giménez, Guillermo Matera, Mariana Pérez, and Melina Privitelli. Average-case complexity of the Euclidean algorithm with a fixed polynomial over a finite field. Combinatorics, Probability and Computing, 31(1):166–183, January 2022. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/averagecase-
complexity-of-the-euclidean-algorithm-with-a-fixed-polynomial-
over-a-finite-field/85632A7CD68C5E0CAA1243BE27D8E5A5.

Gionfriddo:1994:BS

[GMT94] Mario Gionfriddo, Salvatore Milici, and Zsolt Tuza. Blocking sets in SQS(2v). Combinatorics, Probability and Computing, 3 (1):77–86, March 1994. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Grosz:2020:SFG

[GMT20] Dániel Grósz, Abhishek Methuku, and Casey Tompkins. On subgraphs of C_{2k}-free graphs and a problem of Kühn and Osthus. Combinatorics, Probability and Computing, 29(3):436–454, May 2020. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL https://www.cambridge.org/core/
journals/combinatorics-probability-and-computing/article/
on-subgraphs-of-c2kfree-graphs-and-a-problem-of-kuhn-
and-osthus/73CBED075F1B910C19966FC713D68EE3.

Gimenez:2006:CCT

[GN06] Omer Giménez and Marc Noy. On the complexity of computing the Tutte polynomial of bicircular matroids. Combinatorics, Probability and Computing, 15(3):385–395, May 2006. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=15&issueId=03.

Gnedin:2000:NSS

[Gne00] Alexander V. Gnedin. A note on sequential selection from permutations. Combinatorics, Probability and Computing, 9(1):13–17, January 2000. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL http://journals.cambridge.org/
action/displayIssue?jid=CPC&volumeId=9&issueId=01.

Gnedin:2004:TSF

[Gne04] Alexander V. Gnedin. Three sampling formulas. Combinatorics, Probability and Computing, 13(2):185–193, March 2004. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=13&issueId=02.

Glebov:2017:TPL

[GNS17] Roman Glebov, Humberto Naves, and Benny Sudakov. The threshold probability for long cycles. Combinatorics, Probability and Computing, 26(2):208–247, March 2017. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/product/
281BBEF1A6910F3B2716AE0FC7D0ABF0.

Godsil:1992:WGF

[God92] C. D. Godsil. Walk generating functions, Christoffel–Darboux identities and the adjacency matrix of a graph. Combinatorics, Probability and Computing, 1(1):13–25, March 1992. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Goerdt:2010:RBC

[Goe10] Andreas Goerdt. On random betweenness constraints. Combinatorics, Probability and Computing, 19(5–6):775–790, November 2010. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Gouezel:2015:NLB

[Gou15] S. Gouezel. A numerical lower bound for the spectral radius of random walks on surface groups. Combinatorics, Probability and Computing, 24(6):838–856, November 2015. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Gowers:1996:AWI

[Gow96] W. T. Gowers. An almost m-wise independent random permutation of the cube. Combinatorics, Probability and Computing, 5(2):119–130, June 1996. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Gowers:2006:QCR

[Gow06] W. T. Gowers. Quasirandomness, counting and regularity for 3-uniform hypergraphs. Combinatorics, Probability and Computing, 15(1–2):143–184, January 2006. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=15&issueId=01.

Gowers:2008:QG

[Gow08] W. T. Gowers. Quasirandom groups. Combinatorics, Probability and Computing, 17(3):363–387, May 2008. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=17&issueId=03.

Grabner:1997:MSR

[GP97] Peter J. Grabner and Helmut Prodinger. Maximum statistics of N random variables distributed by the negative binomial distribution. Combinatorics, Probability and Computing, 6(2):179–183, June 1997. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http://journals.

cambridge.org/action/displayIssue?jid=CPC&volumeId=6&
issueId=02.

Gnedin:2007:PRE

[GP07] Alexander Gnedin and Jim Pitman. Poisson representation of a Ewens fragmentation process. Combinatorics, Probability and Computing, 16(6):819–827, November 2007. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=16&issueId=06.

Goldschmidt:2019:PRT

[GP19] Christina Goldschmidt and Michal Przykucki. Parking on a random tree. Combinatorics, Probability and Computing, 28(1):23– 45, January 2019. CODEN CPCOFG. ISSN 0963-5483 (print),

1469-2163 (electronic). URL https://www.cambridge.org/
core/journals/combinatorics-probability-and-computing/
article/parking-on-a-random-tree/176352B7E36AB7862C55051FEDA273A6.

Georgakopoulos:2023:EGR

[GP23] Agelos Georgakopoulos and Christoforos Panagiotis. On the exponential growth rates of lattice animals and interfaces. Combinatorics, Probability and Computing, 32(6):912–955, November 2023. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL https://www.cambridge.org/core/
journals/combinatorics-probability-and-computing/article/
on-the-exponential-growth-rates-of-lattice-animals-and-
interfaces/96D87E5F67E90E4E781BDE8108513657.

Gunderson:1998:EPA

[GR98] David S. Gunderson and Vojtéch Rödl. Extremal problems for affine cubes of integers. Combinatorics, Probability and Computing, 7(1):65–79, March 1998. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http://journals.

cambridge.org/action/displayIssue?jid=CPC&volumeId=7&
issueId=01.

Grable:1998:LDI

[Gra98] D. A. Grable. A large deviation inequality for functions of independent, multi-way choices. Combinatorics, Probability and Computing, 7(1):57–63, March 1998. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=7&issueId=01.

Graham:2015:BDC

[Gra15] Benjamin Graham. A binary deletion channel with a fixed number of deletions. Combinatorics, Probability and Computing, 24 (3):486–489, May 2015. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Greenhalgh:1997:MRR

[Gre97] Andrew S. Greenhalgh. A model for random random-walks on finite groups. Combinatorics, Probability and Computing, 6(1):49–56, March 1997. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http://journals.

cambridge.org/action/displayIssue?jid=CPC&volumeId=6&
issueId=01.

Green:2003:SCI

[Gre03] Ben Green. Some constructions in the inverse spectral theory of cyclic groups. Combinatorics, Probability and Computing, 12(2):127–138, March 2003. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=12&issueId=02.

Green:2012:CMF

[Gre12] Ben Green. On (not) computing the Möbius function using bounded depth circuits. Combinatorics, Probability and Computing, 21(6):942–951, November 2012. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Green:2017:CNR

[Gre17] Ben Green. On the chromatic number of random Cayley graphs. Combinatorics, Probability and Computing, 26(2):248– 266, March 2017. CODEN CPCOFG. ISSN 0963-5483 (print),

1469-2163 (electronic). URL https://www.cambridge.org/
core/product/72DEF972689BE6AC69BAC24775610BAB.

Grinstead:1997:MLD

[Gri97] Charles M. Grinstead. On medians of lattice distributions and a game with two dice. Combinatorics, Probability and Computing, 6(3):273–294, September 1997. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http://journals.

cambridge.org/action/displayIssue?jid=CPC&volumeId=6&
issueId=03.

Grolmusz:2003:NER

[Gro03] Vince Grolmusz. A note on explicit Ramsey graphs and modular sieves. Combinatorics, Probability and Computing, 12(5–6): 565–569, November 2003. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). Special issue on Ramsey theory.

Greenberg:2020:SBM

[GRS20] Sam Greenberg, Dana Randall, and Amanda Pascoe Streib. Sampling biased monotonic surfaces using exponential metrics. Combinatorics, Probability and Computing, 29(5):672–697, September 2020. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL https://www.cambridge.org/core/
journals/combinatorics-probability-and-computing/article/
sampling-biased-monotonic-surfaces-using-exponential-
metrics/3D29F7E2E25ACDBCCC43F7FE7EAFD096.

Goyal:2015:QCS

[GRV15] Navin Goyal, Luis Rademacher, and Santosh Vempala. Query complexity of sampling and small geometric partitions. Combinatorics, Probability and Computing, 24(5):733–753, September 2015. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Greenhill:2003:CDB

[GRW03] Catherine Greenhill, Andrzej Ruciński, and Nicholas C. Wormald. Connectedness of the degree bounded star process. Combinatorics, Probability and Computing, 12(3):269–283, May 2003. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=12&issueId=03. Combina-

torics, probability and computing (Oberwolfach, 2001).

Gao:2021:FRM

[GRWW21] Pu Gao, Reshma Ramadurai, Ian M. Wanless, and Nick Wormald. Full rainbow matchings in graphs and hypergraphs. Combinatorics, Probability and Computing, 30(5):762–780, September 2021. CODEN CPCOFG. ISSN 0963-5483 (print),

1469-2163 (electronic). URL https://www.cambridge.org/
core/journals/combinatorics-probability-and-computing/
article/full-rainbow-matchings-in-graphs-and-hypergraphs/
1BC41B2F1E233C134DF7F5DDFF5CFA2B.

Goh:1993:RMB

[GS93] William M. Y. Goh and Eric Schmutz. Random matrices and Brownian motion. Combinatorics, Probability and Computing, 2(2):157–180, June 1993. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Gyarfas:2011:CRN

[GS11] András Gyárfás and Gábor N. Sárközy. The 3-colour Ramsey number of a 3-uniform Berge cycle. Combinatorics, Probability and Computing, 20(1):53–71, January 2011. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Ge:2012:GPI

[GS12a] Q. Ge and D. Stefankovic. A graph polynomial for independent sets of bipartite graphs. Combinatorics, Probability and Computing, 21(5):695–714, September 2012. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Gyarfas:2012:SVT

[GS12b] András Gyárfás and Gábor N. Sárközy. Star versus two stripes Ramsey numbers and a conjecture of Schelp. Combinatorics, Probability and Computing, 21(1–2):179–186, ???? 2012. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Gishboliner:2023:MCS

[GS23] Lior Gishboliner and Benny Sudakov. Maximal chordal subgraphs. Combinatorics, Probability and Computing, 32(5):724– 741, September 2023. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL https://www.cambridge.

org/core/journals/combinatorics-probability-and-computing/
article/maximal-chordal-subgraphs/494F018278BEBED90228478D9813680D.

Godbole:1996:CAU

[GSS96] Anant P. Godbole, Daphne E. Skipper, and Rachel A. Sunley. t-covering arrays: upper bounds and Poisson approximations. Combinatorics, Probability and Computing, 5(2):105–117, June 1996. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Glebov:2014:CFC

[GST14] Roman Glebov, Tibor Szabó, and Gábor Tardos. Conflict-free colouring of graphs. Combinatorics, Probability and Computing, 23(3):434–448, May 2014. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Galanis:2016:IPF

[GSV16] Andreas Galanis, Daniel Stefankovic, and Eric Vigoda. Inapproximability of the partition function for the antiferromagnetic Ising and hard-core models. Combinatorics, Probability and Computing, 25(4):500–559, July 2016. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Gerke:2011:PGY

[GSW11] Stefanie Gerke, Angelika Steger, and Nicholas Wormald. Pegging graphs yields a small diameter. Combinatorics, Probability and Computing, 20(2):239–248, March 2011. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Green:2009:FTF

[GT09] Ben Green and Terence Tao. Freiman’s theorem in finite fields via extremal set theory. Combinatorics, Probability and Computing, 18(3):335–355, May 2009. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Gubser:1996:CAP

[Gub96] Bradley S. Gubser. A characterization of almost-planar graphs. Combinatorics, Probability and Computing, 5(3):227– 245, September 1996. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Georgakopoulos:2014:NBE

[GW14] Agelos Georgakopoulos and Peter Winkler. New bounds for edgecover by random walk. Combinatorics, Probability and Computing, 23(4):571–584, July 2014. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Gao:2015:OTR

[GW15] Pu Gao and Nicholas Wormald. Orientability thresholds for random hypergraphs. Combinatorics, Probability and Computing, 24(5):774–824, September 2015. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Georgakopoulos:2018:SGC

[GW18] Agelos Georgakopoulos and Stephan Wagner. Subcritical graph classes containing all planar graphs. Combinatorics, Probability and Computing, 27(5):763–773, September 2018. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/subcritical-
graph-classes-containing-all-planar-graphs/7A03A98403E1CA61823753232426ECFC.

Green:2021:EPG

[GW21] Ben Green and Aled Walker. Extremal problems for GCDs. Combinatorics, Probability and Computing, 30(6):922–929, November 2021. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL https://www.cambridge.org/core/
journals/combinatorics-probability-and-computing/article/
extremal-problems-for-gcds/D3BE6C493BF68649458DDED6D20AEB55.

Gyori:2006:TFH

[Gyo06] Ervin Gyori. Triangle-free hypergraphs. Combinatorics, Probability and Computing, 15(1–2):185–191, January 2006. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic).  URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=15&issueId=01.

Hadjicostas:1998:APS

[Had98] Petros Hadjicostas. The asymptotic proportion of subdivisions of a 2 × 2 table that result in Simpson’s Paradox. Combinatorics, Probability and Computing, 7(4):387–396, December 1998. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=7&issueId=04. See [Had99].

Hadjicostas:1999:CAP

[Had99] Petros Hadjicostas. Corrigendum: “The asymptotic proportion of subdivisions of a 2 × 2 table that result in Simpson’s paradox” [Combin. Probab. Comput. 7 (1998), no. 4, 387– 396; MR1680088 (99m:05008)]. Combinatorics, Probability and Computing, 8(6):599, November 1999. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=8&issueId=06. See [Had98].

Haggkvist:1992:SNH

[Häg92] Roland Häggkvist. On the structure of non-Hamiltonian graphs. I. Combinatorics, Probability and Computing, 1(1):27–34, March 1992. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Haggkvist:1993:HCO

[Häg93] Roland Häggkvist. Hamilton cycles in oriented graphs. Combinatorics, Probability and Computing, 2(1):25–32, March 1993. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Haggkvist:1996:REC

[Häg96] Roland Häggkvist. Restricted edge-colourings of bipartite graphs. Combinatorics, Probability and Computing, 5(4):385– 392, December 1996. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Haggstrom:1998:CBB

[Häg98] Olle Häggström. On a conjecture of Bollobás and Brightwell concerning random walks on product graphs. Combinatorics, Probability and Computing, 7(4):397–401, December 1998. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=7&issueId=04.

Halin:1993:MPI

[Hal93] R. Halin. Minimization problems for infinite n-connected graphs. Combinatorics, Probability and Computing, 2(4):417–436, December 1993. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Hamidoune:1995:WSS

[Ham95] Yahya Ould Hamidoune. On weighted sequence sums. Combinatorics, Probability and Computing, 4(4):363–367, December 1995. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Hamidoune:1998:ADC

[Ham98] Y. O. Hamidoune. Adding distinct congruence classes. Combinatorics, Probability and Computing, 7(1):81–87, March 1998. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=7&issueId=01. See com-

ments [Baj16].

Hamidoune:2003:SS

[Ham03] Yahya Ould Hamidoune. Subsequence sums. Combinatorics, Probability and Computing, 12(4):413–425, July 2003. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic).  URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=12&issueId=04.

Hamidoune:2008:WGG

[Ham08a] Yahya O. Hamidoune. A weighted generalization of Gao’s n + D−1 theorem. Combinatorics, Probability and Computing, 17(6): 793–798, November 2008. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Hamidoune:2008:IIS

[Ham08b] Yahya Ould Hamidoune. On iterated image size for pointsymmetric relations. Combinatorics, Probability and Computing, 17(1):61–66, January 2008. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=17&issueId=01.

Hamidoune:2011:CCP

[Ham11] Y. O. Hamidoune. Counting certain pairings in arbitrary groups. Combinatorics, Probability and Computing, 20(6):855– 865, November 2011. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Hansen:1993:FBB

[Han93] Jennie C. Hansen. Factorization in F_{q}[x] and Brownian motion. Combinatorics, Probability and Computing, 2(3):285–299, September 1993. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Hansen:1997:LLO

[Han97] Jennie C. Hansen. Limit laws for the optimal directed tree with random costs. Combinatorics, Probability and Computing, 6 (3):315–335, September 1997. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http://journals.

cambridge.org/action/displayIssue?jid=CPC&volumeId=6&
issueId=03.

Han:2015:NPM

[Han15] Jie Han. Near perfect matchings in k-uniform hypergraphs. Combinatorics, Probability and Computing, 24(5):723–732, September 2015. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Hatami:2009:DTI

[Hat09] Hamed Hatami. Decision trees and influences of variables over product probability spaces. Combinatorics, Probability and Computing, 18(3):357–369, May 2009. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Haxell:2001:NVL

[Hax01] P. E. Haxell. A note on vertex list colouring. Combinatorics, Probability and Computing, 10(4):345–347, July 2001. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=10&issueId=04.

Haxell:2004:SCN

[Hax04] P. E. Haxell. On the strong chromatic number. Combinatorics, Probability and Computing, 13(6):857–865, November 2004. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=13&issueId=06.

Heckel:2020:SCE

[Hec20] Annika Heckel. Sharp concentration of the equitable chromatic number of dense random graphs. Combinatorics, Probability and Computing, 29(2):213–233, March 2020. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/sharp-
concentration-of-the-equitable-chromatic-number-of-dense-
random-graphs/5C01926669E65A06E90B76D68C2393A6.

Hegyvari:1996:SSB

[Heg96] Norbert Hegyvári. Subset sums in N^{2}. Combinatorics, Probability and Computing, 5(4):393–402, December 1996. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Hegyvari:2005:IPP

[Heg05] Norbert Hegyvári. On intersecting properties of partitions of integers. Combinatorics, Probability and Computing, 14(3):319– 323, May 2005. CODEN CPCOFG. ISSN 0963-5483 (print),

1469-2163 (electronic). URL http://journals.cambridge.
org/action/displayIssue?jid=CPC&volumeId=14&issueId=
03.

Hatami:2012:NTC

[HHK^{+}12] Hamed Hatami, Jan Hladký, Daniel Král’, Serguei Norine, and Alexander Razborov. Non-three-colourable common graphs exist. Combinatorics, Probability and Computing, 21(5):734–742, September 2012. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Hegyvari:2007:AQB

[HHP07] Norbert Hegyvári, François Hennecart, and Alain Plagne. Answer to a question by Burr and Erdős on restricted addition, and related results. Combinatorics, Probability and Computing, 16(5):747–756, September 2007. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=16&issueId=05.

Harutyunyan:2021:IDS

[HHV21] Ararat Harutyunyan, Paul Horn, and Jacques Verstraete. Independent dominating sets in graphs of girth five. Combinatorics, Probability and Computing, 30(3):344–359, May 2021. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/independent-
dominating-sets-in-graphs-of-girth-five/6FC6BCA3C2FFAB26ABC8BBB40E1EFE2B.

Han:2023:CFG

[HHWY23] Jie Han, Ping Hu, Guanghui Wang, and Donglei Yang. Cliquefactors in graphs with sublinear ℓ-independence number. Combinatorics, Probability and Computing, 32(4):665–681, July 2023. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/cliquefactors-
in-graphs-with-sublinear-boldsymbolell-independence-
number/75D1FCB8E81339202B9EB21F44077966.

Hirth:1997:GBD

[Hir97a] Ulrich Martin Hirth. From GEM back to Dirichlet via Hoppe’s Urn. Combinatorics, Probability and Computing, 6(2):185–195, June 1997. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL http://journals.cambridge.org/
action/displayIssue?jid=CPC&volumeId=6&issueId=02.

Hirth:1997:PNT

[Hir97b] Ulrich Martin Hirth. Probabilistic number theory, the GEM/ Poisson–Dirichlet distribution and the Arc-sine Law. Combinatorics, Probability and Computing, 6(1):57–77, March 1997. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=6&issueId=01.

Haggkvist:1994:FGE

[HJ94] Roland Häggkvist and Anders Johansson. (1, 2)-factorizations of general Eulerian nearly regular graphs. Combinatorics, Probability and Computing, 3(1):87–95, March 1994. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Haggkvist:1997:NBL

[HJ97] Roland Häggkvist and Jeannette Janssen. New bounds on the list-chromatic index of the complete graph and other simple graphs. Combinatorics, Probability and Computing, 6(3): 295–313, September 1997. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http://journals.

cambridge.org/action/displayIssue?jid=CPC&volumeId=6&
issueId=03.

Haggkvist:2008:OLR

[HJ08] Roland Häggkvist and Anders Johansson. Orthogonal latin rectangles. Combinatorics, Probability and Computing, 17(4):519– 536, July 2008. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Hitczenko:2014:WRS

[HJ14] Pawel Hitczenko and Svante Janson. Weighted random staircase tableaux. Combinatorics, Probability and Computing, 23 (6):1114–1147, November 2014. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). Special issue on Honouring the Memory of Philippe Flajolet - Part 2.

Huynh:2022:SDS

[HJW22] Tony Huynh, Gwenaël Joret, and David R. Wood. Subgraph densities in a surface. Combinatorics, Probability and Computing, 31(5):812–839, September 2022. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL https://www.

cambridge.org/core/journals/combinatorics-probability-
and-computing/article/subgraph-densities-in-a-surface/
D2A8D2326DB06835D536A303DAE30338.

Hammer:1994:UTG

[HK94] P. L. Hammer and A. K. Kelmans. On universal threshold graphs. Combinatorics, Probability and Computing, 3(3):327– 344, September 1994. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Heuberger:2014:ABA

[HK14] Clemens Heuberger and Sara Kropf. Analysis of the binary asymmetric joint sparse form. Combinatorics, Probability and Computing, 23(6):1087–1113, November 2014. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). Special issue on Honouring the Memory of Philippe Flajolet - Part 2.

Hamm:2019:EKRb

[HK19a] A. Hamm and J. Kahn. On Erdős–Ko–Rado for random hypergraphs I. Combinatorics, Probability and Computing, 28 (6):881–916, November 2019. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL /core/journals/

combinatorics-probability-and-computing/article/on-erdoskorado-
for-random-hypergraphs-i/6469A6C52B5948B2DFB8B991DD05EE45.

Hamm:2019:EKRa

[HK19b] A. Hamm and J. Kahn. On Erdős–Ko–Rado for random hypergraphs II. Combinatorics, Probability and Computing, 28(1):61– 80, January 2019. CODEN CPCOFG. ISSN 0963-5483 (print),

1469-2163 (electronic). URL https://www.cambridge.org/
core/journals/combinatorics-probability-and-computing/
article/on-erdoskorado-for-random-hypergraphs-ii/6272DC7EDD61A2885455E287AE6CE59C.

Haxell:1995:ISR

[HKŁ95] P. E. Haxell, Y. Kohayakawa, and T. Łuczak. The induced size-Ramsey number of cycles. Combinatorics, Probability and Computing, 4(3):217–239, September 1995. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Hoppen:2012:ECG

[HKL12] Carlos Hoppen, Yoshiharu Kohayakawa, and Hanno Lefmann. Edge colourings of graphs avoiding monochromatic matchings of a given size. Combinatorics, Probability and Computing, 21(1–2): 203–218, ???? 2012. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Hoppen:2020:EPA

[HKL^{+}20] Carlos Hoppen, Yoshiharu Kohayakawa, Richard Lang, Hanno Lefmann, and Henrique Stagni. Estimating parameters associated with monotone properties. Combinatorics, Probability and Computing, 29(4):616–632, July 2020. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/estimating-
parameters-associated-with-monotone-properties/F01EC7C7E509F4F1F0A7FFD395F4DF3F.

Hutchcroft:2023:BCH

[HKNN23] Tom Hutchcroft, Alexander Kent, and Petar Nizić-Nikolac. The bunkbed conjecture holds in the p ↑ 1 limit. Combinatorics, Probability and Computing, 32(3):363–369, May 2023. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/bunkbed-
conjecture-holds-in-the-puparrow-1-limit/DE7E423495703BCB28EFEEB981896B3B.

Han:2021:NPC

[HKP21] Jie Han, Yoshiharu Kohayakawa, and Yury Person. Nearperfect clique-factors in sparse pseudorandom graphs. Combinatorics, Probability and Computing, 30(4):570–590, July 2021. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/nearperfect-
cliquefactors-in-sparse-pseudorandom-graphs/D47868F653ED1CB993807B8A49507FE4.

Hindman:1993:IPR

[HL93] Neil Hindman and Imre Leader. Image partition regularity of matrices. Combinatorics, Probability and Computing, 2(4):437–463, December 1993. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Hutson:2007:ELM

[HL07] Kevin Hutson and Thomas M. Lewis. The expected length of a minimal spanning tree of a cylinder graph. Combinatorics, Probability and Computing, 16(1):63–83, January 2007. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=16&issueId=01.

Hlineny:2006:TPM

[Hli06] Petr Hlinený. The Tutte polynomial for matroids of bounded branch-width. Combinatorics, Probability and Computing, 15(3): 397–409, May 2006. CODEN CPCOFG. ISSN 0963-5483 (print),

1469-2163 (electronic). URL http://journals.cambridge.
org/action/displayIssue?jid=CPC&volumeId=15&issueId=

Holzman:2008:PDS

[HLP08] Ron Holzman, Vsevolod F. Lev, and Rom Pinchasi. Projecting difference sets on the positive orthant. Combinatorics, Probability and Computing, 17(5):681–688, September 2008. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Haxell:2009:RNU

[HLP^{+}09] P. E. Haxell, T. Luczak, Y. Peng, V. Rödl, A. Ruciński, and J. Skokan. The Ramsey number for 3-uniform tight hypergraph cycles. Combinatorics, Probability and Computing, 18(1–2):165– 203, March 2009. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Hamidoune:1999:SSS

[HLS99] Y. O. Hamidoune, A. S. Lladó, and O. Serra. On sets with a small subset sum. Combinatorics, Probability and Computing, 8 (5):461–466, September 1999. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http://journals.

cambridge.org/action/displayIssue?jid=CPC&volumeId=8&
issueId=05.

Hamidoune:2000:RS

[HLS00] Y. O. Hamidoune, A. S. Lladó, and O. Serra. On restricted sums. Combinatorics, Probability and Computing, 9 (6):513–518, November 2000. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http://journals.

cambridge.org/action/displayIssue?jid=CPC&volumeId=9&
issueId=06.

Hindman:2003:OPP

[HLS03] Neil Hindman, Imre Leader, and Dona Strauss. Open problems in partition regularity. Combinatorics, Probability and Computing, 12(5–6):571–583, November 2003. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). Special issue on Ramsey theory.

Han:2021:CTH

[HLSM21] Jie Han, Allan Lo, and Nicolás Sanhueza-Matamala. Covering and tiling hypergraphs with tight cycles. Combinatorics, Probability and Computing, 30(2):288–329, March 2021. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/

combinatorics-probability-and-computing/article/covering-
and-tiling-hypergraphs-with-tight-cycles/66FC8D739C465D1EF2CA472C1A644284.

Han:2017:EMC

[HLTZ17] Jie Han, Allan Lo, Andrew Treglown, and Yi Zhao. Exact minimum codegree threshold for K − 4-factors. Combinatorics, Probability and Computing, 26(6):856–885, November 2017. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL https://www.cambridge.org/core/
product/B9ECF83CA9AF217848B7579842445CD5.

Holroyd:2019:ALG

[HLW19] Alexander E. Holroyd, Lionel Levine, and Peter Winkler. Abelian logic gates. Combinatorics, Probability and Computing, 28(3): 388–422, May 2019. CODEN CPCOFG. ISSN 0963-5483 (print),

1469-2163 (electronic). URL https://www.cambridge.org/
core/journals/combinatorics-probability-and-computing/
article/abelian-logic-gates/D427801293A8606B14D447136774D5B5.

Hammersley:1994:PLE

[HM94] J. M. Hammersley and G. Mazzarino. Properties of large Eden clusters in the plane. Combinatorics, Probability and Computing, 3(4):471–505, December 1994. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Huillet:2015:ASC

[HM15] Thierry Huillet and Martin Möhle. Asymptotics of symmetric compound Poisson population models. Combinatorics, Probability and Computing, 24(1):216–253, January 2015. CODEN CP-COFG. ISSN 0963-5483 (print), 1469-2163 (electronic). Special issue on Honouring the Memory of Philippe Flajolet - Part 3.

Hartarsky:2022:RUC

[HM22a] Ivailo Hartarsky and Laure Marêché. Refined universality for critical KCM: lower bounds. Combinatorics, Probability and Computing, 31(5):879–906, September 2022. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/refined-
universality-for-critical-kcm-lower-bounds/13671D8F48631729372A4F920DA830DC.

Heydenreich:2022:ECP

[HM22b] Markus Heydenreich and Kilian Matzke. Expansion for the critical point of site percolation: the first three terms. Combinatorics, Probability and Computing, 31(3):430–454, May 2022. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/expansion-
for-the-critical-point-of-site-percolation-the-first-
three-terms/4F13BE121491D96EFBF8E273B7023D01.

Hemmecke:2008:TCE

[HMS^{+}08] Raymond Hemmecke, Jason Morton, Anne Shiu, Bernd Sturmfels, and Oliver Wienand. Three counter-examples on semi-graphoids. Combinatorics, Probability and Computing, 17(2):239–257, March 2008. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=17&issueId=02.

Huang:2013:LFA

[HMS^{+}13] Hao Huang, Jie Ma, Asaf Shapira, Benny Sudakov, and Raphael Yuster. Large feedback arc sets, high minimum degree subgraphs, and long cycles in Eulerian Digraphs. Combinatorics, Probability and Computing, 22(6):859–873, November 2013. CODEN CP-COFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Hatami:2013:ERF

[HN13] Hamed Hatami and Serguei Norine. The entropy of randomfree graphons and properties. Combinatorics, Probability and Computing, 22(4):517–526, July 2013. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Hod:2014:CGR

[HN14] Rani Hod and Alon Naor. Component games on regular graphs. Combinatorics, Probability and Computing, 23(1):75–89, January 2014. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Haxell:2018:STM

[HN18] Penny Haxell and Lothar Narins. A stability theorem for matchings in tripartite 3-graphs. Combinatorics, Probability and Computing, 27(5):774–793, September 2018. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/stability-
theorem-for-matchings-in-tripartite-3graphs/30912180E71C820784B5422A06196E20.

Haythorpe:2020:CFC

[HN20] M. Haythorpe and A. Newcombe. Constructing families of cospectral regular graphs. Combinatorics, Probability and Computing, 29(5):664–671, September 2020. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL https://

www.cambridge.org/core/journals/combinatorics-probability-
and-computing/article/constructing-families-of-cospectral-
regular-graphs/C84AF47A6EDD90F5566BB58CDCFA07EB.

Hwang:2019:P

[HNZ19] Hsien-Kuei Hwang, Ralph Neininger, and Marek Zaionc. Preface. Combinatorics, Probability and Computing, 28(4):483–484, July 2019. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL https://www.cambridge.org/core/
journals/combinatorics-probability-and-computing/article/
preface/F88C99871D2FB77579C932B48716CDF9.

Holst:2004:BRL

[Hol04] Lars Holst. Book review: “Logarithmic Combinatorial Structures: A Probabilistic Approach”, by Richard Arratia, A. D. Barbour and Simon Tavaré, European Mathematical Society Monographs in Mathematics, 2003, 375 pp., EU 69. Combinatorics, Probability and Computing, 13(6):916–917, November 2004. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=13&issueId=06.

Holmgren:2010:RRC

[Hol10] Cecilia Holmgren. Random records and cuttings in binary search trees. Combinatorics, Probability and Computing, 19(3):391–424, May 2010. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Hamidoune:1998:CTE

[HOO98] Yahya Ould Hamidoune, Oscar Ordaz, and Asdrubal Ortuño. On a combinatorial theorem of Erdős, Ginzburg and Ziv. Combinatorics, Probability and Computing, 7(4):403–412, December 1998. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=7&issueId=04.

Howard:1996:OMI

[How96] C. Douglas Howard. Orthogonality of measures induced by random walks with scenery. Combinatorics, Probability and Computing, 5(3):247–256, September 1996. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Hueter:1997:SAC

[HP97] Irene Hueter and Yuval Peres. Self-affine carpets on the square lattice. Combinatorics, Probability and Computing, 6(2):197–204, June 1997. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL http://journals.cambridge.org/
action/displayIssue?jid=CPC&volumeId=6&issueId=02.

Hundack:1993:EGP

[HPS93] Christoph Hundack, Hans Jürgen Prömel, and Angelika Steger. Extremal graph problems for graphs with a color-critical vertex. Combinatorics, Probability and Computing, 2(4):465–477, December 1993. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Harant:1999:DSI

[HPV99] Jochen Harant, Anja Pruchnewski, and Margit Voigt. On dominating sets and independent sets of graphs. Combinatorics, Probability and Computing, 8(6):547–553, November 1999. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=8&issueId=06.

Hamidoune:2005:SWP

[HQ05] Y. O. Hamidoune and D. Quiroz. On subsequence weighted products. Combinatorics, Probability and Computing, 14(4):485–489, July 2005. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL http://journals.cambridge.org/
action/displayIssue?jid=CPC&volumeId=14&issueId=04.

Hurst:1995:SSC

[HR95] Fair Barbour Hurst and Talmage James Reid. Some small circuit–cocircuit Ramsey numbers for matroids. Combinatorics, Probability and Computing, 4(1):67–80, March 1995. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Hamidoune:2011:LBS

[HR11] Y. O. Hamidoune and J. Rué. A lower bound for the size of a Minkowski sum of dilates. Combinatorics, Probability and Computing, 20(2):249–256, March 2011. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Han:2021:RTN

[HRRS21] Hiêp Hàn, Troy Retter, Vojtêch Rödl, and Mathias Schacht. Ramsey-type numbers involving graphs and hypergraphs with large girth. Combinatorics, Probability and Computing, 30(5): 722–740, September 2021. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL https://www.cambridge.

org/core/journals/combinatorics-probability-and-computing/
article/ramseytype-numbers-involving-graphs-and-hypergraphs-
with-large-girth/3886DD369AED32470F9F71B411F8A22D.

Hara:1995:SAW

[HS95] Takashi Hara and Gordon Slade. The self-avoiding-walk and percolation critical points in high dimensions. Combinatorics, Probability and Computing, 4(3):197–215, September 1995. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Hindman:1998:APD

[HS98] Neil Hindman and Dona Strauss. An algebraic proof of Deuber’s Theorem. Combinatorics, Probability and Computing, 7(2):167–180, June 1998. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http://journals.

cambridge.org/action/displayIssue?jid=CPC&volumeId=7&
issueId=02.

Haggstrom:2000:PWA

[HS00a] Olle Häggström and Jeffrey E. Steif. Propp–Wilson algorithms and finitary codings for high noise Markov random fields. Combinatorics, Probability and Computing, 9(5):425–439, September 2000. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=9&issueId=05.

Hitczenko:2000:END

[HS00b] P. Hitczenko and G. Stengle. Expected number of distinct part sizes in a random integer composition. Combinatorics, Probability and Computing, 9(6):519–527, November 2000. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=9&issueId=06.

Hush:2004:FSA

[HS04] Don Hush and Clint Scovel. Fat-shattering of affine functions. Combinatorics, Probability and Computing, 13(3):353–360, May 2004. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=13&issueId=03.

Haxell:2006:OIT

[HS06] Penny Haxell and Tibor Szabó. Odd independent transversals are odd. Combinatorics, Probability and Computing, 15(1–2):193– 211, January 2006. CODEN CPCOFG. ISSN 0963-5483 (print),

1469-2163 (electronic). URL http://journals.cambridge.
org/action/displayIssue?jid=CPC&volumeId=15&issueId=
01.

Huang:2012:SHM

[HS12] Po-Shen Loh Hao Huang and Benny Sudakov. The size of a hypergraph and its matching number. Combinatorics, Probability and Computing, 21(3):442–450, May 2012. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Holmes:2016:FCD

[HS16] Mark Holmes and Thomas S. Salisbury. Forward clusters for degenerate random environments. Combinatorics, Probability and Computing, 25(5):744–765, September 2016. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/product/
6119F454D9638160F641AC120919A375.

Hujter:1996:PEI

[HT96] M. Hujter and Zs. Tuza. Precoloring extension. III. Classes of perfect graphs. Combinatorics, Probability and Computing, 5(1): 35–56, March 1996. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Houdre:2001:CMP

[HT01] C. Houdré and P. Tetali. Concentration of measure for products of Markov kernels and graph products via functional inequalities. Combinatorics, Probability and Computing, 10(1):1–28, January 2001. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=10&issueId=01.

Hwang:2002:QDF

[HT02] Hsien-Kuei Hwang and Tsung-Hsi Tsai. Quickselect and the Dickman function. Combinatorics, Probability and Computing, 11(4):353–371, July 2002. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:/

/journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=11&issueId=04.

Hefetz:2016:UGF

[HT16] Dan Hefetz and Mykhaylo Tyomkyn. Universality of graphs with few triangles and anti-triangles. Combinatorics, Probability and Computing, 25(4):560–576, July 2016. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Huber:2016:NOB

[Hub16] Mark Huber. Nearly optimal Bernoulli factories for linear functions. Combinatorics, Probability and Computing, 25(4):577–591, July 2016. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Hooghiemstra:2008:WHS

[HV08] Gerard Hooghiemstra and Piet Van Mieghem. The weight and hopcount of the shortest path in the complete graph with exponential weights. Combinatorics, Probability and Computing, 17(4):537–548, July 2008. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Hoppen:2008:IFR

[HW08] Carlos Hoppen and Nicholas Wormald. Induced forests in regular graphs with large girth. Combinatorics, Probability and Computing, 17(3):389–410, May 2008. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=17&issueId=03.

Hendrey:2019:DCC

[HW19] Kevin Hendrey and David R. Wood. Defective and clustered choosability of sparse graphs. Combinatorics, Probability and Computing, 28(5):791–810, September 2019. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL /core/journals/combinatorics-probability-
and-computing/article/defective-and-clustered-choosability-
of-sparse-graphs/97CE80DA437D31AFDBCBFBEADB59E4B0.

Hwang:1998:PGC

[Hwa98] Hsien-Kuei Hwang. A Poisson* geometric convolution law for the number of components in unlabelled combinatorial structures. Combinatorics, Probability and Computing, 7(1):89–110, March 1998. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=7&issueId=01.

Hyde:2023:TSK

[Hyd23] Joseph Hyde. Towards the 0-statement of the Kohayakawa– Kreuter conjecture. Combinatorics, Probability and Computing, 32(2):225–268, March 2023. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL https://www.cambridge.

org/core/journals/combinatorics-probability-and-computing/
article/towards-the-0statement-of-the-kohayakawakreuter-
conjecture/5F955F23EEC4843C2660B05918EE5FFF.

Hammer:2005:CMS

[HZ05] Peter L. Hammer and Igor E. Zverovich. Construction of a maximum stable set with k-extensions. Combinatorics, Probability and Computing, 14(3):311–318, May 2005. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=14&issueId=03.

Hegarty:2014:DCG

[HZ14] Peter Hegarty and Dmitrii Zhelezov. On the diameters of commuting graphs arising from random skew-symmetric matrices. Combinatorics, Probability and Computing, 23(3):449–459, May 2014. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Hou:2017:JPH

[HZ17] Jianfeng Hou and Qinghou Zeng. Judicious partitioning of hypergraphs with edges of size at most 2. Combinatorics, Probability and Computing, 26(2):267–284, March 2017. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/product/
6217A8D829FE7112F910002DAC4199BB.

Ilinca:2009:NSF

[IK09] L. Ilinca and J. Kahn. On the number of 2-SAT functions. Combinatorics, Probability and Computing, 18(5):749–764, September 2009. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Illingworth:2022:CPL

[Ill22] Freddie Illingworth. The chromatic profile of locally colourable graphs. Combinatorics, Probability and Computing, 31(6):976– 1009, November 2022. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL https://www.cambridge.

org/core/journals/combinatorics-probability-and-computing/
article/chromatic-profile-of-locally-colourable-graphs/
8CD12196550420E5653F23CB527E24A1.

Ingrassia:1993:GAE

[Ing93] Salvatore Ingrassia. Geometric approaches to the estimation of the spectral gap of reversible Markov chains. Combinatorics, Probability and Computing, 2(3):301–323, September 1993. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Isaev:2022:DTP

[ISZ22] Mikhail Isaev, Angus Southwell, and Maksim Zhukovskii. Distribution of tree parameters by martingale approach. Combinatorics, Probability and Computing, 31(5):737–764, September 2022. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL https://www.cambridge.org/core/
journals/combinatorics-probability-and-computing/article/
distribution-of-tree-parameters-by-martingale-approach/
32DFE2BAD7DFC3309F3FAD131FEFC451.

Inkmann:2011:MMP

[IT11] Torsten Inkmann and Robin Thomas. Minor-minimal planar graphs of even branch-width. Combinatorics, Probability and Computing, 20(1):73–82, January 2011. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Jackson:1993:ZFI

[Jac93] Bill Jackson. A zero-free interval for chromatic polynomials of graphs. Combinatorics, Probability and Computing, 2(3):325– 336, September 1993. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Jackson:2007:ZFI

[Jac07] Bill Jackson. Zero-free intervals for flow polynomials of near-cubic graphs. Combinatorics, Probability and Computing, 16(1):85–108, January 2007. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=16&issueId=01.

Jagger:1998:DVP

[Jag98] Chris Jagger. Distant vertex partitions of graphs. Combinatorics, Probability and Computing, 7(4):413–422, December 1998. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=7&issueId=04.

Janson:1994:NST

[Jan94] Svante Janson. The numbers of spanning trees, Hamilton cycles and perfect matchings in a random graph. Combinatorics, Probability and Computing, 3(1):97–126, March 1994. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Janson:1995:RRG

[Jan95] Svante Janson. Random regular graphs: asymptotic distributions and contiguity. Combinatorics, Probability and Computing, 4(4): 369–405, December 1995. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Janson:1999:OTT

[Jan99] Svante Janson. One, two and three times log n/n for paths in a complete graph with random weights. Combinatorics, Probability and Computing, 8(4):347–361, July 1999. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&

volumeId=8&issueId=04. Random graphs and combinatorial structures (Oberwolfach, 1997).

Janson:2003:CUC

[Jan03] Svante Janson. Cycles and unicyclic components in random graphs. Combinatorics, Probability and Computing, 12(1):27–52, January 2003. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL http://journals.cambridge.org/
action/displayIssue?jid=CPC&volumeId=12&issueId=01.

Janson:2005:FER

[Jan05] Svante Janson. The first eigenvalue of random graphs. Combinatorics, Probability and Computing, 14(5–6):815–828, November 2005. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=14&issueId=05.

Janson:2007:RGR

[Jan07] Svante Janson. On a random graph related to quantum theory. Combinatorics, Probability and Computing, 16(5):757–766, September 2007. CODEN CPCOFG. ISSN 0963-5483 (print),

1469-2163 (electronic). URL http://journals.cambridge.
org/action/displayIssue?jid=CPC&volumeId=16&issueId=
05.

Janson:2008:IDH

[Jan08] Svante Janson. Individual displacements in hashing with coalesced chains. Combinatorics, Probability and Computing, 17(6): 799–814, November 2008. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Janson:2009:PRM

[Jan09] Svante Janson. The probability that a random multigraph is simple. Combinatorics, Probability and Computing, 18(1–2):205– 225, March 2009. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Janson:2017:PRP

[Jan17] Svante Janson. Patterns in random permutations avoiding the pattern 132. Combinatorics, Probability and Computing, 26(1): 24–51, January 2017. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL https://www.cambridge.

org/core/product/BB28F694BDC4A021F11DDE41D75EB30E.

Janson:2019:RRT

[Jan19] Svante Janson. Random recursive trees and preferential attachment trees are random split trees. Combinatorics, Probability and Computing, 28(1):81–99, January 2019. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/random-
recursive-trees-and-preferential-attachment-trees-are-
random-split-trees/02A08DA234287E378EC087B424FF8E52.

Janson:2021:TLL

[Jan21] Svante Janson. Tree limits and limits of random trees. Combinatorics, Probability and Computing, 30(6):849–893, November 2021. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL https://www.cambridge.org/core/
journals/combinatorics-probability-and-computing/article/
tree-limits-and-limits-of-random-trees/79B2934A78D3B54A0077E178688FEE56.

Jin:1992:CSP

[Jin92] Guo Ping Jin. Complete subgraphs of r-partite graphs. Combinatorics, Probability and Computing, 1(3):241–250, September 1992. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Jin:1993:TFG

[Jin93] Guo Ping Jin. Triangle-free graphs with high minimal degrees. Combinatorics, Probability and Computing, 2(4):479–490, December 1993. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Javadi:2019:SRN

[JKOP19] R. Javadi, F. Khoeini, G. R. Omidi, and A. Pokrovskiy. On the size-Ramsey number of cycles. Combinatorics, Probability and Computing, 28(6):871–880, November 2019. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL /core/journals/combinatorics-probability-
and-computing/article/on-the-sizeramsey-number-of-cycles/
44661F2AAFFB3FA3984538300F7DB119.

Johannsen:2013:EUC

[JKS13] Daniel Johannsen, Michael Krivelevich, and Wojciech Samotij. Expanders are universal for the class of all spanning trees. Combinatorics, Probability and Computing, 22(2):253–281, ???? 2013. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Jacquet:2014:NPP

[JKS14] Philippe Jacquet, Charles Knessl, and Wojciech Szpankowski. A note on a problem posed by D. E. Knuth on a satisfiability recurrence. Combinatorics, Probability and Computing, 23(5): 829–841, September 2014. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). Special issue on Honouring the Memory of Philippe Flajolet - Part 1.

Jaworski:1992:CUG

[JŁ92] Jerzy Jaworski and Tomasz Łuczak. Cycles in a uniform graph process. Combinatorics, Probability and Computing, 1(3):223– 239, September 1992. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Jungic:2003:RAP

[JLM^{+}03] Veselin Jungić, Jacob Licht, Mohammad Mahdian, Jaroslav Nešetřil, and Radoš Radoičić. Rainbow arithmetic progressions and anti-Ramsey results. Combinatorics, Probability and Computing, 12(5–6):599–620, November 2003. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). Special issue on Ramsey theory.

Jiang:2023:BNU

[JLM23] Tao Jiang, Sean Longbrake, and Jie Ma. Bipartite-ness under smooth conditions. Combinatorics, Probability and Computing, 32(4):546–558, July 2023. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL https://www.cambridge.

org/core/journals/combinatorics-probability-and-computing/
article/bipartiteness-under-smooth-conditions/01A73AA670BEDEBE10CA4D2AD616A570.

Johnson:2015:SSC

[JLR15] J. Robert Johnson, Imre Leader, and Paul A. Russell. Set systems containing many maximal chains. Combinatorics, Probability and Computing, 24(3):480–485, May 2015. CODEN CP-COFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Johnson:2013:TRP

[JM13] J. Robert Johnson and Klas Markström. Turán and Ramsey properties of subcube intersection graphs. Combinatorics, Probability and Computing, 22(1):55–70, ???? 2013. CODEN CP-COFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Juvan:1998:LTC

[JMŠ98] M. Juvan, B. Mohar, and R. Škrekovski. List total colourings of graphs. Combinatorics, Probability and Computing, 7(2):181–188, June 1998. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http://journals.

cambridge.org/action/displayIssue?jid=CPC&volumeId=7&
issueId=02.

Jiang:2022:TEB

[JMY22] Tao Jiang, Jie Ma, and Liana Yepremyan. On Turán exponents of bipartite graphs. Combinatorics, Probability and Computing, 31(2):333–344, March 2022. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL https://www.cambridge.

org/core/journals/combinatorics-probability-and-computing/
article/on-turan-exponents-of-bipartite-graphs/939E454E5E6B3E2BE414B820D3B36B56.

Johnson:2004:DFF

[Joh04] J. Robert Johnson. A disproof of the Fon-der-Flaass Conjecture. Combinatorics, Probability and Computing, 13(2):195–201, March 2004. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL http://journals.cambridge.org/
action/displayIssue?jid=CPC&volumeId=13&issueId=02.

Johnson:2022:CPT

[Joh22] Tobias Johnson. Continuous phase transitions on Galton–Watson trees. Combinatorics, Probability and Computing, 31(2):198–228, March 2022. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL https://www.cambridge.org/core/
journals/combinatorics-probability-and-computing/article/
continuous-phase-transitions-on-galtonwatson-trees/309705D18DDF6EC6C45404F07BC95171.

Jonasson:1998:CTR

[Jon98] Johan Jonasson. On the cover time for random walks on random graphs. Combinatorics, Probability and Computing, 7(3): 265–279, September 1998. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http://journals.

cambridge.org/action/displayIssue?jid=CPC&volumeId=7&
issueId=03.

Johnson:2017:SSH

[JP17] J. Robert Johnson and Trevor Pinto. Saturated subgraphs of the hypercube. Combinatorics, Probability and Computing, 26 (1):52–67, January 2017. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL https://www.cambridge.

org/core/product/52DAFCA1426933D10068D1F4B144FCF0.

Jenssen:2022:ISG

[JPP22] Matthew Jenssen, Will Perkins, and Aditya Potukuchi. Independent sets of a given size and structure in the hypercube. Combinatorics, Probability and Computing, 31(4):702–720, July 2022.

CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/independent-
sets-of-a-given-size-and-structure-in-the-hypercube/
110CCBBAC5E904780FA0CB5D3167A591.

Jiang:2023:MTE

[JQ23] Tao Jiang and Yu Qiu. Many Turán exponents via subdivisions. Combinatorics, Probability and Computing, 32(1):134–150, January 2023. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL https://www.cambridge.org/core/
journals/combinatorics-probability-and-computing/article/
many-turan-exponents-via-subdivisions/3AF62F3C0AAEA4C1EFE0C7CC5D41CA24.

Jelenkovic:2004:OLC

[JR04] Predrag R. Jelenkovic and Ana Radovanovic. Optimizing LRU caching for variable document sizes. Combinatorics, Probability and Computing, 13(4–5):627–643, July 2004. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=13&issueId=04.

Jukna:2002:TFH

[JS02] S. Jukna and G. Schnitger. Triangle-freeness is hard to detect. Combinatorics, Probability and Computing, 11(6):549–569, November 2002. CODEN CPCOFG. ISSN 0963-5483 (print),

1469-2163 (electronic). URL http://journals.cambridge.
org/action/displayIssue?jid=CPC&volumeId=11&issueId=
06.

Janson:2007:PPD

[JS07] Svante Janson and Joel Spencer. A point process describing the component sizes in the critical window of the random graph evolution. Combinatorics, Probability and Computing, 16(4):631– 658, July 2007. CODEN CPCOFG. ISSN 0963-5483 (print),

1469-2163 (electronic). URL http://journals.cambridge.
org/action/displayIssue?jid=CPC&volumeId=16&issueId=
04.

Jaworski:2008:VDD

[JS08] Jerzy Jaworski and Dudley Stark. The vertex degree distribution of passive random intersection graph models. Combinatorics, Probability and Computing, 17(4):549–558, July 2008. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Jin:2020:GLR

[JS20] Emma Yu Jin and Benedikt Stufler. Graph limits of random unlabelled k-trees. Combinatorics, Probability and Computing, 29(5): 722–746, September 2020. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL https://www.cambridge.

org/core/journals/combinatorics-probability-and-computing/
article/graph-limits-of-random-unlabelled-ktrees/AB91DDA376B3097BB7E0EFA712D0E02C.

Johnston:2021:LBB

[JS21] Tom Johnston and Alex Scott. Lipschitz bijections between boolean functions. Combinatorics, Probability and Computing, 30(4):513–525, July 2021. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL https://www.cambridge.

org/core/journals/combinatorics-probability-and-computing/
article/lipschitz-bijections-between-boolean-functions/
72DE1E5F5E34349DC397F5508C076098.

Jain:2022:SSV

[JSS22] Vishesh Jain, Ashwin Sah, and Mehtaab Sawhney. On the smallest singular value of symmetric random matrices. Combinatorics, Probability and Computing, 31(4):662–683, July 2022. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/on-the-
smallest-singular-value-of-symmetric-random-matrices/
D918772E3671A55DC3798B5D1502DF58.

Janson:2008:DSR

[JT08a] Svante Janson and Andrew Thomason. Dismantling sparse random graphs. Combinatorics, Probability and Computing, 17(2):259–264, March 2008. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:/

/journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=17&issueId=02.

Johnson:2008:ITM

[JT08b] J. Robert Johnson and John Talbot. G-intersection theorems for matchings and other graphs. Combinatorics, Probability and Computing, 17(4):559–575, July 2008. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Jukna:2006:GC

[Juk06] S. Jukna. On graph complexity. Combinatorics, Probability and Computing, 15(6):855–876, November 2006. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=15&issueId=06.

Jiang:2003:ESS

[JW03] Tao Jiang and Douglas B. West. On the Erdős–Simonovits–Sós conjecture about the anti-Ramsey number of a cycle. Combinatorics, Probability and Computing, 12(5–6):585–598, November 2003. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). Special issue on Ramsey theory.

Jiang:2020:SEL

[JY20] Tao Jiang and Liana Yepremyan. Supersaturation of even linear cycles in linear hypergraphs. Combinatorics, Probability and Computing, 29(5):698–721, September 2020. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/supersaturation-
of-even-linear-cycles-in-linear-hypergraphs/240493C1F42618819DA9B16CE7F3E4BD.

Kahale:1997:LDB

[Kah97] Nabil Kahale. Large deviation bounds for Markov chains. Combinatorics, Probability and Computing, 6(4):465–474, December 1997. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=6&issueId=04.

Kahn:2001:EAH

[Kah01] Jeff Kahn. An entropy approach to the hard-core model on bipartite graphs. Combinatorics, Probability and Computing, 10(3): 219–237, May 2001. CODEN CPCOFG. ISSN 0963-5483 (print),

1469-2163 (electronic). URL http://journals.cambridge.
org/action/displayIssue?jid=CPC&volumeId=10&issueId=
03.

Kang:2019:SHC

[Kan19] Dong Yeap Kang. Sparse highly connected spanning subgraphs in dense directed graphs. Combinatorics, Probability and Computing, 28(3):423–464, May 2019. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/sparse-
highly-connected-spanning-subgraphs-in-dense-directed-
graphs/FD39F597B3D0D8FAF715AA78924BA504.

Kawarabayashi:2008:WOH

[Kaw08] Ken-Ichi Kawarabayashi. A weakening of the odd Hadwiger’s conjecture. Combinatorics, Probability and Computing, 17(6): 815–821, November 2008. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Kraetzl:1993:TCG

[KC93] Miro Kraetzl and Charles J. Colbourn. Threshold channel graphs. Combinatorics, Probability and Computing, 2(3):337– 349, September 1993. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Kramer:2011:NDS

[KCR11] Josh Brown Kramer, Jonathan Cutler, and A. J. Radcliffe. Negative dependence and Srinivasan’s sampling process. Combinatorics, Probability and Computing, 20(3):347–361, May 2011. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Keller:2011:IVB

[Kel11] Nathan Keller. On the influences of variables on Boolean functions in product spaces. Combinatorics, Probability and Computing, 20(1):83–102, January 2011. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Kim:1995:BTS

[Kim95] Jeong Han Kim. On Brooks’ theorem for sparse graphs. Combinatorics, Probability and Computing, 4(2):97–132, June 1995. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Kahn:2001:SPR

[KK01] Jeff Kahn and János Komlós. Singularity probabilities for random matrices over finite fields. Combinatorics, Probability and Computing, 10(2):137–157, March 2001. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=10&issueId=02.

Kahn:2007:TET

[KK07a] Jeff Kahn and Gil Kalai. Thresholds and expectation thresholds. Combinatorics, Probability and Computing, 16(3):495–502, May 2007. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=16&issueId=03.

Kaul:2007:EGG

[KK07b] Hemanshu Kaul and Alexandr Kostochka. Extremal graphs for a graph packing theorem of Sauer and Spencer. Combinatorics, Probability and Computing, 16(3):409–416, May 2007. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=16&issueId=03.

Kierstead:2008:SPH

[KK08] H. A. Kierstead and A. V. Kostochka. A short proof of the Hajnal–Szemerédi Theorem on equitable colouring. Combinatorics, Probability and Computing, 17(2):265–270, March 2008. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=17&issueId=02.

Kierstead:2009:EGP

[KK09] H. A. Kierstead and A. V. Kostochka. Efficient graph packing via game colouring. Combinatorics, Probability and Computing, 18(5):765–774, September 2009. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Kaiser:2014:DCI

[KK14] Tomás Kaiser and Ross J. Kang. The distance-t chromatic index of graphs. Combinatorics, Probability and Computing, 23(1):90– 101, January 2014. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Katona:2012:MPI

[KKK12] Gyula O. H. Katona, Gyula Y. Katona, and Zsolt Katona. Most probably intersecting families of subsets. Combinatorics, Probability and Computing, 21(1–2):219–227, ???? 2012. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Kostochka:2012:CFC

[KKL12] A. Kostochka, M. Kumbhat, and T. Luczak. Conflict-free colourings of uniform hypergraphs with few edges. Combinatorics, Probability and Computing, 21(4):611–622, July 2012. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Kierstead:2018:SDE

[KKM18] H. A. Kierstead, A. V. Kostochka, and A. McConvey. A sharp Dirac–Erdős type bound for large graphs. Combinatorics, Probability and Computing, 27(3):387–397, May 2018. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/sharp-
diracerdos-type-bound-for-large-graphs/3F599749BE2E301BC44E4EC46F38A006.

Kelly:2008:DTR

[KKO08] Luke Kelly, Daniela Kühn, and Deryk Osthus. A Dirac-type result on Hamilton cycles in oriented graphs. Combinatorics, Probability and Computing, 17(5):689–709, September 2008. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Korandi:2015:DRG

[KKS15] Dániel Korándi, Michael Krivelevich, and Benny Sudakov. Decomposing random graphs into few cycles and edges. Combinatorics, Probability and Computing, 24(6):857–872, November 2015. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Krivelevich:2016:CMR

[KKS16] Michael Krivelevich, Matthew Kwan, and Benny Sudakov. Cycles and matchings in randomly perturbed digraphs and hypergraphs. Combinatorics, Probability and Computing, 25(6): 909–927, November 2016. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL https://www.cambridge.

org/core/product/0B9450323E83AEF77A40AA701CC40EBF.

Kawarabayashi:2006:SDC

[KKY06] Ken-Ichi Kawarabayashi, Alexandr Kostochka, and Gexin Yu. On sufficient degree conditions for a graph to be k-linked. Combinatorics, Probability and Computing, 15(5):685–694, September 2006. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=15&issueId=05.

Keevash:2016:NSC

[KL16] Peter Keevash and Eoin Long. On the normalized Shannon capacity of a union. Combinatorics, Probability and Computing, 25(5):766–767, September 2016. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL https://www.

cambridge.org/core/product/AEF9EA053456FB91107EC17EE1852F64.

Keevash:2017:STE

[KL17] Peter Keevash and William Lochet. The structure of typical eye-free graphs and a Turán-type result for two weighted colours. Combinatorics, Probability and Computing, 26(6):886– 910, November 2017. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL https://www.cambridge.

org/core/product/F8AB63B1BA385414160F5FFD89FBE568.

Kubicki:2002:RIB

[KLM02] Grzegorz Kubicki, Jeno Lehel, and Michal Morayne. A ratio inequality for binary trees and the best secretary. Combinatorics, Probability and Computing, 11(2):149–161, March 2002. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=11&issueId=02.

Kannan:2006:BCM

[KLM06] R. Kannan, L. Lovász, and R. Montenegro. Blocking conductance and mixing in random walks. Combinatorics, Probability and Computing, 15(4):541–570, July 2006. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=15&issueId=04.

Kral:2019:DGE

[KLMP19] Daniel Král’, Bernard Lidický, Taísa L. Martins, and Yanitsa Pehova. Decomposing graphs into edges and triangles. Combinatorics, Probability and Computing, 28(3):465–472, May 2019. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/decomposing-
graphs-into-edges-and-triangles/E4C8C50C5038046E38E191A4C791EAD7.

Kuhn:2013:OPH

[KLO13] Daniela Kühn, John Lapinskas, and Deryk Osthus. Optimal packings of Hamilton cycles in graphs of high minimum degree. Combinatorics, Probability and Computing, 22(3):394–416, ???? 2013. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Kral:2013:NBC

[KLS^{+}13] Daniel Král’, Chun-Hung Liu, Jean-Sébastien Sereni, Peter Whalen, and Zelealem B. Yilma. A new bound for the 2/3 conjecture. Combinatorics, Probability and Computing, 22(3):384–393, ???? 2013. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Kamcev:2018:AFC

[KLS18] Nina Kamcev, Tomasz Luczak, and Benny Sudakov. Anagramfree colourings of graphs. Combinatorics, Probability and Computing, 27(4):623–642, July 2018. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL https://www.

cambridge.org/core/journals/combinatorics-probability-
and-computing/article/anagramfree-colourings-of-graphs/
CB8968416886E00DC91797AF42E0A4C0.

Klartag:2005:RSS

[KM05] B. Klartag and V. Milman. Rapid Steiner symmetrization of most of a convex body and the slicing problem. Combinatorics, Probability and Computing, 14(5–6):829–843, November 2005. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=14&issueId=05.

Kostochka:2008:WAM

[KM08] Alexandr Kostochka and Dhruv Mubayi. When is an almost monochromatic K_{4} guaranteed? Combinatorics, Probability and Computing, 17(6):823–830, November 2008. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Kang:2010:ICN

[KM10] Ross J. Kang and Colin McDiarmid. The t-improper chromatic number of random graphs. Combinatorics, Probability and Computing, 19(1):87–98, January 2010. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Kurauskas:2011:RGF

[KM11] Valentas Kurauskas and Colin McDiarmid. Random graphs with few disjoint cycles. Combinatorics, Probability and Computing, 20(5):763–775, September 2011. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Keevash:2012:TNF

[KM12] Peter Keevash and Dhruv Mubayi. The Turán number of F_{3,3}. Combinatorics, Probability and Computing, 21(3):451–456, May 2012. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Kortchemski:2017:SGN

[KM17] Igor Kortchemski and Cyril Marzouk. Simply generated noncrossing partitions. Combinatorics, Probability and Computing, 26(4):560–592, July 2017. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL https://www.cambridge.

org/core/product/C9AF65A51ADBBBB14FC5D02730C4B4B5.

Kamcev:2023:UPL

[KM23] Nina Kamcev and Alp Müyesser. Unavoidable patterns in locally balanced colourings. Combinatorics, Probability and Computing, 32(5):796–808, September 2023. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/unavoidable-
patterns-in-locally-balanced-colourings/8C1CC0D03AE2F08C2F13117AC6E86ED3.

Kupavskii:2019:BSA

[KMS19] Andrey Kupavskii, Nabil H. Mustafa, and Konrad J. Swanepoel. Bounding the size of an almost-equidistant set in Euclidean space. Combinatorics, Probability and Computing, 28(2):280– 286, March 2019. CODEN CPCOFG. ISSN 0963-5483 (print),

1469-2163 (electronic). URL https://www.cambridge.org/
core/journals/combinatorics-probability-and-computing/
article/bounding-the-size-of-an-almostequidistant-set-
in-euclidean-space/DD5A2CE5B5BD832561D0E7BD5B744636.

Kaplan:2012:UDT

[KMSS12] Haim Kaplan, Jiří Matousek, Zuzana Safernová, and Micha Sharir. Unit distances in three dimensions. Combinatorics, Probability and Computing, 21(4):597–610, July 2012. CODEN CP-COFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Keevash:2007:RTP

[KMSV07] Peter Keevash, Dhruv Mubayi, Benny Sudakov, and Jacques Verstraëte. Rainbow Turán problems. Combinatorics, Probability and Computing, 16(1):109–126, January 2007. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic).  URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=16&issueId=01.

Kinnersley:2012:DRN

[KMW12] William B. Kinnersley, Kevin G. Milans, and Douglas B. West. Degree Ramsey numbers of graphs. Combinatorics, Probability and Computing, 21(1–2):229–253, ???? 2012. CODEN CP-COFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Kostochka:1999:PDC

[KN99] A. V. Kostochka and J. Nešetřil. Properties of Descartes’ construction of triangle-free graphs with high chromatic number. Combinatorics, Probability and Computing, 8(5):467– 472, September 1999. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http://journals.

cambridge.org/action/displayIssue?jid=CPC&volumeId=8&
issueId=05.

Kostochka:2003:ECD

[KN03] A. V. Kostochka and K. Nakprasit. Equitable colourings of d-degenerate graphs. Combinatorics, Probability and Computing, 12(1):53–60, January 2003. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=12&issueId=01.

Kittipassorn:2014:CRT

[KN14a] Teeradej Kittipassorn and Bhargav P. Narayanan. A canonical Ramsey theorem for exactly m-coloured complete subgraphs. Combinatorics, Probability and Computing, 23(1):102–115, January 2014. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Knape:2014:PUC

[KN14b] Margarete Knape and Ralph Neininger. Pólya urns via the contraction method. Combinatorics, Probability and Computing, 23 (6):1148–1186, November 2014. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). Special issue on Honouring the Memory of Philippe Flajolet - Part 2.

Kostochka:2016:AEI

[KN16] Alexandr Kostochka and Jaroslav Nesetril. Adding edges to increase the chromatic number of a graph. Combinatorics, Probability and Computing, 25(4):592–594, July 2016. CODEN CP-COFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Kohayakawa:2003:HPT

[KNR03] Yoshiharu Kohayakawa, Brendan Nagle, and Vojtěch Rödl. Hereditary properties of triple systems. Combinatorics, Probability and Computing, 12(2):155–189, March 2003. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic).  URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=12&issueId=02.

Kuhn:2004:LTC

[KO04a] Daniela Kühn and Deryk Osthus. Large topological cliques in graphs without a 4-cycle. Combinatorics, Probability and Computing, 13(1):93–102, January 2004. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=13&issueId=01.

Kuhn:2004:SGA

[KO04b] Daniela Kühn and Deryk Osthus. Subdivisions of K_{r+2} in graphs of average degree at least r + ε and large but constant girth. Combinatorics, Probability and Computing, 13(3):361–371, May 2004. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=13&issueId=03.

Kuhn:2005:PDR

[KO05] Daniela Kühn and Deryk Osthus. Packings in dense regular graphs. Combinatorics, Probability and Computing, 14(3):325– 337, May 2005. CODEN CPCOFG. ISSN 0963-5483 (print),

1469-2163 (electronic). URL http://journals.cambridge.
org/action/displayIssue?jid=CPC&volumeId=14&issueId=

Kuuhn:2007:MSC

[KO07] Daniela Küuhn and Deryk Osthus. Maximizing several cuts simultaneously. Combinatorics, Probability and Computing, 16(2):277–283, March 2007. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=16&issueId=02.

Kuhn:2008:LOC

[KO08] Daniela Kühn and Deryk Osthus. Linkedness and ordered cycles in digraphs. Combinatorics, Probability and Computing, 17(3): 411–422, May 2008. CODEN CPCOFG. ISSN 0963-5483 (print),

1469-2163 (electronic). URL http://journals.cambridge.
org/action/displayIssue?jid=CPC&volumeId=17&issueId=
03.

Kang:2019:ICG

[KO19] Dong Yeap Kang and Sang-Il Oum. Improper colouring of graphs with no odd clique minor. Combinatorics, Probability and Computing, 28(5):740–754, September 2019. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL /core/journals/combinatorics-probability-
and-computing/article/improper-colouring-of-graphs-with-
no-odd-clique-minor/F0C6091DDB46096417FFDAF60FCEC87C.

Komlos:1999:BL

[Kom99] János Komlós. The blow-up lemma. Combinatorics, Probability and Computing, 8(1–2):161–176, January 1999. CODEN CP-COFG. ISSN 0963-5483 (print), 1469-2163 (electronic). Recent trends in combinatorics (Mátraháza, 1995).

Kordecki:1996:SSR

[Kor96] Wojciech Kordecki. Small submatroids in random matroids. Combinatorics, Probability and Computing, 5(3):257–266, September 1996. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Kotek:2012:CIP

[Kot12] Tomer Kotek. Complexity of Ising polynomials. Combinatorics, Probability and Computing, 21(5):743–772, September 2012. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Kirschenhofer:1998:CHF

[KP98] Peter Kirschenhofer and Helmut Prodinger. Comparisons in Hoare’s find algorithm. Combinatorics, Probability and Computing, 7(1):111–120, March 1998. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http://journals.

cambridge.org/action/displayIssue?jid=CPC&volumeId=7&
issueId=01.

Kuba:2011:ASG

[KP11] Markus Kuba and Alois Panholzer. Analysis of statistics for generalized Stirling permutations. Combinatorics, Probability and Computing, 20(6):875–910, November 2011. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Kang:2018:DCO

[KP18] Ross J. Kang and FranÇOis Pirot. Distance colouring without one cycle length. Combinatorics, Probability and Computing, 27(5):794–807, September 2018. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL https://www.

cambridge.org/core/journals/combinatorics-probability-
and-computing/article/distance-colouring-without-one-
cycle-length/02F6AF04625109193C7FA9CF1CCD0322.

Kuba:2019:CAG

[KP19] Markus Kuba and Alois Panholzer. Combinatorial analysis of growth models for series-parallel networks. Combinatorics, Probability and Computing, 28(4):574–599, July 2019. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/combinatorial-
analysis-of-growth-models-for-seriesparallel-networks/
EEA10F8E281E0D62189A5D6DBE8FB217.

Kuba:2022:BIT

[KP22] Markus Kuba and Alois Panholzer. On bucket increasing trees, clustered increasing trees and increasing diamonds. Combinatorics, Probability and Computing, 31(4):629–661, July 2022. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/on-bucket-
increasing-trees-clustered-increasing-trees-and-increasing-
diamonds/A777F49CA62EB9EEFB57775469935BF2.

Krieger:1994:SCA

[KR94] Abba M. Krieger and Paul R. Rosenbaum. A stochastic comparison for arrangement increasing functions. Combinatorics, Probability and Computing, 3(3):345–348, September 1994. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Krivelevich:1994:FGL

[Kri94] Michael Krivelevich. K^{s}-free graphs without large K^{r}-free subgraphs. Combinatorics, Probability and Computing, 3(3):349– 354, September 1994. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Krivelevich:1997:TFR

[Kri97] Michael Krivelevich. Triangle factors in random graphs. Combinatorics, Probability and Computing, 6(3):337–347, September 1997. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=6&issueId=03.

Krivelevich:2000:CND

[Kri00] Michael Krivelevich. The choice number of dense random graphs. Combinatorics, Probability and Computing, 9(1):19–26, January 2000. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=9&issueId=01.

Kriesell:2005:TDC

[Kri05] Matthias Kriesell. Triangle density and contractibility. Combinatorics, Probability and Computing, 14(1–2):133–146, January 2005. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=14&issueId=01.

Klimosova:2023:SBR

[KRRS23] Tereza Klimosová, Christian Reiher, Andrzej Ruci’nski, and Matas Sileikis. Sandwiching biregular random graphs. Combinatorics, Probability and Computing, 32(1):1–44, January 2023. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/sandwiching-
biregular-random-graphs/D42DB620D6D4B8FAD17F671CDA104E14.

Kohayakawa:2004:TTR

[KRS04] Yoshiharu Kohayakawa, Vojtech Rödl, and Mathias Schacht. The Turán theorem for random graphs. Combinatorics, Probability and Computing, 13(1):61–91, January 2004. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=13&issueId=01.

Kostochka:1999:SSS

[KRT99] A. V. Kostochka, V. Rödl, and L. A. Talysheva. On systems of small sets with no large Δ-subsystems. Combinatorics, Probability and Computing, 8(3):265–268, May 1999. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=8&issueId=03.

Krushkal:2011:GLD

[Kru11] Vyacheslav Krushkal. Graphs, links, and duality on surfaces. Combinatorics, Probability and Computing, 20(2):267–287, March 2011. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Komlos:1994:TCG

[KS94] János Komlós and Endre Szemerédi. Topological cliques in graphs. Combinatorics, Probability and Computing, 3(2):247– 256, June 1994. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Kilakos:1996:EMC

[KS96a] K. Kilakos and B. Shepherd. Excluding minors in cubic graphs. Combinatorics, Probability and Computing, 5(1):57–78, March 1996. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Komlos:1996:TCG

[KS96b] János Komlós and Endre Szemerédi. Topological cliques in graphs. II. Combinatorics, Probability and Computing, 5(1):79– 90, March 1996. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Kezdy:2002:DSM

[KS02] André E. Kézdy and Hunter S. Snevily. Distinct sums modulo n and tree embeddings. Combinatorics, Probability and Computing, 11(1):35–42, January 2002. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=11&issueId=01.

Keevash:2003:LDG

[KS03a] Peter Keevash and Benny Sudakov. Local density in graphs with forbidden subgraphs. Combinatorics, Probability and Computing, 12(2):139–153, March 2003. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=12&issueId=02.

Komjath:2003:PTS

[KS03b] Péter Komjáth and Saharon Shelah. A partition theorem for scattered order types. Combinatorics, Probability and Computing, 12(5–6):621–626, November 2003. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). Special issue on Ramsey theory.

Kostochka:2003:RNS

[KS03c] Alexandr Kostochka and Benny Sudakov. On Ramsey numbers of sparse graphs. Combinatorics, Probability and Computing, 12 (5–6):627–641, November 2003. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). Special issue on Ramsey theory.

Krause:2003:DOC

[KS03d] Matthias Krause and Hans Ulrich Simon. Determining the optimal contrast for secret sharing schemes in visual cryptography. Combinatorics, Probability and Computing, 12(3):285–299, May 2003. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=12&issueId=03. Combina-

torics, probability and computing (Oberwolfach, 2001).

Krivelevich:2003:LES

[KS03e] Michael Krivelevich and Benny Sudakov. The largest eigenvalue of sparse random graphs. Combinatorics, Probability and Computing, 12(1):61–72, January 2003. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=12&issueId=01.

Kezdy:2004:PVD

[KS04] André E. Kézdy and Hunter S. Snevily. Polynomials that vanish on distinct nth roots of unity. Combinatorics, Probability and Computing, 13(1):37–59, January 2004. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=13&issueId=01.

Khalfalah:2006:NMS

[KS06] Ayman Khalfalah and Endre Szemerédi. On the number of monochromatic solutions of boldmath x + y = z^{2}. Combinatorics, Probability and Computing, 15(1–2):213–227, January 2006. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=15&issueId=01.

Korner:2007:CSF

[KS07] János Körner and Blerina Sinaimeri. On cancellative set families. Combinatorics, Probability and Computing, 16(5):767–773, September 2007. CODEN CPCOFG. ISSN 0963-5483 (print),

1469-2163 (electronic). URL http://journals.cambridge.
org/action/displayIssue?jid=CPC&volumeId=16&issueId=
05.

Kang:2008:CPR

[KS08] M. Kang and T. G. Seierstad. The critical phase for random graphs with a given degree sequence. Combinatorics, Probability and Computing, 17(1):67–86, January 2008. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=17&issueId=01.

Kiwi:2009:SRB

[KS09] M. Kiwi and J. Soto. On a speculated relation between Chvátal– Sankoff constants of several sequences. Combinatorics, Probability and Computing, 18(4):517–532, July 2009. CODEN CP-COFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Kiwi:2015:LIS

[KS15] Marcos Kiwi and José A. Soto. Longest increasing subsequences of randomly chosen multi-row arrays. Combinatorics, Probability and Computing, 24(1):254–293, January 2015. CODEN CP-COFG. ISSN 0963-5483 (print), 1469-2163 (electronic). Special issue on Honouring the Memory of Philippe Flajolet - Part 3.

Kupavskii:2018:CUH

[KS18] Andrey Kupavskii and Dmitry Shabanov. Colourings of uniform hypergraphs with large girth and applications. Combinatorics, Probability and Computing, 27(2):245–273, March 2018. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/colourings-
of-uniform-hypergraphs-with-large-girth-and-applications/
38534EFBB8634F71119DAA068DB42AAB.

Komlos:1995:PPC

[KSS95] János Komlós, Gábor N. Sárközy, and Endre Szemerédi. Proof of a packing conjecture of Bollobás. Combinatorics, Probability and Computing, 4(3):241–255, September 1995. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Komlos:2001:STD

[KSS01] János Komlós, Gábor N. Sárkózy, and Endre Szemerédi. Spanning trees in dense graphs. Combinatorics, Probability and Computing, 10(5):397–416, September 2001. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=10&issueId=05.

Kahn:2011:DBI

[KSS11] Jeff Kahn, Michael Saks, and Clifford Smyth. The dual BKR inequality and Rudich’s conjecture. Combinatorics, Probability and Computing, 20(2):257–266, March 2011. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Karonski:1999:RIG

[KSSC99] Michał Karoński, Edward R. Scheinerman, and Karen B. Singer-Cohen. On random intersection graphs: the subgraph problem. Combinatorics, Probability and Computing, 8(1–2):131–159, January 1999. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). Recent trends in combinatorics (Mátraháza, 1995).

Krivelevich:2002:STN

[KSV02] Michael Krivelevich, Benny Sudakov, and Van H. Vu. A sharp threshold for network reliability. Combinatorics, Probability and Computing, 11(5):465–474, September 2002. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=11&issueId=05.

Krivelevich:2009:RSP

[KSV09] Michael Krivelevich, Benny Sudakov, and Dan Vilenchik. On the random satisfiable process. Combinatorics, Probability and Computing, 18(5):775–801, September 2009. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Knox:2013:ESB

[KT13] Fiachra Knox and Andrew Treglown. Embedding spanning bipartite graphs of small bandwidth. Combinatorics, Probability and Computing, 22(1):71–96, ???? 2013. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Korandi:2020:IRT

[KT20] Dániel Korándi and István Tomon. Improved Ramsey-type results for comparability graphs. Combinatorics, Probability and Computing, 29(5):747–756, September 2020. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/improved-
ramseytype-results-for-comparability-graphs/48958A9ECB076D15FC1A9045E2D4ED9C.

Kubicka:1996:EME

[Kub96] Ewa Kubicka. An efficient method of examining all trees. Combinatorics, Probability and Computing, 5(4):403–413, December 1996. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Kung:1993:SCI

[Kun93] Joseph P. S. Kung. Sign-coherent identities for characteristic polynomials of matroids. Combinatorics, Probability and Computing, 2(1):33–51, March 1993. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Kung:2000:CEC

[Kun00] Joseph P. S. Kung. Critical exponents, colines, and projective geometries. Combinatorics, Probability and Computing, 9(4):355–362, July 2000. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http://journals.

cambridge.org/action/displayIssue?jid=CPC&volumeId=9&
issueId=04.

Kurecka:2022:LBS

[Kur22] Martin Kurecka. Lower bound on the size of a quasirandom forcing set of permutations. Combinatorics, Probability and Computing, 31(2):304–319, March 2022. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/lower-
bound-on-the-size-of-a-quasirandom-forcing-set-of-permutations/
8D738E506AB2A44BB7578B2C00425C1D.

Kutin:2002:CLS

[Kut02] Samuel Kutin. Constructing large set systems with given intersection sizes modulo composite numbers. Combinatorics, Probability and Computing, 11(5):475–486, September 2002. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=11&issueId=05.

Kuzjurin:2000:ECR

[Kuz00] N. N. Kuzjurin. Explicit constructions of Rödl’s asymptotically good packings and coverings. Combinatorics, Probability and Computing, 9(3):265–276, May 2000. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=9&issueId=03.

Kuba:2010:DEW

[KW10] Markus Kuba and Stephan Wagner. Deterministic edge weights in increasing tree families. Combinatorics, Probability and Computing, 19(1):99–119, January 2010. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Kostochka:2007:OTG

[KY07] A. V. Kostochka and G. Yu. Ore-type graph packing problems. Combinatorics, Probability and Computing, 16(1):167–169, January 2007. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL http://journals.cambridge.org/
action/displayIssue?jid=CPC&volumeId=16&issueId=01.

Kostochka:2012:LRM

[KY12] Alexandr Kostochka and Matthew Yancey. Large rainbow matchings in edge-coloured graphs. Combinatorics, Probability and Computing, 21(1–2):255–263, ???? 2012. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Keevash:2020:NSF

[KY20] Peter Keevash and Liana Yepremyan. On the number of symbols that forces a transversal. Combinatorics, Probability and Computing, 29(2):234–240, March 2020. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/on-the-
number-of-symbols-that-forces-a-transversal/91F5A09EF35627EBE62B43AF220BC2C5.

Lamaison:2023:RUD

[Lam23] Ander Lamaison. Ramsey upper density of infinite graphs. Combinatorics, Probability and Computing, 32(5):703–723, September 2023. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL https://www.cambridge.org/core/
journals/combinatorics-probability-and-computing/article/
ramsey-upper-density-of-infinite-graphs/6CFD00E4384E6A64E51C1C383B914084.

Larsson:2021:MPM

[Lar21] Joel Larsson. The minimum perfect matching in pseudodimension 0 < q < 1. Combinatorics, Probability and Computing, 30(3):374–397, May 2021. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL https://www.cambridge.

org/core/journals/combinatorics-probability-and-computing/
article/minimum-perfect-matching-in-pseudodimension-
0-q-1/6BAAA80D6E991F2CF818F0EFEBA188F1.

Lawler:1993:DAT

[Law93] Gregory F. Lawler. A discrete analogue of a theorem of Makarov. Combinatorics, Probability and Computing, 2(2):181–199, June 1993. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Lawler:1995:RTC

[Law95] Gregory F. Lawler. Recurrence and transience for a card shuffling model. Combinatorics, Probability and Computing, 4(2):133–142, June 1995. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Loi:2014:SMA

[LCF14] C. Loi, P.-H. Cournède, and J. Françon. A symbolic method to analyse patterns in plant structure. Combinatorics, Probability and Computing, 23(5):842–860, September 2014. CODEN CP-COFG. ISSN 0963-5483 (print), 1469-2163 (electronic). Special issue on Honouring the Memory of Philippe Flajolet - Part 1.

Leader:2004:BRE

[Lea04] Imre Leader. Book review: “Extremal Combinatorics: with Applications in Computer Science” by Stasys Jukna, Springer, 2001, xvii + 375 pp. £32.50; $49.95, ISBN 3-540-66313-4. Combinatorics, Probability and Computing, 13(1):119–121, January 2004. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=13&issueId=01.

Lefmann:2005:SPC

[Lef05] Hanno Lefmann. Sparse parity-check matrices over GF (q). Combinatorics, Probability and Computing, 14(1–2):147–169, January 2005. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=14&issueId=01.

Lehner:2013:RCA

[Leh13] Florian Lehner. Random colourings and automorphism breaking in locally finite graphs. Combinatorics, Probability and Computing, 22(6):885–909, November 2013. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Lemos:1994:NBM

[Lem94] Manoel Lemos. Non-binary matroids having at most three nonbinary elements. Combinatorics, Probability and Computing, 3 (3):355–369, September 1994. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Lemoine:2022:LBT

[Lem22] Thibaut Lemoine. Large N behaviour of the two-dimensional Yang–Mills partition function. Combinatorics, Probability and Computing, 31(1):144–165, January 2022. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/large-
n-behaviour-of-the-twodimensional-yangmills-partition-
function/68E2F00A42AF7D162D81879A8E80B664.

Letzter:2016:PRN

[Let16] Shoham Letzter. Path Ramsey number for random graphs. Combinatorics, Probability and Computing, 25(4):612–622, July 2016. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Lewis:2008:HTH

[Lew08] Thomas M. Lewis. The hitting time for the height of a random recursive tree. Combinatorics, Probability and Computing, 17(6): 831–835, November 2008. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Linial:1993:LGP

[Lin93] Nathan Linial. Local-global phenomena in graphs. Combinatorics, Probability and Computing, 2(4):491–503, December 1993. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Linusson:2011:PBC

[Lin11] Svante Linusson. On percolation and the bunkbed conjecture. Combinatorics, Probability and Computing, 20(1):103–117, January 2011. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Linusson:2019:EPB

[Lin19] Svante Linusson. Erratum to ‘on percolation and the bunkbed conjecture’. Combinatorics, Probability and Computing, 28(6): 917–918, November 2019. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL /core/journals/

combinatorics-probability-and-computing/article/erratum-
to-on-percolation-and-the-bunkbed-conjecture/4D9D52D8F716980FC1E43A664873D4B5.

Liskovets:1996:SAE

[Lis96] Valery A. Liskovets. Some asymptotical estimates for planar Eulerian maps. Combinatorics, Probability and Computing, 5 (2):131–138, June 1996. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Liu:2022:HNT

[Liu22] Xizhi Liu. Hypergraphs without non-trivial intersecting subgraphs. Combinatorics, Probability and Computing, 31(6):1076– 1091, November 2022. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL https://www.cambridge.

org/core/journals/combinatorics-probability-and-computing/
article/hypergraphs-without-nontrivial-intersecting-
subgraphs/FBF9B8E169BF81314F80F16DBB17791D.

Lang:2021:MCP

[LL21] Richard Lang and Allan Lo. Monochromatic cycle partitions in random graphs. Combinatorics, Probability and Computing, 30 (1):136–152, January 2021. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL https://www.cambridge.

org/core/journals/combinatorics-probability-and-computing/
article/monochromatic-cycle-partitions-in-random-graphs/
880B178D98CC67FB4D353847851CF3A6.

Loera:2009:ECP

[LLMO09] J. A. Loera, J. Lee, S. Margulies, and S. Onn. Expressing combinatorial problems by systems of polynomial equations and Hilbert’s nullstellensatz. Combinatorics, Probability and Computing, 18(4):551–582, July 2009. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Lillo:2001:CRB

[LM01] R. E. Lillo and M. Martin. Characterization results based on a functional derivative approach. Combinatorics, Probability and Computing, 10(5):417–434, September 2001. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=10&issueId=05.

Law:2013:ISG

[LM13a] Hiu-Fai Law and Colin McDiarmid. On independent sets in graphs with given minimum degree. Combinatorics, Probability and Computing, 22(6):874–884, November 2013. CODEN CP-COFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Lo:2013:MVH

[LM13b] Allan Lo and Klas Markström. A multipartite version of the Hajnal–Szemerédi theorem for graphs and hypergraphs. Combinatorics, Probability and Computing, 22(1):97–111, ???? 2013.

CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Lima:2014:MTS

[LM14] Yuri Lima and Carlos Gustavo Moreira. A Marstrand theorem for subsets of integers. Combinatorics, Probability and Computing, 23(1):116–134, January 2014. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Lenz:2016:PPQ

[LM16] John Lenz and Dhruv Mubayi. Perfect packings in quasirandom hypergraphs II. Combinatorics, Probability and Computing, 25 (4):595–611, July 2016. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Li:2019:CAS

[LM19] Jiange Li and Mokshay Madiman. A combinatorial approach to small ball inequalities for sums and differences. Combinatorics, Probability and Computing, 28(1):100–129, January 2019. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/combinatorial-
approach-to-small-ball-inequalities-for-sums-and-differences/
D3AD98B958B4D3D4F804C06452A9432D.

Loth:2023:ENF

[LM23] Jesse Campion Loth and Bojan Mohar. Expected number of faces in a random embedding of any graph is at most linear. Combinatorics, Probability and Computing, 32(4):682–690, July 2023. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/expected-
number-of-faces-in-a-random-embedding-of-any-graph-is-
at-most-linear/6A74B34121E95670FB3E8191EBA4D74C.

Lieby:2009:SRE

[LMMW09] Paulette Lieby, Brendan D. McKay, Jeanette C. McLeod, and Ian M. Wanless. Subgraphs of random k-edge-coloured k-regular graphs. Combinatorics, Probability and Computing, 18(4):533– 549, July 2009. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Linial:2008:GCN

[LMST08] Nathan Linial, Jiří Matousek, Or Sheffet, and Gábor Tardos. Graph colouring with no large monochromatic components. Combinatorics, Probability and Computing, 17(4):577–589, July 2008. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Luo:2023:MNE

[LMY23] Cong Luo, Jie Ma, and Tianchi Yang. On the maximum number of edges in k-critical graphs. Combinatorics, Probability and Computing, 32(6):900–911, November 2023. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/on-the-
maximum-number-of-edges-in-kcritical-graphs/9D540DEE2EE90A8D99963128E8250BCC.

Lin:2021:ETG

[LNW21]  Huiqiu Lin, Bo Ning, and Baoyindureng Wu.  Eigen-

values and triangles in graphs. Combinatorics, Probability and Computing, 30(2):258–270, March 2021. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/eigenvalues-
and-triangles-in-graphs/C1C4C121DA20DB524354B1FDB8AF0F7C.

Lo:2009:TRG

[Lo09] Allan Siu Lun Lo. Triangles in regular graphs with density below one half. Combinatorics, Probability and Computing, 18(3):435– 440, May 2009. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Lo:2012:CGB

[Lo12] Allan Lo. Cliques in graphs with bounded minimum degree. Combinatorics, Probability and Computing, 21(3):457–482, May 2012. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Long:2017:LUS

[Lon17] Eoin Long. Large unavoidable subtournaments. Combinatorics, Probability and Computing, 26(1):68–77, January 2017. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/product/
BDD88AC11AE37435DAA175DEB1FC3879.

Long:2020:NBE

[Lon20] Jason Long. A note on the Brown–Erdős–Sós conjecture in groups. Combinatorics, Probability and Computing, 29(4):633– 640, July 2020. CODEN CPCOFG. ISSN 0963-5483 (print),

1469-2163 (electronic). URL https://www.cambridge.org/
core/journals/combinatorics-probability-and-computing/
article/note-on-the-brownerdossos-conjecture-in-groups/
49D3255FFDE30567D42A31F16362C886.

Lozin:2008:BCP

[Loz08] Vadim Lozin. Boundary classes of planar graphs. Combinatorics, Probability and Computing, 17(2):287–295, March 2008. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=17&issueId=02.

Luczak:1992:CRF

[ŁP92] Tomasz Łuczak and Boris Pittel. Components of random forests. Combinatorics, Probability and Computing, 1(1):35–52, March 1992. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Luczak:1993:RGS

[ŁP93] Tomasz Łuczak and László Pyber. On random generation of the symmetric group. Combinatorics, Probability and Computing, 2 (4):505–512, December 1993. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Lawler:2000:IES

[LP00] Gregory F. Lawler and Emily E. Puckette. The intersection exponent for simple random walk. Combinatorics, Probability and Computing, 9(5):441–464, September 2000. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=9&issueId=05.

Lennon:2009:LNS

[LP09] Craig Lennon and Boris Pittel. On the likely number of solutions for the stable marriage problem. Combinatorics, Probability and Computing, 18(3):371–421, May 2009. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Liu:2018:ERN

[LP18] Shiping Liu and Norbert Peyerimhoff. Eigenvalue ratios of non-negatively curved graphs. Combinatorics, Probability and Computing, 27(5):829–850, September 2018. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/eigenvalue-
ratios-of-nonnegatively-curved-graphs/FCE0A41C41E681254D89D8989864B8F7.

Liebenau:2020:AVJ

[LP20]  Anita Liebenau and Yanitsa Pehova.  An approximate

version of Jackson’s conjecture. Combinatorics, Probability and Computing, 29(6):886–899, November 2020. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/an-approximate-
version-of-jacksons-conjecture/BED9014DAFBA84CBD5B36AD5918CA0F6.

Long:2023:BVE

[LP23] Eoin Long and Laurentiu Ploscaru. A bipartite version of the Erdős–McKay conjecture. Combinatorics, Probability and Computing, 32(3):465–477, May 2023. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL https://www.

cambridge.org/core/journals/combinatorics-probability-
and-computing/article/bipartite-version-of-the-erdosmckay-
conjecture/A70EA6DA42FBA58B81E27EDEE331839A.

Liu:2021:NMP

[LPP21] Hong Liu, Péter Pál Pach, and Richárd Palincza. The number of maximum primitive sets of integers. Combinatorics, Probability and Computing, 30(5):781–795, September 2021. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/number-
of-maximum-primitive-sets-of-integers/FB2FA8BEAF0E124CC670674E4763B797.

Lefmann:2009:CHM

[LPRS09] Hanno Lefmann, Yury Person, Vojtech Rödl, and Mathias Schacht. On colourings of hypergraphs without monochromatic Fano planes. Combinatorics, Probability and Computing, 18(5): 803–818, September 2009. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Lyons:2008:GNS

[LPS08] Russell Lyons, Ron Peled, and Oded Schramm. Growth of the number of spanning trees of the Erdős–Rényi giant component. Combinatorics, Probability and Computing, 17(5):711–726, September 2008. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Luczak:1998:PTC

[ŁRS98] Tomasz Łuczak, Vojtěch Rödl, and Endre Szemerédi. Partitioning two-coloured complete graphs into two monochromatic cycles. Combinatorics, Probability and Computing, 7 (4):423–436, December 1998. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http://journals.

cambridge.org/action/displayIssue?jid=CPC&volumeId=7&
issueId=04.

Louchard:1995:PAS

[LS95] Guy Louchard and Wojciech Szpankowski. A probabilistic analysis of a string editing problem and its variations. Combinatorics, Probability and Computing, 4(2):143–166, June 1995. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Loh:2008:SCN

[LS08] Po-Shen Loh and Benny Sudakov. On the strong chromatic number of random graphs. Combinatorics, Probability and Computing, 17(2):271–286, March 2008. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=17&issueId=02.

Linial:2009:LCV

[LS09a] Nati Linial and Adi Shraibman. Learning complexity vs communication complexity. Combinatorics, Probability and Computing, 18(1–2):227–245, March 2009. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Loh:2009:CRN

[LS09b] Po-Shen Loh and Benny Sudakov. Constrained Ramsey numbers. Combinatorics, Probability and Computing, 18(1–2):247– 258, March 2009. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Lengler:2018:DAE

[LS18a] J. Lengler and A. Steger. Drift analysis and evolutionary algorithms revisited. Combinatorics, Probability and Computing, 27(4):643–666, July 2018. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL https://www.cambridge.

org/core/journals/combinatorics-probability-and-computing/
article/drift-analysis-and-evolutionary-algorithms-revisited/
264EEBAC1C92951B2E7B5BE1B4A1D6DE.

Linial:2018:MSH

[LS18b] Nathan Linial and Michael Simkin. Monotone subsequences in high-dimensional permutations. Combinatorics, Probability and Computing, 27(1):69–83, January 2018. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/product/
239407E0538A5D9C3430A7D8F71F4B76.

Lo:2019:ABS

[LSM19] Allan Lo and Nicolás Sanhueza-Matamala. An asymptotic bound for the strong chromatic number. Combinatorics, Probability and Computing, 28(5):768–776, September 2019. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL /core/journals/combinatorics-probability-
and-computing/article/an-asymptotic-bound-for-the-strong-
chromatic-number/90047FC8683FBB3DB2B0ADB3F68BDE37.

Liu:2017:LCE

[LSS17] Hong Liu, Maryam Sharifzadeh, and Katherine Staden. Local conditions for exponentially many subdivisions. Combinatorics, Probability and Computing, 26(3):423–430, May 2017. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/product/
3397C266BA57D031933707B973206FE6.

Lee:2012:GDH

[LSV12] Choongbum Lee, Benny Sudakov, and Dan Vilenchik. Getting a directed Hamilton cycle two times faster. Combinatorics, Probability and Computing, 21(5):773–801, September 2012. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Luczak:2010:VG

[LT10] Tomasz Luczak and Andrew Thomason. Vicarious graphs. Combinatorics, Probability and Computing, 19(5–6):819–827, November 2010. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Loh:2018:ITN

[LTTZ18] Po-Shen Loh, Michael Tait, Craig Timmons, and Rodrigo M. Zhou. Induced Turán numbers. Combinatorics, Probability and Computing, 27(2):274–288, March 2018. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/induced-
turan-numbers/CD98094837510D606F43931567F759D2.

Lo:2022:CSM

[LTZ22] Allan Lo, Andrew Treglown, and Yi Zhao. Complete subgraphs in a multipartite graph. Combinatorics, Probability and Computing, 31(6):1092–1101, November 2022. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/complete-
subgraphs-in-a-multipartite-graph/DC9088C5B343AD3630623632833C5C63.

Lukacs:1998:LEV

[Luk98] András Lukács. On local expansion of vertex-transitive graphs. Combinatorics, Probability and Computing, 7(2):205–209, June 1998. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=7&issueId=02.

Lovasz:2013:NDG

[LV13] László Lovász and Katalin Vesztergombi. Non-deterministic graph property testing. Combinatorics, Probability and Computing, 22(5):749–762, September 2013. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Levin:2018:ACN

[LV18] Leonid A. Levin and Ramarathnam Venkatesan. An average case NP-complete graph colouring problem. Combinatorics, Probability and Computing, 27(5):808–828, September 2018. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/an-average-
case-npcomplete-graph-colouring-problem/4ADD0DD208ED318A79996D8037AE5179.

Lovasz:1998:RMC

[LW98] László Lovász and Peter Winkler. Reversal of Markov chains and the forget time. Combinatorics, Probability and Computing, 7(2):189–204, June 1998. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http://journals.

cambridge.org/action/displayIssue?jid=CPC&volumeId=7&
issueId=02.

Lee:2002:CCP

[LW02] Orlando Lee and Yoshiko Wakabayashi. On the circuit cover problem for mixed graphs. Combinatorics, Probability and Computing, 11(1):43–59, January 2002. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=11&issueId=01.

Linusson:2007:CA

[LW07] Svante Linusson and Johan Wästlund. Completing a (k − 1)- assignment. Combinatorics, Probability and Computing, 16(4): 621–629, July 2007. CODEN CPCOFG. ISSN 0963-5483 (print),

1469-2163 (electronic). URL http://journals.cambridge.
org/action/displayIssue?jid=CPC&volumeId=16&issueId=
04.

Lehre:2021:TBH

[LW21] P. K. Lehre and C. Witt. Tail bounds on hitting times of randomized search heuristics using variable drift analysis. Combinatorics, Probability and Computing, 30(4):550–569, July 2021. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/tail-
bounds-on-hitting-times-of-randomized-search-heuristics-
using-variable-drift-analysis/F30499B93D4498D404E00206005A999A.

Lynch:1993:TFM

[Lyn93] James F. Lynch. Threshold functions for Markov chains: a graphtheoretic approach. Combinatorics, Probability and Computing, 2(3):351–362, September 1993. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Lyons:2005:AES

[Lyo05] Russell Lyons. Asymptotic enumeration of spanning trees. Combinatorics, Probability and Computing, 14(4):491–522, July 2005. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=14&issueId=04.

Lyons:2010:IIT

[Lyo10] Russell Lyons. Identities and inequalities for tree entropy. Combinatorics, Probability and Computing, 19(2):303–313, March 2010. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Lyons:2017:CGD

[Lyo17a] Russell Lyons. Comparing graphs of different sizes. Combinatorics, Probability and Computing, 26(5):681–696, September 2017. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL https://www.cambridge.org/core/
product/34ED4563F8621AF08187867711C045C2.

Lyons:2017:FIT

[Lyo17b] Russell Lyons. Factors of IID on trees. Combinatorics, Probability and Computing, 26(2):285–300, March 2017. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/product/
F503CF939B1070804F7C81EC719C4293.

Li:2009:DEL

[LZ09] Wenbo V. Li and Xinyi Zhang. On the difference of expected lengths of minimum spanning trees. Combinatorics, Probability and Computing, 18(3):423–434, May 2009. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Marchand:2007:LLL

[MA07] R. Marchand and E. Zohoorian Azad. Limit law of the length of the standard right factor of a Lyndon word. Combinatorics, Probability and Computing, 16(3):417–434, May 2007. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=16&issueId=03.

Maazoun:2020:BSP

[Maa20] Mickaël Maazoun. On the Brownian separable permuton. Combinatorics, Probability and Computing, 29(2):241–266, March 2020.

CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/on-the-
brownian-separable-permuton/653EC2FF3B4944F50751766331E2A340.

Mader:1994:VEC

[Mad94] W. Mader. On vertex-edge-critically n-connected graphs. Combinatorics, Probability and Computing, 3(2):257–271, June 1994. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Mader:1995:VDM

[Mad95] W. Mader. On vertices of degree n in minimally n-edge-connected graphs. Combinatorics, Probability and Computing, 4(1):81–95, March 1995. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Maansson:2000:CPA

[Mån00] Marianne Månsson. On compound Poisson approximation for sequence matching. Combinatorics, Probability and Computing, 9(6):529–548, November 2000. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http://journals.

cambridge.org/action/displayIssue?jid=CPC&volumeId=9&
issueId=06.

Maansson:2002:PAO

[Mån02a] Marianne Månsson. Pattern avoidance and overlap in strings. Combinatorics, Probability and Computing, 11(4):393–402, July 2002. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=11&issueId=04.

Manstavicius:2002:MDC

[Man02b] E. Manstavicius. Mappings on decomposable combinatorial structures: Analytic approach. Combinatorics, Probability and Computing, 11(1):61–78, January 2002. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=11&issueId=01.

Mani:2010:SIW

[Man10] Arun P. Mani. Some inequalities for Whitney–Tutte polynomials. Combinatorics, Probability and Computing, 19(3):425–439, May 2010. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Marchal:2001:CAT

[Mar01] Philippe Marchal. A combinatorial approach to the twosided exit problem for left-continuous random walks. Combinatorics, Probability and Computing, 10(3):251–266, May 2001. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=10&issueId=03.

Marchant:2013:LMC

[Mar13] Edward Marchant. The largest minimum codegree of a 3-graph without a generalized 4-cycle. Combinatorics, Probability and Computing, 22(1):112–117, ???? 2013. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Martinsson:2019:LTU

[Mar19a] Anders Martinsson. A linear threshold for uniqueness of solutions to random jigsaw puzzles. Combinatorics, Probability and Computing, 28(2):287–302, March 2019. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/linear-
threshold-for-uniqueness-of-solutions-to-random-jigsaw-
puzzles/A7B4B084923555FA362C457AD4E25CB7.

Marton:2019:LSI

[Mar19b] Katalin Marton. Logarithmic Sobolev inequalities in discrete product spaces. Combinatorics, Probability and Computing, 28 (6):919–935, November 2019. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL /core/journals/

combinatorics-probability-and-computing/article/logarithmic-
sobolev-inequalities-in-discrete-product-spaces/E7174E4BECFC9C6A2741E16FB55BDF5E.

Matus:1995:CIAb

[Mat95] F. Matúš. Conditional independences among four random variables. II. Combinatorics, Probability and Computing, 4(4):407– 417, December 1995. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Matthias:1997:CUB

[Mat97] Ulrich Matthias. Constructive upper bounds for the Turán number. Combinatorics, Probability and Computing, 6(4): 475–479, December 1997. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http://journals.

cambridge.org/action/displayIssue?jid=CPC&volumeId=6&
issueId=04.

Matus:1999:CIA

[Mat99] F. Matúš. Conditional independences among four random variables. III. Final conclusion. Combinatorics, Probability and Computing, 8(3):269–276, May 1999. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http://journals.

cambridge.org/action/displayIssue?jid=CPC&volumeId=8&
issueId=03.

Matousek:2003:LBS

[Mat03] Jiří Matoušek. A lower bound on the size of Lipschitz subsets in dimension 3. Combinatorics, Probability and Computing, 12(4): 427–430, July 2003. CODEN CPCOFG. ISSN 0963-5483 (print),

1469-2163 (electronic). URL http://journals.cambridge.
org/action/displayIssue?jid=CPC&volumeId=12&issueId=
04.

Mathe:2007:APW

[Mát07] András Máthé. The angel of power 2 wins. Combinatorics, Probability and Computing, 16(3):363–374, May 2007. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic).  URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=16&issueId=03.

Matiyasevich:2010:OMP

[Mat10] Yu. Matiyasevich. One more probabilistic reformulation of the four colour conjecture. Combinatorics, Probability and Computing, 19(5–6):829–834, November 2010. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Matousek:2014:NOS

[Mat14] Jirí Matousek. Near-optimal separators in string graphs. Combinatorics, Probability and Computing, 23(1):135–139, January 2014. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Mayhew:2005:IRB

[May05] Dillon Mayhew. Inequivalent representations of bias matroids. Combinatorics, Probability and Computing, 14(4):567–583, July 2005. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=14&issueId=04.

Milenkovic:2004:PTC

[MC04] O. Milenkovic and K. J. Compton. Probabilistic transforms for combinatorial urn models. Combinatorics, Probability and Computing, 13(4–5):645–675, July 2004. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=13&issueId=04.

McCutcheon:2003:STF

[McC03] Randall McCutcheon. A Sárközy theorem for finite fields. Combinatorics, Probability and Computing, 12(5–6):643–651, November 2003. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). Special issue on Ramsey theory.

McColm:2004:TFR

[McC04] Gregory L. McColm. Threshold functions for random graphs on a line segment. Combinatorics, Probability and Computing, 13(3): 373–387, May 2004. CODEN CPCOFG. ISSN 0963-5483 (print),

1469-2163 (electronic). URL http://journals.cambridge.
org/action/displayIssue?jid=CPC&volumeId=13&issueId=
03.

McDiarmid:1992:CIF

[McD92] Colin McDiarmid. On a correlation inequality of Farr. Combinatorics, Probability and Computing, 1(2):157–160, June 1992. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

McDiarmid:1993:RRM

[McD93] Colin McDiarmid. A random recolouring method for graphs and hypergraphs. Combinatorics, Probability and Computing, 2(3): 363–365, September 1993. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

McDiarmid:1997:CSB

[McD97] Colin McDiarmid. Centering sequences with bounded differences. Combinatorics, Probability and Computing, 6(1):79–86, March 1997. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=6&issueId=01.

McDiarmid:2002:CIP

[McD02] Colin McDiarmid. Concentration for independent permutations. Combinatorics, Probability and Computing, 11(2):163–178, March 2002. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL http://journals.cambridge.org/
action/displayIssue?jid=CPC&volumeId=11&issueId=02.

McDiarmid:2009:RGM

[McD09] Colin McDiarmid. Random graphs from a minor-closed class. Combinatorics, Probability and Computing, 18(4):583–599, July 2009. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

McKay:2011:SDR

[McK11] Brendan D. McKay. Subgraphs of dense random graphs with specified degrees. Combinatorics, Probability and Computing, 20 (3):413–433, May 2011. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

McMullen:2005:BRC

[McM05] Peter McMullen. Book review: “Convex Polytopes”, by Branko Grünbaum, second edition (first edition (1967) written with the cooperation of V. L. Klee, M. Perles and G. C. Shephard; second edition (2003) prepared by V. Kaibel, V. L. Klee and G. M. Ziegler), Graduate Texts in Mathematics, Vol. 221, Springer 2003, 568 pp., £61.50/EU 85.55/$79.97. Combinatorics, Probability and Computing, 14(4):623–626, July 2005. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=14&issueId=04.

Mecay:2001:MSM

[Mec01] Stefan T. Mecay. Maximum-sized matroids with no minors isomorphic to U_{2,5}, F_{7}, F_{7}^{−} or P_{7}. Combinatorics, Probability and Computing, 10(6):531–542, November 2001. CODEN CPCOFG.

ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=10&issueId=06.

Mehrabian:2011:LBC

[Meh11] Abbas Mehrabian. Lower bounds for the cop number when the robber is fast. Combinatorics, Probability and Computing, 20(4): 617–621, July 2011. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Menard:2018:VUI

[Mén18] Laurent Ménard. Volumes in the uniform infinite planar triangulation: From skeletons to generating functions. Combinatorics, Probability and Computing, 27(6):946–973, November 2018. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/volumes-
in-the-uniform-infinite-planar-triangulation-from-skeletons-
to-generating-functions/1673969865AE26E0475120F28AA9C2AD.

Mercer:2006:ARB

[Mer06] Idris David Mercer. Autocorrelations of random binary sequences. Combinatorics, Probability and Computing, 15 (5):663–671, September 2006. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:/

/journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=15&issueId=05.

Meszaros:2023:BIR

[Més23] András Mészáros. A BK inequality for random matchings. Combinatorics, Probability and Computing, 32(1):151–157, January 2023. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL https://www.cambridge.org/core/
journals/combinatorics-probability-and-computing/article/
bk-inequality-for-random-matchings/60FD42CCE9AD321CBEDBCFC33CF736DB.

Mills:1994:MRR

[Mil94] Allan D. Mills. A matroid reconstruction result. Combinatorics, Probability and Computing, 3(3):371–373, September 1994. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Milicevic:2017:SAG

[Mil17] Luka Milićević. Sets in almost general position. Combinatorics, Probability and Computing, 26(5):720–745, September 2017. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL https://www.cambridge.org/core/
product/9FC2FBE275BABC6274D86B5FBB91095E.

Milicevic:2023:ASF

[Mil23] Luka Milićević. Approximately symmetric forms far from being exactly symmetric. Combinatorics, Probability and Computing, 32(2):299–315, March 2023. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL https://www.cambridge.

org/core/journals/combinatorics-probability-and-computing/
article/approximately-symmetric-forms-far-from-being-
exactly-symmetric/6B1D0C3825B7CCBE760E26A278A5090D.

Mitzenmacher:1999:SBA

[Mit99] Michael Mitzenmacher. Studying balanced allocations with differential equations. Combinatorics, Probability and Computing, 8(5):473–482, September 1999. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http://journals.

cambridge.org/action/displayIssue?jid=CPC&volumeId=8&
issueId=05.

Maia:2001:MHS

[ML01] Bráulio Maia and Manoel Lemos. Matroids having small circumference. Combinatorics, Probability and Computing, 10(4): 349–360, July 2001. CODEN CPCOFG. ISSN 0963-5483 (print),

1469-2163 (electronic). URL http://journals.cambridge.
org/action/displayIssue?jid=CPC&volumeId=10&issueId=
04.

Melczer:2014:SAI

[MM14] Stephen Melczer and Marni Mishna. Singularity analysis via the iterated kernel method. Combinatorics, Probability and Computing, 23(5):861–888, September 2014. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). Special issue on Honouring the Memory of Philippe Flajolet - Part 1.

Merino:2009:ETG

[MN09] Criel Merino and Steven D. Noble. The equivalence of two graph polynomials and a symmetric function. Combinatorics, Probability and Computing, 18(4):601–615, July 2009. CODEN CP-COFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Morrison:2017:SHB

[MNS17] Natasha Morrison, Jonathan A. Noel, and Alex Scott. Saturation in the hypercube and bootstrap percolation. Combinatorics, Probability and Computing, 26(1):78–98, January 2017. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/product/
EB3A47A0D526EE16F13FFCBE0249A416.

Mousset:2020:TKK

[MNS20] Frank Mousset, Rajko Nenadov, and Wojciech Samotij. Towards the kohayakawa-kreuter conjecture on asymmetric Ramsey properties. Combinatorics, Probability and Computing, 29(6): 943–955, November 2020. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL https://www.cambridge.

org/core/journals/combinatorics-probability-and-computing/
article/towards-the-kohayakawakreuter-conjecture-on-
asymmetric-ramsey-properties/B0E78A36835B2949D44605C3190B90EB.

Maclean:2001:RFT

[MO01] C. F. Maclean and Neil O’Connell. Random finite topologies and their thresholds. Combinatorics, Probability and Computing, 10(3):239–249, May 2001. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=10&issueId=03.

Mohar:1994:ODC

[Moh94] Bojan Mohar. Obstructions for the disk and the cylinder embedding extension problems. Combinatorics, Probability and Computing, 3(3):375–406, September 1994. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Mohar:2006:BRG

[Moh06] Bojan Mohar. Book review: “Graphs on Surfaces and Their Applications”, by Sergei K. Lando and Alexander K. Zvonkin, Encyclopaedia of Mathematical Sciences 141, Springer-Verlag, 2004, 455pp., £73.00/$109.00/94.95 Euros. Combinatorics, Probability and Computing, 15(6):939–941, November 2006. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=15&issueId=06.

Mohle:2007:CNR

[Möh07] Martin Möhle. On a class of non-regenerative sampling distributions. Combinatorics, Probability and Computing, 16(3):435–444, May 2007. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL http://journals.cambridge.org/
action/displayIssue?jid=CPC&volumeId=16&issueId=03.

Mohar:2016:MEB

[Moh16] Bojan Mohar. Median eigenvalues of bipartite subcubic graphs. Combinatorics, Probability and Computing, 25(5):768– 790, September 2016. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL https://www.cambridge.

org/core/product/7D2DCEE8D6737ED6EF05820EB802B1A3.

Montenegro:2014:ICC

[Mon14] Ravi Montenegro. Intersection conductance and canonical alternating paths: Methods for general finite Markov chains. Combinatorics, Probability and Computing, 23(4):585–606, July 2014. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Montgomery:2017:FCD

[Mon17] Richard Montgomery. Fractional clique decompositions of dense partite graphs. Combinatorics, Probability and Computing, 26(6):911–943, November 2017. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL https://www.

cambridge.org/core/product/DB3EC831C9CCB6DDE226F12905764CB0.

Montgomery:2018:AED

[Mon18] Aaron M. Montgomery. Asymptotic enumeration of difference matrices over cyclic groups. Combinatorics, Probability and Computing, 27(1):84–109, January 2018. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/product/
0F9C95BF74792BD2100C0D58541A3944.

Montgomery:2020:HRD

[Mon20] Richard Montgomery. Hamiltonicity in random directed graphs is born resilient. Combinatorics, Probability and Computing, 29(6): 900–942, November 2020. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL https://www.cambridge.

org/core/journals/combinatorics-probability-and-computing/

article/hamiltonicity-in-random-directed-graphs-is-born-
resilient/E9F926DE4CBF255532C9839677528CB5.

Mori:1992:MWT

[Mór92] Tamás F. Móri. Maximum waiting times are asymptotically independent. Combinatorics, Probability and Computing, 1(3): 251–264, September 1992. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Mori:2005:MDB

[Mór05] Tamás F. Móri. The maximum degree of the Barabási–Albert random tree. Combinatorics, Probability and Computing, 14(3): 339–348, May 2005. CODEN CPCOFG. ISSN 0963-5483 (print),

1469-2163 (electronic). URL http://journals.cambridge.
org/action/displayIssue?jid=CPC&volumeId=14&issueId=
03.

Morris:2013:IMT

[Mor13] Ben Morris. Improved mixing time bounds for the Thorp shuffle. Combinatorics, Probability and Computing, 22(1):118–132, ???? 2013. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Mosca:1998:REC

[Mos98] Michele Mosca. Removing edges can increase the average number of colours in the colourings of a graph. Combinatorics, Probability and Computing, 7(2):211–216, June 1998. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=7&issueId=02.

Mossel:2002:MGP

[Mos02] Elchanan Mossel. The Minesweeper Game: Percolation and complexity. Combinatorics, Probability and Computing, 11(5):487–499, September 2002. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=11&issueId=05.

Mount:2000:FUC

[Mou00] John Mount. Fast unimodular counting. Combinatorics, Probability and Computing, 9(3):277–285, May 2000. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic).  URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=9&issueId=03.

Malon:2006:PFC

[MP06a] Christopher Malon and Igor Pak. Percolation on finite Cayley graphs. Combinatorics, Probability and Computing, 15(4):571– 588, July 2006. CODEN CPCOFG. ISSN 0963-5483 (print),

1469-2163 (electronic). URL http://journals.cambridge.
org/action/displayIssue?jid=CPC&volumeId=15&issueId=
04.

Matousek:2006:MIN

[MP06b] Jirí Matoušek and Aleš Prívetivý. The minimum independence number of a Hasse diagram. Combinatorics, Probability and Computing, 15(3):473–475, May 2006. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=15&issueId=03.

Mitsche:2013:RSR

[MP13] Dieter Mitsche and Pawel Pralat. Revolutionaries and spies on random graphs. Combinatorics, Probability and Computing, 22 (3):417–432, ???? 2013. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Methuku:2017:FHI

[MP17] Abhishek Methuku and Dömötör Pálvölgyi. Forbidden hypermatrices imply general bounds on induced forbidden subposet problems. Combinatorics, Probability and Computing, 26(4):593–602, July 2017. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL https://www.cambridge.org/core/
product/B201CDEB3B7AA15CE5E8AEAB3B443726.

Mphako:2000:TPP

[Mph00] Eunice Gogo Mphako. Tutte polynomials of perfect matroid designs. Combinatorics, Probability and Computing, 9(4):363–367, July 2000. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL http://journals.cambridge.org/
action/displayIssue?jid=CPC&volumeId=9&issueId=04.

Maruoka:1998:CNR

[MPK98] Akira Maruoka, Mike Paterson, and Hirotaka Koizumi. Consistency of natural relations on sets. Combinatorics, Probability and Computing, 7(3):281–293, September 1998. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=7&issueId=03.

Morris:2004:SPH

[MPP04] Kate Morris, Alois Panholzer, and Helmut Prodinger. On some parameters in heap ordered trees. Combinatorics, Probability and Computing, 13(4–5):677–696, July 2004. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=13&issueId=04.

Morales:2022:ANS

[MPT22] Alejandro H. Morales, Igor Pak, and Martin Tassy. Asymptotics for the number of standard tableaux of skew shape and for weighted lozenge tilings. Combinatorics, Probability and Computing, 31(4):550–573, July 2022. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/asymptotics-
for-the-number-of-standard-tableaux-of-skew-shape-and-
for-weighted-lozenge-tilings/CB6F609E5CB2718E9D9AED13D0F5D308.

McDiarmid:1995:AEG

[MR95] Colin McDiarmid and Bruce Reed. Almost every graph can be covered by ⌈Δ/2⌉ linear forests. Combinatorics, Probability and Computing, 4(3):257–268, September 1995. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

McKay:1998:AEE

[MR98a] Brendan D. McKay and Robert W. Robinson. Asymptotic enumeration of Eulerian circuits in the complete graph. Combinatorics, Probability and Computing, 7(4):437–449, December 1998. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=7&issueId=04.

Molloy:1998:SGC

[MR98b] Michael Molloy and Bruce Reed. The size of the giant component of a random graph with a given degree sequence. Combinatorics, Probability and Computing, 7(3):295–305, September 1998. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=7&issueId=03.

Martin:2006:DDM

[MR06] Russell Martin and Dana Randall. Disjoint decomposition of Markov chains and sampling circuits in Cayley graphs. Combinatorics, Probability and Computing, 15(3):411–448, May 2006. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=15&issueId=03.

McDiarmid:2008:MDR

[MR08] Colin McDiarmid and Bruce Reed. On the maximum degree of a random planar graph. Combinatorics, Probability and Computing, 17(4):591–601, July 2008. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Mubayi:2009:SS

[MR09] Dhruv Mubayi and Reshma Ramadurai. Simplex stability. Combinatorics, Probability and Computing, 18(3):441–454, May 2009. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Moore:2011:GIF

[MR11] Cristopher Moore and Alexander Russell. A graph integral formulation of the circuit partition polynomial. Combinatorics, Probability and Computing, 20(6):911–920, November 2011. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Mustafa:2019:PD

[MR19] Nabil H. Mustafa and Saurabh Ray. On a problem of Danzer. Combinatorics, Probability and Computing, 28(3):473–482, May 2019. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL https://www.cambridge.org/core/
journals/combinatorics-probability-and-computing/article/
on-a-problem-of-danzer/E71A9BF684CA793D79201CF9B49FF261.

Martin:2022:EDF

[MR22] Ryan R. Martin and Alex W. N. Riasanovsky. On the edit distance function of the random graph. Combinatorics, Probability and Computing, 31(2):345–367, March 2022. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/

combinatorics-probability-and-computing/article/on-the-
edit-distance-function-of-the-random-graph/6B8999F6CCEF86CC1A6DF043EACD6439.

Matus:1995:CIAa

[MS95] F. Matúš and M. Studený. Conditional independences among four random variables. I. Combinatorics, Probability and Computing, 4(3):269–278, September 1995. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Muller:2009:ILD

[MS09] T. Müller and J.-S. Sereni. Identifying and locating-dominating codes in (random) geometric networks. Combinatorics, Probability and Computing, 18(6):925–952, November 2009. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Miranda:2013:EHD

[MS13] Yuri Mejía Miranda and Gordon Slade. Expansion in high dimension for the growth constants of lattice trees and lattice animals. Combinatorics, Probability and Computing, 22(4):527–565, July 2013. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Martin:2020:MFG

[MS20] James B. Martin and Roman Stasiński. Minimax functions on Galton–Watson trees. Combinatorics, Probability and Computing, 29(3):455–484, May 2020. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL https://www.

cambridge.org/core/journals/combinatorics-probability-
and-computing/article/minimax-functions-on-galtonwatson-
trees/7318A926F1AF00465DF51811B9B91246.

Marciniszyn:2009:ORG

[MSS09a] Martin Marciniszyn, Reto Spöhel, and Angelika Steger. Online Ramsey games in random graphs. Combinatorics, Probability and Computing, 18(1–2):271–300, March 2009. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Marciniszyn:2009:UBO

[MSS09b] Martin Marciniszyn, Reto Spöhel, and Angelika Steger. Upper bounds for online Ramsey games in random graphs. Combinatorics, Probability and Computing, 18(1–2):259–270, March 2009. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Mubayi:2019:EGP

[MT19] Dhruv Mubayi and Caroline Terry. An extremal graph problem with a transcendental solution. Combinatorics, Probability and Computing, 28(2):303–324, March 2019. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/an-extremal-
graph-problem-with-a-transcendental-solution/57961022486FF3BC337889DDFC108824.

Mubayi:2002:SER

[Mub02] Dhruv Mubayi. Some exact results and new asymptotics for hypergraph Turán numbers. Combinatorics, Probability and Computing, 11(3):299–309, May 2002. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=11&issueId=03.

Mubayi:2016:IBR

[Mub16] Dhruv Mubayi. Improved bounds for the Ramsey number of tight cycles versus cliques. Combinatorics, Probability and Computing, 25(5):791–796, September 2016. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/product/
5980C527685C7E3ADBC2D4B7B3D9F2FB.

Mundici:2014:IMU

[Mun14] Daniele Mundici. Invariant measure under the affine group over invariant measure under the affine group over Z. Combinatorics, Probability and Computing, 23(2):248–268, March 2014. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Mutafchiev:2013:SLP

[Mut13] Ljuben Mutafchiev. The size of the largest part of random weighted partitions of large integers. Combinatorics, Probability and Computing, 22(3):433–454, ???? 2013. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Mohar:2006:PEC

[MV06] Bojan Mohar and Andrej Vodopivec. On polyhedral embeddings of cubic graphs. Combinatorics, Probability and Computing, 15(6):877–893, November 2006. CODEN CPCOFG.

ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=15&issueId=06.

Moreno:2023:RWS

[MV23] Elvira Moreno and Mauricio Velasco. On random walks and switched random walks on homogeneous spaces. Combinatorics, Probability and Computing, 32(3):398–421, May 2023. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/on-random-
walks-and-switched-random-walks-on-homogeneous-spaces/
8441DE2544CF7CC95326F861C112F506.

Madras:2005:SAW

[MW05a] Neal Madras and C. Chris Wu. Self-avoiding walks on hyperbolic graphs. Combinatorics, Probability and Computing, 14(4):523– 548, July 2005. CODEN CPCOFG. ISSN 0963-5483 (print),

1469-2163 (electronic). URL http://journals.cambridge.
org/action/displayIssue?jid=CPC&volumeId=14&issueId=
04.

May:2005:USI

[MW05b] William D. May and John C. Wierman. Using symmetry to improve percolation threshold bounds. Combinatorics, Probability and Computing, 14(4):549–566, July 2005. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=14&issueId=04.

May:2007:ANC

[MW07] William D. May and John C. Wierman. The application of noncrossing partitions to improving percolation threshold bounds. Combinatorics, Probability and Computing, 16(2):285–307, March 2007. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL http://journals.cambridge.org/
action/displayIssue?jid=CPC&volumeId=16&issueId=02.

Moseman:2008:FCP

[MW08] Elizabeth R. Moseman and Peter Winkler. On a form of coordinate percolation. Combinatorics, Probability and Computing, 17(6):837–845, November 2008. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Micek:2011:GGG

[MW11] Piotr Micek and Bartosz Walczak. A graph-grabbing game. Combinatorics, Probability and Computing, 20(4):623–629, July 2011. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Mathieu:2013:MMW

[MW13] Claire Mathieu and David B. Wilson. The min mean-weight cycle in a random network. Combinatorics, Probability and Computing, 22(5):763–782, September 2013. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

McDiarmid:2017:BAE

[MW17] Colin McDiarmid and Kerstin Weller. Bridge-addability, edge-expansion and connectivity. Combinatorics, Probability and Computing, 26(5):697–719, September 2017. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/product/
F8DD9F4294A07FB20A31C4D6EBD696E0.

Mubayi:2018:MS

[MW18] Dhruv Mubayi and Lujia Wang. Multicolour sunflowers. Combinatorics, Probability and Computing, 27(6):974–987, November 2018. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL https://www.cambridge.org/core/
journals/combinatorics-probability-and-computing/article/
multicolour-sunflowers/83E49E5AF1F0A67BCA1819A6A11A86B7.

Mohar:2022:TFS

[MW22] Bojan Mohar and Hehui Wu. Triangle-free subgraphs with large fractional chromatic number. Combinatorics, Probability and Computing, 31(1):136–143, January 2022. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/trianglefree-
subgraphs-with-large-fractional-chromatic-number/8E6011978AB76FCB7D4762508379D542.

McKay:2002:AEG

[MWW02] Brendan D. McKay, Ian M. Wanless, and Nicholas C. Wormald. Asymptotic enumeration of graphs with a given upper bound on the maximum degree. Combinatorics, Probability and Computing, 11(4):373–392, July 2002. CODEN CPCOFG.

ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=11&issueId=04.

Myers:2002:GLC

[Mye02] Joseph Samuel Myers. Graphs without large complete minors are quasi-random. Combinatorics, Probability and Computing, 11(6):571–585, November 2002. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=11&issueId=06.

Nacu:2006:IRP

[Nac06] Serban Nacu. Increments of random partitions. Combinatorics, Probability and Computing, 15(4):589–595, July 2006. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=15&issueId=04.

Nagy:2017:NEP

[Nag17] DÁNiel T. Nagy. On the number of 4-edge paths in graphs with given edge density. Combinatorics, Probability and Computing, 26(3):431–447, May 2017. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL https://www.cambridge.

org/core/product/0EF887D0410C48310D1F5ED3EF69B7D1.

Naor:2012:BSV

[Nao12] Assaf Naor. On the Banach-space-valued Azuma inequality and small-set isoperimetry of Alon–Roichman graphs. Combinatorics, Probability and Computing, 21(4):623–634, July 2012. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Narayanan:2018:CRO

[Nar18] Bhargav Narayanan. Connections in randomly oriented graphs. Combinatorics, Probability and Computing, 27(4):667–671, July 2018. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL https://www.cambridge.org/core/
journals/combinatorics-probability-and-computing/article/
connections-in-randomly-oriented-graphs/D8C33C2578EB4B3B21B37AF729EBE8A1.

Neininger:2002:WIR

[Nei02] Ralph Neininger. The Wiener index of random trees. Combinatorics, Probability and Computing, 11(6):587–597, November 2002. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=11&issueId=06.

Nesetril:1999:HSC

[Neš99] Jaroslav Nešetřil. The homomorphism structure of classes of graphs. Combinatorics, Probability and Computing, 8(1–2): 177–184, January 1999. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). Recent trends in combinatorics (Mátraháza, 1995).

Nesetril:2005:RCH

[Neš05] Jaroslav Nešetril. Ramsey classes and homogeneous structures. Combinatorics, Probability and Computing, 14(1–2):171–189, January 2005. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL http://journals.cambridge.org/
action/displayIssue?jid=CPC&volumeId=14&issueId=01.

Neuhauser:1996:PTD

[Neu96] Claudia Neuhauser. A phase transition for the distribution of matching blocks. Combinatorics, Probability and Computing, 5 (2):139–159, June 1996. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Newman:2023:AGR

[New23] Andrew Newman. Abelian groups from random hypergraphs. Combinatorics, Probability and Computing, 32(4):654–664, July 2023. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL https://www.cambridge.org/core/
journals/combinatorics-probability-and-computing/article/
abelian-groups-from-random-hypergraphs/C147A716F674EF3D24871747063B7813.

Nguyen:2011:DTT

[Ngu11] Hoi H. Nguyen. On the distribution of three-term arithmetic progressions in sparse subsets of F_{p}^{n}. Combinatorics, Probability and Computing, 20(5):777–791, September 2011. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Nica:2023:PSG

[Nic23] Bogdan Nica. Polynomials over structured grids. Combinatorics, Probability and Computing, 32(2):284–298, March 2023.

CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/polynomials-
over-structured-grids/ADCA0913E60648143E68EB7FC7D9A3CD.

Nikiforov:2001:EDG

[Nik01a] V. Nikiforov. On the edge distribution of a graph. Combinatorics, Probability and Computing, 10(6):543–555, November 2001. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=10&issueId=06.

Nikiforov:2001:MNC

[Nik01b] V. Nikiforov. On the minimum number of k-cliques in graphs with restricted independence number. Combinatorics, Probability and Computing, 10(4):361–366, July 2001. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=10&issueId=04.

Nikiforov:2002:SIL

[Nik02] V. Nikiforov. Some inequalities for the largest eigenvalue of a graph. Combinatorics, Probability and Computing, 11(2):179– 189, March 2002. CODEN CPCOFG. ISSN 0963-5483 (print),

1469-2163 (electronic). URL http://journals.cambridge.
org/action/displayIssue?jid=CPC&volumeId=11&issueId=
02.

Nikiforov:2005:CCG

[Nik05] V. Nikiforov. The cycle-complete graph Ramsey numbers. Combinatorics, Probability and Computing, 14(3):349–370, May 2005. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=14&issueId=03.

Nikiforov:2006:EDG

[Nik06] V. Nikiforov. Edge distribution of graphs with few copies of a given graph. Combinatorics, Probability and Computing, 15(6):895–902, November 2006. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=15&issueId=06.

Nikiforov:2009:SES

[Nik09] Vladimir Nikiforov. A spectral Erdős–Stone–Bollobás theorem. Combinatorics, Probability and Computing, 18(3):455–458, May 2009. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Nilli:1994:PHP

[Nil94] A. Nilli. Perfect hashing and probability. Combinatorics, Probability and Computing, 3(3):407–409, September 1994. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Noble:1998:ETP

[Nob98] S. D. Noble. Evaluating the Tutte polynomial for graphs of bounded tree-width. Combinatorics, Probability and Computing, 7(3):307–321, September 1998. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http://journals.

cambridge.org/action/displayIssue?jid=CPC&volumeId=7&
issueId=03.

Noble:2006:ERG

[Nob06] S. D. Noble. Evaluating the rank generating function of a graphic 2-polymatroid. Combinatorics, Probability and Computing, 15(3):449–461, May 2006. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=15&issueId=03.

Nguyen:2020:SNS

[NP20a] Hoi. H. Nguyen and Elliot Paquette. Surjectivity of near-square random matrices. Combinatorics, Probability and Computing, 29 (2):267–292, March 2020. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL https://www.cambridge.

org/core/journals/combinatorics-probability-and-computing/
article/surjectivity-of-nearsquare-random-matrices/D1909693321D0F774F53A748BFE8D656.

Ning:2020:EEG

[NP20b] Bo Ning and Xing Peng. Extensions of the Erdős–Gallai theorem and Luo’s theorem. Combinatorics, Probability and Computing, 29(1):128–136, January 2020. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL https://www.

cambridge.org/core/journals/combinatorics-probability-
and-computing/article/extensions-of-the-erdosgallai-
theorem-and-luos-theorem/84938C868FE59C5F7521F31F88E40FA6.

Norris:2017:SPM

[NPZ17] James Norris, Yuval Peres, and Alex Zhai. Surprise probabilities in Markov chains. Combinatorics, Probability and Computing, 26(4):603–627, July 2017. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL https://www.cambridge.

org/core/product/EFB9F5F7775CDED628A8F28D47DACC2C.

Nikiforov:2005:BRN

[NRS05] V. Nikiforov, C. C. Rousseau, and R. H. Schelp. Book Ramsey numbers and quasi-randomness. Combinatorics, Probability and Computing, 14(5–6):851–860, November 2005. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic).  URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=14&issueId=05.

Nolte:2001:SAG

[NS01] Andreas Nolte and Rainer Schrader. Simulated annealing and graph colouring. Combinatorics, Probability and Computing, 10(1):29–40, January 2001. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=10&issueId=01.

Nenadov:2016:SPR

[NS16] Rajko Nenadov and Angelika Steger. A short proof of the Random Ramsey Theorem. Combinatorics, Probability and Computing, 25(1):130–144, January 2016. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). Special issue on Oberwolfach Special Issue Part 2.

Nenadov:2020:KTT

[NS20] Rajko Nenadov and Nemanja Skorić. On Komlós’ tiling theorem in random graphs. Combinatorics, Probability and Computing, 29 (1):113–127, January 2020. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL https://www.cambridge.

org/core/journals/combinatorics-probability-and-computing/
article/on-komlos-tiling-theorem-in-random-graphs/37E38B81AE4356478E95035AB36911B2.

Narayanan:2022:MTC

[NS22] Hariharan Narayanan and Piyush Srivastava. On the mixing time of coordinate hit-and-run. Combinatorics, Probability and Computing, 31(2):320–332, March 2022. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/on-the-
mixing-time-of-coordinate-hitandrun/8CFBD4ABE132EA14CA27F9E2D6B69DEE.

Norin:2023:CCG

[NSW23] Sergey Norin, Alex Scott, and David R. Wood. Clustered colouring of graph classes with bounded treedepth or pathwidth. Combinatorics, Probability and Computing, 32(1):122–133, January 2023. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL https://www.cambridge.org/core/
journals/combinatorics-probability-and-computing/article/
clustered-colouring-of-graph-classes-with-bounded-treedepth-
or-pathwidth/AFB47818C14D268E4CB83FF78C0B3C12.

Ngoc:1993:LTA

[NT93] Nguy ˜ ên Văn Ngo.c and Zsolt Tuza. Linear-time approximation algorithms for the max cut problem. Combinatorics, Probability and Computing, 2(2):201–210, June 1993. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Nagy:2016:CPD

[NT16] Gábor V. Nagy and Vilmos Totik. A convexity property of discrete random walks. Combinatorics, Probability and Computing, 25(6):928–940, November 2016. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/product/
17B043E91AAB1025E0149F637E48EEB1.

Narayanan:2018:ISM

[NT18] Bhargav Narayanan and István Tomon. Induced subgraphs with many distinct degrees. Combinatorics, Probability and Computing, 27(1):110–123, January 2018. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL https://www.

cambridge.org/core/product/61BD845CFC0B7D2AFCA32103FF07FCA5.

Nesetril:1994:RTT

[NV94] Jaroslav Nešetřil and Pavel Valtr. A Ramsey-type theorem in the plane. Combinatorics, Probability and Computing, 3(1):127– 135, March 1994. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Naor:2005:NBG

[NV05] Assaf Naor and Jacques Verstraëte. A note on bipartite graphs without 2^{k}-cycles. Combinatorics, Probability and Computing, 14(5–6):845–849, November 2005. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=14&issueId=05.

onChan:2012:CF

[oCM12] Siu on Chan and Michael Molloy. (k + 1)-cores have kfactors. Combinatorics, Probability and Computing, 21(6):882– 896, November 2012. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Odell:2006:BRR

[Ode06] Edward Odell. Book review: “Ramsey Methods in Analysis”, by Spiros A. Argyros and Stevo Todorcevic, Birkhäuser 2005, 257pp.,£29.00/$49.95/38 Euros. Combinatorics, Probability and Computing, 15(4):631–636, July 2006. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=15&issueId=04.

Ofek:2007:EGC

[Ofe07] Eran Ofek. On the expansion of the giant component in percolated (n, d, λ) graphs. Combinatorics, Probability and Computing, 16(3):445–457, May 2007. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=16&issueId=03.

Oliveira:2008:BBP

[Oli08] Roberto Oliveira. Balls-in-bins processes with feedback and Brownian motion. Combinatorics, Probability and Computing, 17(1):87–110, January 2008. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=17&issueId=01.

Omidi:2021:PCG

[Omi21] G. R. Omidi. A proof of a conjecture of Gyárfás, Lehel, Sárközy and Schelp on Berge-cycles. Combinatorics, Probability and Computing, 30(5):654–669, September 2021. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/proof-
of-a-conjecture-of-gyarfas-lehel-sarkozy-and-schelp-
on-bergecycles/C0AAD4BC48E93C061E3CBEB667618723.

Ok:2018:DRZ

[OP18] Seongmin Ok and Thomas J. Perrett. Density of real zeros of the Tutte polynomial. Combinatorics, Probability and Computing, 27(3):398–410, May 2018. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL https://www.cambridge.

org/core/journals/combinatorics-probability-and-computing/
article/density-of-real-zeros-of-the-tutte-polynomial/
C24BF0FB7FBE22C32EB0AD95BE74F9A0.

Oporowski:2002:PMO

[Opo02] Bogdan Oporowski. Partitioning matroids with only small cocircuits. Combinatorics, Probability and Computing, 11(2):191–197, March 2002. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL http://journals.cambridge.org/
action/displayIssue?jid=CPC&volumeId=11&issueId=02.

Orenshtein:2014:GRW

[OS14] Tal Orenshtein and Igor Shinkar. Greedy random walk. Combinatorics, Probability and Computing, 23(2):269–289, March 2014. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Oberkampf:2020:SDG

[OS20] Heiner Oberkampf and Mathias Schacht. On the structure of dense graphs with bounded clique number. Combinatorics, Probability and Computing, 29(5):641–649, September 2020. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/on-the-
structure-of-dense-graphs-with-bounded-clique-number/
8C4ACF57DC47E04C63859D1AA84E99B6.

Oxley:2002:CFR

[OW02] James Oxley and Dominic Welsh. Chromatic, flow and reliability polynomials: The complexity of their coefficients. Combinatorics, Probability and Computing, 11(4):403–426, July 2002. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=11&issueId=04.

Panholzer:2004:DSD

[Pan04] Alois Panholzer. Distribution of the Steiner distance in generalized M-ary search trees. Combinatorics, Probability and Computing, 13(4–5):717–733, July 2004. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=13&issueId=04.

Parviainen:2004:RAI

[Par04] Robert Parviainen. Random assignment with integer costs. Combinatorics, Probability and Computing, 13(1):103–113, January 2004. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=13&issueId=01.

Parzanchevski:2017:MHD

[Par17] Ori Parzanchevski. Mixing in high-dimensional expanders. Combinatorics, Probability and Computing, 26(5):746–761, September 2017. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL https://www.cambridge.org/core/
product/5317C713ED80768DA23606A9B0161CEA.

Paulsen:2006:BOF

[Pau06] W. Paulsen. Best odds for finding a perfect matching in a bipartite graph. Combinatorics, Probability and Computing, 15(5):753–763, September 2006. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=15&issueId=05.

Pach:2017:DDA

[PD17] JÁNos Pach and Frank De Zeeuw. Distinct distances on algebraic curves in the plane. Combinatorics, Probability and Computing, 26(1):99–117, January 2017. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL https://www.cambridge.

org/core/product/0C7AC997A40A263E521284A61804CE46.

Pebody:2004:RFA

[Peb04] Luke Pebody. The reconstructibility of finite Abelian groups. Combinatorics, Probability and Computing, 13(6):867–892, November 2004. CODEN CPCOFG. ISSN 0963-5483 (print),

1469-2163 (electronic). URL http://journals.cambridge.
org/action/displayIssue?jid=CPC&volumeId=13&issueId=
06.

Pebody:2007:RON

[Peb07] Luke Pebody. Reconstructing odd necklaces. Combinatorics, Probability and Computing, 16(4):503–514, July 2007. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic).  URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=16&issueId=04.

Pekec:1996:WSR

[Pek96] Aleksandar Pekeč. A winning strategy for the Ramsey graph game. Combinatorics, Probability and Computing, 5(3):267–276, September 1996. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Pelc:1994:ASG

[Pel94] Andrzej Pelc. Almost safe group testing with few tests. Combinatorics, Probability and Computing, 3(3):411–419, September 1994. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Petridis:2011:PI

[Pet11] Giorgis Petridis. Plünnecke’s inequality. Combinatorics, Probability and Computing, 20(6):921–938, November 2011. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Petersson:2013:MDL

[Pet13] Niclas Petersson. The maximum displacement for linear probing hashing. Combinatorics, Probability and Computing, 22(3):455– 476, ???? 2013. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Picollelli:2011:FSF

[Pic11] Michael E. Picollelli. The final size of the C_{4}-free process. Combinatorics, Probability and Computing, 20(6):939–955, November 2011. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Pikhurko:1999:MSS

[Pik99] Oleg Pikhurko. The minimum size of saturated hypergraphs. Combinatorics, Probability and Computing, 8(5): 483–492, September 1999. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http://journals.

cambridge.org/action/displayIssue?jid=CPC&volumeId=8&
issueId=05.

Pikhurko:2001:WSH

[Pik01] Oleg Pikhurko. Weakly saturated hypergraphs and exterior algebra. Combinatorics, Probability and Computing, 10(5):435–451, September 2001. CODEN CPCOFG. ISSN 0963-5483 (print),

1469-2163 (electronic). URL http://journals.cambridge.
org/action/displayIssue?jid=CPC&volumeId=10&issueId=
05.

Pikhurko:2010:FUA

[Pik10] Oleg Pikhurko. Finding an unknown acyclic orientation of a given graph. Combinatorics, Probability and Computing, 19(1): 121–131, January 2010. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Pimentel:2011:AFP

[Pim11] Leandro P. R. Pimentel. Asymptotics for first-passage times on Delaunay triangulations. Combinatorics, Probability and Computing, 20(3):435–453, May 2011. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Pinto:2017:PTD

[Pin17] Trevor Pinto. The proofs of two directed paths conjectures of Bollobás and Leader. Combinatorics, Probability and Computing, 26(5):762–774, September 2017. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/product/
BEB3CE235FFE09D1E2A1597B76070F36.

Pinsky:2020:ILR

[Pin20] Ross G. Pinsky. The infinite limit of random permutations avoiding patterns of length three. Combinatorics, Probability and Computing, 29(1):137–152, January 2020. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/infinite-

limit-of-random-permutations-avoiding-patterns-of-length-
three/793B9C66C85B2A43CF01862E91B8CD16.

Pittel:1993:RIS

[Pit93] Boris Pittel. On a random instance of a “stable roommates” problem: likely behavior of the proposal algorithm. Combinatorics, Probability and Computing, 2(1):53–92, March 1993. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Pitman:2002:PDG

[Pit02] Jim Pitman. Poisson–Dirichlet and GEM invariant distributions for split-and-merge transformations of an interval partition2. Combinatorics, Probability and Computing, 11(5):501–514, September 2002. CODEN CPCOFG. ISSN 0963-5483 (print),

1469-2163 (electronic). URL http://journals.cambridge.
org/action/displayIssue?jid=CPC&volumeId=11&issueId=
05.

Pittel:2005:DRS

[Pit05] Boris Pittel. On dimensions of a random solid diagram. Combinatorics, Probability and Computing, 14(5–6):873–895, November 2005. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=14&issueId=05.

Plagne:2011:SDG

[Pla11a] Alain Plagne. Sums of dilates in groups of prime order. Combinatorics, Probability and Computing, 20(6):867–873, November 2011. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Plagne:2011:YOH

[Pla11b] Alain Plagne. Yahya Ould Hamidoune the Mauritanian mathematician: 1948–11 march 2011. Combinatorics, Probability and Computing, 20(5):641–645, September 2011. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Pontiveros:2013:FHR

[Pon13a] Gonzalo Fiz Pontiveros. Freiman homomorphisms of random subsets of Z_{N}. Combinatorics, Probability and Computing, 22 (4):592–611, July 2013. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Pontiveros:2013:SD

[Pon13b] Gonzalo Fiz Pontiveros. Sums of dilates in Z_{p}. Combinatorics, Probability and Computing, 22(2):282–293, ???? 2013. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Percus:1992:ESC

[PP92] Ora Engelberg Percus and Jerome K. Percus. An expanded set of correlation tests for linear congruential random number generators. Combinatorics, Probability and Computing, 1(2):161–168, June 1992. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Pemantle:2014:CLF

[PP14] Robin Pemantle and Yuval Peres. Concentration of Lipschitz functionals of determinantal and other strong Rayleigh measures. Combinatorics, Probability and Computing, 23(1):140–160, January 2014. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Panagiotou:2015:RRS

[PPGSS15] Konstantinos Panagiotou, Xavier Pérez-Giménez, Thomas Sauerwald, and He Sun. Randomized rumour spreading: The effect of the network topology. Combinatorics, Probability and Computing, 24(2):457–479, March 2015. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Perarnau:2017:ESF

[PR17a] G. Perarnau and B. Reed. Existence of spanning F -free subgraphs with large minimum degree. Combinatorics, Probability and Computing, 26(3):448–467, May 2017. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/product/
699F50FECEE9E9DDA2984E82748368C8.

Pikhurko:2017:ASG

[PR17b] Oleg Pikhurko and Alexander Razborov. Asymptotic structure of graphs with the minimum number of triangles. Combinatorics, Probability and Computing, 26(1):138–160, January 2017. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL https://www.cambridge.org/core/
product/AE7E297950A0F8C2822C338653DB437A.

Promel:2005:CDI

[Prö05a] Hans Jürgen Prömel. Complete disorder is impossible: The mathematical work of Walter Deuber. Combinatorics, Probability and Computing, 14(1–2):3–16, January 2005. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=14&issueId=01.

Promel:2005:SIM

[Prö05b] Hans Jürgen Prömel. Special issue in memory of Walter Deuber. Combinatorics, Probability and Computing, 14(1–2):1, January 2005. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=14&issueId=01.

Peng:2005:CSC

[PRS05] Y. Peng, V. Rödl, and J. Skokan. Counting small cliques in 3-uniform hypergraphs. Combinatorics, Probability and Computing, 14(3):371–413, May 2005. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=14&issueId=03.

Pach:2016:RTC

[PRT16] János Pach, Natan Rubin, and Gábor Tardos. On the Richter–Thomassen conjecture about pairwise intersecting closed curves. Combinatorics, Probability and Computing, 25(6):941– 958, November 2016. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL https://www.cambridge.

org/core/product/56468ECD9D9540FEC6B9E58C1FB5E22A.

Promel:1992:AAB

[PS92] Hans Jürgen Prömel and Angelika Steger. Almost all Berge graphs are perfect. Combinatorics, Probability and Computing, 1(1):53–79, March 1992. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Pach:1998:NIB

[PS98] János Pach and Micha Sharir. On the number of incidences between points and curves. Combinatorics, Probability and Computing, 7(1):121–127, March 1998. CODEN CPCOFG.

ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=7&issueId=01.

Perarnau:2013:RPM

[PS13] Guillem Perarnau and Oriol Serra. Rainbow perfect matchings in complete bipartite graphs: Existence and counting. Combinatorics, Probability and Computing, 22(5):783–799, September 2013. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Pittel:2016:STX

[PS16] Boris Pittel and Gregory B. Sorkin. The satisfiability threshold for k-XORSAT. Combinatorics, Probability and Computing, 25 (2):236–268, March 2016. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Piga:2023:CTD

[PSS23] Simón Piga, Marcelo Sales, and Bjarne Schülke. The codegree Turán density of tight cycles minus one edge. Combinatorics, Probability and Computing, 32(6):881–884, November 2023. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/codegree-
turan-density-of-tight-cycles-minus-one-edge/3D058F73B8050A7B96E1BEC53E4253A4.

Panagiotou:2013:EPE

[PSST13] Konstantinos Panagiotou, Reto Spöhel, Angelika Steger, and Henning Thomas. Explosive percolation in Erdős–Rényi-like random graph processes. Combinatorics, Probability and Computing, 22(1):133–145, ???? 2013. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Promel:2001:CPO

[PST01] Hans Jürgen Prömel, Angelika Steger, and Anusch Taraz. Counting partial orders with a fixed number of comparable pairs. Combinatorics, Probability and Computing, 10(2):159–177, March 2001. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=10&issueId=02.

Peled:2013:LFE

[PSY13] Ron Peled, Wojciech Samotij, and Amir Yehudayoff. Lipschitz functions on expanders are typically flat. Combinatorics, Probability and Computing, 22(4):566–591, July 2013. CODEN CP-COFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Pach:2009:CFC

[PT09] János Pach and Gábor Tardos. Conflict-free colourings of graphs and hypergraphs. Combinatorics, Probability and Computing, 18 (5):819–834, September 2009. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Perrett:2018:DCR

[PT18] Thomas J. Perrett and Carsten Thomassen. Density of chromatic roots in minor-closed graph families. Combinatorics, Probability and Computing, 27(6):988–998, November 2018. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/density-
of-chromatic-roots-in-minorclosed-graph-families/8C481CED690A7405B98F3B679F946C46.

Pach:2021:DGS

[PTT21] János Pach, Gábor Tardos, and Géza Tóth. Disjointness graphs of segments in the space. Combinatorics, Probability and Computing, 30(4):498–512, July 2021. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/disjointness-
graphs-of-segments-in-the-space/6406E2D53C3A30F2172234ADE22AEA66.

Pelc:1998:RFD

[PU98] Andrzej Pelc and Eli Upfal. Reliable fault diagnosis with few tests. Combinatorics, Probability and Computing, 7(3): 323–333, September 1998. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http://journals.

cambridge.org/action/displayIssue?jid=CPC&volumeId=7&
issueId=03.

Pudlak:2003:AHT

[Pud03] Pavel Pudlák. An application of Hindman’s theorem to a problem on communication complexity. Combinatorics, Probability and Computing, 12(5–6):661–670, November 2003. CODEN CP-COFG. ISSN 0963-5483 (print), 1469-2163 (electronic). Special issue on Ramsey theory.

Pikhurko:2013:MNC

[PV13] Oleg Pikhurko and Emil R. Vaughan. Minimum number of kcliques in graphs with bounded independence number. Combinatorics, Probability and Computing, 22(6):910–934, November 2013. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Poblete:2019:ARH

[PV19] P. V. Poblete and A. Viola. Analysis of Robin Hood and other hashing algorithms under the random probing model, with and without deletions. Combinatorics, Probability and Computing, 28(4):600–617, July 2019. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/analysis-
of-robin-hood-and-other-hashing-algorithms-under-the-
random-probing-model-with-and-without-deletions/933D4F203E3C70EF15053287412242E0.

Pemantle:2004:AMS

[PW04] Robin Pemantle and Mark C. Wilson. Asymptotics of multivariate sequences II: Multiple points of the singular variety. Combinatorics, Probability and Computing, 13(4–5):735–761, July 2004. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=13&issueId=04.

Pach:2016:DMP

[PW16] JÁNos Pach and Bartosz Walczak. Decomposition of multiple packings with subquadratic union complexity. Combinatorics, Probability and Computing, 25(1):145–153, January 2016. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). Special issue on Oberwolfach Special Issue Part 2.

Panova:2017:PFS

[PW17] Greta Panova and David B. Wilson. Pfaffian formulas for spanning tree probabilities. Combinatorics, Probability and Computing, 26(1):118–137, January 2017. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/product/
B0A5D209E60D2AC0E3ACA4233C21279F.

Pronzato:1997:SAC

[PWZ97] L. Pronzato, H. P. Wynn, and A. A. Zhigljavsky. Stochastic analysis of convergence via dynamic representation for a class of line-search algorithms. Combinatorics, Probability and Computing, 6(2):205–229, June 1997. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http://journals.

cambridge.org/action/displayIssue?jid=CPC&volumeId=6&
issueId=02.

Qin:2001:CMS

[Qin01] Hongxun Qin. Connected matroids with symmetric Tutte polynomials. Combinatorics, Probability and Computing, 10(2):179– 186, March 2001. CODEN CPCOFG. ISSN 0963-5483 (print),

1469-2163 (electronic). URL http://journals.cambridge.
org/action/displayIssue?jid=CPC&volumeId=10&issueId=
02.

Qin:2009:REM

[QSZ09] Hongxun Qin, Daniel C. Slilaty, and Xianqqian Zhou. The regular excluded minors for signed-graphic matroids. Combinatorics, Probability and Computing, 18(6):953–978, November 2009. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Rassmann:2019:NSR

[Ras19] Felicia Rassmann. On the number of solutions in random graph k-colouring. Combinatorics, Probability and Computing, 28(1): 130–158, January 2019. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL https://www.cambridge.

org/core/journals/combinatorics-probability-and-computing/
article/on-the-number-of-solutions-in-random-graph-kcolouring/
44C61122BBA4F9C17BD4C9E8ECE9CAD8.

Razborov:2008:MDT

[Raz08] Alexander A. Razborov. On the minimal density of triangles in graphs. Combinatorics, Probability and Computing, 17(4):603– 618, July 2008. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Raz:2020:NDD

[Raz20] Orit E. Raz. A note on distinct distances. Combinatorics, Probability and Computing, 29(5):650–663, September 2020.

CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/note-
on-distinct-distances/2C4D32DA9558DB9F61E8FC5D9FE418F5.

Reed:1996:PSN

[Ree96] Bruce Reed. Paths, stars and the number three. Combinatorics, Probability and Computing, 5(3):277–295, September 1996. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Reimer:2000:PVB

[Rei00] David Reimer. Proof of the Van den Berg–Kesten conjecture. Combinatorics, Probability and Computing, 9(1):27–32, January 2000. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=9&issueId=01.

Reidys:2002:DRI

[Rei02] C. M. Reidys. Distances in random induced subgraphs of generalized n-cubes. Combinatorics, Probability and Computing, 11(6):599–605, November 2002. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=11&issueId=06.

Reimer:2003:ASS

[Rei03] David Reimer. An average set size theorem. Combinatorics, Probability and Computing, 12(1):89–93, January 2003. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=12&issueId=01.

Renlund:2010:FPP

[Ren10] Henrik Renlund. First-passage percolation with exponential times on a ladder. Combinatorics, Probability and Computing, 19(4):593–601, July 2010. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Riordan:2000:SSR

[Rio00] Oliver Riordan. Spanning subgraphs of random graphs. Combinatorics, Probability and Computing, 9(2):125–148, March 2000. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=9&issueId=02.

Riordan:2005:SGC

[Rio05] Oliver Riordan. The small giant component in scale-free random graphs. Combinatorics, Probability and Computing, 14(5–6):897–938, November 2005. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=14&issueId=05.

Riordan:2008:CBP

[Rio08] Oliver Riordan. The k-core and branching processes. Combinatorics, Probability and Computing, 17(1):111–136, January 2008. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=17&issueId=01.

Riordan:2012:PTC

[Rio12] Oliver Riordan. The phase transition in the configuration model. Combinatorics, Probability and Computing, 21(1–2):265– 299, ???? 2012. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Reed:1992:SCC

[RM92] B. Reed and Colin McDiarmid. The strongly connected components of 1-in, 1-out. Combinatorics, Probability and Computing, 1(3):265–274, September 1992. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Roberto:2003:PML

[Rob03] Cyril Roberto. A path method for the logarithmic Sobolev constant. Combinatorics, Probability and Computing, 12(4):431–455, July 2003. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL http://journals.cambridge.org/
action/displayIssue?jid=CPC&volumeId=12&issueId=04.

Ross:2009:SSS

[Ros09] Nathan Ross. Step size in Stein’s method of exchangeable pairs. Combinatorics, Probability and Computing, 18(6):979– 1017, November 2009. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Rodl:1997:BCG

[RR97] Vojtech Rödl and Andrzej Ruciński. Bipartite coverings of graphs. Combinatorics, Probability and Computing, 6(3): 349–352, September 1997. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http://journals.

cambridge.org/action/displayIssue?jid=CPC&volumeId=6&
issueId=03.

Rahmann:2003:DNM

[RR03] Sven Rahmann and Eric Rivals. On the distribution of the number of missing words in random texts. Combinatorics, Probability and Computing, 12(1):73–87, January 2003. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=12&issueId=01.

Rasendrahasina:2019:EMB

[RRR19] V. Rasendrahasina, A. Rasoanaivo, and V. Ravelomanana. Expected maximum block size in critical random graphs. Combinatorics, Probability and Computing, 28(4):638–655, July 2019. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/expected-
maximum-block-size-in-critical-random-graphs/A3F859EA3B9AAB85D7AC299C90442CAD.

Rodl:2006:DTT

[RRS06] Vojtech Rödl, Andrzej Rucinski, and Endre Szemerédi. A Dirac-type theorem for 3-uniform hypergraphs. Combinatorics, Probability and Computing, 15(1–2):229–251, January 2006. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=15&issueId=01.

Rodl:1999:HPG

[RRT99] V. Rödl, A. Ruciński, and A. Taraz. Hypergraph packing and graph embedding. Combinatorics, Probability and Computing, 8(4):363–376, July 1999. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http://journals.

cambridge.org/action/displayIssue?jid=CPC&volumeId=8&

issueId=04. Random graphs and combinatorial structures (Oberwolfach, 1997).

Riordan:2000:MDR

[RS00] Oliver Riordan and Alex Selby. The maximum degree of a random graph. Combinatorics, Probability and Computing, 9 (6):549–572, November 2000. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http://journals.

cambridge.org/action/displayIssue?jid=CPC&volumeId=9&
issueId=06.

Rousseau:2003:MRN

[RS03] C. C. Rousseau and S. E. Speed. Mixed Ramsey numbers revisited. Combinatorics, Probability and Computing, 12(5–6): 653–660, November 2003. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). Special issue on Ramsey theory.

Rodl:2007:RPHb

[RS07a] Vojtech Rödl and Mathias Schacht. Regular partitions of hypergraphs: Counting lemmas. Combinatorics, Probability and Computing, 16(6):887–901, November 2007. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=16&issueId=06.

Rodl:2007:RPHa

[RS07b] Vojtech Rödl and Mathias Schacht. Regular partitions of hypergraphs: Regularity lemmas. Combinatorics, Probability and Computing, 16(6):833–885, November 2007. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=16&issueId=06.

Rao:2018:LBB

[RS18] Shravas Rao and Igor Shinkar. On Lipschitz bijections between Boolean functions. Combinatorics, Probability and Computing, 27(3):411–426, May 2018. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL https://www.cambridge.

org/core/journals/combinatorics-probability-and-computing/
article/on-lipschitz-bijections-between-boolean-functions/
58FB13FAA4D02D5BDA9692680E2A5B48.

Rodl:2022:BVV

[RS22] Vojtech Rödl and Marcelo Sales. A blurred view of Van der Waerden type theorems. Combinatorics, Probability and Computing, 31(4):684–701, July 2022. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL https://www.cambridge.

org/core/journals/combinatorics-probability-and-computing/
article/blurred-view-of-van-der-waerden-type-theorems/
23A6C57B88AE5A62A353778612E1B460.

Remy:2007:OST

[RSS07] Jan Remy, Alexander Souza, and Angelika Steger. On an online spanning tree problem in randomly weighted graphs. Combinatorics, Probability and Computing, 16(1):127–144, January 2007. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=16&issueId=01.

Rivera:2023:MRW

[RSS23] Nicolás Rivera, Thomas Sauerwald, and John Sylvester. Multiple random walks on graphs: mixing few to cover many. Combinatorics, Probability and Computing, 32(4):594–637, July 2023. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/multiple-
random-walks-on-graphs-mixing-few-to-cover-many/22B76C31532C71701E21D60A598012A1.

Razborov:1993:CSS

[RSW93] A. Razborov, E. Szemerédi, and A. Wigderson. Constructing small sets that are uniform in arithmetic progressions. Combinatorics, Probability and Computing, 2(4):513–518, December 1993. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Rodl:2005:CGD

[RT05] V. Rödl and L. Thoma. On cover graphs and dependent arcs in acyclic orientations. Combinatorics, Probability and Computing, 14(4):585–617, July 2005. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=14&issueId=04.

Russell:2012:CPI

[Rus12] Paul A. Russell. Compressions and probably intersecting families. Combinatorics, Probability and Computing, 21(1–2):301– 313, ???? 2012. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Rucinski:1992:RGP

[RW92] A. Ruciński and N. C. Wormald. Random graph processes with degree restrictions. Combinatorics, Probability and Computing, 1(2):169–180, June 1992. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Robalewska:2000:RSP

[RW00] H. D. Robalewska and N. C. Wormald. Random star processes. Combinatorics, Probability and Computing, 9(1):33–43, January 2000. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=9&issueId=01.

Reid:2001:MCB

[RW01] Talmage James Reid and Haidong Wu. On minimally 3connected binary matroids. Combinatorics, Probability and Computing, 10(5):453–461, September 2001. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=10&issueId=05.

Riordan:2010:DSR

[RW10] Oliver Riordan and Nicholas Wormald. The diameter of sparse random graphs. Combinatorics, Probability and Computing, 19 (5–6):835–926, November 2010. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Russell:2013:PIF

[RW13] Paul A. Russell and Mark Walters. Probably intersecting families are not nested. Combinatorics, Probability and Computing, 22 (1):146–160, ???? 2013. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Reed:2016:FSMa

[RW16a] Bruce Reed and David R. Wood. Forcing a sparse minor. Combinatorics, Probability and Computing, 25(2):300–322, March 2016. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). See corrigendum [RW16b].

Reed:2016:FSMb

[RW16b] Bruce Reed and David R. Wood. ‘Forcing a sparse minor’ — corrigendum. Combinatorics, Probability and Computing, 25(2):

323, March 2016. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). See [RW16a].

Riordan:2016:CAP

[RW16c] Oliver Riordan and Lutz Warnke. Convergence of Achlioptas processes via differential equations with unique solutions. Combinatorics, Probability and Computing, 25(1):154–171, January 2016. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). Special issue on Oberwolfach Special Issue Part 2.

Ralaivaosaona:2019:CLT

[RW19] Dimbinaina Ralaivaosaona and Stephan Wagner. A central limit theorem for additive functionals of increasing trees. Combinatorics, Probability and Computing, 28(4):618–637, July 2019. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/central-
limit-theorem-for-additive-functionals-of-increasing-
trees/784149F341FD074D3636B6899ECB0779.

Saad:1996:FLA

[Saa96] Rachid Saad. Finding a longest alternating cycle in a 2-edge-coloured complete graph is in RP. Combinatorics, Probability and Computing, 5(3):297–306, September 1996. CODEN CP-COFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Salez:2016:IMR

[Sal16] Justin Salez. The interpolation method for random graphs with prescribed degrees. Combinatorics, Probability and Computing, 25(3):436–447, May 2016. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Samorodnitsky:2017:IFH

[Sam17] Alex Samorodnitsky. An inequality for functions on the Hamming cube. Combinatorics, Probability and Computing, 26(3): 468–480, May 2017. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL https://www.cambridge.

org/core/product/3770E7921E125EF4A73A413E396348CB.

Sanders:2008:NFT

[San08] T. Sanders. A note on Freiman’s Theorem in vector spaces. Combinatorics, Probability and Computing, 17(2):297–305, March 2008. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=17&issueId=02.

Sauer:2003:CVP

[Sau03] N. W. Sauer. Canonical vertex partitions. Combinatorics, Probability and Computing, 12(5–6):671–704, November 2003. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). Special issue on Ramsey theory.

Savicky:1998:CPS

[Sav98] Petr Savický. Complexity and probability of some Boolean formulas. Combinatorics, Probability and Computing, 7(4): 451–463, December 1998. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http://journals.

cambridge.org/action/displayIssue?jid=CPC&volumeId=7&
issueId=04.

Saloff-Coste:1995:IID

[SC95] L. Saloff-Coste. Isoperimetric inequalities and decay of iterated kernels for almost-transitive Markov chains. Combinatorics, Probability and Computing, 4(4):419–442, December 1995. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Schnorr:1994:BRL

[Sch94] C.-P. Schnorr. Block reduced lattice bases and successive minima. Combinatorics, Probability and Computing, 3(4):507–522, December 1994. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Schoen:1999:DUS

[Sch99] Tomasz Schoen. On the density of universal sum-free sets. Combinatorics, Probability and Computing, 8(3):277–280, May 1999. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=8&issueId=03.

Schwartz:2002:AGD

[Sch02] Gary K. Schwartz. On the automorphism groups of Dowling geometries. Combinatorics, Probability and Computing, 11(3):311– 321, May 2002. CODEN CPCOFG. ISSN 0963-5483 (print),

1469-2163 (electronic). URL http://journals.cambridge.
org/action/displayIssue?jid=CPC&volumeId=11&issueId=

Schachinger:2004:CSP

[Sch04] Werner Schachinger. Concentration of size and path length of tries. Combinatorics, Probability and Computing, 13(4–5):763– 793, July 2004. CODEN CPCOFG. ISSN 0963-5483 (print),

1469-2163 (electronic). URL http://journals.cambridge.
org/action/displayIssue?jid=CPC&volumeId=13&issueId=
04.

Schoen:2007:DC

[Sch07] Tomasz Schoen. Difference covers. Combinatorics, Probability and Computing, 16(5):775–787, September 2007. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic).  URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=16&issueId=05.

Schmutz:2011:PLI

[Sch11] Eric Schmutz. Period lengths for iterated functions. Combinatorics, Probability and Computing, 20(2):289–298, March 2011. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Schulke:2023:PDC

[Sch23] Bjarne Schülke. A pair degree condition for Hamiltonian cycles in 3-uniform hypergraphs. Combinatorics, Probability and Computing, 32(5):762–781, September 2023. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/pair-
degree-condition-for-hamiltonian-cycles-in-3uniform-
hypergraphs/EB22D077566A1ED830A2C646AC50BAF6.

Scott:1992:LIS

[Sco92] A. D. Scott. Large induced subgraphs with all degrees odd. Combinatorics, Probability and Computing, 1(4):335–349, December 1992. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Scott:2004:BRT

[Sco04] Alexander Scott. Book review: “Topics in Graph Automorphisms and Reconstruction” by Josef Lauri and Raffaele Scapellato, Cambridge University Press, 2003, 172 pp. £47.50; $65.00 hardback, £16.95; $23.00, paperback, ISBN 0-521-82151-7 (hardback), 0-521-52903-4 (paperback). Combinatorics, Probability and Computing, 13(1):121–122, January 2004. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=13&issueId=01.

Scott:2011:SRL

[Sco11] Alexander Scott. Szemerédi’s regularity lemma for matrices and sparse graphs. Combinatorics, Probability and Computing, 20(3): 455–466, May 2011. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Saloff-Coste:1996:CNG

[SCW96] Laurent Saloff-Coste and Wolfgang Woess. Computing norms of group-invariant transition operators. Combinatorics, Probability and Computing, 5(2):161–178, June 1996. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Sedgewick:2014:MPF

[Sed14] Robert Sedgewick. In memory of Philippe Flajolet. Combinatorics, Probability and Computing, 23(5):631–634, September 2014. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). Special issue on Honouring the Memory of Philippe Flajolet - Part 1.

Seginer:2000:ENR

[Seg00] Yoav Seginer. The expected norm of random matrices. Combinatorics, Probability and Computing, 9(2):149–166, March 2000. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=9&issueId=02.

Seysen:1999:MNO

[Sey99] M. Seysen. A measure for the non-orthogonality of a lattice basis. Combinatorics, Probability and Computing, 8(3):281–291, May 1999. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=8&issueId=03.

Stark:2002:ILR

[SGO02] Dudley Stark, A. Ganesh, and Neil O’Connell. Information loss in riffle shuffling. Combinatorics, Probability and Computing, 11(1):79–95, January 2002. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=11&issueId=01.

Sharir:2003:CST

[Sha03] Micha Sharir. The Clarkson–Shor technique revisited and extended. Combinatorics, Probability and Computing, 12(2):191– 201, March 2003. CODEN CPCOFG. ISSN 0963-5483 (print),

1469-2163 (electronic). URL http://journals.cambridge.
org/action/displayIssue?jid=CPC&volumeId=12&issueId=
02.

Sharir:2011:IBS

[Sha11] Micha Sharir. An improved bound for k-sets in four dimensions. Combinatorics, Probability and Computing, 20(1):119–129, January 2011. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Shakan:2016:SMD

[Sha16] George Shakan. Sum of many dilates. Combinatorics, Probability and Computing, 25(3):460–469, May 2016. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Shalev:2017:CMC

[Sha17a] Aner Shalev. Corrigendum to ‘Mixing, Communication Complexity and Conjectures of Gowers and Viola’. Combinatorics, Probability and Computing, 26(6):954–955, November 2017. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/product/
140F12D529CF255A617C2ED3D1775D50. See [Sha17b].

Shalev:2017:MCC

[Sha17b] Aner Shalev. Mixing, communication complexity and conjectures of Gowers and Viola. Combinatorics, Probability and Computing, 26(4):628–640, July 2017. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL https://www.cambridge.

org/core/product/5D1FBB5807440C48E5267B69A499A9EA. See

corrigendum [Sha17a].

Shiraishi:1998:MCL

[Shi98] Shuji Shiraishi. On the maximum cut of line graphs. Combinatorics, Probability and Computing, 7(3):335–337, September 1998. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=7&issueId=03.

Siegrist:1998:EVE

[Sie98] Kyle Siegrist. Expected value expansions in random subgraphs with applications to network reliability. Combinatorics, Probability and Computing, 7(4):465–483, December 1998. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=7&issueId=04.

Sinclair:1992:IBM

[Sin92] Alistair Sinclair. Improved bounds for mixing rates of Markov chains and multicommodity flow. Combinatorics, Probability and Computing, 1(4):351–370, December 1992. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Sivakoff:2014:SPD

[Siv14] David Sivakoff. Site percolation on the d-dimensional Hamming torus. Combinatorics, Probability and Computing, 23(2):290– 315, March 2014. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Skrekovski:1999:GTT

[Škr99a] R. Škrekovski. A Grötzsch-type theorem for list colourings with impropriety one. Combinatorics, Probability and Computing, 8 (5):493–507, September 1999. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http://journals.

cambridge.org/action/displayIssue?jid=CPC&volumeId=8&
issueId=05.

Skrekovski:1999:LIC

[Škr99b] R. Škrekovski. List improper colourings of planar graphs. Combinatorics, Probability and Computing, 8(3):293–299, May 1999. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=8&issueId=03.

Skupien:1996:SSL

[Sku96] Z. Skupień. Smallest sets of longest paths with empty intersection. Combinatorics, Probability and Computing, 5(4):429–436, December 1996. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Slivnik:1996:SPG

[Sli96] Tomaž Slivnik. Short proof of Galvin’s theorem on the listchromatic index of a bipartite multigraph. Combinatorics, Probability and Computing, 5(1):91–94, March 1996. CODEN CP-COFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Slilaty:2002:MDT

[Sli02] Daniel C. Slilaty. Matroid duality from topological duality in surfaces of nonnegative Euler characteristic. Combinatorics, Probability and Computing, 11(5):515–528, September 2002. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=11&issueId=05.

Smith:1993:ORG

[Smi93] D. H. Smith. Optimally reliable graphs for both vertex and edge failures. Combinatorics, Probability and Computing, 2(1):93–100, March 1993. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Smyth:2013:RIF

[Smy13] Clifford Smyth. Reimer’s inequality on a finite distributive lattice. Combinatorics, Probability and Computing, 22(4):612–621, July 2013. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Soberon:2018:RTC

[Sob18] Pablo Soberón. Robust Tverberg and colourful Carathéodory results via random choice. Combinatorics, Probability and Computing, 27(3):427–440, May 2018. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL https://www.

cambridge.org/core/journals/combinatorics-probability-
and-computing/article/robust-tverberg-and-colourful-
caratheodory-results-via-random-choice/0F924628DD8F0CF7F40C685A6200E9D8.

Sokal:2001:BCZ

[Sok01] Alan D. Sokal. Bounds on the complex zeros of (di)chromatic polynomials and Potts-model partition functions. Combinatorics, Probability and Computing, 10(1):41–77, January 2001. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=10&issueId=01.

Sokal:2004:CRD

[Sok04] Alan D. Sokal. Chromatic roots are dense in the whole complex plane. Combinatorics, Probability and Computing, 13(2):221– 261, March 2004. CODEN CPCOFG. ISSN 0963-5483 (print),

1469-2163 (electronic). URL http://journals.cambridge.
org/action/displayIssue?jid=CPC&volumeId=13&issueId=
02.

Solymosi:2004:NQE

[Sol04] J. Solymosi. A note on a question of Erdős and Graham. Combinatorics, Probability and Computing, 13(2):263–267, March 2004. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=13&issueId=02.

Solymosi:2006:APS

[Sol06] József Solymosi. Arithmetic progressions in sets with small sumsets. Combinatorics, Probability and Computing, 15(4):597–603, July 2006. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL http://journals.cambridge.org/
action/displayIssue?jid=CPC&volumeId=15&issueId=04.

Solymosi:2015:CFG

[Sol15] József Solymosi. The (7, 4)-conjecture in finite groups. Combinatorics, Probability and Computing, 24(4):680–686, July 2015. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). Special issue on Oberwolfach Special Issue Part 1.

Specker:2005:MCS

[Spe05] E. Specker. Modular counting and substitution of structures. Combinatorics, Probability and Computing, 14(1–2):203–210, January 2005. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL http://journals.cambridge.org/
action/displayIssue?jid=CPC&volumeId=14&issueId=01.

Spink:2016:LMQ

[Spi16] Hunter Spink. Local maxima of quadratic Boolean functions. Combinatorics, Probability and Computing, 25(4):633–640, July 2016. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Spiro:2023:SNS

[Spi23] Sam Spiro. A smoother notion of spread hypergraphs. Combinatorics, Probability and Computing, 32(5):809–818, September 2023. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL https://www.cambridge.org/core/
journals/combinatorics-probability-and-computing/article/
smoother-notion-of-spread-hypergraphs/F086D4773B6655E62CF5AB462D2181FC.

Schwerdtfeger:2010:ALL

[SRT10] U. Schwerdtfeger, C. Richard, and B. Thatte. Area limit laws for symmetry classes of staircase polygons. Combinatorics, Probability and Computing, 19(3):441–461, May 2010. CODEN CP-COFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Shahrokhi:1994:CCF

[SS94] Farhad Shahrokhi and László A. Székely. On canonical concurrent flows, crossing number and graph expansion. Combinatorics, Probability and Computing, 3(4):523–543, December 1994. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Sali:1997:RSS

[SS97] Attila Sali and Gábor Simonyi. Recovering set systems and graph entropy. Combinatorics, Probability and Computing, 6 (4):481–491, December 1997. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http://journals.

cambridge.org/action/displayIssue?jid=CPC&volumeId=6&
issueId=04.

Simonovits:2003:HEP

[SS03] Miklós Simonovits and Vera T. Sós. Hereditary extended properties, quasi-random graphs and induced subgraphs. Combinatorics, Probability and Computing, 12(3):319–344, May 2003. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=12&issueId=03. Combina-

torics, probability and computing (Oberwolfach, 2001).

Smorodinsky:2004:SPH

[SS04] Shakhar Smorodinsky and Micha Sharir. Selecting points that are heavily covered by pseudo-circles, spheres or rectangles.

Combinatorics, Probability and Computing, 13(3):389–411, May 2004. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=13&issueId=03.

Scott:2006:DGL

[SS06a] Alexander D. Scott and Alan D. Sokal. On dependency graphs and the lattice gas. Combinatorics, Probability and Computing, 15(1–2):253–279, January 2006. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=15&issueId=01.

Scott:2006:SSR

[SS06b] Alexander D. Scott and Gregory B. Sorkin. Solving sparse random instances of max cut and max 2-CSP in linear expected time. Combinatorics, Probability and Computing, 15(1–2):281–315, January 2006. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL http://journals.cambridge.org/
action/displayIssue?jid=CPC&volumeId=15&issueId=01.

Steel:2007:TAT

[SS07] M. A. Steel and L. A. Székely. Teasing apart two trees. Combinatorics, Probability and Computing, 16(6):903–922, November 2007. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=16&issueId=06.

Schoen:2011:SCS

[SS11] Tomasz Schoen and Ilya D. Shkredov. On sumsets of convex sets. Combinatorics, Probability and Computing, 20(5):793–798, September 2011. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Sharir:2013:CPG

[SS13] Micha Sharir and Adam Sheffer. Counting plane graphs: Crossgraph charging schemes. Combinatorics, Probability and Computing, 22(6):935–954, November 2013. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Skokan:2014:CSR

[SS14] J. Skokan and M. Stein. Cycles are strongly Ramsey-unsaturated. Combinatorics, Probability and Computing, 23(4):607–630, July 2014. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Sharir:2016:DDT

[SS16] Micha Sharir and József Solymosi. Distinct distances from three points. Combinatorics, Probability and Computing, 25(4):623– 632, July 2016. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Sapir:2020:IRL

[SS20] Shachar Sapir and Asaf Shapira. The induced removal lemma in sparse graphs. Combinatorics, Probability and Computing, 29 (1):153–162, January 2020. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL https://www.cambridge.

org/core/journals/combinatorics-probability-and-computing/
article/induced-removal-lemma-in-sparse-graphs/651B476F7AFFB50ED7B62C44522CC09D.

Salvy:2011:PFD

[SSS^{+}11] Bruno Salvy, Bob Sedgewick, Michele Soria, Wojciech Szpankowski, and Brigitte Vallée. Philippe Flajolet 1 December 1948–22 March 2011. Combinatorics, Probability and Computing, 20(5):647–649, September 2011. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Sah:2021:CBR

[SSTZ21] Ashwin Sah, Mehtaab Sawhney, Jonathan Tidor, and Yufei Zhao. A counterexample to the Bollobás–Riordan conjectures on sparse graph limits. Combinatorics, Probability and Computing, 30(5):796–799, September 2021. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/counterexample-
to-the-bollobasriordan-conjectures-on-sparse-graph-limits/
5D13139270D258DCE816FB1EAFEB36F8.

Sharir:2015:IBI

[SSZ15] Micha Sharir, Adam Sheffer, and Joshua Zahl. Improved bounds for incidences between points and circles. Combinatorics, Probability and Computing, 24(3):490–520, May 2015. CODEN CP-COFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Shawe-Taylor:1996:FSM

[ST96] John Shawe-Taylor. Fast string matching in stationary ergodic sources. Combinatorics, Probability and Computing, 5(4):415– 427, December 1996. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Schelp:1998:RNC

[ST98] Richard H. Schelp and Andrew Thomason. A remark on the number of complete and empty subgraphs. Combinatorics, Probability and Computing, 7(2):217–219, June 1998. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=7&issueId=02.

Szabo:2001:MGE

[ST01] Tibor Szabó and Gábor Tardos. A multidimensional generalization of the Erdős–Szekeres lemma on monotone subsequences. Combinatorics, Probability and Computing, 10(6):557–565, November 2001. CODEN CPCOFG. ISSN 0963-5483 (print),

1469-2163 (electronic). URL http://journals.cambridge.
org/action/displayIssue?jid=CPC&volumeId=10&issueId=
06.

Schlaghoff:2005:IRC

[ST05] Jens Schlaghoff and Eberhard Triesch. Improved results for competitive group testing. Combinatorics, Probability and Computing, 14(1–2):191–202, January 2005. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=14&issueId=01.

Solymosi:2008:QBA

[ST08] József Solymosi and Csaba D. Tóth. On a question of Bourgain about geometric incidences. Combinatorics, Probability and Computing, 17(4):619–625, July 2008. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Sammer:2009:CDT

[ST09] M. Sammer and P. Tetali. Concentration on the discrete torus using transportation. Combinatorics, Probability and Computing, 18(5):835–860, September 2009. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Saxton:2012:LCR

[ST12] David Saxton and Andrew Thomason. List colourings of regular hypergraphs. Combinatorics, Probability and Computing, 21(1– 2):315–322, ???? 2012. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Saxton:2016:SCS

[ST16] David Saxton and Andrew Thomason. Simple containers for simple hypergraphs. Combinatorics, Probability and Computing, 25 (3):448–459, May 2016. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Stark:1995:FOP

[Sta95] Dudley Stark. First occurrence in pairs of long words: a Penneyante conjecture of Pevzner. Combinatorics, Probability and Computing, 4(3):279–285, September 1995. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Stark:1997:ELT

[Sta97] Dudley Stark. Explicit limits of total variation distance in approximations of random logarithmic assemblies by related Poisson processes. Combinatorics, Probability and Computing, 6(1):87–105, March 1997. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http://journals.

cambridge.org/action/displayIssue?jid=CPC&volumeId=6&
issueId=01.

Stacey:1999:SBT

[Sta99a] Alan Stacey. Searches on a binary tree with random edgeweights. Combinatorics, Probability and Computing, 8(6): 555–565, November 1999. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http://journals.

cambridge.org/action/displayIssue?jid=CPC&volumeId=8&
issueId=06.

Stark:1999:TVA

[Sta99b] Dudley Stark. Total variation asymptotics for refined Poisson process approximations of random logarithmic assemblies. Combinatorics, Probability and Computing, 8(6):567–598, November 1999. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=8&issueId=06.

Stadje:2002:RMP

[Sta02a] Wolfgang Stadje. The residues modulo m of products of random integers. Combinatorics, Probability and Computing, 11(5):529–540, September 2002. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=11&issueId=05.

Stark:2002:ILT

[Sta02b] Dudley Stark. Information loss in top to random shuffling. Combinatorics, Probability and Computing, 11(6):607–627, November 2002. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=11&issueId=06.

Stacey:2003:BRW

[Sta03] Alan Stacey. Branching random walks on quasi-transitive graphs. Combinatorics, Probability and Computing, 12(3):345–358, May 2003. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=12&issueId=03. Combina-

torics, probability and computing (Oberwolfach, 2001).

Steger:2005:ETF

[Ste05a] Angelika Steger. On the evolution of triangle-free graphs. Combinatorics, Probability and Computing, 14(1–2):211–224, January 2005. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=14&issueId=01.

Steif:2005:BRP

[Ste05b] Jeffrey E. Steif. Book review: “Probability on Discrete Structures”, edited by H. Kesten, Encyclopaedia of Mathematical Sciences, Vol. 110, Springer 2004, 351 pp., ISBN 3-540-00845-4, EU 94.95/£73.00/$109.00. Combinatorics, Probability and Computing, 14(4):619–623, July 2005. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=14&issueId=04.

Steiner:2022:ILB

[Ste22] Raphael Steiner. Improved lower bound for the list chromatic number of graphs with no K_{t} minor. Combinatorics, Probability and Computing, 31(6):1070–1075, November 2022. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/improved-
lower-bound-for-the-list-chromatic-number-of-graphs-
with-no-kt-minor/8132D67EC14954FEFEF399184D7C9A8B.

Steiner:2023:IBI

[Ste23] Raphael Steiner. Improved bound for improper colourings of graphs with no odd clique minor. Combinatorics, Probability and Computing, 32(2):326–333, March 2023. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/improved-
bound-for-improper-colourings-of-graphs-with-no-odd-
clique-minor/D1379954D4625F660F6C935A2B6DFB72.

Stufler:2020:UGP

[Stu20] Benedikt Stufler. Unlabelled Gibbs partitions. Combinatorics, Probability and Computing, 29(2):293–309, March 2020. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/unlabelled-
gibbs-partitions/64D5647200F23911B839DC7E0DDB7C64.

Subramanian:2000:ACR

[Sub00] C. R. Subramanian. Algorithms for colouring random kcolourable graphs. Combinatorics, Probability and Computing, 9(1):45–77, January 2000. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http://journals.

cambridge.org/action/displayIssue?jid=CPC&volumeId=9&
issueId=01.

Subramanian:2007:LSC

[Sub07] C. R. Subramanian. List set colouring: Bounds and algorithms. Combinatorics, Probability and Computing, 16(1):145–158, January 2007. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL http://journals.cambridge.org/
action/displayIssue?jid=CPC&volumeId=16&issueId=01.

Sanders:2003:TBE

[SV03] Peter Sanders and Berthold Vöcking. Tail bounds and expectations for random arc allocation and applications. Combinatorics, Probability and Computing, 12(3):301–318, May 2003. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=12&issueId=03. Combina-

torics, probability and computing (Oberwolfach, 2001).

Spencer:1992:TTL

[SW92] Joel Spencer and Peter Winkler. Three thresholds for a liar. Combinatorics, Probability and Computing, 1(1):81–93, March 1992. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Steger:1999:GRR

[SW99] A. Steger and N. C. Wormald. Generating random regular graphs quickly. Combinatorics, Probability and Computing, 8(4):377–396, July 1999. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http://journals.

cambridge.org/action/displayIssue?jid=CPC&volumeId=8&

issueId=04. Random graphs and combinatorial structures (Oberwolfach, 1997).

Sharir:2004:PLI

[SW04] Micha Sharir and Emo Welzl. Point–line incidences in space. Combinatorics, Probability and Computing, 13(2):203–220, March 2004. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL http://journals.cambridge.org/
action/displayIssue?jid=CPC&volumeId=13&issueId=02.

Shi:2007:CRRa

[SW07a] Lingsheng Shi and Nicholas Wormald. Colouring random 4-regular graphs. Combinatorics, Probability and Computing, 16(2):309–344, March 2007. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=16&issueId=02.

Shi:2007:CRRb

[SW07b] Lingsheng Shi and Nicholas Wormald. Colouring random regular graphs. Combinatorics, Probability and Computing, 16(3):459– 494, May 2007. CODEN CPCOFG. ISSN 0963-5483 (print),

1469-2163 (electronic). URL http://journals.cambridge.
org/action/displayIssue?jid=CPC&volumeId=16&issueId=

Semple:2008:NCG

[SW08] Charles Semple and Dominic Welsh. Negative correlation in graphs and matroids. Combinatorics, Probability and Computing, 17(3):423–435, May 2008. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=17&issueId=03.

Sousi:2014:MTM

[SW14] Perla Sousi and Peter Winkler. Mixing times and moving targets. Combinatorics, Probability and Computing, 23(3):460–476, May 2014. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Stark:2018:PNE

[SW18] Dudley Stark and Nick Wormald. The probability of nonexistence of a subgraph in a moderately sparse random graph. Combinatorics, Probability and Computing, 27(4):672–715, July 2018. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL https://www.cambridge.org/core/
journals/combinatorics-probability-and-computing/article/
probability-of-nonexistence-of-a-subgraph-in-a-moderately-
sparse-random-graph/B654A9E2D34901659E29D31B6D5E7663.

Szuster:1996:RSR

[SWŻ96] Janusz Szuster, Paweł Wlaź, and Jerzy Żurawiecki. On recognition of shift registers. Combinatorics, Probability and Computing, 5(3):307–315, September 1996. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Szekely:1997:CNH

[Szé97] László A. Székely. Crossing numbers and hard Erdős problems in discrete geometry. Combinatorics, Probability and Computing, 6 (3):353–358, September 1997. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http://journals.

cambridge.org/action/displayIssue?jid=CPC&volumeId=6&
issueId=03.

Tanushev:1997:NDE

[TA97] Miroslav Tanushev and Richard Arratia. A note on distributional equality in the cyclic tour property for Markov chains. Combinatorics, Probability and Computing, 6(4):493–496, December 1997. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=6&issueId=04.

Talbot:2002:LH

[Tal02] J. M. Talbot. Lagrangians of hypergraphs. Combinatorics, Probability and Computing, 11(2):199–216, March 2002. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic).  URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=11&issueId=02.

Tao:2010:SIS

[Tao10] Terence Tao. Sumset and inverse sumset theory for Shannon entropy. Combinatorics, Probability and Computing, 19(4):603– 639, July 2010. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Telcs:2002:NRI

[Tel02] András Telcs. A note on rough isometry invariance of resistance. Combinatorics, Probability and Computing, 11(4):427–432, July 2002. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=11&issueId=04.

Tetali:1994:EFN

[Tet94] Prasad Tetali. An extension of Foster’s network theorem. Combinatorics, Probability and Computing, 3(3):421–427, September 1994. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Teufl:2003:ADS

[Teu03] Elmar Teufl. The average displacement of the simple random walk on the Sierpiński graph. Combinatorics, Probability and Computing, 12(2):203–222, March 2003. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=12&issueId=02.

The:2004:AIM

[The04] Michel Nguyen The. Area and inertial moment of Dyck paths. Combinatorics, Probability and Computing, 13(4–5):697–716, July 2004. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL http://journals.cambridge.org/
action/displayIssue?jid=CPC&volumeId=13&issueId=04.

Thiel:2016:INM

[Thi16] Marko Thiel. The inversion number and the major index are asymptotically jointly normally distributed on words. Combinatorics, Probability and Computing, 25(3):470–483, May 2016. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Thomassen:1992:PCG

[Tho92] Carsten Thomassen. Plane cubic graphs with prescribed face areas. Combinatorics, Probability and Computing, 1(4):371–381, December 1992. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Thorup:1995:SPD

[Tho95] Mikkel Thorup. Shortcutting planar digraphs. Combinatorics, Probability and Computing, 4(3):287–315, September 1995. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Thomassen:1996:SG

[Tho96a] Carsten Thomassen. K_{5}-subdivisions in graphs. Combinatorics, Probability and Computing, 5(2):179–189, June 1996. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Thomassen:1996:NHC

[Tho96b] Carsten Thomassen. On the number of Hamiltonian cycles in bipartite graphs. Combinatorics, Probability and Computing, 5 (4):437–442, December 1996. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Thomassen:1997:ZFI

[Tho97] Carsten Thomassen. The zero-free intervals for chromatic polynomials of graphs. Combinatorics, Probability and Computing, 6(4):497–506, December 1997. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http://journals.

cambridge.org/action/displayIssue?jid=CPC&volumeId=6&
issueId=04.

Thomason:2004:TMP

[Tho04] Andrew Thomason. Two minor problems. Combinatorics, Probability and Computing, 13(3):413–414, May 2004. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic).  URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=13&issueId=03.

Timar:2007:CIG

[Tim07] Ádám Timár. Cutsets in infinite graphs. Combinatorics, Probability and Computing, 16(1):159–166, January 2007. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic).  URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=16&issueId=01.

Timar:2012:ACD

[Tim12] Ádám Timár. Approximating Cayley diagrams versus Cayley graphs. Combinatorics, Probability and Computing, 21(4):635– 641, July 2012. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Timar:2023:URO

[Tim23] Ádám Timár. Unimodular random one-ended planar graphs are sofic. Combinatorics, Probability and Computing, 32(6):851–858, November 2023. CODEN CPCOFG. ISSN 0963-5483 (print),

1469-2163 (electronic). URL https://www.cambridge.org/
core/journals/combinatorics-probability-and-computing/
article/unimodular-random-oneended-planar-graphs-are-
sofic/0C1BF8E92302BB0ED603B9FBA5FA7706.

Tokushige:2013:CII

[Tok13] Norihide Tokushige. Cross t-intersecting integer sequences from weighted Erdős–Ko–Rado. Combinatorics, Probability and Computing, 22(4):622–637, July 2013. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Traldi:2004:SEC

[Tra04] Lorenzo Traldi. A subset expansion of the coloured Tutte polynomial. Combinatorics, Probability and Computing, 13(2):269–275, March 2004. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL http://journals.cambridge.org/
action/displayIssue?jid=CPC&volumeId=13&issueId=02.

Traldi:2010:WIP

[Tra10] Lorenzo Traldi. Weighted interlace polynomials. Combinatorics, Probability and Computing, 19(1):133–157, January 2010. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Treglown:2015:DVH

[Tre15] Andrew Treglown. On directed versions of the Hajnal–Szemerédi theorem. Combinatorics, Probability and Computing, 24(6):873– 928, November 2015. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Thomas:2000:EMN

[TT00] Robin Thomas and Jan McDonald Thomson. Excluding minors in nonplanar graphs of girth at least five. Combinatorics, Probability and Computing, 9(6):573–585, November 2000. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=9&issueId=06.

Turova:2011:LCS

[Tur11] Tatyana S. Turova. The largest component in subcritical inhomogeneous random graphs. Combinatorics, Probability and Computing, 20(1):131–154, January 2011. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Toth:1997:TFE

[TW97] B. Tóth and W. Werner. Tied favourite edges for simple random walk. Combinatorics, Probability and Computing, 6(3): 359–369, September 1997. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http://journals.

cambridge.org/action/displayIssue?jid=CPC&volumeId=6&
issueId=03.

Trotter:1998:RTS

[TW98] William T. Trotter and Peter Winkler. Ramsey theory and sequences of random variables. Combinatorics, Probability and Computing, 7(2):221–238, June 1998. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=7&issueId=02.

Tyomkyn:2021:MDT

[Tyo21] Mykhaylo Tyomkyn. Many disjoint triangles in co-triangle-free graphs. Combinatorics, Probability and Computing, 30(1):153– 162, January 2021. CODEN CPCOFG. ISSN 0963-5483 (print),

1469-2163 (electronic). URL https://www.cambridge.org/
core/journals/combinatorics-probability-and-computing/
article/many-disjoint-triangles-in-cotrianglefree-graphs/
292CDBFEB9E252D3636177E54166ADAB.

Tyros:2015:CSV

[Tyr15] Konstantinos Tyros. Combinatorial structures on van der Waerden sets. Combinatorics, Probability and Computing, 24(6): 929–953, November 2015. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Tillich:2000:DII

[TZ00] Jean-Pierre Tillich and Gilles Zémor. Discrete isoperimetric inequalities and the probability of a decoding error. Combinatorics, Probability and Computing, 9(5):465–479, September 2000. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=9&issueId=05.

Tomon:2021:TTR

[TZ21] István Tomon and Dmitriy Zakharov. Turán-type results for intersection graphs of boxes. Combinatorics, Probability and Computing, 30(6):982–987, November 2021. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/turantype-
results-for-intersection-graphs-of-boxes/F2BC191D56DC97328ECA7E7785960DEA.

Uzzell:2019:IUB

[Uzz19] Andrew J. Uzzell. An improved upper bound for bootstrap percolation in all dimensions. Combinatorics, Probability and Computing, 28(6):936–960, November 2019. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL /core/journals/combinatorics-probability-
and-computing/article/an-improved-upper-bound-for-bootstrap-
percolation-in-all-dimensions/83A88608CA5334707D423B38535E58E0.

VanDenHeuvel:2001:AAC

[Van01] Jan Van Den Heuvel. Algorithmic aspects of a chip-firing game. Combinatorics, Probability and Computing, 10(6):505–529, November 2001. CODEN CPCOFG. ISSN 0963-5483 (print),

1469-2163 (electronic). URL http://journals.cambridge.
org/action/displayIssue?jid=CPC&volumeId=10&issueId=
06.

Vatter:2008:ESR

[Vat08] Vincent Vatter. Enumeration schemes for restricted permutations. Combinatorics, Probability and Computing, 17(1):137–159, January 2008. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL http://journals.cambridge.org/
action/displayIssue?jid=CPC&volumeId=17&issueId=01.

Verbitsky:1995:HAS

[Ver95] Oleg Verbitsky. On the hardness of approximating some optimization problems that are supposedly easier than MAX CLIQUE. Combinatorics, Probability and Computing, 4(2):167– 180, June 1995. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Verstraete:2000:APC

[Ver00] Jacques Verstraëte. On arithmetic progressions of cycle lengths in graphs. Combinatorics, Probability and Computing, 9 (4):369–373, July 2000. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http://journals.

cambridge.org/action/displayIssue?jid=CPC&volumeId=9&
issueId=04.

Verstraete:2002:NVD

[Ver02] Jacques Verstraëte. A note on vertex-disjoint cycles. Combinatorics, Probability and Computing, 11(1):97–102, January 2002. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=11&issueId=01.

VanDerHofstad:2006:SWS

[VHV06] Remco Van Der Hofstad, Gerard Hooghiemstra, and Piet Van Mieghem. Size and weight of shortest path trees with exponential link weights. Combinatorics, Probability and Computing, 15(6):903–926, November 2006. CODEN CPCOFG.

ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=15&issueId=06.

VanBatenburg:2018:PGB

[VK18] Wouter Cames Van Batenburg and Ross J. Kang. Packing graphs of bounded codegree. Combinatorics, Probability and Computing, 27(5):725–740, September 2018. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/packing-
graphs-of-bounded-codegree/95F60E36C02E097BF6C15DB74E6CA903.

Voigt:2000:AAP

[Voi00] M. Voigt. Algorithmic aspects of partial list colourings. Combinatorics, Probability and Computing, 9(4):375–380, July 2000. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=9&issueId=04.

VonZurGathen:2015:CDU

[Von15] Joachim Von Zur Gathen. Counting decomposable univariate polynomials. Combinatorics, Probability and Computing, 24(1): 294–328, January 2015. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). Special issue on Honouring the Memory of Philippe Flajolet - Part 3.

VanDerHofstad:2006:EBP

[VS06] Remco Van Der Hofstad and Gordon Slade. Expansion in n^{−1} for percolation critical values on the n-cube and Z^{n}: the first three terms. Combinatorics, Probability and Computing, 15(5):695–713, September 2006. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=15&issueId=05.

Vu:1997:TG

[Vu97] Van H. Vu. On a theorem of Ganter. Combinatorics, Probability and Computing, 6(2):247–254, June 1997. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=6&issueId=02.

Vu:2000:CNR

[Vu00] Van H. Vu. On the choice number of random hypergraphs. Combinatorics, Probability and Computing, 9(1):79–95, January 2000. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=9&issueId=01.

Vu:2001:LDR

[Vu01] Van H. Vu. A large deviation result on the number of small subgraphs of a random graph. Combinatorics, Probability and Computing, 10(1):79–94, January 2001. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=10&issueId=01.

Vu:2002:GUB

[Vu02] Van H. Vu. A general upper bound on the list chromatic number of locally sparse graphs. Combinatorics, Probability and Computing, 11(1):103–111, January 2002. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=11&issueId=01.

Vu:2018:SSA

[Vu18] Van Vu. A simple SVD algorithm for finding hidden partitions. Combinatorics, Probability and Computing, 27(1):124–140, January 2018. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL https://www.cambridge.org/core/
product/3691FFE262617493CEC469769154EA19.

Vogt:1997:SCD

[VV97] Frank Vogt and Bernd Voigt. Symmetric chain decompositions of linear lattices. Combinatorics, Probability and Computing, 6(2):231–245, June 1997. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http://journals.

cambridge.org/action/displayIssue?jid=CPC&volumeId=6&
issueId=02.

Vertigan:1992:CCT

[VW92] D. L. Vertigan and D. J. A. Welsh. The computational complexity of the Tutte plane: the bipartite case. Combinatorics, Probability and Computing, 1(2):181–187, June 1992. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Vertigan:1993:RPA

[VW93] Dirk Vertigan and Geoff Whittle. Recognizing polymatroids associated with hypergraphs. Combinatorics, Probability and Computing, 2(4):519–530, December 1993. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Voigt:2006:SMH

[VZ06] Thomas Voigt and Günter M. Ziegler. Singular 0/1-matrices, and the hyperplanes spanned by random 0/1-vectors. Combinatorics, Probability and Computing, 15(3):463–471, May 2006. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=15&issueId=03.

Wagner:2000:ZRP

[Wag00] David G. Wagner. Zeros of reliability polynomials and fvectors of matroids. Combinatorics, Probability and Computing, 9(2):167–190, March 2000. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http://journals.

cambridge.org/action/displayIssue?jid=CPC&volumeId=9&
issueId=02.

Wagner:2006:WCG

[Wag06a] Donald K. Wagner. Weakly 3-connected graphs. Combinatorics, Probability and Computing, 15(4):605–617, July 2006. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=15&issueId=04.

Wagner:2006:UBC

[Wag06b] Peter Wagner. An upper bound for constrained Ramsey numbers. Combinatorics, Probability and Computing, 15(4):619–626, July 2006. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=15&issueId=04.

Wagner:2008:TES

[Wag08] Roy Wagner. Tail estimates for sums of variables sampled by a random walk. Combinatorics, Probability and Computing, 17(2):307–316, March 2008. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=17&issueId=02.

Wagner:2015:CLT

[Wag15] Stephan Wagner. Central limit theorems for additive tree parameters with small toll functions. Combinatorics, Probability and Computing, 24(1):329–353, January 2015. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). Special issue on Honouring the Memory of Philippe Flajolet - Part 3.

Walters:2007:EPH

[Wal07] Mark Walters. Extensions of the polynomial Hales–Jewett theorem. Combinatorics, Probability and Computing, 16(5):789–803, September 2007. CODEN CPCOFG. ISSN 0963-5483 (print),

1469-2163 (electronic). URL http://journals.cambridge.
org/action/displayIssue?jid=CPC&volumeId=16&issueId=
05.

Wanless:2010:CMT

[Wan10] Ian. M. Wanless. Counting matchings and tree-like walks in regular graphs. Combinatorics, Probability and Computing, 19(3): 463–480, May 2010. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Warnke:2016:MTB

[War16] Lutz Warnke. On the method of typical bounded differences. Combinatorics, Probability and Computing, 25(2):269– 299, March 2016. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Wastlund:2011:BPA

[Wäs11] Johan Wästlund. The blind passenger and the assignment problem. Combinatorics, Probability and Computing, 20(3):467–480, May 2011. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Weineck:1997:LCS

[Wei97] Ekkehard Weineck. Log-concavity and the spectral gap of stochastic matrices. Combinatorics, Probability and Computing, 6(3):371–379, September 1997. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http://journals.

cambridge.org/action/displayIssue?jid=CPC&volumeId=6&
issueId=03.

Welsh:1994:RAT

[Wel94] D. J. A. Welsh. Randomised approximation in the Tutte plane. Combinatorics, Probability and Computing, 3(1):137–143, March 1994. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Watanabe:2011:NGP

[WF11] Y. Watanabe and K. Fukumizu. New graph polynomials from the Bethe approximation of the Ising Partition Function. Combinatorics, Probability and Computing, 20(2):299–320, March 2011. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Whittle:1992:DPS

[Whi92] Geoff Whittle. Duality in polymatroids and set functions. Combinatorics, Probability and Computing, 1(3):275–280, September 1992. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

White:1994:IRT

[Whi94] Arthur T. White. An introduction to random topological graph theory. Combinatorics, Probability and Computing, 3(4):545–555, December 1994. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Wierman:1992:EBP

[Wie92] John C. Wierman. Equality of the bond percolation critical exponents for two pairs of dual lattices. Combinatorics, Probability and Computing, 1(1):95–105, March 1992. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Wierman:1995:SMC

[Wie95] John C. Wierman. Substitution method critical probability bounds for the square lattice site percolation model. Combinatorics, Probability and Computing, 4(2):181–188, June 1995. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Wierman:2002:IUB

[Wie02] John C. Wierman. An improved upper bound for the hexagonal lattice site percolation critical probability. Combinatorics, Probability and Computing, 11(6):629–643, November 2002. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=11&issueId=06.

Wierman:2003:ULB

[Wie03] John C. Wierman. Upper and lower bounds for the Kagomé lattice bond percolation critical probability. Combinatorics, Probability and Computing, 12(1):95–111, January 2003. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic).  URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=12&issueId=01.

Wilson:2015:DAP

[Wil15] Mark C. Wilson. Diagonal asymptotics for products of combinatorial classes. Combinatorics, Probability and Computing, 24 (1):354–372, January 2015. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). Special issue on Honouring the Memory of Philippe Flajolet - Part 3.

Witt:2013:TBO

[Wit13] Carsten Witt. Tight bounds on the optimization time of a randomized search heuristic on linear functions. Combinatorics, Probability and Computing, 22(2):294–318, ???? 2013. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Weifan:2001:SPE

[WL01] Wang Weifan and Ko-Wei Lih. Structural properties and edge choosability of planar graphs without 6-cycles. Combinatorics, Probability and Computing, 10(3):267–276, May 2001. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=10&issueId=03.

Woess:2005:LDL

[Woe05] Wolfgang Woess. Lamplighters, Diestel–Leader graphs, random walks, and harmonic functions. Combinatorics, Probability and Computing, 14(3):415–433, May 2005. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=14&issueId=03.

Wollan:2006:EFS

[Wol06] Paul Wollan. Extremal functions for shortening sets of paths. Combinatorics, Probability and Computing, 15(6):927–932, November 2006. CODEN CPCOFG. ISSN 0963-5483 (print),

1469-2163 (electronic). URL http://journals.cambridge.
org/action/displayIssue?jid=CPC&volumeId=15&issueId=
06.

Wolfovitz:2011:SGT

[Wol11] Guy Wolfovitz. Sub-Gaussian tails for the number of triangles in G(n, p). Combinatorics, Probability and Computing, 20(1):155– 160, January 2011. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Wu:1997:UBN

[Wu97] Pou-Lin Wu. An upper bound on the number of edges of a 2connected graph. Combinatorics, Probability and Computing, 6(1):107–113, March 1997. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http://journals.

cambridge.org/action/displayIssue?jid=CPC&volumeId=6&
issueId=01.

Wu:1998:VTC

[Wu98] Haidong Wu. On vertex-triads in 3-connected binary matroids. Combinatorics, Probability and Computing, 7(4):485–497, December 1998. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL http://journals.cambridge.org/
action/displayIssue?jid=CPC&volumeId=7&issueId=04.

Wu:2003:CEG

[Wu03] Haidong Wu. Contractible elements in graphs and matroids. Combinatorics, Probability and Computing, 12(4):457–465, July 2003. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=12&issueId=04.

Wegener:2005:OMP

[WW05] Ingo Wegener and Carsten Witt. On the optimization of monotone polynomials by simple randomized search heuristics. Combinatorics, Probability and Computing, 14(1–2):225–247, January 2005. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=14&issueId=01.

Wang:2023:LES

[WW23] Ke Wang and Philip Matchett Wood. Limiting empirical spectral distribution for the non-backtracking matrix of an Erdős–Rényi random graph. Combinatorics, Probability and Computing, 32(6):956–973, November 2023. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/journals/
combinatorics-probability-and-computing/article/limiting-
empirical-spectral-distribution-for-the-nonbacktracking-
matrix-of-an-erdosrenyi-random-graph/64FDC86CE96AE6C22DA20C5CF4209E79.

Xu:2011:BBP

[XY11] Baogang Xu and Xingxing Yu. Better bounds for k-partitions of graphs. Combinatorics, Probability and Computing, 20(4):631– 640, July 2011. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Yehudayoff:2020:PCT

[Yeh20] Amir Yehudayoff. Pointer chasing via triangular discrimination. Combinatorics, Probability and Computing, 29(4):485–494, July 2020. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL https://www.cambridge.org/core/
journals/combinatorics-probability-and-computing/article/
pointer-chasing-via-triangular-discrimination/1792C00D506C602217BB778835A990FD.

Yuster:1997:ITI

[Yus97] Raphael Yuster. Independent transversals and independent coverings in sparse partite graphs. Combinatorics, Probability and Computing, 6(1):115–125, March 1997. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=6&issueId=01.

Yuster:2004:FTD

[Yus04] Raphael Yuster. Families of trees decompose the random graph in an arbitrary way. Combinatorics, Probability and Computing, 13(6):893–910, November 2004. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=13&issueId=06.

Yuster:2007:PCG

[Yus07] Raphael Yuster. Packing cliques in graphs with independence number 2. Combinatorics, Probability and Computing, 16(5):805–817, September 2007. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=16&issueId=05.

Yuster:2012:DGL

[Yus12] Raphael Yuster. Dense graphs with a large triangle cover have a large triangle packing. Combinatorics, Probability and Computing, 21(6):952–962, November 2012. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Yuster:2017:MNS

[Yus17] Raphael Yuster. On the maximum number of spanning copies of an orientation in a tournament. Combinatorics, Probability and Computing, 26(5):775–796, September 2017. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (elec-

tronic). URL https://www.cambridge.org/core/product/
2DD0D54C5486DB2D9410DEBE50CBFE71.

Zahl:2020:CHO

[Zah20] Joshua Zahl. Counting higher order tangencies for plane curves. Combinatorics, Probability and Computing, 29(2):310–317, March 2020. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL https://www.cambridge.org/core/
journals/combinatorics-probability-and-computing/article/
counting-higher-order-tangencies-for-plane-curves/3537675844B65CE8958B7A9B4587D755.

Zhang:2008:LDA

[ZD08] Jim X. Zhang and Paul Dupuis. Large-deviation approximations for general occupancy models. Combinatorics, Probability and Computing, 17(3):437–470, May 2008. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL http:

//journals.cambridge.org/action/displayIssue?jid=CPC&
volumeId=17&issueId=03.

Zhao:2010:NIS

[Zha10] Yufei Zhao. The number of independent sets in a regular graph. Combinatorics, Probability and Computing, 19(2):315– 320, March 2010. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Zhao:2017:LTV

[Zha17] Yufei Zhao. On the lower tail variational problem for random graphs. Combinatorics, Probability and Computing, 26(2): 301–320, March 2017. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic). URL https://www.cambridge.

org/core/product/B678B0FE9950922161CFD66278A45187.

Zorin-Kranich:2017:CQG

[ZK17] Pavel Zorin-Kranich. Corners over quasirandom groups. Combinatorics, Probability and Computing, 26(6):944–951, November 2017. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-

2163 (electronic). URL https://www.cambridge.org/core/
product/60B98255EC52112ED40F387057FA0CA3.